A Complete Bibliography of IEEE Security & Privacy

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

24 May 2024 Version 1.91

Title word cross-reference

#SP18 [Ano17-75].

4 [BS13a]. $80 [GK11]. = [Sch14b].

.NET [For04a, Chu08, TG04]. .onion [Boy16].

0-13-008215-5 [Pfl04]. 0-13-100851-X [For04a]. 0-Day [RSM13]. 024-Bit [GS07].

1 [Ano03a, AAG15, Han12a, Mar15c, Nar13a, SD16a]. 101 [CDFW14]. 16 [SSW21]. 160 [MMKP16]. 18th [TP06]. 19 [PCK21, Pei20, SLKJ21, Sch20a, TXL^{+}21].

2 [AA15, Han12b, Nar13b, VVSL12a]. 2.0 [Dre16, SP07]. 20-Year [LRW^{+}23]. 2003 [Mea03b]. 2004 [Wei05]. 2005 [Gat05, Kei05]. 2006 [TE06]. 2007 [Ano06g]. 2008 [BBB^{+}08, Mic08b]. 2010 [Vig11]. 2014 [Ano13b]. 2016 [Ano15a, Ano15b, Ano16-27, Ben17, CGL^{+}16, SZ17b, SZ17c]. 2017 [Ano17a, BP18]. 2018 [Ano18e, EB19]. 2019 [Ano19a, Ano19b, Ano19-62]. 2020 [Ano20r, YB21]. 2021 [BH22]. 2022 [Ano22-61]. 2023 [Ano22a]. 2024 [Ano23b]. 2025 [May15a]. 2036 [Koh22a]. 20th [BC23a].

39th [Ano15c]. 3rd [Ano15d].

4.1 [DeM15]. 40th [Ano19c]. 42nd [Ano21a]. 4th [Ano16a, Ano19d, Ano19e].

5G [Ber23, RLB^{+}23].

6 [FG08].

800 [MMKP16]. 800-160 [MMKP16]. 802.16 [JW04].

9/11 [GM12].

Aargh [Les13c]. Abandonment [Gee13a]. Abstract [And06a, Gee06a]. Abuse [Ede10, HMA04, MOT^{+}17]. Abuses [AC21]. Abusing [VWC19]. Academia [KJC^{+}12]. Academic [Bre12, FB04a, Nic20b, SDC^{+}17, TA23, YB05]. Academy [BV06, DR05]. Accelerating [ADE^{+}22]. Access [eDIATK22, AR24, Ano20l, Ano21-39, BV11b, Cam03a, Cat10a, Eva14, FKS07, GLM11, HHK18, LBB07, Mas19, NBLC09, NSSS08, SS10, VdCA07, Ano22-60]. Account [Arc04b]. Accountability [MRF^{+}24, Sch09a]. Accountable [Ede14, Lan09b]. Accreditation [KG17]. Achieve [Ano17b]. Achieving [BF07]. Acknowledging [HvO12]. ACM [Pop04, SZ17b, SZ17c, SZ17a]. ACNS [SMGK04]. Acquisition [SMK20a]. Across [BS20a, CZ23, DOL13, In 07]. ACSAC [YB21]. Act [Fid11, MC12, BBD^{+}08, Fra07b, KH12, Sta20]. Acting [PT05b]. Action [Ano14-39, DEKM^{+}21, DMB^{+}22, Wil16a, Win03]. Active [DBD11]. Acts [Mil03]. Actually [Pei19]. Ad [Car08, Ede14, HP04]. Adam [McG08a]. Adapting [KvS14]. Adaptive [GCY^{+}22]. Address [Bel15b, KvS14, O’R17]. Address-Sharing [O’R17]. Addresses [Alf05, SJ03]. Addressing [AC21, Amo16, Ber20a, JP11, PS09, TBBR17]. Administration [vO22a]. Adolescent [Wis18]. Adopting [Ste06, TM05]. Adoption [Ded17, Gup17, JKSM19, Mei18, Sho23]. Advance [CK10]. Advanced [Bai12b, BS14, DR10, LTB15, SE13, Bur03, MS21a]. Advancements [MCW17, MCW19]. Advances [Ano17-56, PB04]. Advancing [Dar14]. Advantage [May15a].

Adversarial [Cel16, DSJK19, DBR22, FOM23, MBA^{+}22, YG19]. Adversary [DFG^{+}11, Gor06]. Advertisement [Ano13b, Ano13c, Ano13f, Ano14a, Ano14b, Ano14d, Ano14c, Ano14n, Ano14i, Ano14j, Ano14k, Ano14l, Ano14m, Ano14o, Ano14w, Ano14p, Ano14y, Ano14-27, Ano14z, Ano14-28, Ano14-29, Ano14-31, Ano14-32, Ano15d, Ano15c, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano15r, Ano15s, Ano15t, Ano15w, Ano15u, Ano15v, Ano15x, Ano15y, Ano15z, Ano15-27, Ano15-28, Ano15-29, Ano15a, Ano15b, Ano15-36, Ano15-37, Ano15-38, Ano15-39, Ano15-41, Ano15-43, Ano16j, Ano16c, Ano16f, Ano16e, Ano16r, Ano16s, Ano16t, Ano16w, Ano16u, Ano16z, Ano16x, Ano16y, Ano16v, Ano16-32, Ano16-29, Ano16-30, Ano16-31, Ano16-41, Ano16-43, Ano16-44, Ano17a, Ano17f, Ano17k, Ano17-29, Ano17-42, Ano17-58, Ano17-64, Ano17-65, Ano17-84, Ano20o, iC15, Ano08d, Ano08f]. advertisement [Ano08a, Ano08g, Ano17-49, Ano17-63, Ano21g, Ano21p]. Advertisements [MRF^{+}24]. Advertising [Cra12]. Advice [Ano17-72, RIC17]. Advisories [LS05b]. Advocates [HLJ21]. Aerial [BKS18]. Aerospace [DDAN12]. AES [Mar10]. Affects [Sch12a]. Afford [MM14]. Afraid [Pfl06]. After [Gif10, MHF19, KH12, McL06, Sch18a]. Again [Les13d, Sch10a]. Against [ACPZ20, LDZ^{+}22, SG18, AH17, BFK16, Bon16, Che06, EGB18, KRK18, KS12a, KS12b, SND14]. Age [ASC15, Bai12b, BGMP19, CP09a, KDC20]. Agencies [Bel14a, CFH05, DOP^{+}21]. Agenda [FWFA23, HKMR24, HvO12, Mul05]. Agent [PGNRV24]. Agent-Based [PGNRV24]. Agents [KT20]. Aggregated [Sta20]. Aggression [Hut24]. Agile [Mar13a, RAAB20]. Agility [CUCD14]. Ahead [Ano17-36, Les15b, LC13, WL18].

AI [Ano17c, Ano17l, Ano17m, Ano20a, Bel24, Don03a, JKE18, Kro18, Nis24, Sch23, SG18, SW18b, SCHC24]. Aid [BB17, JZ04, LGO06]. Ain’t [Don05c]. Air [May15a]. Airplane [Sch03b]. Airport [KTT19, Sas07]. Alan [Ano21-39]. Alchemy [Cyb04b]. Alert [HRTT03, VT12, Vie12d]. Algorithm [EV18, PZTE18, SW18a]. Algorithmic [AWV18, Bin18]. Algorithms [MR14, SZ16]. Alice [Opp17]. Alien [GN07a]. Aliens [Sch18a]. Aligning [DeL14, Yee04]. Alliance [Alf05]. Alliances [Kal12]. Allies [Bel05]. Alone [DFJ^{+}20]. along [Mei18]. Alternative [AS09, Gar03, LS06, MC12]. Always [RP10]. Am [PB07]. Ambitious [Cyb03a]. Amendment [Bel06a]. America [BBD^{+}08, KH12, Ami03, BS20b]. Among [MŽ22]. Amoroso [McG08c]. Amplification [Bla03, Hut24]. Analog [Gra13]. Analogy [Gee08b]. Analysis [ADG23, AMM19, AIKR13, And12, Ano17-59, AL03, ABK^{+}04, AEP^{+}19, BM11, BCG06, CAL^{+}24, CPC^{+}18, CCS06, CM04, CBE^{+}12, HQD^{+}03, HBN09, HFT^{+}19, Her06, JKSM19, KN13, LH07, MBP^{+}24, Opp15, PMA^{+}20, Puz16, RKA15, SLW22b, SGCG^{+}11, SCHC24, TW21, VM04, WHF07, ZOC^{+}22]. Analytics [ADE^{+}22, Ano14y, Ano16-42, Ano17-68, CMR13, CL12, Sch11c, SBA^{+}18, SW18a, VKM^{+}15, ZRM14]. Analyze [SND14]. Analyzing [Hu16, PKBS15, RBBK04a, RBBK04b]. Anchored [Han12a, Han12b]. Anderson [MA07, Spa08]. Android [EOM09, LCJ21, SFK^{+}10, SFE10]. Android-Powered [SFE10]. Angels [JJV^{+}09]. Angle [BLG^{+}15]. Annals [Ano19i, Ano19j, Ano20c, Ano21c, Ano21b, Ano21g, Ano21d, Ano21e, Ano21f, Ano22c, Ano24b, Ano24c, Ano23d]. Annie [McG07a]. Anniversary [BC23a, Vie12c]. Annual [Ano03a, Ano04a, Ano05, Ano06c, Ano08b, Ano09a, Ano15d, Ano15c, Ano16a].

Anomaly [AMM19, Ano17h, SWL21]. Anomaly-Based [SWL21]. Anon [LDK^{+}14]. Anon-Pass [LDK^{+}14]. ANONIZE [HMPS15]. Anonymity [AK03, BPB^{+}04, DMS07, IAM03, Kos15, PKBS15, PRC17]. Anonymization [AR19, Sim21]. Anonymous [AC21, HW13, HMPS15, LDK^{+}14]. ANSI [FKS07, LBB07]. Answer [Her14]. Answers [WCV^{+}04]. Ant [FHMF14]. Ant-Based [FHMF14]. Anthropological [SMO^{+}14]. Anthropologist [Ort14]. Anti [GTG07, Hol15b, Kra04]. Anti-Debugging [GTG07]. Anti-fragility [Hol15b]. Anti-Honeypot [Kra04]. Antiphishing [ACPZ20]. Antón [McG07a]. Any [Bel14a, Pei22]. Anycast [AS09]. Anyone [Ros07]. Anything [Ano18-29, O’D08]. Anyway [Don04e]. APIs [GS16]. Apocalypse [Don03c]. App [BC14, MHF19]. Apple [Ano17d, KS12b]. Applicants [Ano23-39]. Application [Ano17-74, CP10a, CA06, IKP^{+}07, KCL^{+}23, ME10, Mei06, NP07, Roe17, Sho22, Ste08, Tho05, WBT^{+}18]. Application-Aware [IKP^{+}07]. Applications [Ano15c, Ano17u, Ano19l, Bos16, Fet16, GL04, JSS20, LWL^{+}21, Mar15c, Sas18, Sty04c, Whi03, Ano19m, Ano20d, Ano20e, Ano21j, Ano21h, Ano21i, Ano22e, Ano22f, Ano24d]. Applied [DM15a, SVR^{+}19, GNP05]. Apply [Cal04f]. Applying [Cha12, Her06, KS06, LBF^{+}19, VE06]. Apprentice [IN10]. Approach [AA12, AC21, BLCDK11, CSSvdH14, Dia11, DGU^{+}12, KWRK13, Kos15, LS06, LS05b, Mic22, SLM^{+}23, SMO^{+}14, Yan16]. Approaches [OMB17, OY16, RRAK14]. Approaching [Lev04a]. Appropriate [WC04]. Apps [Ano17-66, BBS19, MA15a, RSOD22]. Arce [McG10f]. Architectural [ASK07, HWF^{+}10, PEC^{+}14, RKA15]. Architecture [Ber21b, DOP^{+}21, Dia11, Pet10, Sch09c, JB05, VVYY11]. Architectures [BMMM10, MGH^{+}15, NS11]. Areas [BKS18, vO22b]. Aren’t [Nan09]. Argument [DMB^{+}22]. Ariadne [FD10]. Armed [Mic19]. Arms [McC03b, Pei20, Som04]. Army [KW12]. ARP [Zdr09]. Art [And07, BGG^{+}22, BLM17, Sen17, Sty04a]. Articles [Ano19r, Ano19-36, Ano19u, Ano19v, Ano19-27, Ano20i, Ano20j, Ano20k, Ano20m, Ano20n, Ano21t, Ano21u, Ano21v, Ano21w, Ano21-29, Ano21-30, Ano21-31, Ano22m, Ano22n, Ano22s, Ano22t, Ano23j, Ano23k, Ano24j]. Artifact [Her22]. Artifacts [Ben23]. Artificial [AACG19, Ano23r, Ber21a, Far21, FOM23, Lan08b, MBP^{+}24, Mic21, Mic22, MMR24, Raa19, Sch18b, TCC^{+}24, vO22a]. Arun [For04a]. Arya [LRKR18]. Ashcroft [Les03c]. Asia [SAB^{+}19]. Ask [Pfl06]. Aspects [San21]. Assertion [HJK08]. Assessing [BGLEP15, HKN^{+}09, Hoo16, KLL17, XL16]. Assessment [BCD^{+}15, CA06, HKN^{+}09, HN10, PR08, SFK^{+}10, Ste08, TW21, WTML17]. Assessments [Hoo16]. Assist [AEP^{+}19]. Assistant [LWL^{+}21]. Assisted [ECG^{+}07, MA20]. Assumptions [KZZ17, SP09, Oeh05]. Assurance [Bar15, BBB^{+}08, BF08, BC22, BGM^{+}06, Blo13, CW12, GHK^{+}06, KMP^{+}11, KR10, Lan07a, LJZ12, MA12, Mea03b, MLM^{+}07, OC12, RWD03, SFH05, SFH06, WL11, ITA05]. Assured [May15a]. Atlantic [In 07]. ATM [HMKT07]. Attack [AK03, Ahm08b, Arc04b, Bel16a, BCM^{+}15b, BCLM09, BGL^{+}23, CCR09, CRBM06, Gee11a, GAA04, Gro08, GW07, JJV^{+}09, KS13a, Por09, PRCC10, RVK05, Sch18b, Sto04, Wil24, Zan09, Zdr09]. Attack/Defense [Sch18b]. Attacker [CP09b, Her13]. Attacker-Targeted [CP09b]. Attackers [AAG15, CLS07, Zur23]. Attacking [AM04, Hu16]. Attacks [ACPZ20, Ano17e, Ano21-41, AS09, AH17, BGC^{+}14, Ber21a, BFK16, Bil10, BO14, BKMN09, DSJK19, DBR22, Der13, DDAN12, EHK^{+}04, GY21, GHG14, HQD^{+}03, Hay13, HMR06, JS07, Jak19, JFK21, KS12b, LPS^{+}23, Law09, Mil11, O’D08, Per07, Pie23, PvO24, RSM13, SND14, SG20, Syt05, TA23, VWC19, YE09, Zdr09, CA05, Lev05, Sch05a]. Attaining [Mic22]. Attractors [GC08e]. Attribution [Gee17, Gro20]. audience [Ano18-28]. Audit [Mar15a]. Auditing [BPVS04, JZ04, KAAEa17, RRAK14, TCJ^{+}19]. Audits [BFS^{+}21, LS12, NL23]. Augmented [OGM^{+}24]. Augmenting [CAL^{+}24, LRKR18]. Australia [BW12, HDB08, ST06]. Authenticate [HM12]. Authenticated [MR14, Smi11]. Authentication [AS22, ACAT^{+}15, BCE^{+}12, Boy16, DeF11, FD11, Gar03, GHS14, GU13, Hay13, HLS^{+}21, HKW06, JCM12, Lys07, MR03, May15b, OdH12, PWVT12, RS11, Sch05c, SWCS21, SE09, SCMS18, TGC16, WKB08, WDL21]. Authority [Pet19]. Authorization [CEK^{+}10, DK10]. Authorship [Ano17e, KF13]. Auto [Gee21a]. Auto-Update [Gee21a]. Automated [BDTJ19, BD18, CRBM06, Coh10a, DRS16, HFT^{+}19, KTT19, McC03a, PS23, RVK05, SyPKZ22, Ste15, Tom16, VM21, WHF07, ZKH^{+}23, ZRM14]. Automatic [KS13a, LBF15, VNC^{+}06, SK05]. Automating [JKE18, OJR^{+}23]. Automation [AA12, Kam14]. Automotive [TW21]. Autonomic [BPVS04]. Autonomous [KT20, MV22, Mic19, NTMD^{+}18, SBB^{+}18]. Autonomy [Blo15]. Auxiliaries [Mar13b]. Availability [O’R17, KN13]. Available [Les03d]. Aversion [Gee12d]. Averting [PT05a]. Avi [McG06a]. Aviation [DMS^{+}15, SLM^{+}23]. Avoid [EJ06].

Avoiders [Ede10]. Avoiding [BS13c]. Award [Ano15g, Ano15u, Ano16j, Ano16u, Ano16v, Ano17-42, Ano17-41, Ano15a, Ano15b, Ano17a, Ano17-43, Ano19f, Ano22a, Ano23b, Ano23-66]. Awards [Ano16c, Ano17k, Ano17i, Ano17j, Ano18-40]. Aware [BCD^{+}15, BKS18, IKP^{+}07, NBLC09, PMNT12]. Awareness [AS15, CPFJ14, CW08a, CW09, LTB15, MBA^{+}22, vWS06]. Away [Lan10b]. AWK [McL05b].

B [Ano17a, Sch17b]. B. [Ano19f]. Baase [Pfl04]. Bace [McG07e]. Back [FA07b, FB04b, KS08b, SEKR18, Sho22, Vie12b, Ano17-75, Ano18-43]. Backdoors [LDZ^{+}22]. Backwards [LRW^{+}23]. Bad [Av18, Arc05a, BKS13, Les03e, Sch19a, TWC^{+}15, VT12]. Badly [PPHB08]. Bake [Boy16]. Balance [Sch18b, SAA04]. Balancing [GG04, Mic10a, Red24, San21, Sol07]. Ballot [Cha04b, SCM08, vdG17]. Bandwidth [GHG14]. Bang [Arc04b, Lan03]. Banking [HKW06, HMT06]. Barken [Sty04b]. Barrier [JR14, McG13b]. Barriers [CPS^{+}16]. Bart [McG14a, McG15a]. Based [eDIATK22, AZ20, Ano16-43, Ano17-67, AR19, BS13b, But17, BMMM10, DK10, DXS21, Dun10, El 10, FKS07, FHMF14, Gar03, GM09, GY13, GSLA15, GLM11, Han12a, Han12b, HPSR18, KP10, Kos15, LKL^{+}19, LBB07, LL11, Mar15c, Mar08, NBLC09, NKJ^{+}14, PGNRV24, PvDS08, Por09, RBB19, RSM13, SCM08, SMJM11, Sen17, SPG^{+}19, SWL21, SW12, SBF^{+}15, WDL21, Wil16b, Ada05, EHK^{+}04, GSHU08]. Basic [CCR09, Chu08, Hen09, How09a, How09b, YZA08]. basics [Mar05]. Basing [FTCS12]. Battery [JMD06]. Batteryless [TJS^{+}22]. Battle [Les06, Mil03, NTMD^{+}18]. Battlefield [Gra13]. Battleground [Gif10]. Bazaar [McK07]. Be [Ano15v, Ano16w, Ano16x, Ano16y, Ano17-39, Ano17-72, Ano18u, BCM^{+}15b, CBOS20, Eps12a, FA07a, FA07b, Gee15b, Les03d, Les06, MM11, Mos18, Ohm14, Sch06b, TP11]. Beacons [TJS^{+}22]. Beast [Dun10]. Beat [Gee20]. Beautiful [Gee11f]. Becky [McG07e]. Become [Lev04b]. Becoming [How08]. BedRock [MSeF^{+}22]. Beef [RBB19]. Before [NCA19, Ost20]. Behaving [PPHB08]. Behavior [HLJ21, HWF^{+}10, LXS22]. Behavioral [Acq09, AMM19, Bor13b, Cra12, Far20, Lan09a]. Behavioral-Anomaly [AMM19]. Behaviors [McL13]. Behind [Ano17f, FRA10]. Being [YRX^{+}22]. Bellovin [McG13e, McG15e]. Benchmarking [vdKHA^{+}20]. Bendy [Sas07]. Benefit [GC09a]. Benefits [Ano20-34, Ano21-60, Ano21-61, Ano21-62, Ano21-63, Ano21-64, Ano21-65, Ano22-49, Ano23v, Ano23-34, Ano23-35, Ano23-36, Ano23-37, Ano24t, Ano24u, Fel03, JKSM19]. Best [Ano15d, Ano16a, Don04f, MDR^{+}20, RP09, Ano17-49]. Best-Seller [Don04f]. Betrayal [Kal12]. Better [Ano16i, Ano16-39, Ano16-40, Ano17-60, Ano17-64, Ano17-61, Ano17-62, Bor15, CB14, EV18, EJ06, FF10, HAO^{+}22, Law16, MDR^{+}20, Ano17-63, dAMM13]. Between [HAO^{+}22, MM10, Ksh14, Law16, MK13, vWM05]. Beware [GC08b]. Beyond [AS22, BS14, BF05a, BDL^{+}14, Gor06, HKMR24, JL18, Mar08, PB04, RS18, Sty04b, Win03, Woo23, ZS19, JR14, RLB^{+}23]. Beyond-5G [RLB^{+}23]. Bias [DMB^{+}22, KTT19, Smi12b]. Biblio [Don04a, Don09a]. Bickering [LBS09]. Big [Ano14y, Ano18c, Ano19-28, Ano19-29, Ano19-30, Ano19-31, Ano19-38, Ano20-48, Ano21-33, Ano21-34, Ano22z, Ano24k, CMR13, ES14, Gee11f, LSG^{+}18, Les13a, SW18b, ASC15, Ber23, Lan03, Les13a]. Bigger [Les04a]. Bill [BBB^{+}23, Bel15b, TAGM23, McG08b]. Billion [GK11]. Bills [ZLT^{+}23]. Binary [Ano17-71, Bos16]. Bio [Sag13].

Bio-hacking [Sag13]. biobanking [VB13]. Biology [Ber20a]. Biomanufacturing [GCY^{+}22]. Biometric [AMF15, AFFOG12, LZ16, May15b, PPJ03, RdRdCG10, Sas07, Web05]. Biometrics [MR03, Way08]. Bishop [McG08f]. Bit [Ano17-59, GS07, MOG^{+}20]. Bitcoin [TMU23, GKCC14, GCR18]. Bites [Don03a]. Bits [SHB20, Way04]. Black [Gee07, Sty05a]. BlackBerry [HK09]. Blacklisting [BS14]. Blakley [McG09b]. Blaming [Ren12]. Blankets [Les11c]. Blaster [BCJ^{+}05]. Blended [ACAT^{+}15]. Blessing [MP21]. Blind [Arc05b, DGU^{+}12, SLW22a]. Blindfold [Nac16]. Blink [Sas07]. Bloatware [McD12]. Blockchain [CIL^{+}21, DP18, HHK18, HV20, KC18, KNL22, RBB19, SJZG19]. Blockchain-Based [RBB19]. Blockchains [van20]. Blog [ANC10]. Blue [Dun10]. BlueBag [CMZ07]. Blues [Don10a]. Bluetooth [CMZ07, Dun10, TJS^{+}22]. Bluff [DGU^{+}12]. Blunderbuss [SC04]. Blur [OHBS14]. Board [Ano14o, Pfl15d, Ano03f, Ano19-58, Ano19-59, Ano19-60, Ano20-35, Ano22-55, Ano22-56, Ano22-57, Bel12b, Pal04]. Boards [Ano19-52, Ano19-53, Ano19-54, Ano21-67, Ano22-52, Ano22-54, Ano23-41, Ano23-42]. Boats [Gee11c]. Bob [McG09b, Opp17]. Body [NRH^{+}22, RDC^{+}18, TG07, vO21a]. Bolt [CB15]. Bolt-On [CB15]. Book [Don05b, For04a, Les06, Pfl04, Pfl09, Sty04a, Sty04b, Sty08]. Books [PP06a, Spa06]. Boot [FF10]. Bootstrapping [TP08]. Border [MJF07, Mac06, McL08]. Borders [PGT07]. Bot [MKKP09, Hol05]. Both [Bel11a]. Botnet [Ano17h, PH18, SGCG^{+}11]. Botnets [AJV18, AvEB15, Sch06a]. Bots [KO08, Hol05]. Bottom [Smi11]. Bought [MB16]. Boundaries [PGT07]. Bounding [BV15, DM15b]. Bounties [BDTJ19]. Box [BCGN16, DBR^{+}08, Mic10b, Vét21]. Brain [Koh22b, Sch08b, Sch09d]. Brazil [Ano17q, Ano17r, Av18, BA18, SEKR18, SBA^{+}18]. Brazilian [CPC^{+}18]. Breach [BH19, Hei16, ZS19]. Breaches [ABAGOB10, Eva15, Ren12, Tel15]. Break [Les09a]. Breaking [BGG^{+}22, Sch13a]. Breaking-in [Sch13a]. Breaks [GCR18]. Breakthrough [San14]. Brian [McG15b]. Bridge [DOL13]. Bridges [Jae23]. Bridging [BLCDK11, TMGP13, vWM05]. Briefs [And03b, And04, DK05, DFK05, DFK06, Fig09a, Fig09b, Fig09c, Fig09d, Fig10a, Fig10b, Gar11a, Gar11b, Gar11c, Gar11d, Gar11e, Gar11f, Gar13a, Gar13b, KDF06, KD06a, KD06b, Law10a, Law10b, Ort06a, Ort06b, Ort07a, Ort07b, Ort07c, Ort07d, Ort07e, Ort07f, Ort08a, Ort08b, Ort08c, Ort08d, Ort08e, Ort09a, Ort09b, Pau10b, Pau10c, SAA04]. Brigade [Don08a]. Bright [BBCL13]. Bring [Cyb04a]. Bringing [IAM03, PHR20]. British [Gro14]. Brittleness [Bel06c]. broadcast [Les05]. broader [Ano18-28]. Broken [Ahm08b, GCR18, HLS^{+}21]. Brokerage [VdCA07]. Brokered [SMJ^{+}19]. Brokers [OAB07]. Brother [ASC15, Les13a]. Brought [Ras11]. Browser [Bon16, JS07, PH18]. Browser-Sniffing [JS07]. Browsers [FVJ19]. Bruce [Sty04b, McG18c]. Brute [JR14]. Brute-Force [JR14]. Buddy [vZH^{+}19]. Budgets [CHMO07]. Budweiser [Ras11]. Buffer [PB04]. Buffet [KCL^{+}22]. Bug [Arc04b, BDTJ19]. Buggy [FLDGB19]. Bugs [BDL^{+}14]. Build [Ano14-39, DLMS21, OP15]. Building [Ada05, BFC^{+}17, BPS16, BC08, CP09a, CCLS21, DOL13, DBD11, DCC^{+}09, EJ06, Fet16, FF10, GBK09, HRMS21b, HM13, HBT12, How04, KMP^{+}11, Knu07, KSL^{+}20, Law16, MS20, Mas19, MV22, Mea03a, MH13, Pet09, PEC^{+}14, SP09, Ste08, TKZ21, WL11]. Built [CR20, JJ13]. Built-In [CR20]. Bullet [McG06d, McG07d, McG07c, McG07e, MA07, McG09f, McG10f, McG09a, McG06a, McG06c, McG06b, McG07a, McG07b, McG08d, McG08c, McG08e, McG08a, McG08b, McG08f, McG09d, McG09e, McG09b, McG09c, McG10a, McG10b, McG10e, McG10d, McG10c, McG11e, McG11a, McG11d, McG11b, McG11c, McG11f, McG12a, McG12e, McG12b, McG12f, McG12d, McG12c, McG13d, McG13e, McG13f, McG13c, McG13h, McG13g, McG14b, McG14c, McG14f, McG14d, McG14a, McG14e, McG15b, McG15f, McG15d, McG15c, McG15a, McG15e, McG16e, McG16b, McG16a, McG16d, McG16c, McG17e, McG17d, McG17a, McG17b, McG17c, McG17f, McG18f, McG18d, McG18c, McG18e, McG18b, McG18a, Ran16]. Bullying [Sas15]. Business [Gri04, Law16]. Businesses [Ano17-38]. Butler [McG16b]. Butterfly [Len03]. Buy [Max20, PLW07]. Buying [Phi19]. Buys [Sch08b]. Bypassing [DeM15]. Bystander [CDJS^{+}24, Fla16]. Bystander-Centered [Fla16].

C [AT13, Sea06, vO21b, vO23b]. C-Suite [AT13]. CA [Ano20z]. Cache [DXS21]. Caches [DXS21, LLGJ16]. Caernarvon [KMP^{+}11]. Café [KCL^{+}22]. CAI [Ano23r]. Calculating [Mus21]. CALEA [Gid06]. Calendar [Ano06a, Ano06d]. Calendars [Ano15-29]. California [Kos15, Sta20]. Call [Ano08c, Ano15e, Ano15f, Ano15g, Ano15u, Ano16b, Ano16c, Ano16f, Ano16e, Ano16d, Ano16g, Ano16h, Ano16u, Ano16v, Ano16-44, Ano17a, Ano17k, Ano17i, Ano17j, Ano17o, Ano17n, Ano17l, Ano17p, Ano17m, Ano17q, Ano17r, Ano17-42, Ano17-41, Ano18b, Ano19r, Ano19s, Ano19-36, Ano19u, Ano19v, Ano19w, Ano19-27, Ano20a, Ano20-27, Ano20-28, Ano20i, Ano20j, Ano20k, Ano20m, Ano20n, Ano21-66, Ano21-48, Ano21-49, Ano21-50, Ano21-51, Ano21t, Ano21u, Ano21v, Ano21w, Ano21-29, Ano21-30, Ano21-31, Ano21-35, Ano22-35, Ano22-40, Ano22-36, Ano22-37, Ano22-38, Ano22-39, Ano22m, Ano22n, Ano22s, Ano22t, Ano22u, Ano22-27, Ano23x, Ano23b, Ano23y, Ano23j, Ano23k, Ano24v, Ano24s, Ano24w, Ano24j, Hor19, IP10, Pei20, SNS10, Win03, VWW05]. Call-Filtering [SNS10]. Callas [McG14b]. Caller [Les14a]. Camouflage [JS07]. Camp [FF10, McG15d]. Campaigns [CGC19, Gra23]. Campus [Ano06b]. Campuses [HNE^{+}08]. Can [ACL07, Ano16i, Ano17d, Ano17-38, Ano17-72, Bar15, BH19, Bin18, Bor15, Cra12, Eps12a, GM14, GSLA15, Kau10b, Kau10a, KNBV17, Lan16, MM14, Mar13b, MJP21, MOG^{+}20, Pal04, Pay04, Ros07, RXJ^{+}21, Kri13]. Canadian [Fra07b]. Canal [SE15]. Candidates [Ano18p]. Candy [Way04]. Candy-Coated [Way04]. Canning [PB05]. Cannot [Cal03e]. Can’t [Cyb06b, Kni17]. Capability [Bab16, Ded17]. Capital [Bac13]. CAPTCHA [YE09]. CAPTCHAs [KO08]. Capture [Irv11, MBC^{+}18, Vig11]. Capture-the-Flag [Irv11, MBC^{+}18]. Card [Ano17w, CL07, KST^{+}21]. Cards [LA10, PWVT12]. Care [Les13b, RC06]. Career [Ano15-28, Ano23s, Ano23z, Ano24q, Ano24r, Ano24x, MH13, Ano17b]. Careful [Eps12a]. Carhart [McG17d]. Carl [BC23a]. Carterfone [Les10b]. Case [BHJM19, BF07, CCF^{+}20, CPS^{+}16, CW09, CB15, DeF11, DBR22, GS07, Gui11, Gup17, HMT06, Kap07, Kni17, Lev06, LLT14, Mar15b, McL13, MDD^{+}15, NRH^{+}22, PA03, PRS09, Pow07, Sty03, YE09, vO23b, YB05]. Cases [BGM^{+}06, De 24, HMA04, Kni15, MCW08]. Castle [FB04d, Pfl12a]. Categorizing [LGO06]. CATS [SOG^{+}19]. Causal [Far20]. Causation [Gee11b]. Caused [Bye04, PH18]. Causes [Ren12]. Caution [SW03a]. Caveman [Per07]. CCS [SZ17b, SZ17c, SZ17a]. CCTV [KvS14]. CD [Cha04a]. CD-ROM [Cha04a]. Cell [VB13]. Cellular [Ber19, Tra09]. Censorship [DRS16, Les03b, PEL^{+}18, Sel11, DA05]. Census [NH23]. Center [Ano15v, Ano16w, Ano16x, Ano16y, Ano18u, Ano23s, CDF^{+}12, LPS^{+}23, McG14e, Ano23z, Ano24q, Ano24r, Ano24x]. Centered [Fla16, MM09a]. Centers [SDC^{+}17, YB05]. Centric [BHRR07, NSLW20]. Century [BC22, TP06]. Ceremonies [HLS^{+}21]. Certificate [MB16, SDO21]. Certification [CN13, FB04a, HRMS21b, MHRS19, MT23, MBA12, Sun16, Vét21, NRH^{+}22]. Certified [Opp07]. Certifying [SK04]. CFP [Ano22-50, Ano22-64, Ano23-27]. CG&A [Ano19k]. CGA [Ano22d]. Chaff [All07]. Chain [AK15, DXR^{+}23, EW22, FWFA23, LPS^{+}23, Lev03b, MW23, MTA23, RSWO18, Wil22, Wil24]. Chaining [Val12]. Chains [MT23, RALS23]. Chalice [Nis24]. Challenge [Ano17c, Ano17l, Ano17m, Ano17-59, GMB12, GHS14, Hag09, Hav13, Jus04, MC15, NCA19, NTMD^{+}18, PZTE18, Tho18, Tzi18, AF16, SAF15]. Challenge-Question [Jus04]. Challenges [AZ20, AMF15, AM12, AS22, Ano17-36, Ano17-56, Ano17-73, AKN12, BGMP08, BBB^{+}23, BV11b, Bos16, BV15, CR06, CP10b, DSEB12, DD08, DMS07, EW22, FREP17, HBN09, HHK18, HRMS21b, HRMS^{+}21a, JS10, JSS20, KM22, Kau10b, KA13, Lan07b, LXS22, LLGJ16, LAYG16, MN21, MJP21, MM09b, MS21b, NSLW20, OZ09, PGNRV24, Raa19, RBE03, RI12, RLB^{+}23, RRAK14, SS04, SMK^{+}20b, TVR23, TJA10, Tra09, TBBR17, Wai16, WK05, YAZ21, vS19]. Chance [JS10]. Change [CW08a, CW09, Hal10, HLJ21, Les03e, SHH^{+}21, WCV^{+}04]. Changin [GC13]. Changing [Lan05b, Sch12a, VLOS18]. Channel [Law09, PvO24]. Channels [HPSP10, SS19b]. Characteristics [EEA^{+}23]. Characterization [BS13b].

Characterizing [MP22, Sav17]. Charge [Don08a]. Charging [SSJ^{+}23]. Chasing [Gru16]. Chasm [DOL13]. Chat [BC23a]. ChatGPT [Gra23]. Chattering [Dif06]. Cheap [MP10, Sho23]. Cheat [CC11]. Cheaters [Sch11a]. Cheating [PRS09, YR09]. Check [Vie11a, Zan09]. Checklist [Bel08a]. Checkpoints [KTT19]. Chess [BCG06]. Cheswick [McG08b]. Chicken [Les03a]. Chief [Ano21-89, Ano14-30, Ano22-87, Ano23-38, Ano23-39, Ano24-38, Nic21a]. Child [WMDX23]. Children [Gee15a]. Chill [Mei04]. China [Ksh14, WMDX23]. Chips [KDC20, Man13]. Choice [Cat10b]. ChoicePoint [OAB07]. Choices [Lan14c]. Choosing [Gee11g]. Chris [Irv11]. Christma [CCW03]. Christmas [GC08a]. Christofer [McG10a]. Chrome [SDO21]. CIPHER [SSW21]. Circumvention [BKS13]. Cities [HRMS^{+}21a, VM21]. Citizens [LXS22, MŽ22]. Civil [DMS^{+}15, San21]. Civilian [WZE18]. CLARK [TKZ21]. Clarke [GA04b, McG10d]. Classic [KRK18]. Classics [Don04b]. Classification [VVSL12a]. Classifiers [CPC^{+}18]. Classroom [HCRS18]. Classrooms [BS24]. Clean [MC12]. Cleaner [TR09a, TR09b]. Cleaning [Dre16]. Clear [Bel08b, Bel09, BC09, Gee08a, Sch04c, Sch09c, Sch09d, WCV^{+}04]. ClearText [Gee09a]. Click [Ksh10]. Client [FD11, RI12, WKB08, Wu10, ZOC^{+}22, IS05]. Client-Side [RI12, ZOC^{+}22]. Clients [JS07]. Clinic [Bis10a]. Clinical [CRK^{+}13]. Clinicians [Hec11]. Cloaking [ZOC^{+}22]. Clock [Pei22]. Close [DM15b, RC17]. Closed [Cow23]. Closer [Pau05]. Closing [PM14]. Cloud [AAC^{+}17, Amo13, Ano14-28, Ano14-29, Ano17z, BHP24, BCD^{+}15, Blu10, CEK^{+}10, FB11, GA10, Gre13, GWS11, GHR^{+}10, HPSP10, Her19, Kau09, Kau10b, KAAEa17, Les14b, Mar15c, Mar13a, PP15, Pau10a, Pet10, RBB19, RRAK14, SCZ^{+}13, SBA^{+}18, Spr11a, Spr11b, SLS18, Sun16, Sur16, TJA10, TCJ^{+}19, Vie12a]. Cloud-Based [Mar15c]. Cloud-Enabled [Mar13a]. Clouds [Bel11a, Les12a, LWLL10, Mic09]. Cloudy [JS10]. Club [BCG06]. Clue [Vie12d]. Clues [Wil19]. Co3 [Ano14b]. Coast [Les11c]. Coated [Way04]. Code [Ano17-71, BMM07, CCS06, FG22, How06, KTS18, KF13, Loc09, MPM22, Pot10a, RW21, RAAB20, Sag13, Sch20b, Sen17, SW03a, ZRM14]. Code-Based [Sen17]. Codebreaking [RS18]. Coding [Sea06, YRX^{+}22]. Coevolution [vO21a]. Cognitive [SS19a, Smi12b]. Coherent [TPPM07]. Collaborating [SDC^{+}17]. Collaboration [RAAB20]. Collaborative [AA12, CHK^{+}05, Che11, SSW21]. Collection [CP10b]. Collective [Cha12, DEKM^{+}21]. Colleges [SFH05, SFH06]. collide [GNP05]. Colluding [MA15a]. Collusion [Lan03]. colorless [Kri13]. Column [Ano22-66, Ano22-88, Koh22e]. Combat [Der13, HB22]. Combating [Cur06, GG11a, GMC^{+}08, WYA06]. Combining [HvO18]. Come [JJ13, RXJ^{+}21, Sty08]. Comes [DFJ^{+}20, KNBV17, McD12]. Coming [Lib05]. Commands [SWCS21]. Comments [FKS07]. Commission [Bre12, Hoo16]. Commitment [Ano17-84]. Common [FOM23, Hea04, Kal12, KS06, MSR06, Mic08b, NSA^{+}23, TG07, Wil16a]. Communicating [CFN13]. Communication [Kap07, Mar07, OP15, OY16, Pop04]. Communications [Bel06a, BBD^{+}08, BBCL13, GN06, Gro14, Kot08, KYEV^{+}18, KNL22, May10, Ree16]. Community [Ben14, Ben16a, BPS16, De 14, DBD11, FB04c, HW13, HB22, Pei20, SFH05, SFH06].

Comp [Ano22-65]. Company [CR20]. Comparative [MBP^{+}24]. Comparison [NS11, WMDX23]. Compasses [DeL14]. Competency [HM13]. Competing [LBS09]. Competition [Ano16a, Bur08, Lan07c, Sch22b, Ano15d, Vig11, WWH10]. Competitions [BLG^{+}15, CBN11, Eag13, PGP16]. Competitive [SLM^{+}23]. Competitor [AF16, SAF15]. Competitors [MM10]. Complementary [Acq13, Dia11]. Completely [Ano22-87]. Completion [Ano17y]. Complex [ADE^{+}22, Ano17c, Gar14e, Bor05]. Complexity [Bel07a, DFJ^{+}20, Gee08a]. Compliance [DeL14, DPP15, HV20, RXJ^{+}21, Sch19a, Sch04e, SDKM20, BLR19]. Composition [LBS09]. Compositional [DFG^{+}11]. Comprehensive [SFK^{+}10]. Compromise [Ahm08b, HPSR18, LB08]. Compromised [RBBK04b]. Compromises [WAF11]. COMPSAC [Ano18e, Ano20r]. Computation [ABPP16, BB17, BS13c]. Computational [Ano17u]. Computations [Ano17z, Nac16]. Computer [AIKR13, All05, AW04, Ano06b, Ano14i, Ano15w, Ano15u, Ano15v, Ano16j, Ano16w, Ano16u, Ano16z, Ano16x, Ano16y, Ano16v, Ano17-42, Ano17-41, Ano17-40, Ano17-43, Ano18u, Ano18q, Ano18r, Ano18s, Ano18v, Ano18t, Ano18w, Ano18-41, Ano18-42, Ano19m, Ano19l, Ano19-46, Ano19-52, Ano19-47, Ano19-53, Ano19-48, Ano19-55, Ano19-49, Ano19-54, Ano19a, Ano19-56, Ano19b, Ano19-50, Ano19-51, Ano20d, Ano20e, Ano20-29, Ano20-30, Ano20-34, Ano20-31, Ano20-27, Ano20-32, Ano20-28, Ano20-33, Ano20l, Ano21j, Ano21h, Ano21i, Ano21-54, Ano21-60, Ano21-55, Ano21-61, Ano21-48, Ano21-56, Ano21-62, Ano21-49, Ano21-57, Ano21-63, Ano21-53, Ano21-50, Ano21-58, Ano21-64, Ano21-52, Ano21-51, Ano21-65, Ano21-59, Ano21x, Ano21y, Ano21z, Ano21-27, Ano21-28, Ano22e, Ano22f, Ano22-49, Ano22-35, Ano22-43]. Computer [Ano22-40, Ano22-36, Ano22-44, Ano22-45, Ano22-42, Ano22-37, Ano22a, Ano22-46, Ano22-38, Ano22-47, Ano22-41, Ano22-39, Ano22-48, Ano22-59, Ano22o, Ano22p, Ano23v, Ano23-38, Ano23x, Ano23-31, Ano23-34, Ano23b, Ano23t, Ano23-35, Ano23u, Ano23w, Ano23-36, Ano23-32, Ano23y, Ano23-28, Ano23-37, Ano23-29, Ano23-30, Ano23z, Ano23-27, Ano23-39, Ano23-33, Ano24d, Ano24y, Ano24t, Ano24v, Ano24-29, Ano24z, Ano24s, Ano24-28, Ano24-30, Ano24x, Ano24-27, Ano24w, Ano24u, Ano24-31, Arc07b, AQP^{+}23, BBC^{+}19, Bis03, BF06, BLCDK11, Cal03a, CF14, DBD11, Du11, Eag13, Far21, GM12, KST^{+}21, Len03, OBS^{+}20, Pei22, Pop04, RBK23, RC06, SMGK04, ST06, Smi03b, For05, OR05, SPH^{+}05, Smi05b, Ano18x, Ano19-61, Ano19-63, Ano20s, Ano22-60, CSZ^{+}14, Hor14b, RDM^{+}14]. Computer-Security-Oriented [BBC^{+}19]. Computers [Ano15c, Ano24e, Don11a, Mos18, Pfl04, Ano20p, Ano21-35, Ano22-27, Ano24l]. Computing [AAC^{+}17, Ano18f, Ano19n, Ano19o, Ano20f, Ano20h, Ano20g, Ano21k, Ano21l, Ano21p, Ano21m, Ano21n, Ano21o, Ano21-66, Ano21r, Ano21q, Ano21s, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22q, Ano23e, Ano23f, Ano23-40, Ano23g, Ano24f, Ano24g, Ano24h, AR15, Baa03, BGMP08, CDS08, Fel03, FB11, GA10, GWS11, GHR^{+}10, HSS11, Kau09, LRW^{+}23, LLGJ16, PCMN08, Pot09, RS18, SCZ^{+}13, Sch07a, SA12, Sma23, Spr11a, Spr11b, SLS18, TJA10, Thi06, Trč11, Vie11b, IS05, OR05, Ano18d, Ano19g, Ano19i, Ano19j, Ano19q, Ano19u, Ano19v, Ano19w, Ano19-32, Ano19-33, Ano19-34, Ano19-35, Ano19-39, Ano20b, Ano20c, Ano20m, Ano20n, Ano20q, Ano21c, Ano21b, Ano21g, Ano21d, Ano21e, Ano21f, Ano21-29, Ano21-30, Ano21-31, Ano21-36].

Computing [Ano21-37, Ano21-38, Ano22b, Ano22c, Ano22s, Ano22r, Ano22t, Ano22u, Ano22-28, Ano23c, Ano23w, Ano24a, Ano24b, Ano24c, Ano24j, Ano24i, Ano24m]. Concentration [GC09c]. Concept [SOG^{+}19]. Concepts [Wha12]. Conceptual [CW08a, CW09]. Concern [Whi15]. Concerned [TXL^{+}21]. Concerns [ASC15, AEY10, Der03, ES14, PPJ03, Wei04]. Conference [Ano06g, Ano13b, Ano23r, Ben16a, SMGK04, Ano15c, Pop04]. Confidence [BRST17, BRS^{+}21]. Confidential [DK10]. Confidentiality [Les13b, PP15, PBW^{+}08, ZHC^{+}20]. Confidentiality-Preserving [ZHC^{+}20]. Configuring [KR10]. Conflict [KMP20, Mic19, AS13]. Conflicts [MK13, NH23]. Conformist [Cal04a]. Confronting [MA20]. Confused [Ahm08a]. Congress [Ano19-62]. Connect [Ano18g, Ano18-38]. Connected [Ano18-45]. Connections [PRCC10]. Connelly [McG12d]. Consensus [Wei04]. Consent [Bel18a, Bor15, JKE18]. Consequence [Gla19]. Consequences [EPH12, Ess17, Pee13, SS19b, Tro06, VWC19, VVSL12a]. Consider [Ive05]. Considerations [AC21, Cur06, Hei07, Sch07d]. Considered [Gee21a]. Considering [HDB08, Sty04c, WA07]. Constraining [Cal04c]. Constraints [JMD06]. Construction [Rog16, Ada05]. Constructive [Irv03]. Consultants [Mat12]. Consumer [Ano04c, Ano17v, MRF^{+}24, Sta20]. Consumers [DEKM^{+}21, Sch03e, ZS19]. Contact [PCK21, San21, TJS^{+}22]. Contact-Tracing [San21]. Containerization [OMB17]. Containing [Fra07a]. containment [CHK^{+}05]. Contains [SW03a]. Contemporary [Ahm07, Smi12a]. Content [Bon16, FD10, GSLA15, MSW09]. Content-Based [GSLA15]. Contents [Ano13e, Ano14-33, Ano14-34, Ano14-35, Ano15-42, Ano16-45, Ano16-46, Ano16-47, Ano16-48, Ano16-49, Ano16-50, Ano17-76, Ano17-77, Ano17-78, Ano17-80, Ano17-81, Ano18-46, Ano18-48, Ano18-49, Ano18-50, Ano18-51, Ano19-78, Ano19-79, Ano19-80, Ano19-82, Ano20-49, Ano20-50, Ano20-51, Ano20-52, Ano20-53, Ano20-54, Ano21-82, Ano21-83, Ano21-85, Ano21-86, Ano21-87, Ano22-81, Ano22-82, Ano22-83, Ano22-84, Ano22-85, Ano22-86, Ano23-60, Ano23-61, Ano23-62, Ano23-63, Ano23-64, Ano23-65, Ano24-40, Ano24-41, Ano24-42, MMR24, Ano14-36, Ano14-37, Ano14-38, Ano17-79, Ano18-47, Ano19-81, Ano19-83, Ano21-84]. Context [BS09, Eps17, KG17, RS06, Sch07a, SBA^{+}18, SLS18]. Contextual [BS24, Mal23, WYA06]. Continual [PEL^{+}18]. Continuing [Hea03a]. Continuous [ACAT^{+}15, GL04, KF13, KA13]. Contracts [CXG^{+}18, ZHC^{+}20]. Contribution [Gel14]. Control [eDIATK22, AAB13, AJQA23, BW07, BGC^{+}14, BN08, Cal04d, Coh10a, CB15, Cra12, DXR^{+}23, FKS07, KvdHK^{+}14, LBB07, McL13, MPS14, Mil05, NBLC09, NSSS08, NSA^{+}23, PR08, PM14, PMB^{+}14, PHS^{+}08, RORL22, SS10, TP08, ZRM14, KS13b]. Controller [Ano17-69]. Controllers [AJQA23]. Controls [AED^{+}21, Mar13a]. Conundrum [Can14]. Convenience [Cal04b]. Convenient [Ren12]. Convergence [AR24, Cat11, Gee06a, Gee21b, Gee23a, RS06]. Conversations [Ano17h]. Cookie [FVJ19]. Coordination [Mor11]. Copyright [Les03b, Les03c, Les04b]. Core [Ano15-43, Dim07]. Corner [BKMN09, CFN13, Mar08, TR09a, TR09b, KN13]. Corporate [Ano08d, SAA04]. Correction [LN08]. Correctness [Bel07a, CXG^{+}18]. Correlation [Gee11b]. Correlators [HRTT03]. Cosimulation [SSJ^{+}23].

Cosmetology [Cal04e]. Cosmology

[Lan03]. Cost

[Cal04b, Ede10, Man13, PD07, SC07, ZC09]. Cost-Effective [PD07]. Could [CBOS20, MB16, Mei04, Per07]. Count [YE09]. Counter [GC06]. Counter-Forensic [GC06]. Counterattack [Cal03d]. Counterfactuals [Far20]. Countering [CUCD14, PP06b, SK05, CA05]. Countermeasures [CHMO07, RSOD22, Lin05]. Counterpoint [McG13b, Pee13]. Counting [YB08]. County [BFS^{+}21]. Course [TA23]. Courseware [RAAB20]. Cover [Ano08e, Ano13a, Ano14e, Ano14f, Ano15l, Ano15m, Ano15n, Ano15o, Ano15p, Ano15q, Ano16k, Ano16l, Ano16m, Ano16n, Ano16o, Ano16p, Ano17-30, Ano17-31, Ano17-32, Ano17-33, Ano17-34, Ano17-35, Ano18j, Ano18k, Ano18l, Ano18m, Ano18n, Ano18o, Ano19-40, Ano19-41, Ano19-42, Ano19-44, Ano20v, Ano20w, Ano20x, Ano20y, Ano21-42, Ano21-43, Ano21-46, Ano21-47, Ano22-30, Ano22-31, Ano22-32, Ano22-33, Ano22-34, Ano23l, Ano23m, Ano23n, Ano23o, Ano23p, Ano23q, Ano24n, Ano24o, Ano24p, Ano14g, Ano14h, Ano19-43, Ano19-45, Ano21-44, Ano21-45, Ano22-29]. Covered [Ano23-30, Ano24-28]. Covertly [CFN13]. COVID [PCK21, Pei20, SLKJ21, Sch20a, TXL^{+}21]. COVID-19 [PCK21, Pei20, SLKJ21, Sch20a, TXL^{+}21]. Cowbell [SPT19]. CPI [ECG^{+}07]. CPU [CFN13]. Craftspeople [Bel07b]. Craig [McG18d]. Crashes [Pay19]. CrawlPhish [ZOC^{+}22]. Cray [Ano18-40]. Craze [Arc04a]. Create [EJ06, Lan09b, MOG^{+}20, SE07]. Creating [All07, PMNT12, RSWO18, SOG^{+}19]. Creativity [Les04b, Les04d, Les11c]. Credentials [CLN12, DN07]. Credit [KST^{+}21]. Creep [Les15a, Sch10b]. Crime [Cam03b, Gra13, Hut24]. Crimeware [Sty08]. Criminal [O’R17]. Criminals [Lev04b]. Crisis [Ark11, Les03a, MS20, OdH12, LS05a]. Criteria [Hea04, Kal12, KS06, LLT14]. Critical [AIKR13, Ami05, BSC^{+}08, BL13b, CZ23, Cyb03a, De 21, Fet16, Mat23, MWM13, Nic20b, OJR^{+}23, PR08, RBK23, RALS23, SE09, Wil16b, YAZ21]. Critique [FKS07, LBB07]. Cross [Ano18h, CCF^{+}20, Mac06, McL08]. Cross-Border [Mac06, McL08]. Cross-pollinate [Ano18h]. Cross-Site [CCF^{+}20]. Crossing [BL13a, BOR^{+}13a, MBLT13]. Crossover [Lev03a]. Crowd [ADG23]. Crowd-Sourced [ADG23]. Crowdsourcing [Cox11]. Crucial [Sad17]. Crutial [BSC^{+}08]. Crying [MA15b]. Cryptanalysis [ASK07, GS07, JL18]. Crypto [Ahm08b, Bel14c, BKMN09, CFN13, GCR18, Mar08, Nar13a, Nar13b, Pfl10, TR09a, TR09b, WL11, KN13, Wei05]. Cryptocurrencies [GCR18, Tzi18]. Cryptocurrency [LLKA18]. Cryptographers [BPS16, Tsa07]. Cryptographic [Bar15, BCM^{+}15a, BCDS22, Bur06, Ess17, GG11b, HV20, Law09, LN08, MR22, RRS06, Tom16]. Cryptography [Ano16g, Ano16h, Ano16-44, Ano21-39, Bar15, BCGN16, BGG^{+}22, BLM17, BLM18, Che17, Cor06, DP17, Dod22, Ell04, ECG^{+}07, Gen06, GN06, KL23, MO12, MS21b, Mic10b, MMB17, PY06, Rog16, Sch16a, Sch18a, Sch22b, Sen17, Sim15, Ste15, TG04, Tro08, TR09a, TR09b, YAZ21, GNP05, For04a]. Cryptojacking [CBOS20]. CS [Ano22-50, Ano22-51, Ano22-52, Ano17-82, Ano18-52]. CS1 [Nan09]. CSIRTs [SMO^{+}14]. Cult [Don04b]. Culture [Pow07]. Currency [GKCC14]. Current [DP17, NKJ^{+}14, RALS23]. Curricula [TKZ21, vO21a]. Curriculum [Bra07, MLM^{+}07]. Curse [GG11b]. Curves [CT11]. Customers [Sch04d]. Customizable [Gro08]. Customization [OdH12]. CVSS [SHH^{+}21]. CWE [How09a]. CWSandbox [WHF07]. Cybenko [BC23a]. Cyber [AF16, Ano15-36, ABD^{+}18, BSS10, Ber20a, BS20b, BD18, CV10, CZ23, CW08b, FHMF14, GB09, GMC^{+}08, Gro20, GFJ^{+}18, Kel10, KT20, KJC^{+}12, LL16, Lin09, MV22, May15a, MWM13, MLM^{+}07, NRH^{+}22, NTMD^{+}18, PE12, PMN^{+}14, PW17, PHR20, PZTE18, RDC^{+}18, SI09, SAF15, Tho18, TH13, WL18, WDP23, WWH10, LS05a, IP10, NRH^{+}22, SE15]. Cyber-Defense [SI09]. Cyber-Physical [BSS10, CV10, GFJ^{+}18, MWM13, PMN^{+}14]. Cyber-Threat [Kel10]. Cyberassault [Don07, Les07a]. Cyberassured [CAB^{+}22]. Cyberattack [BGLEP15, Ell11, Fid11, Hor20]. Cyberattack-Resilient [Hor20]. Cyberattacks [KLL17, SE13, TWC^{+}15, Bor05]. CyberCIEGE [ITA05]. Cyberconflict [Mul05]. Cybercontrol [Ksh14]. Cybercrime [Ano17-38, AAG15, AA15, Gla19, SLKJ21, Mei04]. Cybercrimes [Ksh06]. Cyberdefense [BS14, Nic20b]. Cyberdyne [GD18]. Cyberexercises [WC04]. CyberHEAD [ADG23]. Cyberincident [Bel12b, HW13]. Cyberinsurance [BP07, Far20, MTS15]. Cybermilitias [App11]. Cyberoperations [Irv13]. Cyberpandemics [MVL09]. CyberPatriot [WWH10]. Cyberphysical [AZ20, BGL^{+}23, CR20, Hor20, Mas19, TB19]. Cyberresiliency [SLS18]. Cybersecurity [ADG23, AR18, Amo16, Ano14-27, Ano14z, Ano15d, Ano15-37, Ano15-38, Ano16a, Ano16-27, Ano16-41, BHJM19, BLG^{+}15, Bel09, Bel15b, Bel18b, Bel23, Ben15b, Ben21, Bis16, BVB22, CK20, CSSvdH14, CR20, Dar14, DM15a, Dar15, DHR^{+}04, DMS^{+}15, Ded17, DLMS21, DMB^{+}22, Ell14, Eps19, Far21, FWBC15, FO08, FTCS12, GMB12, Gee10b, GA04b, HLJ21, HRGK^{+}20, HRMS21b, HRDR05, HBT12, Koh22a, Ksh15, Lan08b, Les11a, LXS22, MHRS19, MBLT13, McG13a, MB12, MDD^{+}15, Mos18, Nic20a, Okh21, Ort09b, Ost20, PMNT12, PW17, PHR20, PLW07, PGP16, RP09, RM22, SSW21, Sch19a, Sch20a, Sch13b, Sch19b, SEKR18, SDC^{+}17, SOG^{+}19, SPG^{+}19, SBF^{+}15, SMK^{+}20b, SLM^{+}23, TFB10, TKZ21, TI18, TA23, VKM^{+}15, WMD20, WTML17, Wil16b, WM20, YK16, Yan16, YG19, Wil16a]. cybersomethings [Bel13a]. Cyberspace [And03a, Don10b, Lan09b, LTB15, May15a, Say08, Les13c]. Cyberwar [Lew11]. Cyberwarfare [App11, BD11, Lan11, PD11]. Cyberworthiness [Tro04]. Cyborg [Hee11]. CybSI [YK16]. Cycle [Don11b, GLM11, Yav24].

D [Ano17-43, CBOS20]. Dan [McG06c]. dance [Cam05]. Dancing [Hea03c, MJP21, ZYG15]. Danger [Ano17s, EMM06]. Dangerous [DMB^{+}22]. Dangers [PA03]. DARPA [AF16, SAF15]. Dartmouth [GB09]. Daswani [McG12e]. Data [AACG19, All07, Ano14y, Ano17v, AR15, BPB^{+}04, BH19, BBL^{+}17, Bel14a, Ber20b, BW10, Bye04, CMR13, Cat10a, CK10, CP10b, De 24, Don04e, ES14, El 08, EK09, El 10, El 20, Eva14, GS03, GS06, GC09b, Gee14b, GC06, Gru16, GSLA15, HRGK^{+}20, How09b, HKM17, HCC09, In 07, JW11, KM22, Kau09, KWRK13, KAAEa17, KE09, Kot08, Kro18, Ksh14, KJ15, Lan09a, LSG^{+}18, Les03d, Les13a, Lyo24, McD11, McG13b, McL08, NROS24, O’R17, OAB07, PMA^{+}20, Pet19, dQSzL13, RBC14, RDK19, Rou09, Sch06c, Sch10c, SyPKZ22, SSH22, SBA^{+}18, SC19, SLW22b, Sma23, SW18b, SB21, SS22, Tel15, Ten16, TPPM07, VC04, VKM^{+}15, WMDX23, YG19, ZS19, Lin05, PT05b, VWW05, BLR19, Ber23, CM23, Ano18c, Ano19-28, Ano19-29, Ano19-30, Ano19-31, Ano19-38, Ano20-48, Ano21-33, Ano21-34]. Data [Ano22z, Ano24k]. Data-Driven [HRGK^{+}20]. Data-Protection [HRGK^{+}20]. Data-Sharing [CK10]. Database [ADG23, WAF11]. DataPort [Ano19-57]. Dating [EH13]. David [McG11a]. Dawn [Ano17t]. Day [MMB17, RSM13]. De-Anonymization [Sim21]. De-Identification [El 10]. De-identifying [El 08]. Dead [Ano17t]. Deal [Cal04b, Sag13, VVYY11]. Dealing [SLKJ21]. dear [dAMM13]. Death [MBLT13, Sch07c, VC08, Lan13a]. Debate [Don06a]. Debian [Ahm08b]. Deborah [McG12a]. Debugging [GTG07]. Debunking [SSH^{+}16]. Dec [Ano20z]. Decade [AvEB15, GM12, GGBM21]. Decades [Av18]. Decency [Mol13]. Decentralized [GKCC14, KNL22]. Deception [WL18]. Deceptive [EH13]. Decide [Rya03]. Decision [Ano17u, CRK^{+}13, CP10b, KTT19, SG18, AG05, Sah05]. decision-tree [Sah05]. Decisions [Ano21-88, CP15, EV18, EGB18, MT24, PR12a, Sch07d]. Declaration [TK15]. Decoding [Fid11]. Deconstructing [Gro20, Mas09]. Dedicated [KDC20]. Deduplication [HPSP10]. Deep [CGR^{+}22, DO22, RDK19, SB19]. Deeper [AJQA23]. Deepfake [MP22, RVG24, Yav24]. Default [Mas19, Mol13]. Defeating [DSJK19]. Defend [Bor15]. Defenders [AA15, FA07b]. Defending [AJV18, Bon16]. Defense [BPFG24, Cha12, FHMF14, Gui11, LCJ21, Mat23, Mil11, MLM^{+}07, RCK17, SI09, Sch18b, Sty04c, Syt05, WWH10, Yan16]. Defenses [AJV18, CF14, GY21, GHG14, LS09]. Defensive [DSKB22, Pfl10]. Defined [Ano17-73, KYEV^{+}18, LAYG16]. Defining [Mus21, PS06]. Definition [Mal23]. Degenerative [Bel24]. Degradation [Bil10]. Degraded [GD11]. Degree [NRH^{+}22]. Degrees [FB04a]. Deidentified [Sta20]. DelegaTEE [SMJ^{+}19]. Delegation [SMJ^{+}19]. Deletion [RBC14]. Delicate [SAA04]. Delivery [Nic21b, TH13]. Delta [MT23]. Demand [LMW10, ZYG15]. Demand-Response [LMW10]. Demands [GD17]. Demise [Eps15a]. demographics [VWW05]. Demystifying [LCJ21, Tor05]. Demythifying [TFB10]. Denial [Alj03, Che06, GHG14, KS12b, Cam05, CA05]. denial-of-information [CA05]. Denial-of-Service [Alj03, KS12b, Cam05]. Denied [Cam03a]. Denning [McG07c]. Deny [Mas19]. Dependability [ACL07, Ami05, AKN12, Blo15, Gar12f, Gar13a, Gar13b, Gar13c, Gar13d, Gar13f, Gar14a, Gar14b, Gar14c, Gar14d, KvdHK^{+}14, Gar13e]. Dependable [Hav13]. Dependence [GD17]. Dependencies [MP21]. Dependent [Eps17]. Depends [HK09, KvM11, Kau09, MVL09, Sch05b]. Deploying [DMS07, GSB^{+}04]. Deployment [CL07, CR09b]. Deployments [And12]. Depth [BPFG24, Eps15a, LBS09, Sty04c]. Deputy [Ahm08a]. Desensitization [BPB^{+}04]. Desert [Spa06]. Deserve [Ano15t]. Design [AJQA23, CGR^{+}22, DD03, DMS07, EGB18, Fla16, GP19, GW07, Hav13, LPJ^{+}16, MM09a, MN21, McL05c, Pei22, PA03, TVR23, TKZ21, VM04, McG14e, Smi12a]. Designed [PMN^{+}14]. Designed-in [PMN^{+}14]. Designers [TEG^{+}19]. Designing [BSH^{+}09, DOP^{+}21, Gid06, Jus04, PZTE18, Roe17, SSW21, Yan16]. Deskilling [Gee09b]. Desktop [Cur06]. Detectable [Ess17]. Detecting [ARvD12, BNSE21, CMS09, Ess17, Fly11, LGO06, MBP^{+}24, Sch11a, ZRM14]. Detection [ANC10, AJW13, AMF15, AMM19, Ano17h, BSP11, BV11b, CCF^{+}20, CKPP23, CRBM06, Che11, DO22, HCL11, HW13, Hee11, HPSR18, KTT19, LGO06, MP22, MPM22, MNS08, MK13, MC15, MKKP09, RVG24, RSM13, SWL21, TPMT15]. DeterLab [MB12]. Determining [CP09b]. Deterrent [Alj03]. Deterring [Ell11]. Deus [Don04c, TPS18]. Develop [Ano17-39, Pei23b]. Developer [Dhi11, Hal10, Mar15a, Ver06]. Developer-Driven [Dhi11]. Developers [BC14, FG22, GLS23, GS16, TVR23, WDP23, YRX^{+}22]. Developing [Bai12b, BC08, MSeF^{+}22, Pei19, Pow07, RS06, SPT19, SFH05, SFH06, YRX^{+}22]. Development [Ano16-27, AP05, Ben15b, BC23b, CKL^{+}12, CGL^{+}16, FPP12, HRGK^{+}20, How04, KG17, PS06, RDM^{+}14, WA07, vWM05]. Deviancy [Hut24]. Device [GFJ^{+}18, Her06, Men24, MŽ22, NKJ^{+}14, Sch22a, SMCA14, VVSL12a, VVSL12b, VM10, VT12]. Devices [Ada15, AIKR13, ASE21, AKN12, CAL^{+}24, DMA09, EKA14, ENDAC22, Hei07, LLT14, May15b, OKH07, SFE10, VWC19, WDP23, Wha11a]. Devil [MJP21]. Dialect [GG04]. Dialing [CK10]. Dialogues [Law16]. Dictionary [Koh22a]. Did [MHF19]. Dies [MMB17]. Different [Ano22-87, RW21]. Differential [NH23]. Difficult [Whi03]. Difficulties [LS08]. Diffie [McG15f]. Diffusion [AS15]. Digital [Ano17n, Ano17-36, AEP^{+}19, BP08, Cal03c, CMV09, Car09, DP20, EEA^{+}23, EPH12, FH06, Gee09a, Gee09b, Gra13, HKMR24, HCL11, HB22, Ksh17, LPJ^{+}16, Les04b, Les04c, Les08a, Les15b, MB19, MD09a, Mat23, MCW17, MCW19, OR03, PS23, RCK17, RdRdCG10, RBB19, Rou09, Sch04b, SVR^{+}19, SS06, TAP19, TBBR17, TWA08, VVYY11, WCV^{+}04, Les05, Tro05]. Dilemma [Hec11, OAB07]. DIMACS [McG03b]. Dimensions [CZ23]. Dinosaur [Len03]. Directions [BSP11, BV11b, CHK^{+}20, HHK18, HRMS^{+}21a, Ker09, OMB17, Por09]. Director [Tro05]. Dirty [GT22, MP10]. Discipline [LL11]. Disciplines [Pfl15c]. Disclosure [AK03, BC23b, McK08, PCK21, Sch03d, AT05]. Discovery [AM12, CFH05]. Discrimination [Acq08]. Discussion [Ano21-39, BC21]. Discussions [PLK23, Wha12]. Dish [Nic20a]. Disillusioning [Opp17]. Disinformation [AS11, Gra23, PGNRV24, Zur22]. Disk [BM14, BMMM10, GS03, HCC09, Thi06]. Disks [Mar10]. Dispelling [MP22]. Disposable [SJ03]. Disposal [HCC09]. Dissecting [DXR^{+}23, Lan11]. Distance [Ano06b, BV15, DM15b]. Distance-Bounding [DM15b]. Distinguished [Ano15a, Ano15b]. Distributed [eDIATK22, EGB18, GHG14, HW13, Kuf13, KS12b, MN21, Mei18, PCMN08, RBB19]. Distribution [Bon16, MGH^{+}15, Sas18]. Divergent [PCK21]. Diverse [FPP12]. Diversity [Ano18i, Ano21-52, Ano22-41, Ano22-53, Ano23-28, Ano23-29, Ano24y, Ano24z, Ano24-27, Hol15a, LBF15, WHD^{+}09]. Divide [BL13a, Sch04b, BOR^{+}13a]. DMCA [Les03b]. Dmitrieva [McG17c]. Dmitrieva-Peguero [McG17c]. DNA [Phi19]. DNP3 [CB15]. DNS [AS09, CR09b, WO09]. DNSKEYs [OZ09]. DNSSEC [WO09]. Do [Bel22b, BKS13, Bor15, BMS08, Bur06, CP10a, HAO^{+}22, Les10a, MS21c, Per07, SA12, Str12, WDP23, Wha11a, VWW05]. do-not-call [VWW05]. Do-Over [SA12]. Doctor [YI13]. Doctrinal [SM11]. Document [Gar14e]. Documents [Bye04, Ste11]. Doers [KT20]. Does [Ano17v, Ano17w, Ber21b, Gee11c, Hea04, Her13, Les14b, LB04, OR05, PRB^{+}12, WM20]. Doesn’t [RS11]. Doing [JCM12, Vie12d]. Domain [Ahm08a, Ano17t, Sel11, SZ16, TW21, CR06, Che06, MD09b]. Domain-Generation [SZ16]. Domains [AJW13, Ano17t]. Done [Her22, Sty07b, VB13]. Dongle [PSB^{+}07]. Dongle-Protected [PSB^{+}07]. Don’t [Bra07, Cyb04a, Les14c, Pet10, Sch19a]. Door [Mar15b]. Doors [Lan14a]. Dorothy [McG07c]. Doubt [Cal03c, GC09a]. Down [Sty05b]. Download [JJV^{+}09]. Dr. [Bel14b]. Draft [And03a]. Dragons [Sch06b]. Drawbacks [Fel03]. Drawing [Lan10a]. Dream [Nar13a, Nar13b, OR03]. Dress [Ahm08b]. Drive [NP07]. Driven [Dhi11, Gui11, HRGK^{+}20, KLL17, Str10]. Drives [Ano17s, Thi06]. Driving [BBL^{+}17, ES14, FPP12, NH23, WMD20]. DRM [Bel07a, Les03e]. Drone [WZE18]. Drones [BNSE21]. Due [DXS21]. During [DSKB22, Pei20, Sch20a, HL13, WA07]. Dust [Smi03a]. Dutch [CSSvdH14]. duties [Tro05]. Dynamic [SB06, Sun16, WHF07, HJK08]. Dynamically [KCC07, WBT^{+}18]. Dynamics [JKSM19, YB08].

E-biobanking [VB13]. E-Government [MCW08]. E-Messaging [Sol07]. E-Passport [Vau07]. E-passports [CT11]. E-Voting [Av18, HWB^{+}20, BPR^{+}04, BF07, Cha04b, DR04, EP04, GSHU08, GG04, BW12, KZZ17, SY12]. E2E [KZZ17]. Eagle [Irv11]. Early [CCW03, CCF14, Sto04]. Earth [Koh23a]. Easy [Bel16b]. Eat [Gee18b]. Ebb [HS17]. EBIA [FBG04]. econometric [Sch05a]. Economic [BCD^{+}15, DMS^{+}15, KJ15, RP09, Wil16b]. Economic-Aware [BCD^{+}15]. Economically [Bor05]. Economics [Acq09, Acq13, AS05, AT05, AvEB15, Bis16, Far20, FWBC15, Ksh06, Ksh10, Les09b, Les09a, Les11a, Les13c, Les13d, MM09a, Ver16, Yan16, DA05]. Economy [Gru16, Ksh17]. Ecosystem [ZKH^{+}23]. Ecosystem-Wide [ZKH^{+}23]. Ed [McG06d, McG08c]. ED288l.6 [LM08].

Edge [AR24, Ano23-40, LTB15, Mic21, Mic22, Nic20a, RORL22, Ano19g, Ano19n, Ano19o, Ano20b, Ano20f, Ano20h, Ano20g, Ano21k, Ano21l, Ano21p, Ano21m, Ano21n, Ano21o, Ano22b, Ano22g, Ano22h, Ano22i, Ano23c, Ano23e, Ano23f, Ano23g, Ano23w, Ano24f, Ano24g, Ano24h]. Editor [Ano03d, Ano06e, Ano09b, Ano14-30, FBG04, LB04, SC04, CL13, Ami05, And06b, Ano03c, Ano04b, Ano21-89, Ano22-87, Arb04, Cyb03b, Lan13a, Lan03, MHV18, Nic21a, PKBS15]. Editor-in-Chief [Ano22-87]. Editorial [Ano06a, Ano06d, Pfl15d, Ben14, Ano03f, Pal04]. Editors [BC23a, Cyb04a, Don04e, Don04d, Don09b, Mic09, Pei21, AS05, Ano23-38, Ano23-39, Ano24-38, AM04, CMV09, CV10, CG04, Cyb03c, DR06, DR04, Don09c, Eps13, ES11, GS06, GA10, GL14, IR11, KS08b, LS09, Lan08b, Lan09b, MW23, MD09b, MC09, Mic08a, MKF21, OP15, PM14, PTP07, PIK12, SS16, Sch04b, Sch09a, SEKR18, TP11, TP08, VM10, WCV^{+}04, Wil21, Zur23, Pfl16b]. Educating [IN10, KJC^{+}12, Lan14a, SE07]. Education [ADG23, AR18, Ano16c, Ano17k, Ano17i, Ano17j, BF08, BV06, CW08a, CW09, CBN11, Dar14, DM15a, DLMS21, Du11, Far21, FB04c, Gri04, Hag09, Irv11, KS12a, Loc09, Mat23, MA12, McG13a, MDD^{+}15, Nan09, PHR20, RC06, Rya07, RTL09, Sch13b, SMK20a, ST06, WTML17, Zat16, DR05]. Educational [Eag13, Gri04]. Educators [Fri03, RH06]. EDURange [WTML17]. Effect [AEV^{+}07, McG03a]. Effective [Mar13a, McG13a, PD07, SWL21]. Effectively [Cra12]. Effectiveness [SBF^{+}15]. Effects [Ano17y, Lew11, Tho18]. Efficacy [Amo16]. Efficiency [BB17, LB04]. Efficient [KZZ17, SLW22b]. Effortless [KP15]. eHealth [TMGP13]. EIC [Ano18-30]. EICAR [Gat05]. Eight [GLS23]. Eisenhower [Gee11d]. Ekiden [ZHC^{+}20]. Eldred [Les03c]. Election [Ano19a, Ano19b, Ano21-39, Ano22-42, Ess17, EP04, Mus21, RS17]. Elections [Ano21-53, BP12, Cha04b, CEL^{+}19, EH12, SW12]. Electoral [CEL^{+}19, GG04]. Electric [BGLEP15, Gra23, SSJ^{+}23, VM21]. Electricity [Ami03]. Electrification [Don08b]. Electronic [Ano16e, Ano16d, ASH^{+}08, BCM03, BPR^{+}04, BCS16, CLN12, KH12, Kot08, Les13b, Pee13, PWVT12, RJ08, SCM08]. Electronics [TA23]. Element [Tho13]. Eliminating [How09a]. Elinor [McG11b]. Elliptic [CT11]. Email [ARvD12, Bel16b, Gar03, Kap07, RS19, SJ03, PB05]. Email-Based [Gar03]. Embedded [BSS10, CPFJ14, CV10, HSTV06, KR10, MB19, VT12, KS13b]. Embedded-Device [VT12]. Embedding [JG07, TVR23]. Embracing [CC11]. Emergency [CRK^{+}13, Gro14]. Emerging [Bel06a, BMMM10, JKMS23, KPM^{+}19, Lan08a, Men24, RRAK14, Sch06a, SMCA14, dA16]. EMET [DeM15]. Emotional [MA15b]. Empathy [Sch11b]. Empirical [BHJM19, FWBC15, Pau10a, YBAG04]. Employees [Eps14]. employment [HL13]. Emulation [KKK^{+}21]. EMV [BCM^{+}15b]. Enable [McG13b]. Enabled [Amo13, HRMS^{+}21a, Mar13a, Por06]. Enabling [KKK^{+}21, SPH^{+}05, MR05]. Enclaves [CCX^{+}20, KDC20]. Encoded [ADE^{+}22]. Encouraging [Ano18i]. Encrypted [ADE^{+}22, BHP24, LH07, Sma23]. Encrypting [CC10, Mar10]. Encryption [Bel16b, Bur03, Con18, DR10, Gue16, Hei07, Hor19, JR14, Lau17, MR14, Mar07, Mar08, MS21a, Nac16, Smi11]. End [BHP24, BH19, Bra06, CEC^{+}08, DEKM^{+}21, Gee07, HWB^{+}20, MS21c, RST15, Ste13]. End-to-End [BHP24, Bra06, CEC^{+}08, HWB^{+}20, MS21c, RST15]. End-User [DEKM^{+}21]. Endeavor [BF05a]. Ending [Hag09]. Endosymbiosis [Gee09a]. Enduring [Ben15a, Mul14]. Enemies [Bel05]. Enemy [Don09c, Gee08a, GS16]. Energy [Nic21b, PM14, TJS^{+}22, THM15]. Energy-Delivery [Nic21b]. Enforcement [GN06, KvS14, Les03b, McL08, Nah08, PHS^{+}08, Raa19, Tzi18]. Engage [SDC^{+}17]. Engagement [Amo16]. Engineer [IN10]. Engineering [Ano18f, Bay11, Cal03d, CAB^{+}22, DBR^{+}08, EHK^{+}04, Far21, GC08d, IR11, KS06, MMKP16, Mei06, OC12, PKB^{+}24, Sch12b, Str10, TR09a, TR09b, dA16, Smi05b, Ano21r, Ano21q, Ano21s, Ano22j, Ano22k, Ano22l, Ano24a, Ano24e]. Engineers [CS13, Lan14a, TEG^{+}19, BG13]. Engines [RI12]. Enhanced [AACG19, Var09]. Enhancing [Cal04a, EEA^{+}23, LSG^{+}18, PS21, SZ06, SE09, Wei06]. Enigma [KM16]. Enlightened [Pfl13a]. Enough [DM15b, GK11, JCM12, Kni15, Sho23, VE06, Wil24]. Enslaving [EV18]. Ensure [RM22]. Enter [TEG^{+}19]. Enterprise [Amo13, LGO04, MJF07, Mar13a, MP10, SLW22a, Ste06, TGC16]. Enterprises [RM22, TPMT15]. Entrepreneur [Ano16j]. Entrepreneurs [Bac13]. Entropy [LH07]. Entry [DMA09]. Enumerations [NSA^{+}23]. Environment [BCGN16, DeL07, DN07, FPP12, Kau10a, KMP20, Kuf13]. Environments [EKA14, JSS20, Ker09, KSL^{+}20, LL23, MRS14, PCMN08, SG20, TJA10, TMGP13, Ada05]. EPC [HCL11]. Epidemic [Kel10]. Epidemiology [Les13b]. Episode [BC23a]. Equipped [ASH^{+}08]. Equitable [SAB^{+}19]. Equity [PM21]. Era [AAC^{+}17, Ber23, BS14, GM12, Gro14, Gru16, HK09, HFT^{+}19, Men24, MR22, Mos18, TCC^{+}24, Trč11, YG19]. Erickson [Sty04a]. Errata [Ano03b, Ano20t, Ano20u, Ano21-40]. Erratum [Ano21-41, SC04]. Error [Les15c]. Errors [Ano17-59, SP09, ZC09, vO23b, vO23a, TCM05]. Escape [BBC^{+}19]. Essential [VKM^{+}15, vWS06]. Est [Don04c]. Estimating [Jon07, LS06, LB08]. Estonia [Don07, Les07a]. Eternal [SPWS14]. Ethical [Baa03, BJD22, KvS14, Lan14c, Pfl04, SG18]. Ethics [AACG19, Ano17c, Ano17l, Ano17m, Ano17x, BV06, DBD11, JKE18, Kro18, Len03, MCKS10, Res16, Sch11c, SW18b]. Eugene [McG08d]. Euphoria [KNL22]. Euro [BS20a]. Europe [GOPB12, SBA^{+}18]. European [Ano19d, Ano19e, ADG23, Bor13b, DLMS21, GOPB12, SMK^{+}20b]. Evaluate [Cyb06b, El 20]. Evaluating [ASH^{+}08, Bab16, BPFG24, Jus04, MDD^{+}15]. Evaluation [All07, DM15a, FE11, FVJ19, GC06, Her22, HRK21, LK17, MT24, Nic05, Pau10a, PR12a]. Evaluations [MGH^{+}15]. Evasion [CRBM06]. Even [Kni17]. Event [BMZ14, Kuf13, PMA^{+}20, Tho18]. Events [HN10]. Ever [Her22, Les10a, MJP21, Sad17]. Everyday [Mol13, OGM^{+}24]. Everything [FH12, Pfl06]. Everywhere [Pfl16b]. Evidence [Cal03c, EPH12, OR03, SW12]. Evidence-Based [SW12]. Evidentiary [Gee06b]. Evidently [Gee06b]. Evolution [Cha09, RDM^{+}14]. Evolutionary [O’D07, Som04]. Evolved [AEY10, Pfl14b]. Examination [AEP^{+}19, MMKP16]. Examining [CSZ^{+}14, YG19]. Example [AP05, Val12]. Excellence [SDC^{+}17, YB05]. Exceptional [Ano21-39]. Excerpts [Koh22a]. Exchanges [MN21]. Executables [LKL^{+}19]. Execution [CCX^{+}20, EKA14, JSS20, KSL^{+}20, LL23, Pie23, SG20]. Executive [TH13]. Exercise [HRDR05, MLM^{+}07]. Exercises [Du11, Irv11, TI18]. Expanding [Eag13, Pfl14a]. Expectations [Gla19, ME10, Roe17]. Experience [CDJS^{+}24, FPP12, GLM11, Mas09, Mus21, NCA19, Sas07, SK04, YRX^{+}22]. Experience-Based [GLM11]. Experienced [Ano17u]. Experiences [MOT^{+}17, MRF^{+}24, Pei19, Ras11, SLM^{+}23].

Experimenting [DHP14]. Experiments [JJF08]. Expert [How08]. Experts [Ano14-31, Ano15-39, Ano15-43, Sty07b]. Expiration [Sch05c]. Explained [Mal23, Ste11]. Explanation [EV18]. Exploitation [Ano17t, Sty04a]. Exploiting [PB04]. Exploits [MJF07, Val12]. Exploration [Ano17y]. Exploring [CPFJ14, CFH05, HRDR05, MRF^{+}24, Smi12b, WMW22, DEKM^{+}21]. Expo [Kei05]. Exposing [FVJ19, Sel11]. Expression [TMU23]. Extended [CAL^{+}24, NROS24, NB23]. Extending [WO09]. Extension [Les03c, PLK23]. Extensions [CB15, FVJ19, PH18]. Extensive [FVJ19]. Eye [Sas07]. Eyed [Cyb05]. Eyes [Koh23b, Pfl14b].

Facial [dAMM13]. Facilitate [Ano17w]. Facing [Lan14c]. Factor [DeF11, Sas07]. Factoring [BGG^{+}22]. Factors [DMS07, FWFA23, vWS06]. Facts [Les03d, PLK23, RBE03]. Faculty [GB09]. Fail [Ano21-88, MBA12, Mas19]. Fail-Safe [Mas19]. Failed [ME10]. Failing [RØTM08]. Failings [MP22]. Failure [AALS21, GP19, Les15c, LL11, McG03a, Res16]. Failure-Based [LL11]. Failures [AIKR13, DMA09, JW11, VVSL12a]. Fair [Gel14, Les06]. Fairness [AWV18, Bin18, DMS^{+}15, KTT19, LLKA18]. Fairy [Opp15, Smi03a]. fancy [Don05b]. Fannie [Der13]. Fast [Sho23]. Faster [Ano17z, Ano17-39]. Fatigue [Res16]. Fault [CN13]. Faults [HQD^{+}03, Nac05]. Faustian [Cal04b]. Favorite [Ano18-42, Ano18-41]. Fear [JJV^{+}09, Sty04b]. Features [Gro08]. Federal [HV20, Bre12, Hoo16]. Federated [Ano21-41, BS13b, DeL07, JFK21, MR08, JJ13]. Feist [Les03d]. Felten [McG06d]. Fernbach [Ano18-40]. Fi [GM14, Sty04b, Ohm14]. Fick [McG14d]. fiction [Don05b]. fiduciary [Tro05]. Field [BDGS04, OY16, SMCA14, TGC16, dA16].

Fifteen [Orm03]. Fifteen-Year [Orm03]. Fight [FA07b]. Fighting [AvEB15, Bel12a]. Figures [Gar08]. Files [GM09, Gar14e, Kra04, MB19]. filler [Ano22-50, Ano22-51, Ano22-59, Ano22-63, Ano22-64, Ano22-65, Ano23d, Ano23-29, Ano23z, Ano23-27, Ano23h]. filler2 [Ano22-62]. Filling [PE12]. Filter [EJ06]. Filtering [KCC07, SNS10, Var10]. Final [Tho18]. Financial [AEH^{+}04, MN21, MA15b, TBBR17]. Find [GC09b, LH07, YB21]. Finding [Nac05, OHBS14, Ste15, Res05]. Fine [NB23]. Fine-Grained [NB23]. Fingerprinting [Ano17h, GFJ^{+}18, NKJ^{+}14, Rou09]. FinTechSec [TBBR17]. Fire [Baa03, Pfl04]. Firing [KMP20]. Firmware [HFT^{+}19, KKK^{+}21]. First [DR10, DP18, LSBB22, Pei23a, BC22]. Fisher [McG18a]. Fit [Ano16-39, Ano16-40, Ano17-60, Ano17-64, Ano17-61, Ano17-62, Ano17b, Ano17-63]. Fitting [Mar06]. Five [AR19, BDGS04, EW22, Gro08, TPS18]. Fix [Kup06, Sch16b]. Fixed [CHMO07, GBK09]. Fixes [DFJ^{+}20, SPL^{+}24]. Fixing [Lan08b, RØTM08]. Flag [Irv11, MBC^{+}18, Vig11, Les05]. Flake [McG11c]. Flash [Ano17-59]. Flaw [Ste15]. Flaw-Finding [Ste15]. Flaws [DD08, FVJ19, PA03, Sel11, vdKHA^{+}20, Ano16q, Ano16r, Ano16s, Ano16t, Ano17-37]. Flow [HPSR18, HS17]. Flow-Based [HPSR18]. Flows [McL08]. Fluxes [MKG09]. Focus [Ano14d, Ano14c, Ano15h, Ano15i, Ano15j, Ano15k, Ano17-27, Ano17-29, Ano17-28, EOM09, KAAEa17, OHBS14, Pfl13b]. Folk [Was12]. Follow [Sch19a]. Followed [Ano17-72]. Following [BS20a]. Food [Lan07a]. Fools [JJV^{+}09]. Footsteps [BS20a]. Force [Don04h, HLJ21, JR14, May15a, WC04].

Force-on-Force [WC04]. Forces [App11]. Forecast [BBB^{+}08, GHK^{+}06]. Forensic [Ano17-59, AEP^{+}19, GC06, MD09a, PMA^{+}20, Tzi18]. Forensics [AMM19, All05, Ano17n, Ano17-36, Ano17-56, Ano17-69, Ano17-68, BP08, Cal09, CMV09, Car09, GM09, HB22, MCW17, MCW19, RBBK04a, RBBK04b, RBB19, Rou09, SLKJ21, Sch22a]. Forever [GC10a]. Forgery [CCF^{+}20]. Forget [BDT17]. Formal [Kob19, MSeF^{+}22, MLS22]. Formally [Dod22]. Format [BM14]. Fortress [FB04d]. Forum [Les08b]. Forward [Ano15-28, Hea03b, Sho22]. Forwards [LRW^{+}23]. Foundation [Mea03a]. foundations [MR05]. Four [Sno05, vO22b]. Fourth [Bel06a]. fragility [Hol15b]. Framework [AK15, AWV18, AEP^{+}19, BDTJ19, CK10, Ded17, FOM23, Gro20, GLM11, Ive05, Kos15, KYEV^{+}18, LWLL10, MHRS19, NP07, PGNRV24, Ste06, Tel15, WYA06, CA05]. Frameworks [Ano17-74, KvS14]. Francisco [Ano17-75, Ano18-43, NL12]. Fratricide [Gee10a]. Fraud [Ano17w, CEL^{+}19, Ede10, Fly11, JJF08, Jon06, KST^{+}21, Ksh10, MA15b, Whi15]. Fred [McG09c]. Free [Boy16, HM22, KNBV17, MSM^{+}22, SC07]. Freedom [Con18, MSM^{+}22]. Freedoms [TPS18]. Frei [Don04a]. Friction [Red24]. Friend [YRX^{+}22]. Frightened [CR09a]. Frincke [McG12a]. Froelich [McG18d]. Front [Ano08e, Ano13a, Ano14e, Ano14f, Ano15l, Ano15m, Ano15n, Ano15o, Ano15p, Ano15q, Ano16k, Ano16l, Ano16m, Ano16n, Ano16o, Ano16p, Ano17-30, Ano17-31, Ano17-32, Ano17-33, Ano17-34, Ano17-35, Ano18j, Ano18k, Ano18l, Ano18m, Ano18n, Ano18o, Ano19-40, Ano19-41, Ano19-42, Ano19-44, Ano20v, Ano20w, Ano20x, Ano20y, Ano21-42, Ano21-43, Ano21-46, Ano21-47, Ano22-30, Ano22-31, Ano22-32, Ano22-33, Ano22-34, Ano23l, Ano23m, Ano23n, Ano23o, Ano23p, Ano23q, Ano24n, Ano24o, Ano24p, CDF^{+}12, Les07a, PMB^{+}14, Ano14g, Ano14h, Ano19-43, Ano19-45, Ano21-44, Ano21-45, Ano22-29]. Frontier [BDL^{+}14]. Frontiers [Ben21, Sad16]. Fuel [SG20]. Fueling [Ben21]. Full [CUCD14, Sch03d]. Fully [Nac16]. Fumbling [Sch10a]. Function [Cal03e, HRK21, Sch10b]. Functionality [BMMM10]. Functions [BH14, RRS06]. Fund [Ano24y, Ano21-52, Ano22-53, Ano24z, Ano24-27]. Funny [SRG03]. furiously [Kri13]. Further [WZE18]. Furthering [YB21]. Fusion [HRK21]. Futility [Ell14]. Future [AS22, Ano17-36, BCE^{+}12, BV11b, BW12, CK20, CS09, DHP14, EEA^{+}23, EGB18, GD18, Hea04, HRMS^{+}21a, Ker09, KS08b, Les07b, LGO06, MC09, Mea03b, OP15, PS23, Pei21, Pfl15b, PHS^{+}08, Sch10a, Sch14a, SDO21, WM20, Yav24]. Fuzzing [Pay19, Oeh05]. Fuzzy [Gar08].

G [For04a, BS13a]. Gadgets [Arc03a].

Gained [GC10b]. Gaining [LTB15]. Game [GG04, LPJ^{+}16, McL05c, Pal04, VLOS18, WCV^{+}04, WMS10]. Games [BCLM09, Fel06, GY21, MH07, MC09, MKKP09, Sad16, Sch22c, YR09]. Gaming [Ano08c, Kan09, ITA05]. Ganesh [For04a]. Gap [BLCDK11, JBZ^{+}15, PM14, TMGP13, vWM05]. Gaps [Amo16]. Garage [Mar15b]. Gartner [Yav24]. Gary [McG09a, McG13c, Ran16]. Gates [Bel13b]. Gateway [WL11]. Gathering [ASC15]. GDPR [MHF19, TEG^{+}19]. Gedanken [Don04a]. geek [Don05b]. Geer [McG06c]. GenCyber [LL16]. Gender [Amo16, MT24, SAB^{+}19]. Gender-Equitable [SAB^{+}19]. General [Gue16, Pet19]. General-Purpose [Gue16]. Générale [Eps08]. Generated [MBP^{+}24, MMR24]. Generating [KM22]. Generation [Arc04c, BL13b, Cam03b, DMR^{+}22, KS13a, LL16, MR14, RLS^{+}08, SZ16, MR05, SK05]. Generative [SCHC24, TCC^{+}24]. Generators [Ano22-74]. Generic [Ano22-64, Lev05]. Genetic [Phi19]. Genome [Ano16f]. Genomic [AH17, BB17, BDT17, De 14, HKM17, Sav17]. Genteel [Fla16]. Gentle [LS12]. George [BC23a]. Georgia [Les12b]. geospatial [PT05b]. Gesture [Ano17-67]. Gesture-Based [Ano17-67]. Get [Ano15r, Ano15s, Ano15t, Lib05, MB16, Pei19]. Gets [GG04]. Getting [Blo12, HMA04, MTA23]. Ghost [Arc07a]. Giant [OBS^{+}20]. Giants [ZYG15]. Gift [Ano18-42, Baa03, Pfl04, Ano18-41]. Gillian [McG10b]. Giovanni [McG12b]. Give [Ano18-42, Ano18-41]. Giving [Eva14, Vie12b]. Glass [Wil21]. Gligor [McG09f]. Global [Ahm08b, Hea04, Hei16, Kel10, NSSS08, PEL^{+}18, Por06]. Globalization [Ano21-81, CWVM22]. Globalized [ABMS22]. Gnana [For04a]. Go [Bur06, CB14, Wha11a, Ano17-82, Ano18-52]. Goal [Lan09b]. goals [Ano17b]. Goes [Gee20, RSWO18]. Going [BBCL13, CPFJ14]. Gold [For04a, Tsa07]. Golden [GM12, SWL21]. Gone [AAB13]. Good [Av18, AM04, BKS13, GC08c, GC08d, GC09a, GC09b, GC09d, GC09c, Kni15, Les03e, Sho23, Res05]. Google [GM14, Les06, McL05b, SFK^{+}10, SDO21]. Got [Ano16q, Ano16r, Ano16s, Ano16t, Ano17-37, GC08a, SA12]. Gotten [Bel14a]. Governance [BW10, Kro18, SMK^{+}20b, TPPM07]. Governing [WM20]. Government [ASC15, Bel09, Cat10a, DOP^{+}21, EW22, Gee14b, KJC^{+}12, MCW08, Don04g]. Governments [Mic08a]. GPS [BO14]. Graduate [Rya07]. Grail [Mas24, San14]. Grained [NB23]. Grand [AF16, NTMD^{+}18, PZTE18, SS04, SAF15, Tho18]. Graphics [Ano19m, Ano19l, Ano20d, Ano20e, Ano21j, Ano21h, Ano21i, Ano22e, Ano22f, Ano24d]. Graphs [GW07, Lau17]. Gratification [Qua03]. Great [BL13a, DFJ^{+}20, ME10, BOR^{+}13a]. Greatly [FA09]. Green [TR09a, TR09b, Kri13, McG14c, McG15e]. Greg [McG10c]. Grid [BGC^{+}14, Coh10b, FRA10, KHLF10, SSJ^{+}23, CDS08, MM09b]. Grids [DHP14]. Grief [GC09e]. Grossman [McG09e]. Ground [Ano22-74, GA12, McG03b]. Ground-Truth [Ano22-74]. Group [BGM^{+}06, Sch09d]. Grover [JL18]. Growing [Ano17-59, Ksh17, Whi15]. Growth [Ben16a]. Guard [IP10]. Guarding [FB04d]. Guessed [THS20]. Guessing [THS20]. Guest [Ben14, Ami05, AS05, And06b, Arb04, AM04, CL13, CMV09, CV10, CG04, DR06, DR04, ES11, GS06, GA10, GL14, IR11, KS08b, LS09, Lan13a, MW23, MD09b, MC09, MKF21, OP15, PKBS15, PM14, PTP07, PIK12, Pfl16b, SS16, SEKR18, TP11, TP08, VM10]. Guide [Mar15a]. Guidelines [BJD22]. Guilty [Sch03c]. Gun [SC04]. Gunfight [Cyb04a]. Gunnar [McG09d]. Guys [Sch19a].

Habits [BBL^{+}17]. Hack [BPR^{+}04]. Hack-a-Vote [BPR^{+}04]. Hackathon [SOG^{+}19]. Hacked [MB16]. Hacker [Bra07]. Hackers [Ano17d, App11, Bra07, PP06a, Sch03b, Spi03]. Hacking [Ano17o, Ano17p, Bel21, CBN11, Don04f, GHS14, Sch20b, Sch22c, SBB^{+}18, Sty04a, Sag13]. Half [GC10a]. Half-Life [GC10a]. Hall [For04a, Pfl04, Eva15]. Halting [Hol15b]. Halvar [McG11c]. Hand [CB14]. Handle [OAB07, RC06]. Handling [DN07, LS05b, Sta20]. Hands [Du11, Kuh16]. Hands-On [Du11, Kuh16]. Happened [Nar13a, Nar13b, SRG03]. Happy [Vie12c, Wil21]. Harbor [SE15]. Hard [GC09b, KCL^{+}23, Mar10, PC10, YB21]. Hardening [Red24, TWA08]. Harder [Sch16a]. Hardware [DMR^{+}22, GS07, Han12a, Han12b, JKMS23, KR10, MA20, Men24, MRO^{+}23, PvDS08, TA23, ZL20, KN13]. Hardware-Anchored [Han12a, Han12b]. Hardware-Assisted [MA20]. Hardware-Based [PvDS08]. Harlan [Ano17-43]. Harm [AWV18, ABAGOB10]. Harm-Reduction [AWV18]. Harmful [Gee21a]. Harms [Sav17]. Hash [Bur06, Bur08, But17, EH12, RRS06]. Hash-Based [But17]. Hash-Only [EH12]. hashes [GNP05]. Hashing [Rou09]. Hathaway [Say08]. Hauling [Sty05b]. Haven’t [ME10]. Having [Gee12b]. Hayes [McG10b]. haystack [Pfl13c]. Hazards [BBD^{+}08]. HbbTV [Wai16]. Heading [HKM17]. Healing [ASE21]. Health [Ano04c, BC21, CL13, Cha12, El 08, El 10, McG13b, Pee13, San21, WDP23]. Healthcare [ACC^{+}13, Hec11, JW11, KJ15]. Healthtech [TK15]. Hear [Arc06]. Heartbleed [GK14, CDFW14]. Heights [Ano15-43]. Hellman [McG16d]. Hello [KTS18]. Help [Ber21b, HB22, Pay04, YI13, KH12]. Helping [Ano17-39, Loc09, SPJ^{+}14, VVYY11]. Hemorrhages [JW11]. Here [AALS21, Bur06, CM22, Nah08, Sch06b, Wha11a]. Herring [Cal04c]. Heterogeneous [DN07]. Heuristic [KKK^{+}21]. Heuristics [El 08]. Hey [Don03b]. Hidden [Bye04, LS08, OSM11, Sim21]. Hide [Blu10, PH03]. Hiding [CLS07, MC15]. High [Bar15, BC22, KMP^{+}11, LJZ12, Max20, Mea03b, Pot09, SR16, TKZ21, WL11, WWH10, YRX^{+}22, KN13]. High-Assurance [Bar15, KMP^{+}11, LJZ12, Mea03b, WL11]. high-availability [KN13]. High-Precision [YRX^{+}22]. High-Quality [TKZ21].

High-Risk [SR16]. Higher

[ADG23, CW09, DLMS21, SMK20a].

Highlights [Lan14b]. Hijacker [Ahm08a]. Hindering [DBR^{+}08]. Hint [GB12]. Hip [DR06]. HIPAA [AEV^{+}07, Nah08]. Historical [MA12]. History [Ano19i, Ano19j, Ano20c, Ano21c, Ano21b, Ano21g, Ano21d, Ano21e, Ano21f, Ano22c, Ano24c, Les04d, MVL09, Rui06, Yav24]. Hoc [Car08, HP04]. Hoff [McG10a]. Holes [Ano17d, Mar06, Res05]. Holistic [McL05a]. Holistically [HBT12]. Holy [Mas24, San14]. Home [Ano17-74, Ano21-81, ABMS22, BJD22, CWVM22, SyPKZ22, ABMS22]. Homes [BVB22, WMDX23]. Homo [TPS18]. Homomorphic [Lau17, Nac16]. Honest [EH13]. Honey [JR14]. Honeybug [BD18]. Honeynet [Kra04, McC03b, Spi03]. Honeynets [Dim07, LGO04]. Honeypot [Kra04, RBBK04a, RBBK04b, Tho18]. Honeywall [Cha04a]. Hong [LZ16]. Hord [McG13g]. Horizon [Ano03b, Cha09, Var09, Win03]. Horse [Eps19, Fra07a, Ten16]. Host [BSH^{+}09, RBBK04b, Ano20z]. Hosts [JMD06]. Hot [GG04, Woo23]. House [Ano13c, Ano13f, Ano14d, Ano14o, Ano14w, Ano14p, Ano14-27, Ano14-28, Ano14-29, Ano14-31, Ano14-32, Ano15d, Ano15c, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano15t, Ano15w, Ano15u, Ano15v, Ano15x, Ano15y, Ano15z, Ano15-27, Ano15-28, Ano15-29, Ano15a, Ano15b, Ano15-37, Ano15-38, Ano15-39, Ano15-41, Ano15-43, Ano16j, Ano16c, Ano16f, Ano16e, Ano16w, Ano16u, Ano16z, Ano16x, Ano16y, Ano16v, Ano16-32, Ano16-29, Ano16-30, Ano16-31, Ano16-41, Ano16-43, Ano16-44, Ano17a, Ano17k, Ano17-29, Ano17-42, Ano17-58, Ano17-64, Ano17-65, iC15, PZTE18, YI13, Ano15r, Ano15s, Ano16r, Ano16s, Ano16t]. Houses [Wil21]. Hover [EH12]. Howard [McG12c]. HTTPS [CCR09, PRCC10]. Hubris [Pfl15b]. Human [BLG^{+}15, BF05a, FWFA23, Hag09, HSS11, Les15c, MBP^{+}24, OGM^{+}24, Sch08b, Sch09d, Smi03b, Tho13]. Human-Computer [Smi03b]. Human-Physiology-Immunity [HSS11]. Humans [Ano17o, Ano17p, Hor14a, KO08, Smi03b, Wil22]. Humphery [Ano23b]. Humphrey [Ano22a, Ano23-66]. Hurdles [McK08]. Hybrid [LBF^{+}19]. Hydra [BDTJ19]. Hype [Pay19, Yav24]. Hype-Train [Pay19]. Hypervisor [LKL^{+}19]. Hypervisor-Based [LKL^{+}19]. Hyppönen [McG07b].

I/O [KS08a, SND14, ZYG15]. IA [BV06]. ICS [DP20, GP12]. ICT [Ksh14]. ID [CL07, Les14a]. Idea [AM04, Les04c, Res05]. Ideas [FREP17, KRK18, Les15b, Ano18h, Kri13]. Identification [Don08b, El 10, Gar03, Lys07, PKW04, Sch04a]. Identify [BCG^{+}09]. identifying [El 08]. Identities [CLN12]. Identity [Acq08, ACAT^{+}15, Bel10a, BSSB07, BS13b, BHRR07, Cam10, Cha09, CCF^{+}10, DD08, DP18, GOPB12, Gee13b, Gee22, GY13, GLM11, HSC08, JJ13, MR08, Mar08, McC03a, MCW08, OdH12, Pet06, PWVT12, Slo14, TMGP13, TAP19, WYA06, Way08]. Identity-Based [GY13, Mar08]. IdM [ACAT^{+}15]. IDS [All07, GC08b]. IEEE [Ano19a, Ano19b, Ano19d, Ano19e, Ano19c, Ano21a, Ano22a, Ano23r, Ano23b, Ben17, BP18, BH22, Ano03a, Ano06b, Ano06d, Ano06c, Ano14i, Ano14j, Ano14k, Ano14l, Ano14m, Ano14-30, Ano15w, Ano15u, Ano15v, Ano15x, Ano15y, Ano15z, Ano15-27, Ano16-27, Ano16w, Ano16u, Ano16z, Ano16x, Ano16y, Ano16v, Ano16-32, Ano16-29, Ano16-30, Ano16-31, Ano16-28, Ano17-42, Ano17-41, Ano17-40, Ano17-43, Ano17-44, Ano17-45, Ano17-46, Ano17-47, Ano18u, Ano18q, Ano18r, Ano18s, Ano18v, Ano18t, Ano18w, Ano18x, Ano18y, Ano18z, Ano18-27, Ano18-41, Ano18-42, Ano19-46, Ano19-52, Ano19-47, Ano19-53, Ano19-48, Ano19-55, Ano19-49, Ano19-54, Ano19-56, Ano19-50, Ano19-51, Ano19-57, Ano19-58, Ano19-59, Ano19-60, Ano19-61, Ano19-62, Ano20-29, Ano20-30, Ano20-34, Ano20-31, Ano20-27, Ano20-32]. IEEE [Ano20-28, Ano20-33, Ano20-35, Ano21-66, Ano21-54, Ano21-60, Ano21-55, Ano21-61, Ano21-48, Ano21-56, Ano21-62, Ano21-49, Ano21-57, Ano21-63, Ano21-53, Ano21-50, Ano21-58, Ano21-64, Ano21-52, Ano21-51, Ano21-65, Ano21-59, Ano21-68, Ano21-67, Ano22-50, Ano22-51, Ano22-52, Ano22-49, Ano22-35, Ano22-43, Ano22-40, Ano22-36, Ano22-44, Ano22-45, Ano22-42, Ano22-37, Ano22-46, Ano22-38, Ano22-47, Ano22-41, Ano22-39, Ano22-48, Ano22-53, Ano22-58, Ano22-55, Ano22-54, Ano22-56, Ano22-57, Ano22-60, Ano22-59, Ano22-61, Ano22-62, Ano22-63, Ano22-64, Ano22-65, Ano23s, Ano23-40, Ano23r, Ano23v, Ano23-38, Ano23x, Ano23-31, Ano23-34, Ano23t, Ano23-35, Ano23u, Ano23w, Ano23-36, Ano23-32, Ano23y, Ano23-28, Ano23-37, Ano23-29, Ano23-30, Ano23z, Ano23-27, Ano23-39, Ano23-33, Ano23-41, Ano23-42, Ano23-43, Ano23-44, Ano24q, Ano24r, Ano24y]. IEEE [Ano24t, Ano24v, Ano24-29, Ano24z, Ano24s, Ano24-28, Ano24-30, Ano24x, Ano24-27, Ano24w, Ano24u, Ano24-31, Ano24-32, Ano24-33, Ano24-34, Ben16b, BS20a, CGL^{+}16, CCF14, iC15, JW04, McG14e, NL12, NPS14, Nic21a, Ano14-32, Ano17g, Ano18b, Ano18a, Ano18c, Ano18d, Ano19i, Ano19j, Ano19n, Ano19o, Ano19m, Ano19l, Ano19k, Ano19q, Ano19r, Ano19s, Ano19p, Ano19t, Ano19u, Ano19v, Ano19w, Ano19-27, Ano19y, Ano19h, Ano19x, Ano19z, Ano19-28, Ano19-29, Ano19-30, Ano19-31, Ano19-32, Ano19-33, Ano19-34, Ano19-35, Ano20c, Ano20f, Ano20h, Ano20g, Ano20d, Ano20e, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano20o, Ano20p, Ano20q, Ano20-48, Ano21c, Ano21b, Ano21g, Ano21d, Ano21e, Ano21f, Ano21k, Ano21l, Ano21p, Ano21m]. IEEE [Ano21n, Ano21o, Ano21j, Ano21h, Ano21i, Ano21r, Ano21q, Ano21s, Ano21t, Ano21u, Ano21v, Ano21w, Ano21x, Ano21y, Ano21z, Ano21-27, Ano21-28, Ano21-29, Ano21-30, Ano21-31, Ano21-32, Ano21-33, Ano21-34, Ano21-35, Ano21-36, Ano21-37, Ano21-38, Ano21-89, Ano22c, Ano22d, Ano22g, Ano22h, Ano22i, Ano22e, Ano22f, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano22-66, Ano22o, Ano22q, Ano22p, Ano22s, Ano22r, Ano22t, Ano22u, Ano22v, Ano22y, Ano22x, Ano22w, Ano22z, Ano22-27, Ano22-28, Ano22-87, Ano23d, Ano23e, Ano23f, Ano23g, Ano23j, Ano23k, Ano23i, Ano23h, Ano24b, Ano24c, Ano24f, Ano24g, Ano24h, Ano24d, Ano24e, Ano24j, Ano24i, Ano24k, Ano24l, Ano24m, BO23, Pei23a]. IETF [Woo23]. If [HAO^{+}22, Les03d, Les03e, Wha11b]. II [FA07b, GC08d, Lan14b, RBBK04b]. I’ll [PLW07]. Illegal [Ohm14]. Illusion [GHS14]. Illusions [Lev04c]. Im [CXG^{+}18, Heb22, LM08]. Image [TCC^{+}24]. Images [MBP^{+}24, dQSzL13]. Immaterial [Tro06]. Immunity [HSS11]. Impact [Ano18-28, Bai12a, Bre12, HQD^{+}03, HN10, Hol15a, JKMS23, KLL17, MA15b, OGM^{+}24, Pfl04]. Impacts [Wil16b]. Imparts [TH13]. Impending [Don06a]. Impersonation [Vil04]. Implement [SCZ^{+}13]. Implementation [CW08b, Ded17, Pei22]. Implementations [AS09, BFK16, Tom16]. Implementing [Mar13a, SZ05, SW18b]. Implications [Ano17-74, BBL^{+}17, GM09, GM14, Gra23, Kan09, MBR^{+}21, PGP16, SC19, SDKM20, vZH^{+}19]. Importance [Kni15, Mul14, Sch12b]. Improve [JMD06, Les09a, MP10, MHF19, Syt05, Sno05]. Improved [How04]. Improvement [Mat12, NP07, TM05]. Improvements [JS10]. Improving [AA12, BC14, BCM^{+}15a, BB17, CP15, Dim07, How09a, Lan07a, SBF^{+}15]. In-Depth [LBS09]. In-VM [LWLL10]. Inaction [Wil16a]. Inc. [MSeF^{+}22]. Incentive [MM09a]. Incentive-Centered [MM09a]. Incentives [Cam10, DHP14, Hal10, Les09a, LLWC11]. Incident [BSP11, CSZ^{+}14, DMB^{+}22, Hor14b, Mor11, PSO^{+}21, RDM^{+}14, Sch14a, SBF^{+}15, Tom06, WB22]. INCITS [CW08b]. Inclusion [Ano22-41, Ano22-53, Ano23-28, Ano24y, Ano24z, Ano24-27]. Inclusive [FD11, Wan18]. Incorporating [LPJ^{+}16]. Increase [FA09, LB04]. Increased [McL08]. Increasing [LWLL10]. Increasingly [KvM11]. Independent [BM14]. Index [Ano03a, Ano04a, Ano05, Ano06c, Ano08b, Ano09a, GC09d, Gee10b]. Indexing [Les06]. India [Ksh15]. Indications [Pet09]. Indicators [EEA^{+}23]. individual [AG05]. Individualized [Les08a]. Individualizing [TI18]. Individuals [Sav17]. Industrial [Ano17c, AJQA23, CB15, DXR^{+}23, KvdHK^{+}14, NSA^{+}23, RM22, ZRM14]. Industry [Ano17v, Ben16a, Ben21, Dod22, EW22, Kup06, MC12, Sch07c, Sch17a]. Inevitability [MVL09]. Inevitable [Gee11f, Pfl15a]. Inexpensive [Fra18]. Infeasibility [Bel06c]. Infection [Gif10]. Infectious [GL12]. Inference [AH17, CGR^{+}22, Far20, PS23]. Inferring [LMW10]. Inflation [Gee11a]. Influencers [Sho23]. Information [Acq09, AA12, AS05, Ano04c, Ano17x, Ano18v, Ano19-46, Ano19-47, Ano19-48, Ano19-49, Ano19-50, Ano19-51, Ano20-29, Ano20-30, Ano20-31, Ano20-32, Ano20-33, Ano21-54, Ano21-55, Ano21-56, Ano21-57, Ano21-58, Ano21-59, Ano22-43, Ano22-44, Ano22-45, Ano22-46, Ano22-47, Ano22-48, Ano23-31, Ano23-32, Ano23-33, Ano24-29, Ano24-30, Ano24-31, ACC^{+}13, ASC15, Bab16, BW07, BMZ14, BF08, BW10, Bye04, Cal03b, DSEB12, DK10, DPP15, Eva14, FE11, FB11, Gar14e, Gel14, GHK^{+}06, Gri04, HV20, JG07, JGP09, JP11, Kam14, KMP20, Les11d, LMW10, LL11, MC15, MA12, Mic22, MLM^{+}07, NSLW20, Opp15, OKH07, OAB07, Pay04, PCMN08, Pet07, PP06b, Pot10a, RWD03, Ren12, RS06, Rui06, Rya07, RR08, Sad16, Sav17, SFH05, SFH06, SB19, SS04, Spa08, Sta20, ST10, TG07]. Information [Tho13, TMU23, WMD20, Wha12, Yan16, vWM05, Ano18w, CA05, ITA05, BBB^{+}08]. Information-Assurance [MLM^{+}07]. Information-Centric [NSLW20]. Information-Processing [HV20]. Information-Sharing [Kam14]. Informative [ENDAC22]. Informed [Bor15]. InfoSec [Ano13b, Hei16, VC08]. Infrastructure [Ami03, Ami05, BBCL13, BSC^{+}08, CL07, CZ23, KNL22, Mat23, MKG09, Mic22, NK12, PR08, Pei23b, RW21, Tra09, Wil16b, HKN^{+}09]. Infrastructures [MM06, Nic20b, OJR^{+}23, RALS23, SLW22a, YAZ21]. Initiative [GB09]. Injection [CN13, RSM13]. Injections [Puz16]. Innocent [Sch03c]. Innovate [Les09a]. Innovation [SW18b]. Input [BDL^{+}14]. Input-Processing [BDL^{+}14]. Insecure [AJQA23, BCGN16]. Insecurely [Pei20]. Insecurity [AF11, Bel19, Ber22, DXS21, Don06b, Pei22, SWYP12]. Insertion [Zdr09]. Insider [BSH^{+}09, CMS09, DCC^{+}09, GMC^{+}08, PS09]. Insiders [PPHB08]. Insight [AJQA23]. Insights [CR20, KvdHK^{+}14, WMD20]. Inspection [TA23, TPMT15]. Inspires [SPG^{+}19]. Inspiring [LL16]. Installation [BV11a, SMJM11]. Instant [Qua03]. Instituting [RAAB20]. Institutional [Ell14]. Institutions [SB21]. Instruction [AR18, Ker09]. Insurance [CIL^{+}21, WM20]. INSuRE [SDC^{+}17]. Integer [BGG^{+}22]. Integers [Sea06]. Integrating [Far21, MA12, OdH12, PHR20]. Integrity [Ano17e, Ano21-52, BS24, KS13b, Mal23]. Intel [CCX^{+}20]. Intelect [Ano14n]. Intellectual [FREP17]. Intelligence [AACG19, Bel14a, Ber21a, BS20b, Far21, FOM23, Gui11, Jon06, Lan08b, MBP^{+}24, Mic21, Mic22, MMR24, Raa19, Sch18b, SSH22, TCC^{+}24, TH13, TMU23, Ano23r]. Intelligence-Driven [Gui11]. Intelligence-Generated [MBP^{+}24, MMR24]. Intelligence/Machine [FOM23]. Intelligent [Ano19p, CUCD14, DPP15, KT20, SB19]. Intensive [CCLS21]. Intention [EHK^{+}04]. Interaction [HAO^{+}22, Smi03b]. Interactions [MWM13]. Interadministrative [OZ09]. Interdependence [DMS^{+}15]. Interest [Sch19b]. Interests [Cat09, CS09, EK09, KE09, Mas09, McL08, Nah08]. Interface [Ano03b, Ano07a, Ano09b, Ano18g, Lev04c, LB04, SC04]. Interfaces [CS09, Her06, SS05]. Interference [NBH08]. Intermediary [Cur06]. Intermediation [Gru16]. Internal [Bon08]. International [Ano15c, Ano17-38, BGM^{+}06, Don10a, Hea03a, Vig11]. Internet [Ano19q, AED^{+}21, ASE21, ABAGOB10, Ano17-39, Ano18-29, AEY10, Baa03, Bel17a, BCG06, CHK^{+}05, Cal13, CHK^{+}20, CMT^{+}19, Cha12, CK10, CEL^{+}19, DRS16, EJ06, ENDAC22, Eps13, Fel06, FREP17, FG08, FD10, Gee15c, GR13, Gra13, HL13, HRMS^{+}21a, HKW06, JKSM19, Kau11, KKK^{+}21, Kuh16, KPM^{+}19, Lan05a, Lan09b, Les08b, LK17, Mar15b, MHRS19, McG03a, MSW09, Mil03, MA15b, MKF21, NK12, NSLW20, Opp07, Pfl04, PLW07, Sch06a, Sch08a, Sch17a, SS19a, SS19b, THM15, TB19, VWC19, Var10, WMD20, vS19]. Internet-of-Things [Kuh16]. Interoperability [CW08b]. Interoperable [VVSL12a, VVSL12b]. Interoperable-Medical-Device [VVSL12a, VVSL12b]. INTERPOL [Tzi18]. interrelationships [Ada05]. Interventions [MDD^{+}15]. Interview [Irv11, McG09a, McG06a, McG06c, McG06b, McG07a, McG07b, McG08d, McG08c, McG08e, McG08a, McG08b, McG08f, McG09d, McG09e, McG09f, McG09b, McG09c, McG10a, McG10b, McG10e, McG10d, McG10c, McG11e, McG11a, McG11d, McG11b, McG11c, McG12a, McG12e, McG12b, McG12f, McG12d, McG12c, McG13d, McG13e, McG13f, McG13c, McG13h, McG13g, McG18a, McL05a, SZ17a, Say08]. Interviews [SD16a, SD16b, SZ17b, SZ17c, SD17]. Intimacy [HKMR24]. Intimate [MOT^{+}17]. Introducing [Ano22-66, Nan09]. Introduction [Ami05, AS05, And06b, Ano18-30, Arb04, AM04, CMV09, CV10, Chu08, CG04, DR06, DR04, ES11, GS06, GA10, IR11, KS08b, LS09, Lan13a, LS12, MW23, MD09b, MC09, MKF21, Pet06, PTP07, PIK12, PH03, SEKR18, TP11, TP08, VM10, CL13, GL14, OP15, PKBS15, PM14, Pfl16b, SS16]. Introspection [NBH08]. Introspections [JBZ^{+}15]. Intrusion [ABK^{+}04, CRBM06, Che11, MK13, NS11, SWL21, VNC^{+}06]. Intrusion-Tolerant [NS11, VNC^{+}06]. Intrusions [Cur06]. Intrusive [AJW13]. Invariants [PEC^{+}14]. Invent [Ano16i]. Inventories [SOG^{+}19]. Investigating [BW07, CGC19]. Investigation [Ano17-38, MB19, PS23, YR09]. Investigations [Bel12b, O’R17]. Investigative [Sim21]. Investigators [Cam03b, MD09a]. Investing [Mic22]. Investments [CHMO07]. Invisible [Don11a]. Invited [SS19b]. Inviting [GK14]. Involving [O’R17]. Inyo [BFS^{+}21]. iOS [dQSzL13]. IoT [Ano17-48, BBS19, Ber23, HFT^{+}19, JKSM19, MDR^{+}20, RORL22, RSWO18, Sch17b]. IP [Alj03, Ker10, MNS08, O’R17, WK05]. iPhone [Ano17d]. IPv6 [DGU^{+}12]. Irregular [App11]. ISBN [For04a, Pfl04]. Island [Spa06]. Isn’t [MPM22, TCC^{+}24]. Isogeny [Lau17]. Isolating [Pei20].

Isolation [BPFG24, NB23, ZYG15]. Issue [Ano16f, Ano16e, Ano16d, Ano16g, Ano16h, Ano16-44, Ano17o, Ano17n, Ano17l, Ano17p, Ano17m, Ano17q, Ano17r, Ano18b, Ano21-81, Cyb03c, Ano22-40]. Issues [AR24, Ano21-81, Baa03, BPR^{+}04, BS13a, CFH05, CWVM22, CR09b, Far20, GG04, KS08a, KHLF10, MR08, MD09a, MBA^{+}22, Pfl04, SEKR18, Slo14, Wai16, WAF11]. Italian [BGLEP15]. Items [SOG^{+}19]. Iterated [BM10]. Itself [Gee15c]. Ivan [McG10f].

Jacob [McG16a]. James [Spa08]. Jamie [McG16b]. Jamming [CUCD14, Ree16]. Janca [McG18b]. Java [BBB^{+}23, vO23a]. JavaScript [Zdr09]. Jean [McG15d]. JELL [GA04b]. JELL-O [GA04b]. Jennifer [Don04g]. Jeremiah [McG09e]. JetBlue [AHB04]. Jewels [MA12]. Jim [McG16c]. Job [Ano14d, Ano14c, Ano15h, Ano15i, Ano15j, Ano15k, Ano17-27, Ano17-29, Ano17-28, Ano17-49, Ano21-67, Ano22-52, Ano22-54, Ano23-41, Ano23-42, HL13]. JobBoards [Ano21-68]. Jobs [Ano14o, Ano19-52, Ano19-53, Ano19-54, Ano19-58, Ano19-59, Ano19-60, Ano20-35, Ano22-55, Ano22-56, Ano22-57]. JobSite [Ano22-58]. Joffe [GM14]. John [McG07d, McG11d]. Johnny [Cyb06b, HM12, RS19]. Joined [DR06]. Joining [FB04c]. Jolly [Sty05b]. Jon [Sty04a, McG08e, McG14b]. Jose [Ano20z]. Journal [Ano21x, Ano21y, Ano21z, Ano21-27, Ano21-28, Ano22o, Ano22q, Ano22p, Ano20l]. Journey [DeL14, LPS^{+}23, RS19]. Judgment [Kan09]. Jurisdiction [Bel17a]. Just [Pfl10, PM21, RI12]. Just-in-Time [RI12].

K-16 [SSW21]. Kafka [BV06]. Karger [SLZ^{+}10]. Kate [McG17a]. Kathleen [McG18a]. Katie [BC23b, McG15c]. Kay [McG12d]. Keep [Ano15-28, FB04d]. Keeping [Cal03a, Man13]. Kelly [McG17b]. Ken [Ano18-40]. Kennedy [Ano18-40]. Kent [McG10e]. Kerf [ABK^{+}04]. Kernel [Arc04a, LGO06, Wu10]. Kernel-Level [LGO06]. Kernels [ZYG15]. Key [Bel15a, BGG^{+}22, GSB^{+}04, KNL22, MGH^{+}15, PLK23, PKW04, Pfl12a, Sas18, SZ06]. Keyed [Bla03]. Keys [CC10, PLK23]. Kin [AH17]. Kind [RW21]. King [Cyb05]. Kingdom [NRH^{+}22]. kingdoms [TCM05]. KISS [KYEV^{+}18]. Knaves [CEL^{+}19]. Knife [Cyb04a]. Knights [CEL^{+}19]. Know [FH12, Pfl06, Ros07, BG13]. Knowing [TAGM23]. Knowledge [Ano18-42, BM05, FOM23, Lyo24, MBC^{+}18, NRH^{+}22, RDC^{+}18, TG07, VKM^{+}15, van19, vO21a, Ano18-37, Ano18-41]. Known [AFO13, SPL^{+}24]. Knows [CM23]. Kobayashi [CC11]. Kocher [McG11e]. Kohno [McG14f]. Kong [LZ16]. Korea [Les07b, PCK21]. Krebs [McG15b]. Ksenia [McG17c]. Kubernetes [MBR^{+}21].

L. [McG15d]. Lab [Du11, TI18]. Label [ENDAC22]. Labeling [Sch09b]. Labeling-in [Sch09b]. Labtainers [TI18]. Labware [BYG^{+}14]. Lack [MRF^{+}24, Pee13]. LAN [BPVS04, Sty04b]. Land [Arc05b, Sch18a]. Landau [BC21]. Landscape [Ahm07, Ano17w, FO08, Ksh15, Lan08a, Ort09b]. Landwehr [BC23a]. Langner [McG11f]. LangSec [AALS21]. Language [Bez24, Hut24, LDZ^{+}22, PGNRV24, vO21b, HJK08]. languages [Sal05]. Laptop [PGT07]. Laptops [TP06]. Large [Ano22-74, AKN12, AEP^{+}19, BCDS22, Bez24, CR20, HQD^{+}03, HMPS15, HN10, KKK^{+}21, LGO04, O’R17, PGNRV24, YCU^{+}21, ZOC^{+}22]. Large-Impact [HN10]. Large-Language-Model-Powered [PGNRV24]. Large-Scale [Ano22-74, AEP^{+}19, BCDS22, HQD^{+}03, HMPS15, KKK^{+}21, O’R17, YCU^{+}21, ZOC^{+}22].

Largely [Kob19]. Last [Ano17-48, Bel12a, Bel13a, Bel13b, Bel18a, Ber21a, Gee13b, Gee14a, Sch13d, Sch23, SW03b]. Lasting [Gel14]. Lately [Sty07b]. Latency [DMS07]. Later [MS21a]. Latest [Ano22-88, Koh22e]. Lattices [Lau17]. Launching [Say08]. Launchpad [Jak19]. Law [Bel21, Cat09, For04b, Gee12c, GN06, Gup17, KvS14, Raa19, Ten16, Tzi18, BG13]. Lawful [Bel21, MV22]. Lawrence [BC23a]. Laws [HDB08, Sch19a]. Lawyers [DR06, WB22]. Layer [Kob19, Spr11a, Spr11b]. Layered [Bel19]. Lead [SP09, Tsa07]. Leadership [Ano16z]. Leading [Ano15-43, Ben15a, Mic22]. Leading-Edge [Mic22]. Leakage [Bye04, GSLA15]. Leaked [And03a]. Leaking [Gar14e]. Leap [OBS^{+}20]. Learn [Bra07, Gee08b, KCL^{+}23]. Learned [AQP^{+}23, DOP^{+}21, Dhi11, Eps08, GSB^{+}04, HPSR18, KMP^{+}11, Pfl15d, Pfl15a, SPT19, Vie11b, WL11]. Learning [ARC19, Ano06b, Ano21-41, AQP^{+}23, BO22, Ber23, BYG^{+}14, BF07, CCF^{+}20, CKPP23, Cel16, DSJK19, De 21, DBR22, DO22, Ell14, FOM23, JFK21, KL23, Kuh16, MPM22, Pfl13c, Pfl15c, RDK19, RØTM08, Rui06, SPG^{+}19, SB19, SLW22b, SS22]. Least [Dre16, Ede10, Sch03a]. Least-Cost [Ede10]. Leaves [Wil19]. Ledger [EGB18]. Ledgers [Mei18]. Lee [Sty04b]. Left [FRA10]. Legacy [Smi11, TP08]. Legal [Baa03, Bor13b, Cur06, EPH12, Hec03, Kan09, Kos15, Lan16, MD09a, Mil03, PS23, Pfl04, RM22, Rya03, SNS10]. Lens [RBK23]. Lesley [McG17d]. Less [Ano15r, Ano17-39, Eps15b, FLDGB19, Gee12a, Gee15c, Les03d, Ano15s]. Less-Buggy [FLDGB19]. Lesson [Don04d, Ell14, Pfl15a, RC04]. Lessons [AQP^{+}23, AvEB15, BDGS04, BF05a, DOP^{+}21, Dhi11, Don08b, Eps08, Far20, GSB^{+}04, HPSR18, HMKT07, KMP^{+}11, Les15c, LJZ12, MBA12, Pay04, Pfl07, Pfl15d, SPT19, SLM^{+}23, TGC16, TP06, Vie11b, WL11]. Lest [BDT17]. Let [Rya03]. Lethal [MV22]. Letter [MHV18]. Letters [Ano03c, Ano03d, Ano04b, Ano06e, Ano09b, Ano18x, Ano19-61, Ano19-63, FBG04, LB04, SC04, WCV^{+}04]. Level [LGO06, OC12, PEL^{+}18, dQSzL13]. Levels [Ded17]. Leverage [MP21]. Leveraging [DO22, SWCS21, WHD^{+}09]. Lexicon [Gro20]. Liability [Bel23, CIL^{+}21, Hec03, Mac06, Rya03]. Liar [Heb22]. Liberties [San21]. Liberty [Mic10a]. Library [Ano17-82, Ano18-52]. License [Les15a]. Licenses [SSH22]. Licensing [MD09a]. Life [Don11b, GC10a, GSB^{+}04, GLM11, Ras11, RTL09, Sag13, SCMS18, VC08]. Life-Cycle [GLM11]. Lifecycle [KG17]. Lift [Gee11c]. Lifting [Lin09]. Light [Don08a, Les04d, Pfl13a]. Lightweight [GLS23, MO12]. Like [BHJM19, Bel14c, DK10]. Limitations [MA20, NB23]. Limiting [BFS^{+}21, LS12]. Limits [AAB13, Cat10b, THM15]. Line [KRK18, Lan10a, Les07a]. Lines [PMB^{+}14]. Lining [Les14b]. Link [Arc03b, BM10, SS06]. Linking [Sav17]. Links [CR09a, Cyb06a]. Linux [SK04]. List [Ano18-44, Don04f, VWW05]. Listing [LKL^{+}19]. Litigation [EJ06]. Little [MOG^{+}20, Les03a]. Live [HBN09, Wil21]. Liveness [AMF15]. Living [AF11]. Location [BBL^{+}17, GL04, Ten16, Wha11a]. Location-Tracking [GL04]. Lock [SHB20]. Locks [Bla03, Sch03d]. Log [SLW22b]. Logging [CP09a, CP10a, Mar15a]. Logic [Ano17-69, AJQA23, HRK21]. Logical [Dat10]. Long [vdG17]. Long-Term [vdG17]. Look [Bel14c, DP18, LC13, LSBB22, Pau05, Smi12a, SDKM20, Win03, Yav24, vO23a]. Looking [Ano17-49, LRW^{+}23, Nan09, Pfl16a, Sho22]. Looks [Sch16a]. Looming [Mil03]. Loop [Hor14a, Smi03b, Wil22]. Lose [Kup06, Ost20]. Loss [GSLA15]. Losses [LS06]. Lost [GM12, MTA23, PY06, TEN12]. Lot [MOG^{+}20]. Love [SE15, Lan13a]. Low [DMS07, Man13, Max20, dQSzL13, TJS^{+}22, THM15]. Low-Cost [Man13]. Low-Energy [TJS^{+}22, THM15]. Low-Latency [DMS07]. Low-Level [dQSzL13]. LTE [BS13a]. LTE/SAE [BS13a]. Lucky [SHB20]. Luddites [Les10a]. Luke [Don04h, SHB20]. Lum [McG17b]. Lurks [SDKM20].

M [McG15e]. M2M [Bai12a]. Machina [Don04c]. Machine [ARC19, Arc07a, AQP^{+}23, BO22, Ber23, CCF^{+}20, CKPP23, Cel16, DSJK19, De 21, DBR22, DO22, Eps15a, FOM23, KS08a, KL23, LWLL10, MBP^{+}24, MPM22, NBH08, Sch13d, SLW22b, SC04, SS22, WHD^{+}09]. Machines [GBK09, Mic19, SZ05, PEC^{+}14]. Mae [Der13]. Magazine [Ano13c, Ano14p, Ano19g, Ano19m, Ano19q, Ano19p, Ano19t, Ano19x, Ano19-37, Ano19-77, Ano20b, Ano20h, Ano22w, Ano23j, Ano23k, Ben14, Cyb03a, Ano19l, Ano19-27]. Magenta [Gee15a]. Mail [Opp07]. Maintaining [MJF07]. Major [Ano15u, Ano16u, Ano16v, Ano17-42, Ano17-41, KS12a, LXS22, Bel12b]. Majority [CXG^{+}18]. Make [Her13, Mar13b, MDR^{+}20, PRB^{+}12]. Makers [Sch20c]. Making [Ano17u, Ben23, Cra03, Fra18, KTT19, Lan13b, Lan14b, LA10, PR12a, RP09, SG18, SB19, ZS19, AG05, Kni17]. Malcode [For05]. Malicious [ARvD12, Bon16, CS09, PH18, Ste11, Zdr09, Hol05]. Malware [BMM07, CMZ07, CLS07, CGC19, CKPP23, CM22, CPC^{+}18, DBR22, Der13, FA09, Gif10, GG11a, Hol15a, Hol15b, LH07, MB19, MC15, MA15a, MRF^{+}24, O’D08, OSM11, RCT06, SCHC24, WHF07, YZA08, Zan09, ZRM14]. Malware-Halting [Hol15b]. Mamilli [McG18f]. Man [CCR09, Cyb05, Don03a, DGU^{+}12, Sch13d, Pfl15b]. Man-in-the-Middle [CCR09]. Man/Machine [Sch13d]. Management [Acq08, AA12, Bac13, BP07, BW07, BSSB07, BMZ14, BS13b, BHRR07, Cam10, CCF^{+}10, DD08, DP18, Eva15, FH06, FE11, FB11, GOPB12, GLM11, HSC08, JGP09, KB06, Les08a, LLWC11, MR08, MCW08, MSW09, Opp15, Pet06, RR08, TMGP13, Trč11, Vis10, Way08, WCV^{+}04, JJ13]. Managers [HAO^{+}22]. Managing [How09b, KF13, OZ09, PTP07, RS06]. Mandatory [ZS19]. Manico [McG16c]. Manipulation [YCU^{+}21]. Manual [CRBM06]. Manufacturing [TWC^{+}15]. Many [Ano18p]. Map [CCLS21]. March [Hea03a, GC08b]. Marcus [McG06b]. Marie [McG17e]. Mark [Ano15-29]. Market [ABAGOB10, BL13a, BOR^{+}13a, Gup17, Lan07a, PS21, PKK05, MM10]. Marketing [MRF^{+}24, Whi15]. Marketplace [PLW07, SRG03]. Marketplaces [Hu16]. Markets [ME10, Pet07, YCU^{+}21]. Markup [HJK08]. Mars [Smi05b, TGS20]. Martin [McG16d]. Maru [CC11]. Masks [IAM03]. Mass [Whi15]. Mass-Marketing [Whi15]. Masses [CW12, Fel06, TB19]. Massively [BCLM09, MKKP09]. Master [Bla03]. Master-Keyed [Bla03]. Masthead [Ano06f, Ano13d, Ano14q, Ano14r, Ano14s, Ano14t, Ano14u, Ano14v, Ano15-30, Ano15-31, Ano15-32, Ano15-33, Ano15-34, Ano15-35, Ano16-33, Ano16-34, Ano16-35, Ano16-36, Ano16-37, Ano16-38, Ano17-50, Ano17-51, Ano17-52, Ano17-53, Ano17-54, Ano17-55, Ano18-31, Ano18-32, Ano18-33, Ano18-34, Ano18-35, Ano19-64, Ano19-65, Ano19-66, Ano19-67, Ano19-68, Ano19-69, Ano20-36, Ano20-37, Ano20-38, Ano20-39, Ano20-40, Ano20-41, Ano21-69, Ano21-70, Ano21-71, Ano21-72, Ano21-73, Ano21-74, Ano22-67, Ano22-68, Ano22-69, Ano22-70, Ano22-71, Ano22-72, Ano23-45, Ano23-46, Ano23-47, Ano23-48, Ano23-49, Ano23-50, Ano24-35, Ano24-36, Ano24-37]. Material [Tro06]. Materials [BBB^{+}23, TAGM23, ZLT^{+}23]. Mathematics [OP15]. Matrix [JZ04]. Matt [McG08f]. Matter [Wis18]. Mattered [Wha11b]. Matters [Ano14w, Cal04d, GFJ^{+}18, Red24]. Matthew [McG14c, McG15e]. Mature [Tro08]. Maturity [ABPP16, SB21]. Maximizing [Gri04]. Mayhem [ABD^{+}18]. McGraw [McG09a, Ran16]. Me [FC14, Heb22, Hor16]. Mean [LB08]. Means [Bel14a, Sch23]. Measure [GC08c, GC08d, GC09a, GC09b, GC09d, GC09c]. Measured [DMB^{+}22, Mic22]. Measurement [HvO18, PEL^{+}18, Pfl12b, PSB^{+}07]. Measurements [CL12]. Measures [ACPZ20, AFFOG12, Gee11e, RM22]. Measuring [AK03, And12, Ano17t, LWLL10, McG03a, PC10, SBE11]. Mechanical [Bla03, Les15c, SBB^{+}18]. Mechanism [ASE21, DEKM^{+}21]. Mechanisms [LK17, MH23, MK13, PCK21]. Media [Hei07]. Mediated [Cal13]. Mediating [McL13]. Medical [AIKR13, Les13b, Sch22a, VVSL12a, VVSL12b]. Meet [Kau10b, Pfl14a, Roe17]. Meets [Kal12]. Melissa [Say08]. Meltdown [GY21]. Member [Ano20-34, Ano21-60, Ano21-61, Ano21-62, Ano21-63, Ano21-64, Ano21-65, Ano22-49, Ano23-34, Ano23-35, Ano23-36, Ano23-37, Pal04]. Members [Ano03f]. Membership [Ano08a, Ano14w, Ano16-39, Ano16-40, Ano17-60, Ano17-64, Ano17-61, Ano17-62, Ano18-42, Ano17-63, Ano18-37, Ano18-41]. Memorability [YBAG04]. Memoriam [SLZ^{+}10]. Memory [Ano17-59, Gue16, Les12d, PLK23, SPWS14, vO23b, vO23a]. MEMS [ECG^{+}07]. MEMS-Assisted [ECG^{+}07]. Menlo [BDKM12]. Mental [BLCDK11]. Merry [CCW03]. Merwin [Ano15a, Ano15b, Ano16z]. Mess [Dre16]. Message [Ano17-68, Nic21a]. Messaging [HLS^{+}21, Sol07]. Metadata [GT22, Sch14b]. Metaphor [FB04d, HSS11, Ver16, Ada05]. Metaverse [RK24]. meter [Sah05]. Methodological [MR05]. Methodology [AR18]. Methods [CRBM06, Dat10, De 24, MSeF^{+}22, MLS22]. Metricon [SP07]. Metrics [Bel06c, NP07, Pet06, RR08, San14, Sur16, ZKH^{+}23]. Mettle [Res16]. Michael [Ano21-39]. Micro [ASK07]. Micro-Architectural [ASK07]. Microarchitectural [GY21, SG20]. Microcosm [O’D07]. Micropayments [Les04c]. Microservices [Fet16]. Microsoft [Chu08, KS12b, Mic08b, Ten16]. Middle [CCR09]. Middleware [VNC^{+}06]. Midlife [MS20]. Might [Ano17-39, Les03e]. Mikko [McG07b]. Mile [Ano17-48]. Military [Bel13a, DR05]. Miller [McG14a]. Mills [Ano17-43, McG11b]. Mind [CM22, Les12b, Smi12b, SE15]. Mindset [BC08]. Mine [CBOS20, For04a]. Minimalist [Tro08]. Mining [Lan09a, VC04]. Misaligned [Cam10]. Misconceptions [MP22]. Mishap [Bai12a]. Misinformation [PGNRV24]. Mission [Ano17-84, Cyb03a, KLL17]. Missteps [Pfl12b, PT05a]. Misuse [HMA04, PS06]. Mitigate [BSH^{+}09]. Mitigating [MTS15, PA03, RSM13]. Mitigation [AvEB15, MBA^{+}22, Puz16]. miTLS [BFK16]. Mixed [CDJS^{+}24, MP21, OJR^{+}23]. Mixing [All07]. MMM [SMGK04]. MMM-ACNS [SMGK04]. Mobile [Ada15, Ano14-28, Ano14-29, Ano17-56, Ano17-66, BYG^{+}14, Car08, DeF11, Dim07, EKA14, FD11, GN06, JMD06, LC13, LAYG16, Mar13a, May15b, MA15a, Mil11, OMB17, OY16, RdRdCG10, SFE10, VM10, Wha11a, WBT^{+}18]. Mobility [Amo13, VM21].

Mobility-Enabled [Amo13]. Mode

[Mar10]. Model [BLR19, BLCDK11, Cha12, CW08b, KLL17, PGNRV24, RVK05, Sah05]. Model-Driven [KLL17]. Modeling [AK15, BM11, Dhi11, HWF^{+}10, Nic05, SGSM22, Ste10, SVR^{+}19, Tor05, YCU^{+}21]. Models [Ano17-74, Bez24, BW10, CP15, CKPP23, DFG^{+}11, Don09b, Roe17, RP09, YG19, Sch05a]. Modern [GG11a, Opp15]. Modest [KS12a]. module [KN13]. Moe [McG17e]. Money [Les13a]. Monitoring [Cal04d, CFN13, KA13, Kuf13, MPS14, Por06, Por09, SLW22a, Spr11a, Spr11b]. Monitors [KS08a]. Monoculture [BS09, GAW03, LS09]. Moonshot [Okh21]. Morris [Orm03]. Morrisett [McG10c]. Most [DMB^{+}22]. Motion [CBE^{+}12, NROS24]. Moussouris [BC23b, McG15c]. Move [FHMF14]. Moving [AJV18, Ano15-28, Bai12a, CF14, GL14, Hea03b, Lan03, NL12, OHBS14]. Moving-Target [CF14, OHBS14]. Much [KNBV17, Str12]. Muddling [Sch15]. Mudge [McG16e]. Mules [Hu16]. Multi [Ano15-43, CEK^{+}10]. Multi-Core [Ano15-43]. Multi-Tenancy [CEK^{+}10]. Multidisciplinary [Yav24]. Multilevel [Say04]. Multimedia [May10, Ano19t]. Multiplayer [BCLM09, MKKP09]. Multiple [DN07]. Multipliers [HLJ21]. Multivariant [Fra18]. Multivariate [DP17]. Music [Les03a]. Must [AS13]. Mutating [BMM07]. Mutual [GD17]. Mx. [Koh22b]. My [Eva14, Les12b, Pot10a, Sch06a, VB13, dAMM13]. myCS [Ano17-58, Ano17-57, Ano18-36]. mysteries [For05]. Myth [RCT06, SS16]. Myths [SSH^{+}16].

N [GSHU08]. N-Version [GSHU08]. Name [CR06, Che06, Don05a, MD09b, Sel11]. NAND [Ano17-59, dQSzL13]. Nanocomputing [AKN12]. Narrowing [Wil24]. Nate [McG14d]. National [Bel18b, DHR^{+}04, HRDR05, IP10, Lan09b, Ste13, WWH10, HKN^{+}09]. Nationwide [GN07b]. Natural [LDZ^{+}22]. Navigation [RSOD22, Ver16]. Near [OY16]. Near-Field [OY16]. Necessary [Pot10b, Sch19a]. Need [ACPZ20, AEH^{+}04, Cal04f, CLN12, CPC^{+}18, Cyb03a, GS16, OGM^{+}24, Sty04b, WZE18]. needle [Pfl13c]. Needs [Pfl14a]. Negotiation [BSSB07]. Negotiations [KP15]. Neighbor [AM12]. Neil [McG12e]. Network [AJW13, Ano08d, BSH^{+}09, CCW03, CRBM06, CGR^{+}22, Dim07, DRS16, HCL11, Hay13, JMD06, Les10b, LAYG16, Mar05, NSSS08, PEL^{+}18, Por09, RBBK04a, SLKJ21, SCZ^{+}13, SMJM11, Sch08a, SMGK04, Sty04b, SK05]. Network-Based [Por09, SMJM11]. Network-Level [PEL^{+}18]. Networkable [OKH07]. Networked [MPS14]. Networking [Ano17-73, Ano19-55, Ano19-56, Ano19-76, ES14, KYEV^{+}18, LLT14, MBR^{+}21, Ros07, Sch10c, AS13]. Networks [ACL07, AS15, Ber19, BS13a, Bon16, Car08, CF14, HQD^{+}03, HNE^{+}08, LGO04, LAYG16, LS08, MK13, NSLW20, PBW^{+}08, RH06, Sch06c, SWYP12, Smi11, GD13, KS13b, Kri13, dAMM13]. Neumann [McL05c]. Neural [CGR^{+}22]. Neutrality [Les10b, Sch08a]. Never [Ark11, CM22, Hag09, RP10, SE15]. Never-Ending [Hag09]. News [And03b, And04, Ano03e, DK05, DFK05, DFK06, FO08, Fig09a, Fig09b, Fig09c, Fig09d, Fig10a, Fig10b, Gar11a, Gar11b, Gar11c, Gar11d, Gar11e, Gar11f, Gar13a, Gar13b, GM03, GA03, GAA04, GA04b, GG04, Got04, GA04a, KDF06, KD06a, KD06b, Law10a, Law10b, Ort06a, Ort06b, Ort07a, Ort07b, Ort07c, Ort07d, Ort07e, Ort07f, Ort08a, Ort08b, Ort08c, Ort08d, Ort08e, Ort09a, Ort09b, Pau10b, Pau10c, SAA04]. Next [BL13b, Cam03b, Gif10, KMP20, LL16, MR14, McL06, Pfl12b, Sch04b, MR05].

Next-Generation [MR14]. NICE

[NCA19, PMNT12]. Nick [McG18e]. Nicole [McG17f]. Ninjas [KW12]. NIST [MMKP16, RRS06, Sch22b]. NIZKCTF [MBC^{+}18]. No [CM23, Don05c, FRA10, GB12, Sty04b, Vie12d, WCV^{+}04, Ano03b, Lan13a]. Nomination [Ano22a]. Nominations [Ano15g, Ano15u, Ano16c, Ano16u, Ano16v, Ano17a, Ano17k, Ano17i, Ano17j, Ano17-42, Ano17-41, Ano20a, Ano23b, Ano23-66]. Nominees [Ano15e, Ano16c, Ano17k, Ano17i, Ano17j, Ano16b]. Non [RIC17]. Non-Tech-Savvy [RIC17]. Noncompliance [Ren12]. Nonconfidential [CRK^{+}13]. Nonfunctional [HWF^{+}10]. Nonidentifiable [KM22]. Noninteractive [MBC^{+}18]. Nonsecurity [Sch07d]. Normal [Mus21]. Normalization [BMM07]. North [Ami03]. Northwest [Les11c]. Norwegian [HMKT07]. NoSQL [Puz16]. Note [BV06, Var09]. Notes [Bra07]. Nothing [GC10b, KNBV17, Sho23]. Notice [Cat10b]. Notifications [ZS19]. NRC [Lan09a]. NSA [Ano17f, Lan13b, Lan14b]. NSF [Eps16, LRW^{+}23]. Nuclear [RSWO18]. Nudging [Acq09, BC14, FG22]. Numbers [Gee12b, Sty07a]. Numerology [GG11b]. Nutrition [ENDAC22].

O [GA04b, KS08a, SND14, ZYG15]. Oakland [BS20a]. Obfuscation [BPB^{+}04, DRS16, OSM11, XL16]. Objectives [BF07]. Oblivious [DK10]. Obscurity [Don06b]. Observable [HvO18]. Observation [NBH08]. Observational [FWBC15]. Observations [EW22]. Obstacles [Mei18]. Octopuses [Hor15]. Off [Ano22-88, GHS14, Koh22e, Ano22-66]. Off-Path [GHS14]. Offense [Lin09]. Officer [CM23, Eps16]. Officers [RVG24]. Offline [Cal03a, Lev03a]. Oh [Sch06a]. OJ [Ano22-59]. Old [Che17, Don09b, FREP17, GY13, Les13d].

Olof [McG13d]. On-Demand [ZYG15]. On-Disk [BM14]. Onboard [DDAN12]. One [Ano18-37, Ano22-66, Ano22-88, CM23, Cyb05, FD11, GMB12, Koh22e, OBS^{+}20, TEG^{+}19]. One-Eyed [Cyb05]. One-Time-Password [FD11]. Ongoing [Gee23a]. Online [Ano17w, BCLM09, BMS08, BGL^{+}23, Cal04d, Cha09, Cra12, Ede10, Ede14, EH13, GT22, HMT06, Hu16, Hut24, Kan09, Kos15, Lev03a, MH07, MC09, MKKP09, Phi19, RIC17, SLW22b, THS20, Vil04, Wis18, YR09, GD13, HL13, JB05, Kri13, dAMM13, Ano08c, VVYY11]. Only [EH12, KNL22]. Ontology [Don03d]. Opaque [Kob19]. Open [AR24, Ano20l, Ano21x, Ano21y, Ano21z, Ano21-27, Ano21-28, Ano22-60, Ano22o, Ano22q, Ano22p, Bel22a, Blo13, CR09b, Cow03, Cow23, HM22, HNE^{+}08, KSL^{+}20, Lan14a, Lan09b, MSW09, OMB17, PGNRV24, PW17, SPL^{+}24, SK04, SSH22, SS19b, YCU^{+}21]. Open-Source [Cow03]. Openers [Mar15b]. OpenSSF [ZKH^{+}23]. OpenTitan [MRO^{+}23]. Operating [Cur06, KMP^{+}11, LRKR18]. Operation [OJR^{+}23, Pei22]. Operational [BMZ14, WA07]. Operationally [Pot10c]. Operations [GD11, PE12]. Opinion [Ksh17]. Opportunities [AZ20, AMF15, AS22, Ano17-56, Ano17-73, LL23, Lau17, LAYG16, Mic10b, NROS24, NB23, PGNRV24, Raa19, RLB^{+}23, THM15, Tra09, YAZ21]. Opportunity [Tzi18]. Oppression [Sch13c]. Optical [CEC^{+}08, GSHU08]. Optical-Scan [CEC^{+}08]. Optimization [Ede14]. Optimize [Les14c]. Optimizing [CHMO07, Mat12]. Options [Ano16-39, Ano16-40, Ano17-60, Ano17-64, Ano17-61, Ano17-62, Kam14, MR08, Ano17-63]. Orbit [Pfl13c]. Order [Sty04b, TH13]. Organization [JG07]. Organizational [CPS^{+}16, CSZ^{+}14, PTP07]. Organizations [EW22]. Oriented [BBC^{+}19, Pet09, PR12b, Rog16]. Origins [SJZG19]. Other [Cur06, Pfl15c, Thi06]. Our [EPH12, Eps14, Koh22c, Pfl13c, Pfl15d, San14, Sch17a, TWC^{+}15, YRX^{+}22]. Outcomes [Eag13, PGP16]. Outgoing [Nic21a]. Outing [GR13]. Outlook [JS10]. Output [SS04]. outsourcing [PT05a]. Outweigh [Fel03]. Overflow [KCL^{+}22]. Overlay [SCZ^{+}13]. Overruns [PB04]. Overview [BN08, De 21, HMPS15, JW04, LS05a, MD09a, OMB17, WL18, GJP05]. Owe [Bel22b]. Own [Loc09]. Owned [GC09d]. Ownership [Gee19a]. Owns [BF06].

P [Spa08, Ano03b]. P3P [Cra03]. Packed [LH07, YZA08]. Packet [TPMT15]. Page [NB23]. Pain [Bac13, TW21]. Pan [GOPB12]. Pan-European [GOPB12]. Panacea [Hor19]. Pandemic [GCY^{+}22]. Pandemic-Adaptive [GCY^{+}22]. Panel [LRW^{+}23]. Paper [Ano15d, Ano16a, ASH^{+}08, Les11b]. Papers [Ano08c, Ano15f, Ano16f, Ano16e, Ano16d, Ano16g, Ano16h, Ano16-44, Ano17o, Ano17n, Ano17l, Ano17p, Ano17m, Ano17q, Ano17r, Ano18b, Ano19s, Ano19w, Ano20-27, Ano20-28, Ano21-66, Ano21-48, Ano21-49, Ano21-50, Ano21-51, Ano21-35, Ano22-35, Ano22-36, Ano22-37, Ano22-38, Ano22-39, Ano22u, Ano22-27, Ano23x, Ano23y, Ano24v, Ano24s, Ano24w, Ben17, BP18, BH22, EB19]. Paradigm [Gee16a, Hea04, KLL17, TAP19]. Paradox [Wis18]. Parental [AED^{+}21]. Parfait [CKL^{+}12]. Part [AF16, BLM18, EK09, MCW19, SD16b, SZ17b, SZ17c, SZ17a, SD17, Sch04f, SFH05, SFH06, Spr11a, Spr11b, TR09b, VVSL12b, AAG15, AA15, Han12a, Han12b, KE09, Lan14b, Mar15c, Nar13a, Nar13b, RBBK04a, RBBK04b, SD16a, VVSL12a]. Participation [CSSvdH14, Hea03a]. Partly [Dre16]. Partner [Mar13b, MOT^{+}17]. Partnership [Ksh15]. Partnerships [Ben21]. Parts [Eps15b, TWC^{+}15]. Party [Arc07b, Mat12]. Pass [LDK^{+}14]. Passed [GS03, Les04c]. Passport [Vau07]. passports [CT11]. Password [FH12, FD11, HAO^{+}22, RS11, THS20, YBAG04]. Passwords [AS22, HvO12, Sch04d, THS20]. Past [Eps12b, GD18, HMA04, Ker09, Les09a, Pei21, SDO21]. Pastures [Gra23]. Patch [GC09e, SPB12, SK05]. Patched [GBK09]. Patching [MJP21]. Path [GHS14, KT20, Mei18, ZKH^{+}23]. Paths [PCK21]. Pathways [CCF^{+}10]. Patient [BCM03, CRK^{+}13, EK09, KJ15]. Patriot [Fra07b]. Patterns [PKW04]. Paul [McG11e, SLZ^{+}10]. Pay [Sch05b]. Payment [Ano17w]. Payments [LLKA18, OY16]. PC [SZ06]. Pc01002 [LM08]. Pc01002/SpringPeeper/ED288l.6 [LM08]. PCs [SND14]. PDF [MB19, Ste11]. Pearce [McG17a]. Pearl [SE15]. Peer [AR18, Ano09c]. Pegs [Mar06]. Peguero [McG17c]. Peisert [Ano21-89, Ano22-87, BC23a]. Peiter [McG16e]. Penetration [Bis07, ASM05, Tho05]. People [BMS08, DEKM^{+}21, Eva14, HM22, Sas15, Wha11b, Wil21, HL13]. Per-Olof [McG13d]. Perceive [WDL21]. Perception [EP04]. Perceptions [Bel10b, FTCS12]. Perceptual [OGM^{+}24]. Perfect [Ami03, Red24]. Perfection [Sch09a]. Perform [JJF08]. Performance [ABPP16, KS08a, MGH^{+}15, RR08]. Performant [ZHC^{+}20]. Performing [How06]. Perils [Ano17-66]. Perimeter [Amo13]. Peripherals [Thi06, Arc05a]. Perlroth [McG17f]. Permission [Ano17-66, Ano17-74, Roe17]. Permissions [Jak20, Mic08b, WBT^{+}18]. pernicious [TCM05]. Persistence [HvO12]. Persistent [Bai12b, Ber22, BS14, SE13]. Personal [Acq09, Gee14b, KWRK13, LZ16, LMW10, OAB07, RORL22, SyPKZ22, Sta20, Sty05b, VdCA07, Wha12]. Personalization [BVB22, CS09, IAM03]. Perspective [AZ20, AF16, Bor13b, BS24, CSZ^{+}14, FB11, HRGK^{+}20, KJ15, Orm03, Pet19, RM22, SAF15, Tzi18]. Perspectives [Acq13, AM12, AAG15, AA15, BJD22, EEA^{+}23, Hea04, Lan09a, PLK23, PSO^{+}21, Pei21, RALS23, Sen17]. Persson [McG13d]. Pervasive [Ano16-42, Ano19u, Ano19v, Ano19w, Ano20m, Ano20n, Ano21-29, Ano21-30, Ano21-31, Ano22s, Ano22r, Ano22t, Ano22u, Ano24j, Ano24i, ACAT^{+}15, Kel10, Trč11]. Pests [Lev03a]. Pet [vZH^{+}19]. Peter [For04a, McL05b]. Peterson [McG09d]. Petri [Nic20a]. Pfleeger [BC23a]. PGP [McL06]. Phagocytes [Don10b]. Philip [McL06]. Philosopher [PB07]. Philosophy [Bin18]. Phish [SBB^{+}18, MM10]. Phish-Market [MM10]. Phishing [ACPZ20, CPFJ14, DO22, Eps14, Gra23, KS12a, MKG09, ZOC^{+}22]. Photography [Les04d]. Photos [Ano17d]. Physical [BSS10, BH14, CMT^{+}19, CV10, CZ23, GFJ^{+}18, McL13, MWM13, PMN^{+}14, TA23]. Physics [AZ20, GFJ^{+}18, MPS14, OP15]. Physics-Based [AZ20]. Physiology [HSS11]. Pie [Les04a]. Pieces [Lan05b]. Pilot [BFS^{+}21]. Pilots [DHP14]. PIN [DMA09]. Pioneer [Spa08]. Piracy [GG04, LB04]. Pirated [FD10]. Pitfalls [AED^{+}21, Phi19, BRS^{+}21]. Pixel [YE09]. Pixel-Count [YE09]. Pixels [Les11b]. PKI [Ahm08b, FBG04, Gar03, HKN^{+}09]. Places [Gee18a]. Plagiarism [KF13, Rya07]. Plaguing [Lev03a]. Plan [Ben15b, Sch17b]. Planted [BDL^{+}14]. Platform [LPJ^{+}16, MBC^{+}18, YK16, ZHC^{+}20]. Play [But17, GPD13]. Playing [BCLM09]. PLC [ZRM14]. Plumbing [Lan08b]. Plundervolt [MOG^{+}20]. Podcast [Ano22-73, Ano23-51, Ano23-52, Ano23-53, Ano24-32, Ano22x]. Point [BV11b, McG13b, Pee13, Res16, SS05, RWD03]. Point/Counterpoint [McG13b, Pee13].

Points [TW21]. Poisoned [Nis24]. Poisoning [Lev03b, RSOD22, Zdr09]. Poker [WMS10]. Polarization [Gee14a]. Police [RVG24]. Policies [And06a, AEH^{+}04, AEV^{+}07, Bel20a, Cra03, HRGK^{+}20, KG17, LWL^{+}21, SMK^{+}20b, Ver06, WMDX23]. Policy [AC21, And04, AHB04, Bel18b, CP15, CK20, FVJ19, Gar12a, Gar12b, Gar12c, Gar12d, Gar12e, Gar12f, Gar13a, Gar13b, Gar13c, Gar13d, Gar13f, Gar14a, Gar14b, Gar14c, Gar14d, KP10, Koh22c, NL23, NSSS08, Pfl13b, Sch07b, Sch20c, SW18a, Tel15, WMD20, Gar13e]. Political [App11, Bin18]. Politics [Lan13a]. pollinate [Ano18h]. Polling [HWB^{+}20]. Pond [BS20a]. Populations [PGP16]. Portable [Hei07, LA10, RH06]. Portal [MM05]. Positive [HMA04]. Possessing [Ada15]. possibility [CXG^{+}18]. Possible [Bel14a, DPP15]. Post [Don03c, MR22, Sch22b]. Post-Apocalypse [Don03c]. Post-Quantum [MR22, Sch22b]. Postel [SPB12]. Postquantum [Ano16g, Ano16h, BLM17, BLM18, Lau17]. Potential [BBD^{+}08, BGLEP15, CBN11, DDAN12, DP20, EKA14, Kal12, LS06, Per07, Phi19, SND14]. Potter [McG18c]. Power [BGC^{+}14, DSEB12, DEKM^{+}21, Fel06, Gee12c, McG03a, RBK23, Way04]. Powered [PGNRV24, SFE10]. PQChain [EGB18]. Practical [CHMO07, DBR22, Fra18, LDK^{+}14, LK17, YB21, Sah05]. Practice [BCS16, CA11, DM15a, HCRS18, Ker10, MBLT13, PKB^{+}24, PY06, PWVT12, Rog16, RST15, dA16]. Practice-Oriented [Rog16]. Practices [GS03, Gel14, KM22, MOT^{+}17, MDR^{+}20, NKJ^{+}14, HL13]. Precision [YRX^{+}22]. Predictable [Hen09]. Predictions [Ano20s, Eps12b, Mas24]. Predictive [Ano16-42, SW18a]. Preferences [SyPKZ22]. Preneel [McG15a]. Prentice [For04a, Pfl04]. Prentice-Hall [For04a, Pfl04]. Prepare [Ano18-38]. Prepared [BCM^{+}15b].

Preparing [BS20b, DLMS21, MH13].

Preplay [BCM^{+}15b]. Prepose [Ano17-67]. Prescriber [KE09]. Prescription [BCM03, EK09, KE09]. Present [GD18, PHS^{+}08, SDO21]. Preserving [ARC19, ABMS22, KP10, KL23, ST10, TJS^{+}22, VC04, ZHC^{+}20]. President [HK09, Koh22b]. Presidential [HK09]. Pretending [Smi05a]. Prevent [Gar14e, GSLA15]. Preventable [SP09]. Prevention [GSLA15, Wis18]. Price [Acq08, GC09d, Les12c]. Pricing [Les08a]. Prime [CKPP23]. Primer [BH14, Lad06]. Primitive [Val12]. Primitive-Chaining [Val12]. Primitives [GCR18]. Principle [KYEV^{+}18, Mas19, SPB12]. Principled [BDTJ19]. Principles [GLS23, MN21, MV22, PKB^{+}24, PD11, SWL21, Smi12a]. prioritization [LS05a]. Prioritizing [CP09b, CGL^{+}16]. Privacy [AG05, Acq08, Acq09, Acq13, AAB13, ARC19, And04, Ano03a, Ano14m, Ano16f, Ano17l, Ano17m, Ano17q, Ano17r, Ano17u, Ano17v, Ano17-48, Ano17-67, Ano18i, Ano21-81, AEH^{+}04, AHB04, AEV^{+}07, AEY10, AR15, ABMS22, AS15, AH17, Bab16, Bai12a, BC14, BCM03, BO22, BA18, BBL^{+}17, Bel05, Bel08b, Bel20a, Bel20b, Ben15a, Ben21, Ben23, Ber23, BS13b, BC21, BKS18, Bor15, BS24, CL13, Cal03d, Cal03b, Cal03e, Cal04d, Cal04c, Cal04f, CIL^{+}21, CFH05, Cat09, Cat11, CWVM22, CDF^{+}12, CAL^{+}24, CK10, CS09, CDJS^{+}24, Cra03, CS13, Cyb03c, DEKM^{+}21, Dat10, De 14, De 21, Der03, Don11b, EK09, ENDAC22, Eps17, Eva15, For04b, Gar12a, Gar12b, Gar12c, Gar12d, Gar12e, Gar12f, Gar13a, Gar13b, Gar13c, Gar13d, Gar13f, Gar14a, Gar14b, Gar14c]. Privacy [Gar14d, Gee13b, Gee16a, GC06, Gel14, GT22, GL04, Gup17, HSC08, Hea03a, HHK18, HRMS^{+}21a, Hoo16, HCL04, HKM17, HAF05, HDB08, JKMS23, Jak20, KP10, KC18, KL23, Kni17, Knu07, Koh22c, KAAEa17, Kos15, KE09, Kot08, Kri13, KP15, Ksh14, LPJ^{+}16, Lan08a, Lan09a, Lan14a, Lan14c, Lan06, Les09b, Les09a, Les12c, Les13c, Les13d, Les14a, LZ16, LWL^{+}21, Lyo24, MM09a, MM14, MH23, Mal23, Mas09, MOT^{+}17, MK13, MM09b, McG13b, McL08, Mic21, MKF21, MHF19, Nah08, NH23, NBLC09, NSLW20, Ort14, OAB07, PKB^{+}24, PCK21, Pee13, Pei20, Pfl06, Pfl15b, PS21, PP06b, Por06, Pow07, PGT07, PPJ03, Pre15, PM21, RBK23, RBE03, RORL22, RS06, RK24, Ros07, RXJ^{+}21, Sad16, Sad17, SD17, SAB^{+}19, San21, Sch09c, Sch11c, SEKR18, SJ03, SyPKZ22]. Privacy [SW18a, Slo14, Sol07, SE09, SW18b, Sta20, ST10, SAA04, SS22, Sty05b, TVR23, TJA10, TJS^{+}22, Tom06, TMGP13, TXL^{+}21, VC04, Wai16, Wan18, Wei04, Wei06, Wha11a, WZE18, Wis18, Woo23, WMDX23, ZH23, dA16, vO20, vdG17, vZH^{+}19, AS13, BG13, Gar13e, GJP05, GD13, HL13, IS05, JB05, Lan13a, LB04, Lin05, SPH^{+}05, SC04, dAMM13, Ano03f, Ano06d, Ano06c, Ano14j, Ano14k, Ano14l, Ano15x, Ano15y, Ano15z, Ano15-27, Ano16-32, Ano16-29, Ano16-30, Ano16-31, Ano16-28, Ano17-44, Ano17-45, Ano17-46, Ano17-47, Ano18y, Ano18z, Ano18-27, Ano19d, Ano19e, Ano19c, Ano21a, Ben16b, Ben17, BP18, BH22, CCF14, ES14, iC15, NL12, NPS14, SD16a, SD16b, Ano14-32, Ano17g, Ano18b, Ano18a, Ano19y, Ano19h, Ano19x, Ano19z, Ano19-37]. Privacy [Ano19-77, Ano20o, Ano21-32, Ano21-89, Ano22-66, Ano22v, Ano22y, Ano22x, Ano22w, Ano22-87, Ano23i, Ano23h, BO23, Pei23a]. Privacy-Aware [BKS18, NBLC09]. Privacy-Based [BS13b, Kos15]. Privacy-Enabled [Por06]. Privacy-Enhancing [PS21, SE09]. Privacy-Preserving [ARC19, KL23, ST10, TJS^{+}22, VC04]. Private [Ano04c, BB17, Cal03a, CLN12, Cat10a, CGR^{+}22, DK10, Kau11, Ksh15, KJC^{+}12, RCV^{+}12, RDK19].

Private-Sector [Cat10a]. Privateers [Les13c]. privately [SBM^{+}13]. Privilege [Sch03a]. Proactive [Ano17-68]. Proactively [SG18]. Problem [Ber22, Gee05b, Hor19, KST^{+}21, LLGJ16, Sch04f, Vie12a, Wil16a]. Problematic [LWL^{+}21]. Problems [Ede14, HAO^{+}22, KvS14, YB21, GJP05, OR05]. Process [BN08, How06, Mil05, NSSS08, PS06, SS04, Tor05, TP08]. Processes [DHR^{+}04, How04]. Processing [BDL^{+}14, HV20, LDZ^{+}22]. Processor [DXS21]. Processors [Gue16]. Procurement [Lad06]. Producing [BBB^{+}23, DHR^{+}04]. Product [CR20, FPP12, Gee05a, KvS14, Mat12, SW03a]. Productive [Ksh14]. Products [Bai12b, JKSM19, Knu07]. Professional [FB04a, Ano22-51, Ano19-36, Ano20i, Ano20j, Ano20k, Ano21t, Ano21u, Ano21v, Ano21w, Ano22m, Ano22n]. Professionals [For04b, KJC^{+}12, VKM^{+}15, Ano19r, Ano19s]. Profile [EH13, Pal04, PW17]. Profiles [Ded17]. Profiling [PRB^{+}12]. Profit [Pei23b]. PROFORMA [Ano17-68]. Program [Eps16, Irv13, NRH^{+}22, TM05]. Programmable [Ano17-69, ABPP16, AJQA23]. Programming [Ano17-67, BF04, BF05b, Bis10a, Fra18, GSHU08, PR12b, Sal05, vO21b]. Programs [ADG23, DLMS21, FLDGB19, SDKM20]. Progress [AR24, Ben23, BF08, GL12, HBN09]. Prohibition [Hor19]. Project [CMZ07, SPG^{+}19, TKZ21, ABMS22, NCA19, Spi03]. Project-Based [SPG^{+}19]. Projects [SPL^{+}24]. Proliferation [ANC10, Kel10]. Promise [EH13, PWVT12]. Promises [JKSM19, KA13]. Promising [KNL22]. Prone [CRBM06]. Proof [Kni15, Ste15]. Proofing [DMA09]. Proofs [Can14]. Propagation [CMZ07, KvS14, Zan09, Lev05]. Properties [JSS20, XL16]. Property [BC09, DMR^{+}22, Wil16a]. Proposal [KS12a]. Proposals [Ano22-40]. Proposed [CK10, Gro20, PB05, GJP05]. Protect [BC14, BCG^{+}09, LGO04, SPJ^{+}14, BBD^{+}08]. Protected [Les03d, PSB^{+}07]. Protecting [Ano04c, AS09, CDD13, CDJS^{+}24, FC14, GL04, IS05, JS07, KRK18, KJ15, SP14, SG18, Sty05b]. Protection [Bab16, BLR19, BSC^{+}08, CM23, ECG^{+}07, GTG07, HRGK^{+}20, PLK23, Pet19, PP06b, Red24, SS19a, SW03b, SS06, TWA08, WMS10, Wu10, Les05]. Protections [Mil03]. Protest [MSM^{+}22]. Protocol [BCM^{+}15a, BFK16, CCR09, CUCD14, FG08, Her06, TB19, MM10]. Protocols [BCDS22, CB15, DM15b, Lyo24]. Provable [DPW11, Rog16]. Prove [Pei23b]. Proven [Sch03c]. Provenance [BH19, Gee16b, McD11]. Proverbs [GC09e]. Provide [Kau10a]. Provider [Pau10a]. Providers [Wil16b]. Providing [DeL07, NCA19, Opp07]. Provocation [DH13]. Proximity [RC17]. Proxy [Ano17z]. Pseudo [Ano22-74]. Pseudonymized [Sta20]. Psychology [CSZ^{+}14]. Public [BC21, BGG^{+}22, Cha12, GSB^{+}04, Kau10b, KNL22, Ksh15, LZ16, MM11, Mar13b, PMNT12, San21, Sch19b, Sch23, SW18a, TCJ^{+}19]. Public-Cloud [Kau10b]. Public-Health [Cha12]. Public-Key [BGG^{+}22, KNL22]. Public-Private [Ksh15]. Publication [MMKP16]. Publications [Ano23-38, Ano24-38]. Published [Bye04]. PUFs [Han12a, Han12b]. PURE [GCY^{+}22]. Purpose [Gar08, GS07, Gue16]. Pursuit [Red24, SLM^{+}23]. Push [HL03, LH23]. Pushed [Nic20a]. Put [BH19, BS09]. Putting [CIL^{+}21, CCS06, MMKP16, Sur16]. Puzzle [Bel08b, Lan05b].

Qmags [Ano16-32]. QoS [AJW13]. Qualitative [PR12a]. Quality [AFFOG12, Ano23b, Ano23-66, BW07, TKZ21, TAGM23]. Quantifiably [Sur16]. quantify [Sah05]. Quantitative [Far20, LS06, Opp15, San14]. Quantum [Ano22-61, Ano23-43, Ano23-54, Che17, Ell04, JL18, MGH^{+}15, MR22, Mos18, RS18, Sas18, Sch22b]. Quest [Ber20b, YB21]. Question [Jus04]. Questions [Wei04]. Quick [MP10]. Quine [GN07a].

Race [McC03b, Say08, Som04]. Railways [Les15c]. Rainbow [Ano22-62, Ano23-51, Ano23-52, Ano23-53, Ano22x, Ano22-73, Ano24-32]. Ralph [McG11f]. Ramakrishna [Ano17a]. Ramifications [SNS10]. Ramkrishna [Ano19f]. Ramsey [Pfl13c]. Random [Pay19]. Randomized [Ker09]. Randomness [Gen06]. Randy [McG12f]. Ransomware [HB22, KRK18, Nic21b, TMU23]. Ransomware-Bitcoin [TMU23]. Ranum [McG06b]. Rare [HN10]. Rational [Sch06a]. Rationale [FKS07]. rationality [AG05]. Rau [Ano17a, Ano19f]. RBAC [CW08b, FKS07]. Reach [San14]. Reaction [MNS08, RSWO18]. Reader [Pfl14a]. Reading [Don09c, Les09b, Les11b]. Ready [Ami03, CKPP23, GOPB12, Koh23a, Mos18]. Real [AJW13, Ano16-44, AJQA23, BCDS22, Bel15b, BFK16, BPS16, DPW11, DOP^{+}21, GSB^{+}04, SSJ^{+}23, SS10, Smi03a, Tom16, Val12, ZC09]. Real-Time [AJW13, SSJ^{+}23]. Real-World [Ano16-44, AJQA23, BCDS22, BFK16, BPS16, DOP^{+}21, Tom16, Val12]. Realities [MMR24, SY12]. Reality [Bel10b, BS24, CAL^{+}24, CDJS^{+}24, EP04, Koh22c, LRKR18, MRF^{+}24, NROS24, OGM^{+}24, OR03, Vie11a, Zan09]. Realization [AR15]. Realizing [OY16]. Really [Cal04c, CM23, GBK09, RC17, Ren12]. Realm [CAL^{+}24]. Realpolitik [Kal12]. Reaper [SPT19]. Reasonable [Cal03c]. Reasoning [ABD^{+}18, BD18]. Rebuilding [LS08]. Receipts [Cha04b]. Recognition [Ano15t, PPJ03, dAMM13]. Recognizers [RVG24]. Recommendations [AM12, Don05b]. Reconciling [PGT07]. Reconfigurable [DFJ^{+}20, Mat23]. Recorded [Les03a]. recording [Les05]. Records [ASH^{+}08, FB11, Les13b]. Recounting [YB08]. Recovery [BGL^{+}23, Gif10, OJR^{+}23, dQSzL13, SJ03]. Rectification [MH23]. Recycling [TR09a, TR09b]. Red [Cal04c, Sas07, RVK05]. Red-Eye [Sas07]. Redaction [BCG^{+}09]. Redefining [LLT14]. Reduces [Hol15a]. Reducing [BCLM09, Dia11]. Reduction [AWV18]. Redundant [BRS^{+}21]. Redux [Cow23, Cyb03b, SC04]. Refining [SOG^{+}19]. Reflect [HvO18]. Reflecting [Eps12b]. Reflections [Eps16, Kob19, Pei21, Zur22]. Reforms [Phi19]. Registers [LLGJ16]. Regulating [Ell14, WBT^{+}18]. REgulation [BLR19, Ano17v, BHRR07, Les10b, Pay04, WZE18, Pet19]. Regulations [CIL^{+}21, DMS^{+}15, Ksh14, Wil16b]. Rehash [FREP17]. Rehearsal [Ahm08b]. Rejuvenation [Bos16]. Relationships [Hag09, Law16]. Relevance [Sim15]. Relevant [BGMP19]. Reliability [Ami05, Ano17-67, Ano20-45, Ano20-46, Ano20-47, Ano21-75, Ano21-78, Ano21-79, Ano22-77, Ano22-78, Ano22-79, Ano22-80, Ano23-44, Ano23-55, Ano23-56, Ano23-57, Ano23-58, Ano23-59, Ano24-33, Ano24-34, Ano24-39, BSS10, BO14, CV10, GC08c, IKP^{+}07, Ano19-70, Ano19-71, Ano19-72, Ano19-73, Ano19-74, Ano19-75, Ano20-42, Ano20-43, Ano20-44, Ano21-76, Ano21-77, Ano21-80, Ano22-75, Ano22-76]. Reliable [MR03, PEC^{+}14]. Rely [ACL07]. Remediation [CP09b]. remedy [OR05]. Remembrance [GS03]. Remote [WKB08, Sch05a]. Reoccurring [GGBM21]. Repeat [BCG^{+}09]. Replicated [SZ05].

Report [BRS^{+}21, Ksh17, Lin09, SMGK04, BDKM12, DHR^{+}04, MBA12]. Reporting [LS05b, Ste13]. Reports [SMGK04]. Reprise [Say04]. Reproducibility [Max20]. Repurposing [GY13]. Request [CCF^{+}20]. Required [O’R17, ZLT^{+}23]. Requirements [MKF21]. Rescue [RXJ^{+}21]. Research [AMF15, Ano17q, Ano17r, Ano17-56, Ano18i, BA18, BHJM19, BL13a, BOR^{+}13a, Ben15a, Ben15b, Ben21, Ben23, CHK^{+}20, CK10, CCLS21, De 14, Far21, FREP17, FWFA23, HKMR24, HvO12, Ker09, Ker10, KRK18, Mat23, MCKS10, MBLT13, Mei04, Mul05, Nic20b, OMB17, PGNRV24, RC06, RJ08, Sch13a, SG20, SDC^{+}17, SPG^{+}19, ST06, SW18b, SBF^{+}15, TKZ21, dA16, vdKHA^{+}20, Ada05]. Residual [Ano17t]. Resilience [KT20, KS12b, LBF^{+}19, Str12, Wis18]. Resiliency [BGC^{+}14]. Resilient [Blo11, CCLS21, Dia11, Hor20, SBB^{+}18, SLS18]. resisting [DA05]. Resource [For04b]. Resources [Ano17-39, GAA04, HM22]. Response [BSP11, CSZ^{+}14, DMB^{+}22, Fra07b, GHS14, Hor14b, LMW10, MVL09, RDM^{+}14, Sch14a, SBF^{+}15, WB22]. Responses [Sch06a]. Responsibilities [TH13]. Responsibility [WMW22]. Responsible [McG13b, SW18b]. Responsibly [KTS18, PT05b]. Rest [Bra07]. Restating [KRK18]. Restricted [BKS18]. Results [RJ08, YBAG04]. Retaining [BC08]. Retention [Kot08]. Rethink [Gee10c, KS12a]. Rethinking [AAC^{+}17, LBS09, Zat16]. Retrospect [NPS14]. Retrospective [LRW^{+}23, LH23]. Return [PR12b]. Return-Oriented [PR12b]. revealed [For05]. Revealing [YZA08]. Revelations [Lan13b, Lan14b]. Reverse [DBR^{+}08, GC08d]. Review [AM12, BCG^{+}09, DO22, KS13a, LDZ^{+}22, OY16, PP06a, PS21, SLW22b]. Reviewer [Ano04d, Ano11, Ano14x, Ano17-70, Ano18-39]. Reviewers [Ano06h, Ano07b, Ano09c]. Reviews [For04a, How06, Pfl04, Pfl09, Sty04a, Sty04b, Sty08]. Revisited [Arc03b, Gee11d]. Revolution [DSKB22, Lan07c, SA12]. RFID [GJP05, MO12, PKK05, RLS^{+}08, RCT06, SE09, Wei04]. RFIDs [Lib05]. Rice [McG11a]. Richard [Ano15a, Ano15b, Ano16z, GA04b, McG10d]. Rid [McG13f]. Riddle [KM16]. Right [CP10a, Gee15b, Gid06, Hec03, Lan16, RCK17, Ano17b, EV18]. Rights [Bla03, FH06, Les04b, Les04c, Les08a, OGM^{+}24, SC19, WCV^{+}04]. Rise [Arc03a, CBOS20, De 14]. Rising [Gee11c, Jak19]. Risk [Ano16-43, AR19, BCD^{+}15, BFS^{+}21, BS09, BRS^{+}21, Cyb06b, El 10, EHK^{+}04, Eva15, FE11, GC09c, Gee12d, HKN^{+}09, HN10, JGP09, JP11, KA13, LS12, LLWC11, McL05c, MTS15, Opp15, PW17, Pet06, RR08, Sag13, SR16, Sol07, TW21, TWC^{+}15, VM04, WDL21, Wil16b, Wis18, WA07, Sah05, Sch05a]. Risk-Based [Ano16-43, AR19, El 10, WDL21, Wil16b, EHK^{+}04]. Risk-Limiting [BFS^{+}21, LS12]. Risking [BBD^{+}08]. Risks [De 24, JKSM19, LS09, Les15c, MRF^{+}24, Mus21, NROS24, Ros07, Sav17, Lin05]. Road [Ano17-36, CCLS21, HM13, RSOD22, RH06, VNC^{+}06, WL18]. Roadmap [PKB^{+}24]. Roads [CIL^{+}21]. Robinson [EMM06]. Robot [Don03b, Hee11, Sch22c]. Robotics [Blo15]. Robust [BF04]. Robustness [SPB12]. Rock [Ano14y, Ano14-27, Ano14z, Ano14-28, Ano14-29, Ano15-36, Ano15-37, Ano15-38, Ano15-41, Ano16-41, Ano16-42, Ano16-43]. Rode0day [FLDGB19]. Roger [BC22, Sty05b]. Rogue [BV11b]. Rogue-Access-Point [BV11b]. Role [BMZ14, BCLM09, Cat09, CM23, FKS07, LBB07, MT24, NBLC09, SFH05, SFH06, Smi12b, Str10, Zat16]. Role-Based [FKS07, LBB07, NBLC09]. Role-Playing [BCLM09]. Roles [Ksh15]. Roll [Les12a].

ROM [Cha04a]. Room [BBC^{+}19, Smi11]. Root [MRO^{+}23, SMJM11]. Rootkits [LGO06, RC04]. Rosetta [PW17]. Ross [MA07]. Round [Mar06]. Roundtable [BSS10, GHR^{+}10]. Roundup [Gar12a, Gar12b, Gar12c, Gar12d, Gar12e, Gar12f, Gar13a, Gar13b, Gar13c, Gar13d, Gar13f, Gar14a, Gar14b, Gar14c, Gar14d, Gar13e]. Route [PCK21]. Route-Disclosure [PCK21]. Routing [AS09, HP04, MM06]. RSA [GS07]. RSA(R) [Ano06g]. Rubin [McG06a]. Rules [BCG^{+}09, PZTE18, Wil16b]. Rules- [Wil16b]. Rump [Eps19]. Run [CEL^{+}19]. Runtime [Ker09]. Rust [vO23a].

S [Ano22a, Ano23b, Ano23-66]. SP [BC23a, Ano06h, Ano07b, Ano09c, BS20a, Nic21a]. Sabett [McG12f]. Sacrifice [KNBV17]. SAE [BS13a]. Safe [DSJK19, RIC17, Mas19]. Safeguarding [MC09, Sty04b]. Safely [HM12]. Safer [AED^{+}21, EJ06, HKMR24]. Safes [AR19]. Safety [AIKR13, Ano21-88, Bai12a, BL13b, CMT^{+}19, Les15c, MWM13, Nis24, Wis18, WMDX23, vO23b, vO23a]. Safety-Critical [AIKR13, BL13b, MWM13]. Sailing [Lan10b]. sales [LB04]. Salespeople [Cal13]. Salmon [Les11c]. Saltzer [Smi12a]. Salute [Les05]. Same [RP10]. Samples [VB13]. Samsung [Lan15]. San [Ano20z, Ano17-75, Ano18-43, NL12]. Sandboxing [GG11a]. Sanitization [GS03, HCC09]. Sapphire [Cyb03b]. Sapphire/Slammer [Cyb03b]. Sara [Pfl04]. SAv5 [CB15]. Savage [McG11d]. Saving [Gee15c]. Saviors [Sch07b]. Savvy [Lev04b, RIC17]. Scalable [HSS11]. Scale [Ano22-74, AEP^{+}19, BCDS22, Bel20b, CAB^{+}22, DMR^{+}22, GN07b, GU13, HQD^{+}03, HMPS15, KKK^{+}21, MLS22, O’R17, YCU^{+}21, ZOC^{+}22]. Scams [Vil04]. Scan [CEC^{+}08, GSHU08]. Scantegrity [CEC^{+}08]. Scapegoats [Sch07b]. Scaring [Sas15]. Scenario [Lev06, Str10]. Scenario-Driven [Str10]. Scene [Cam03b]. Scenes [Ano17f]. Schell [BC22]. Schemes [DP18]. Schmidt [McG12c]. Schneider [McG09c]. Schneier [Sty04b]. Scholarship [SPG^{+}19, Ano16z]. Scholarship-for-Service [SPG^{+}19]. School [FB04b, PHR20, WWH10]. Schroeder [Smi12a]. Schuhmacher [Koh22d]. Science [Ano18f, Ano21r, Ano21q, Ano21s, Ano22j, Ano22k, Ano22l, Ano24a, Ano24e, BHJM19, ES11, Far21, Gra13, HvO18, Kro18, MRS14, PW17, TKZ21, Wil19, YG19, vO22a, Don05b, Smi05b]. Scientific [Ano15d, Ano16a]. Scientist [PB07]. Scope [Lan08c]. Scoped [LCJ21]. Scoping [RDC^{+}18]. Scorecard [LS05b, ZKH^{+}23]. Scores [BRS^{+}21]. Scoring [KA13, MSR06, PZTE18]. Scripts [RW21]. Scrubbing [GC06]. SDRS [MNS08]. Seals [Hu16]. Sean [Ano21-89, Ano22-87, BC23a]. Search [Ano14d, Ano14c, Ano15h, Ano15i, Ano15j, Ano15k, Ano17-27, Ano17-29, Ano17-28, BDGS04, Les06, McL05b, Pre15, RCV^{+}12, HL13]. Searches [PGT07]. Searching [Pfl09]. SecDev [CGL^{+}16, Ano16-27]. Second [RLS^{+}08, RTL09]. Second-Generation [RLS^{+}08]. Secondary [PHR20, RS11]. Seconds [AAB13]. Secret [Cha04b]. Secret-Ballot [Cha04b]. Secrets [CMS09, CCX^{+}20, Man13, Smi03a]. Sector [Ben21, Cat10a, Ksh15, KJC^{+}12, PM14]. Secure [AR24, AM12, Amo13, And03a, Ano17z, Ano17-65, AP05, ABPP16, Bai12b, BV11a, Ber19, BF05b, Bis10a, Bla09, CHK^{+}20, CR09b, CG04, CGL^{+}16, CCLS21, DD03, DHR^{+}04, FPP12, FG22, FD11, GCY^{+}22, HLS^{+}21, HKW06, How04, HP04, HCC09, HRK21, IR11, Kap07, KR10, KDC20, KYEV^{+}18, LRW^{+}23, LLGJ16, Mar07, May10, MM09b, MRS14, McG14e, MM06, MM10, OP15, PRB^{+}12, PEC^{+}14, PRCC10, RCV^{+}12, RBC14, RS19, SCM08, SMJ^{+}19, Sea06, SMK20a, SA12, SBA^{+}18, SS04, Sty04b, TGC16, Tra09, VC08, Whi03, YRX^{+}22, ZL20, ZC09, Ada05, Smi05a, MŽ22]. Securely [Ano17-39, CC10, LRKR18]. SecureWorld [Kei05]. Securing [Ano17-71, ACC^{+}13, BBS19, CT11, CR06, Che11, CZ23, CDJS^{+}24, Gra13, HSTV06, Ksh17, Mar15c, MD09b, MC09, McL13, PCMN08, PM14, RALS23, SFE10, Tra09, WAF11, WO09, YAZ21, WK05]. Securities [Vil04]. Security [Acq13, AA12, Ahm07, AZ20, AAC^{+}17, Alf05, Ami05, AS05, And07, ABAGOB10, And04, AW04, And06b, Ano03a, Ano06d, Ano06c, Ano08d, Ano08f, Ano08g, Ano14j, Ano14k, Ano14l, Ano14m, Ano14-32, Ano14-30, Ano14-39, Ano15x, Ano15y, Ano15z, Ano15-27, Ano15-36, Ano16f, Ano16-32, Ano16-29, Ano16-30, Ano16-31, Ano16-28, Ano16-43, Ano17q, Ano17r, Ano17y, Ano17x, Ano17-44, Ano17-45, Ano17-46, Ano17-47, Ano17g, Ano17-67, Ano17-72, Ano17-73, Ano17-74, Ano18i, Ano18y, Ano18z, Ano18-27, Ano18b, Ano18a, Ano18-29, Ano19d, Ano19e, Ano19y, Ano19h, Ano19x, Ano19z, Ano19c, Ano20o, Ano21-39, Ano21-32, Ano21a, Ano21-81, Ano21-88, Ano21-89, Ano22-66, Ano22v, Ano22y, Ano22x, Ano22w, Ano22-87, Ano23i, Ano23h, AQP^{+}23, Aur06, BP07, BW07, BDGS04, BPR^{+}04, BO22, BA18]. Security [BCM^{+}15a, BFC^{+}17, BM11, Bay11, BBC^{+}19, Bel05, Bel06c, BBD^{+}08, Bel08a, BC09, Bel10a, Bel11b, Bel20b, BL13a, Ben15a, Ben16b, Ben17, BP18, BH22, BO23, Ben23, BCGN16, Bez24, BMZ14, BYG^{+}14, BS13a, Bis03, Bis11, BP12, BCG06, BC23b, BB17, BGM^{+}06, BKS13, BPFG24, BC08, Bon08, BDT17, Bra06, BPVS04, BN08, BLCDK11, BMMM10, CL13, Cal04a, Cal04d, Cal04c, Cal04e, CPS^{+}16, CMR13, Car08, Cat09, CP15, CHMO07, CWVM22, CAL^{+}24, CMT^{+}19, CDS08, CW08a, CW09, CSZ^{+}14, CM04, CA11, Chu08, CP09a, Coh10a, CBN11, CP10b, Cow03, Cow23, CW08b, CB15, CB14, CR20, CA06, Cyb04b, Cyb06b, CL12, CCF14, DSEB12, Dat10, DFG^{+}11, DeL07, DHP14, DPW11, DPP15, DFJ^{+}20, DOP^{+}21, DMR^{+}22, DD08, Dia11, DP20, DR04, Dim07, DBD11, Don03d, DLMS21]. Security [Dre16, Du11, DGU^{+}12, DCC^{+}09, Eag13, EEA^{+}23, iC15, ENDAC22, EOM09, EB19, EW22, EMM06, Eps08, EHK^{+}04, EP04, ES11, FOM23, FH06, FE11, FREP17, For04a, FWFA23, Fri03, FB04c, Gar12a, Gar12b, Gar12c, Gar12d, Gar12e, Gar12f, Gar13a, Gar13b, Gar13c, Gar13d, Gar13e, Gar13f, Gar14a, Gar14b, Gar14c, Gar14d, GBK09, GD11, Gee05a, GC08c, Gee09b, Gee20, GY21, GM12, GCR18, GB09, GPD13, GLS23, GG04, Gra23, GS16, Gri04, Gro08, Gro12, Gru16, GFJ^{+}18, Gup17, GG05, Hag09, Hal10, Han12a, Han12b, HLJ21, HWF^{+}10, HK09, HJK08, Hea03a, Hec11, Hec03, Hei07, Her19, HvO18, HRMS21b, HRMS^{+}21a, Her06, HM22, HM13, HMT06, HKN^{+}09, Hor14b, HL03, How06, How08, How09a, How09b, HCL04, HKM17, Irv03, IN10, IKP^{+}07, JMD06, JS10, JKMS23, JG07]. Security [JGP09, JW04, KG17, Kam14, KC18, KS08a, KvdHK^{+}14, Kau09, Kau10b, Kau10a, KW12, Ker10, KHLF10, KB06, KS08b, KH12, KST^{+}21, KWRK13, KS12a, Kni15, Kob19, KAAEa17, KDC20, KCL^{+}23, Kuf13, Kuh16, KNBV17, Kup05, KJ15, KTT19, Lad06, LS09, Lan05a, Lan08a, Lan14c, Lan03, Law16, LS06, LS05b, Les09b, Les09a, Les13c, Les13d, Les14c, LC13, LZ16, LLT14, LH23, LWLL10, LLWC11, LAYG16, LBS09, LK17, Lyo24, MJF07, MM09a, MS20, MMKP16, MA20, Mas19, MN21, MW23, Mat12, MOT^{+}17, MCKS10, MT24, MM09b, ME10, McD11, MA12, MP10, McG03b, McG04, MP04, MH07, MC09, McG13b, McL05a, MPS14, MM05, MH13, Mei06, MSW09, MTA23, MS21c, Men24, MRO^{+}23, Mic21, Mic10a, MT23, Mil05, MBR^{+}21, MKF21, Mol13, MBA^{+}22, Nah08, Nan09, NL23, NRH^{+}22]. Security [NK12, NL12, NPS14, NP07, Nic05, NSLW20, OBS^{+}20, Opp15, OKH07, Pal04, Pay04, PD07, PMB^{+}14, PMN^{+}14, Pei19, Pei20, Pei23a, PvDS08, Pet09, Pet10, Pet07, PA03, PR12a, PTP07, PC10, Pfl12b, Pfl13a, Pfl15a, Pfl15b, PKK05, PSB^{+}07, Pop04, Por09, PPJ03, RW21, RDC^{+}18, RBK23, RIC17, RS06, RK24, Rog16, RI12, RH06, RC06, RDM^{+}14, Rui06, RLB^{+}23, Rya03, Rya07, RR08, RTL09, RRAK14, RKA15, Sad16, Sad17, SD17, Sah05, SCZ^{+}13, SAB^{+}19, San14, SSJ^{+}23, Sas15, SSH^{+}16, SS16, Say04, Sch09b, Sch03e, Sch04c, Sch04e, Sch06c, Sch07c, Sch07d, Sch08b, Sch09d, Sch10b, Sch11b, Sch12a, Sch12b, Sch13d, Sch15, Sch17b, SMCA14, SR16, SFK^{+}10, SY12, SZ06, Sho22, Sho23, SC19, SB19, SMGK04, SLW22a, ST06, Smi03b, SS04, Smi12b]. Security [Spa08, Ste13, SP09, Ste06, Ste08, SDKM20, SBE11, SAA04, SW03a, SW03b, Sty04b, SB06, SLS18, TJA10, TP11, TG07, Tho13, TG04, TGS20, THM15, TBBR17, TP08, TPPM07, TB19, VM21, Var09, VVSL12a, VVSL12b, Ver16, Ver06, VNC^{+}06, VKM^{+}15, Vét21, VLOS18, VM10, Vie12a, VT12, Wai16, Wan18, Was12, Web05, WMW22, WDP23, Wha11b, WF06, WHD^{+}09, WMS10, Woo23, WA07, Wu10, XL16, YBAG04, YE09, Yee04, ZKH^{+}23, Zat16, Zur22, Zur23, ZH23, vS19, van19, van20, vO20, vO21a, vO22a, vO22b, vWM05, vWS06, vdKHA^{+}20, BM05, BG13, KN13, LS05a, Mar05, MR05, OR05, PT05a, Res05, Sal05, Sch05a, Sno05, TM05, TCM05, VWW05, Ano03f, Ano08c, SD16a, SD16b, Ano19-37, Ano19-77]. Security-Enhancing [EEA^{+}23]. See [TAGM23, dAMM13]. SEED [Du11]. Seek [Ano23-38, Ano24-38, Blu10, PH03]. Seeking [Ano14-30]. Seeks [Ano23-39]. Seemingly [PMA^{+}20]. Seen [FA07a, FA07b]. Seers [Bel07b]. SEHAS [Mea03b]. Selected [Ben17, BP18, BH22, EB19]. Selecting [Bur03]. Selectivity [LSG^{+}18]. Self [eDIATK22, ASE21, Amo16, Ano17v, BMM07, Phi19, RCK17, TAP19]. Self-Defense [RCK17]. Self-Efficacy [Amo16]. Self-Healing [ASE21]. Self-Mutating [BMM07]. Self-Regulation [Ano17v]. Self-Sovereign [TAP19]. Self-Sovereign-Based [eDIATK22]. Selfie [Kob19]. SELinux [SFE10]. Sell [Max20]. Seller [Don04f]. Semantic [CGC19, JBZ^{+}15, KP10, MH23]. Senior [MŽ22]. Sense [Her13, Lan13b, Lan14b]. Sensibly [Sty04b]. Sensitive [Gar14e, ST10]. Sensitivity [PKBS15]. Sensor [HRTT03, MK13, PBW^{+}08]. Sensors [BSH^{+}09]. Sensory [Ano17y]. Separate [SR16]. Server [BB17, Mic08b, MKKP09, IS05]. Server-Side [MKKP09]. Servers [WO09]. Service [AR24, Alj03, Ano15a, Ano15b, Che06, DeL07, GHG14, Hay13, KS12b, Les12d, Pet09, SPG^{+}19, VdCA07, WAF11, WB22, Wu10, Cam05]. Service-Oriented [Pet09]. Services [And06a, Ano19-62, CEK^{+}10, CFH05, CP09a, GR13, HPSP10, Knu07, KPM^{+}19, Opp07, SZ05, Sim21, Sun16, TCJ^{+}19, TBBR17, VM21, VE06]. Session [Vis10]. Sets [BPB^{+}04, Ker09]. Setting [Gid06, PGT07, KN13]. Settings [Cel16]. Setup [KZZ17]. Seven [DD08, El 20, HS17, SWL21, TCM05, TW21]. Sex [HKMR24]. Seymour [Ano18-40]. SGX [CCX^{+}20]. SgxPectre [CCX^{+}20]. Shadows [Bla09]. Shakespeare [BF05a]. Shaking [FO08, Ort09b]. Shall [FA09, Mic09]. Shannon [Smi05b]. Shape [Sty08]. Shaped [MLM^{+}07]. Shaping [dA16]. Share [Ano18-41, Ano18-42, Les04a]. Shared [VM21]. Shari [BC23a]. Sharing [CK10, DK10, GD17, In 07, Kam14, MM10, O’R17, SyPKZ22, ST10, TKZ21, TMU23].

Sharp [Gar08]. Shedding [Les04d, Pfl13a]. Sheep [Gra23]. Shellcode [Arc04c]. Shift [DEKM^{+}21, TAP19, WMW22]. Shockvertising [MRF^{+}24]. Shopping [Les08b]. Shor [JL18]. Short [GC10a, PvO24, Hol05]. Shortage [CS13]. Shostack [McG08a]. Shot [KMP20]. Should [Bel14c, Bel15b, CC11, GA10, Lan16, Les06, Ohm14, OAB07, BG13]. Shoulder [Les09b]. Shouldn’t [Aur06, TP11]. Shuffle [Sas07]. Sibling [BS20a]. Side [HPSP10, Law09, MKKP09, PvO24, RI12, ZOC^{+}22]. Side-Channel [Law09, PvO24]. Sides [Bel11a]. Sidney [Ano18-40]. Siege [SAA04]. SIFIS [ABMS22]. SIFIS-Home [ABMS22]. SIGINT [Dif06]. Sign [HJK08]. Sign-On [HJK08]. Signaling [SCCB05]. Signals [EH13]. Signature [KS13a]. Signatures [But17, GY13]. Significant [Lan13b, Lan14b]. Signing [RdRdCG10]. Silent [TWC^{+}15]. Silver [Les14b, McG06d, McG07d, McG07c, McG07e, MA07, McG09f, McG10f, McG09a, McG06a, McG06c, McG06b, McG07a, McG07b, McG08d, McG08c, McG08e, McG08a, McG08b, McG08f, McG09d, McG09e, McG09b, McG09c, McG10a, McG10b, McG10e, McG10d, McG10c, McG11e, McG11a, McG11d, McG11b, McG11c, McG11f, McG12a, McG12e, McG12b, McG12f, McG12d, McG12c, McG13d, McG13e, McG13f, McG13c, McG13h, McG13g, McG14b, McG14c, McG14f, McG14d, McG14a, McG14e, McG15b, McG15f, McG15d, McG15c, McG15a, McG15e, McG16e, McG16b, McG16a, McG16d, McG16c, McG17e, McG17d, McG17a, McG17b, McG17c, McG17f, McG18f, McG18d, McG18c, McG18e, McG18b, McG18a, Ran16]. Simple [Ksh06, Per07, RIC17]. Simplifying [HJK08]. SIMS [Sch04f]. Simulation [Nic05]. since [AEY10, GM12]. Single [HJK08, Res16]. Single-Point-of-Failure [Res16]. Singularity [SG18]. Site [AEV^{+}07, CCF^{+}20, Gro08]. Sites [Ros07, Sch04d]. Situational [LTB15, MBA^{+}22]. Size [Sch09d]. Skill [SCMS18]. skills [HL13]. Skim [SPT19]. Slammer [Cyb03b]. Slapper [AL03]. sleep [Kri13]. Sloganeering [Cal03d]. Slow [Hea03c, Smi11]. Small [AKN12, Gee11f]. Smaller [Les04a]. Smart [Ano17-74, ABMS22, BMS08, BVB22, BJD22, CIL^{+}21, CL07, CXG^{+}18, HRMS^{+}21a, HCL04, KHLF10, Lan08b, LA10, MŽ22, SSJ^{+}23, SyPKZ22, VM21, WMDX23, ZHC^{+}20, MM09b]. Smart-Grid [KHLF10]. Smart-Home [Ano17-74]. Smarter [Coh10b, GB12, KTT19]. Smartphone [McD12]. Smartphones [BV11a]. Smashing [PB04]. Smells [RW21]. SMTP [Ano17h]. Sniff [GM14]. Sniffing [JS07, Ohm14]. Snowden [Lan13b, Lan14b]. SOA [DN07, EMM06]. Social [Ano19-55, Ano19-56, Ano19-76, AS15, Baa03, BHJM19, Cal04d, DMS07, Pfl04, Rog16, Ros07, Sch10c, AS13, GD13, Kri13, dAMM13]. Social-Science [BHJM19]. Societal [OGM^{+}24]. Société [Eps08]. Society [Ano06b, Ano17-43, Ano19-52, Ano19-53, Ano19-54, Ano20-45, Ano20-46, Ano20-47, Ano21-52, Ano21-75, Ano21-78, Ano21-79, Ano22a, Ano22-77, Ano22-78, Ano22-79, Ano22-80, Ano23b, Ano23z, Ano23-55, Ano23-56, Ano23-57, Ano23-58, Ano23-59, Ano24z, Ano24x, Ano24-27, Ano24-39, Cal03e, HRGK^{+}20, Ano14i, Ano15w, Ano15u, Ano15v, Ano16w, Ano16u, Ano16z, Ano16x, Ano16y, Ano16v, Ano17-42, Ano17-41, Ano17-40, Ano18u, Ano18q, Ano18r, Ano18s, Ano18v, Ano18t, Ano18w, Ano18x, Ano18-41, Ano18-42, Ano19-46, Ano19-47, Ano19-48, Ano19-55, Ano19-49, Ano19a, Ano19-56, Ano19b, Ano19-50, Ano19-51, Ano19-61, Ano19-63, Ano19-70, Ano19-71, Ano19-72, Ano19-73, Ano19-74, Ano19-75, Ano20s, Ano20-29, Ano20-30, Ano20-34, Ano20-31, Ano20-27, Ano20-32, Ano20-28, Ano20-33, Ano20-42, Ano20-43, Ano20-44, Ano21-66, Ano21-54]. Society [Ano21-60, Ano21-55, Ano21-61, Ano21-48, Ano21-56, Ano21-62, Ano21-49, Ano21-57, Ano21-63, Ano21-53, Ano21-50, Ano21-58, Ano21-64, Ano21-51, Ano21-65, Ano21-59, Ano21-76, Ano21-77, Ano21-80, Ano22-49, Ano22-35, Ano22-43, Ano22-40, Ano22-36, Ano22-44, Ano22-45, Ano22-42, Ano22-37, Ano22-46, Ano22-38, Ano22-47, Ano22-41, Ano22-39, Ano22-48, Ano22-60, Ano22-59, Ano22-75, Ano22-76, Ano23v, Ano23-38, Ano23x, Ano23-31, Ano23-34, Ano23t, Ano23-35, Ano23u, Ano23w, Ano23-36, Ano23-32, Ano23y, Ano23-28, Ano23-37, Ano23-29, Ano23-30, Ano23-27, Ano23-39, Ano23-33, Ano23-44, Ano24y, Ano24t, Ano24v, Ano24-29, Ano24s, Ano24-28, Ano24-30, Ano24w, Ano24u, Ano24-31, Ano24-33, Ano24-34, Ano20l, Ano21x, Ano21y, Ano21z, Ano21-27, Ano21-28, Ano22o, Ano22q, Ano22p]. Socio [AK15]. Socio-technical [AK15]. Sociotechnical [JB05]. Soft [Her19]. Software [Ahm07, AK15, And07, Ano14-31, Ano15c, Ano15-39, Ano17-73, Ano19-27, Ano23b, Ano23-66, AP05, ASM05, BBB^{+}23, BV11a, Bar15, Bel06c, Ber22, CA11, CW12, CDD13, CN13, Cow03, Cow23, CCLS21, DHR^{+}04, DXR^{+}23, EW22, EMM06, FPP12, FG22, FWFA23, GTG07, Gee23b, GGBM21, Hec03, HRMS21b, HM22, HS17, How04, How09a, Jae23, JZ04, Jon07, KR10, Knu07, KYEV^{+}18, Lad06, LPS^{+}23, LBF15, Law09, Lev03b, LAYG16, MC12, MW23, McG03b, McG04, MP04, MC09, MM05, MH13, MTA23, MT23, Pei23b, Pfl16a, Pfl16b, PSB^{+}07, PS21, Rya03, SP09, Ste06, Sty03, SW03b, Sty04c, SB06, Syt05, TVR23, TGS20, TAGM23, VM04, Ver06, VLOS18, WMS10, Wil22, Wil24, XL16, ZLT^{+}23, ZL20, ZC09, van19, vWM05, vWS06, AT05, BM05, Hol05, TM05, TCM05, Ano23j]. Software [Ano23k]. Software-Defined [Ano17-73, KYEV^{+}18, LAYG16]. Software-Intensive [CCLS21]. SolarWinds [MJP21, PSO^{+}21]. Solution [Coh10c, Gid06, Sch04f]. Solutions [ARC19, AED^{+}21, BBL^{+}17, BV11b, Ede14, GSLA15, LLGJ16, RBE03, RH06, SNS10, TVR23, Web05, YB21, GJP05, PB05]. Some [DK10, Eps12b, LXS22, Pei19]. Something [Ano22-87]. Son [Les10b]. Songs [Les11c]. Sophisbots [BGMP19]. Sound [McL05c]. Soups [RA05, TE06]. Source [Bel22a, CCS06, Cow03, MPM22, MSW09, SPL^{+}24, SK04, SSH22]. Sourced [ADG23]. South [Les07b, PCK21, SAB^{+}19]. Sovereign [eDIATK22, TAP19]. SP [Ano22-62]. SP18 [Ano18-43]. Space [MM11]. Spaces [CMT^{+}19]. Spafford [McG08d]. Spam [ANC10, Cur06, KCC07, MNS08, O’D07, PB05]. Speak [WDP23]. Speakers [WMDX23]. Speaking [Lan06]. Speaks [McG06a, McG06c, McG06b, McG06d, McG07d, McG07c]. Spear [ACPZ20, CPFJ14]. Spear-Phishing [ACPZ20]. Special [Ano06h, Ano07b, Ano09c, Ano16f, Ano16e, Ano16d, Ano16g, Ano16h, Ano16-44, Ano17o, Ano17n, Ano17l, Ano17p, Ano17m, Ano17q, Ano17r, Ano18b, Ano21-81, Ano22-40, GS07, Lin09, MMKP16]. Special-Purpose [GS07]. Specter [Ano21-39]. Spectrum [SB21]. Speculative [CCX^{+}20]. Speech [Cal04f]. Speed [Ano17-38, CPC^{+}18, Pei22]. Spending [Mat12]. Sphynx [CGR^{+}22]. Spider [Pfl15b]. Spider-Man [Pfl15b]. Spikes [SLKJ21]. Splitting [PRCC10]. Spoiler [VT12, Vie12d]. Sponsor [Ano15-40]. Spooky [Pfl07]. Spotlight [Say08]. Spots [SLW22a]. Spread [SM04]. SpringPeeper [LM08]. Spying [BNSE21, VWC19]. Spyware [Cur06, FH06]. Square [Mar06]. SRAM [Han12a, Han12b]. SSDF [Wil24].

Stability [Mic10a]. Stack [CUCD14, Pet10, PB04]. Staff [Ano18-44, GB09]. Staffing [Les14c]. Stake [Gru16]. Stakeholder [AAG15, AA15, SMK^{+}20b]. Stamping [Les15b]. Stand [DFJ^{+}20, GA12, HKM17, SW03b]. Stand-Alone [DFJ^{+}20]. Standard [CW08b, FKS07, Bur03, LBB07, MS21a]. Standardization [AEH^{+}04, BHRR07, Woo23]. Standards [Ano15g, BCM^{+}15a, Bur06, CL07, Che17, Gid06, Hei16, HV20, KG17, MR22, RS06, Sch22b, VE06, PKK05]. Stars [Ano14y, Ano14-27, Ano14z, Ano14-28, Ano14-29, Ano15-36, Ano15-37, Ano15-38, Ano15-41, Ano16-41, Ano16-42, Ano16-43, LL16]. Startup [Ano15-41]. State [And07, And06b, BGG^{+}22, BLM17, But17, DP17, Gro12, Kup05, SZ05, Sen17, Ste08, VT12]. Statement [Gee05b]. States [WMDX23]. Static [And12, CM04, CBE^{+}12, KCL^{+}23]. Static-Analysis [And12]. Station [HWB^{+}20, Sch04a]. Statistical [Sim15]. Statistics [KCC07]. Stay [Ano18-45, RIC17]. Steal [Ano17d]. Stealing [CCX^{+}20, FH12, VWC19]. Stealth [van20]. Stealthily [Bis11]. Stealthy [AJV18, RC04]. Steganographic [MJF07]. Steganography [PH03]. Stegomalware [CM22]. Steps [AR15, Pfl12b, RIC17]. Steve [McG13e]. Steven [McG10e, McG15e]. Stewardship [AACG19]. Stewart [McG07d]. Stickler [Bon16]. Stimuli [Ano17y]. Stock [O’D07, YCU^{+}21]. Stone [PW17]. Stones [Wil21]. Stop [MJP21, Sch16b]. StopGuessing [THS20]. Stopping [EHK^{+}04]. Storage [BHP24, HPSP10, LCJ21, RBB19, TCJ^{+}19]. Store [KWRK13]. Stores [SyPKZ22]. Stories [Don09a]. Storm [Sch05d]. Storms [Ami03]. Strangecode [Bel14b]. Stranger [DH13]. Strategic [Ben15b, EGB18, Ell11].

Strategies [Ano15-43, HL13]. Strategy [OKH07, RSM13, TPPM07]. Strategy-Based [RSM13]. Stream [PRCC10]. Street [BMS08]. Street-Smart [BMS08]. Strength [PSB^{+}07]. Strengthen [Hal10]. Strengthening [SS06]. Strengths [LL23]. Strings [Sea06]. Strong [Cyb06a, GC08e]. Stronger [Boy16]. Structured [TMU23]. Stubborn [GC06]. Student [Ano16z, Ano18-42, Ano18-41]. Students [BHJM19, BP08, CC11, DD03, DLMS21, For04b, GB09, Loc09, SDC^{+}17, SPG^{+}19, SE07, VKM^{+}15]. Studies [Ano22-74, BF07, CPS^{+}16, FWBC15, Gup17, LLT14, MDD^{+}15, Mul05]. Study [Aur06, CW09, CB15, DBR22, GS03, Gui11, HMT06, Kap07, Mar15b, NRH^{+}22, PA03, PRS09, Pow07, SPG^{+}19, YE09, vO23b, YB05]. Studying [CMZ07, SMO^{+}14, Syt05]. Stuff [For04c]. Stupid [BMS08]. Stuxnet [Fid11, Lan11]. Style [BFS^{+}21, Jon06]. Subgroup [DHR^{+}04]. Subject [SC19, vO22b]. Submit [Ano20-48]. Subscribe [Ano13c, Ano14p, Ano14-32, Ano19k, Ano19y, Ano19h, Ano19x, Ano19z, Ano19-77, Ano22v, Ano22y, Ano22w, Ano23i, Ano23h]. Subscription [Ano16-32, Ano21-32]. Subscriptions [LDK^{+}14]. Subset [GC08c]. Subsystem [Bil10]. Subversion [Bil10, VC08]. Succeed [CCS06]. Success [BFC^{+}17]. Successful [PR12a, vWS06]. Sudanese [DSKB22]. Sufficient [Pot10b, Sch19a]. Suite [TB19, AT13]. Sum [Eps15b]. Summary [DHR^{+}04]. Summer [AALS21]. Summit [Ano14-31, Ano15-39, DHR^{+}04]. Super [RVG24]. Super-Recognizers [RVG24]. Superset [SE13]. Supersingular [Lau17]. Supervision [VVYY11]. Supply [AK15, DXR^{+}23, EW22, FWFA23, LPS^{+}23, Lev03b, MW23, MTA23, MT23, RALS23, Wil22, Wil24]. Support [CRK^{+}13, CP10b, LRKR18, Mat23, O’R17].

Supporting [FOM23]. Suppose [SA12]. Sure [Tom06]. Surface [BCLM09, Gee11a]. Surfaces [Bel16a]. Surprise [Arc07b]. Surprises [Hen09]. Surprising [Per07]. Surveillance [Bla09, GS06, Gee14b, Gro14, Lan09a, Lan13b, Les14b, MM14, Pfl14b, Sch14b]. Survey [DLMS21, Dun10, HMPS15, HP04, LZ16, SLW22a, TB19, Lin05]. Survivors [MOT^{+}17]. Susan [BC21]. Susceptibility [JZ04]. Sustainable [Ano18d, Ano19-32, Ano19-33, Ano19-34, Ano19-35, Ano19-39, Ano20q, Ano21-36, Ano21-37, Ano21-38, Ano22-28, Ano22-65, Ano24m, TJS^{+}22]. Sustaining [SFH05, SFH06]. SVS [LJZ12]. Swartz [McG08e]. Switching [NB23]. Symposium [Ano19d, Ano19e, Ano19c, Ano21a, Ben14, Ben15a, Ben16b, Ben17, BP18, BH22, BO23, NL12, NPS14, Ano08g, EB19]. Syntactic [MH23]. Synthesis [HRK21]. Synthetic [De 24, El 20, MMR24, TCC^{+}24]. System [eDIATK22, ABD^{+}18, BGLEP15, CEK^{+}10, Coh10a, CB15, DSEB12, Don11b, DCC^{+}09, EPH12, GOPB12, GSHU08, GFJ^{+}18, HMPS15, HMKT07, KMP^{+}11, KZZ17, LRKR18, LB08, LK17, MGH^{+}15, MNS08, MSR06, Men24, NTMD^{+}18, NS11, Pei22, RdRdCG10, Rya03, SCM08, SS19a, TP08, WA07, vO22a, CR06, Che06, MD09b]. System-on-Token [RdRdCG10]. Systematic [JKSM19]. Systematizing [van19]. Systemic [KB06]. Systems [AC21, AZ20, AFFOG12, Ami05, Ano14b, ASH^{+}08, AM04, AJQA23, BPR^{+}04, BSS10, BGC^{+}14, Bay11, BC09, BMZ14, BS13b, BKS18, Bla09, BL13b, BN08, BD18, BGL^{+}23, CMR13, Che11, CV10, CAB^{+}22, Cow03, CR20, CCLS21, DD03, DPP15, DDAN12, DXR^{+}23, EHK^{+}04, GY21, GLM11, GW07, Hav13, Hee11, HMK12, Hor20, HSTV06, IN10, IR11, Jus04, KvdHK^{+}14, KS13a, KS06, Kuf13, LJZ12, LMW10, LBS09, LL11, MSeF^{+}22, Mas19, MM09b, McL13, MPS14, Mea03b, MSW09, Mil05, MWM13, MBA12, NSSS08, Nic21b, NSA^{+}23, PR08, Pee13, PM14, PMB^{+}14, PMN^{+}14, RC17, RSM13, RST15, Sas07, Sch13d, SS04, SE07, Tra09, TWC^{+}15, VVSL12a, VVSL12b, VC08, Way08, WDP23, Yan16, YI13, ZC09, Sal05, SCCB05, Smi05a, MMKP16, Ano19p]. Systems-of-Systems [Hav13].

Table [Ano13e, Ano14-33, Ano14-34, Ano14-35, Ano14-36, Ano14-37, Ano14-38, Ano15-42, Ano16-45, Ano16-46, Ano16-47, Ano16-48, Ano16-49, Ano16-50, Ano17-76, Ano17-77, Ano17-78, Ano17-79, Ano17-80, Ano17-81, Ano18-46, Ano18-47, Ano18-48, Ano18-49, Ano18-50, Ano18-51, Ano19-78, Ano19-79, Ano19-80, Ano19-81, Ano19-82, Ano19-83, Ano20-49, Ano20-50, Ano20-51, Ano20-52, Ano20-53, Ano20-54, Ano21-82, Ano21-83, Ano21-84, Ano21-85, Ano21-86, Ano21-87, Ano22-81, Ano22-82, Ano22-83, Ano22-84, Ano22-85, Ano22-86, Ano23-60, Ano23-61, Ano23-62, Ano23-63, Ano23-64, Ano23-65, Ano24-40, Ano24-41, Ano24-42, NB23]. Tackle [YB21]. Tactics [van20]. Tagging [BW10]. Tags [MO12]. Tailored [Kni17]. Take [Ano15-43, Ano17-82, Ano18-52, McL05c, Sch21]. Takeover [SGCG^{+}11]. Takes [Ber19, Lan09a]. Taking [Ano14-39, Bla09, Mic22, RC04, RH06]. Tale [Len03, Opp15, SI09]. tales [GD13]. Talking [FD11, WDP23]. Talks [GA04b, McG09a, McG07a, McG07b, McG07e, MA07, McG08d, McG08c, McG08e, McG08a, McG08b, McG08f, McG09d, McG09e, McG09f, McG09b, McG09c, McG10a, McG10b, McG10e, McG10d, McG10c, McG10f, McG11e, McG11a, McG11d, McG11b, McG11c, McG11f, McG12a, McG12e, McG12b, McG12f, McG12d, McG12c, McG13d, McG13e, McG13c, McG13h, McG13g, McG14b, McG14c, McG14f, McG14d, McG14a, McG14e, McG15b, McG15f, McG15d, McG15c, McG15a, McG15e, McG16e, McG16b, McG16a, McG16d, McG16c, McG17e, McG17d, McG17a, McG17b, McG17c, McG17f, McG18f, McG18d, McG18c, McG18e, McG18b, McG18a, McL05b, Ran16, McG13f]. Tamarin [BCDS22]. Taming [CZL08, Dun10]. Tamper [DMA09, HCL11]. Tamper-Proofing [DMA09]. Tanya [McG18b]. Tape [HCC09]. Tapping [Sag13]. Target [AJV18, CF14, GL14, OHBS14, Red24, TWA08]. Targeted [ARvD12, CP09b, Gra23, SE13]. Targeting [Bor13b, Her13]. Task [Mic08a]. Tasks [Ano17y]. Tax [Sch20b]. Taxonomy [AS11, Ano21-41, CDS08, JFK21, Sch10c, SGSM22, SZ16, Ree16, TCM05]. tbd [Ano22-63]. TC [Ano22-64]. Teach [Bin18, CC11, Nan09]. Teaching [AR18, BHJM19, BP08, BF04, BF05b, Bis11, CW08a, CW09, DD03, FB04d, GY13, Irv03, Lan14a, MB12, SMK20a, SCMS18, van20]. Team [Pfl10, SBF^{+}15, RDM^{+}14]. Teams [RVK05, SBF^{+}15, CSZ^{+}14, Hor14b]. Teams-Based [SBF^{+}15]. Tear [Boy16]. Tear-Free [Boy16]. Tech [Ano17-49, Ben21, Don04a, Don09a, Lev04b, RIC17]. TechIgnite [Ano16-51, Ano17-83]. Technical [AC21, HKMR24, MP21, OY16, Pet19, AK15]. Technique [Hol15b]. Techniques [FOM23, KAAEa17, MBA^{+}22, OHBS14, SMCA14, Sim21, ZOC^{+}22]. Technological [HRGK^{+}20]. Technologies [Ano16i, Bel06a, CIL^{+}21, CGC19, Cur06, EGB18, Lan08a, Men24, O’R17, PS21, PP06b, RR06, San21, SJZG19, SE09, Wei06]. Technologists [DR06, Sch20c]. Technology [Ano20s, ACC^{+}13, BBB^{+}08, BL13a, Bis10b, Cal03b, DSKB22, DeL14, DR05, GHK^{+}06, GSB^{+}04, Kra04, LZ16, Lin05, Lyo24, OJR^{+}23, Pei19, Pfl14c, RXJ^{+}21, Sch07b, Sch04c, Sch12a, WHD^{+}09, Mar13b]. Technology/IT [OJR^{+}23]. Technophobes [Cal04f]. Teen [Amo16]. Telecoms [LL23]. Temperature [BKMN09]. Temporal [BRS^{+}21]. Ten [Mei18, Vie11b, Vie12d]. Tenancy [CEK^{+}10]. Tensors [ADE^{+}22]. Term [vdG17]. Terminarch [Ost20]. Terrorism [PP06b]. Test [All07, LN08, SOG^{+}19, BGMP19]. Testbed [SSJ^{+}23]. Testing [Bis07, BC08, Bon08, KCL^{+}23, MP04, Pay19, Phi19, SB06, ASM05, Tho05]. Tests [Can14, Sim15]. Tests-versus-Proofs [Can14]. Text [Bel08b, Bel09, BC09, Gee08a, Sch04c, Sch09c, Sch09d]. Thanks [Ano04d, Ano06h, Ano07b, Ano09c, Ano11, Ano14x, Ano17-70, Ano18-39]. Theater [Gee20]. Theft [CMS09, McC03a, WYA06]. Their [BGC^{+}14, Bel14a, CBN11, Eva14, FVJ19, Les15b, Loc09]. Them [Nan09, Wil21]. Themes [vO22b]. Theory [DM15a, HvO18, PY06, RST15, Pfl13c]. There [Don05c, MPM22, TCC^{+}24, YI13, ZLT^{+}23]. These [Don04e]. Thesis [SM11]. Thief [BHJM19]. Thing [SRG03]. Things [Blo12, BKS13, BMS08, MP10, MDR^{+}20, ASE21, Ano18-29, Cal13, CHK^{+}20, CMT^{+}19, ENDAC22, FREP17, HRMS^{+}21a, JKSM19, KKK^{+}21, Kuh16, KPM^{+}19, Mar15b, MHRS19, MKF21, NSLW20, Sch17a, SS19a, SS19b, THM15, TB19, VWC19, vS19]. Things-Enabled [HRMS^{+}21a]. Think [Bel11b, Hee11, WF06]. Thinking [Dar15, DBR^{+}08, Lan15, Pot10c, Sty04b, BHJM19]. Third [Mat12]. Third-Party [Mat12]. Thomas [McG13f]. Thorsteinson [For04a]. Thought [Lan07a]. Thousands [Pay19]. Thread [FD10]. Threat [AK15, BSH^{+}09, Dhi11, GT22, Gre13, GMC^{+}08, Jak19, Jon06, Kam14, Kel10, MA15a, Nic21b, PS09, Por06, SGSM22, SSH22, Ste10, SVR^{+}19, Tor05, TW21, TMU23]. Threat-Modeling [Tor05]. Threats [ARC19, Bai12b, Ber20a, BS14, Dun10, EGB18, HMR06, KPM^{+}19, LL23, LBF^{+}19, MBA^{+}22, PH18, RSOD22, Sch06a, SE13, Vau07, VVYY11, vdG17]. Three [MDD^{+}15, SCM08, SI09, CPS^{+}16]. Three-Ballot-Based [SCM08]. Thresholds [Lew11]. Thrown [Wil21]. Thwart [THS20]. Ticking [Vét21]. Tide [Gee11c]. Tiers [Ded17]. Ties [Cyb06a]. Tile [ADE^{+}22]. Time [AJW13, ASC15, BO14, CKPP23, Che17, FD11, GMB12, Gee10c, Gee11g, GR13, HMK12, Les04c, Les15b, LB08, PCK21, Pei20, Pot09, RI12, SSJ^{+}23, Sch04a, SHH^{+}21, TXL^{+}21, Wha12]. Time-Outing [GR13]. Time-to-Compromise [LB08]. Times [GC13, JP11]. Timing [DXS21, Hay13]. Timing-Based [DXS21]. Tiny [Cal13]. TIPPI [SS05]. Tipton [McG13g]. Today [Kel10, Pee13, Vis10, Ano20-48]. Together [CB14, GB09, HAO^{+}22, IAM03]. Token [RdRdCG10]. Tolerant [NS11, VNC^{+}06]. Tomorrow [MA15a]. Too [Ano18p, Eps12a, GFJ^{+}18, Ren12]. Tool [And12, CHMO07, CKL^{+}12, SND14, Sho23, YRX^{+}22]. Toolbox [PE12]. Toolkit [ABK^{+}04]. Tools [Bon08, CCS06, CA06, FOM23, GC06, GG11a, Hec03, Hu16, KP10, SGSM22, SC07, Ste15, vO22a]. Top [EW22, How09a, Mei18]. Topics [Woo23]. Tour [PvO24]. TPM [GY13]. Traceback [Alj03]. Traces [MB19]. Tracing [FD10, PCK21, San21, TJS^{+}22]. Tracking [Ano17-66, GL04]. Trade [Ano13b, Bre12, CMS09, Hoo16]. Tradeoff [Ksh14, PM21, SSH^{+}16, SS16]. Trading [Gee18a]. Traffic [AJW13, DRS16, GSLA15, RSOD22, SLKJ21]. Train [Pay19]. Training [Bis10b, CCR09, CPFJ14, Chu08, FTCS12, Hen09, HM12, HM22, How09a, How09b, KW12, Mat23, PE12, YZA08, vWS06]. Trans [Ano22-65]. Transaction [Mac06]. Transactional [PBW^{+}08]. Transactions [Ano18c, Ano18d, Ano19-28, Ano19-29, Ano19-30, Ano19-31, Ano19-32, Ano19-33, Ano19-34, Ano19-35, Ano20p, Ano20q, Ano20-48, Ano21-33, Ano21-34, Ano21-35, Ano21-36, Ano21-37, Ano21-38, Ano22z, Ano22-27, Ano22-28, Ano24k, Ano24l, Ano24m, Cal13, RdRdCG10, Ano19-38, Ano19-39]. Transatlantic [Cat11]. Transcription [LRW^{+}23]. Transfer [BCM03, PCMN08]. Transferring [BL13a]. Transfers [Tro06]. Transformation [Bis10b]. Transient [Pie23]. Transition [DOL13]. Transitioning [CKL^{+}12, MBLT13]. Translation [MTA23, PY06]. Transparency [AAB13, Ber20b, Mul14, Pau10a, Pfl14c, Pfl16a, SDO21]. Transparent [SW18a]. Transport [DPP15, Kob19]. Trap [EJ06]. Trapping [Spi03]. Tread [JJV^{+}09]. Treasures [TEN12]. Treaty [Mei04]. tree [Sah05]. Trenches [Dhi11]. Trend [GAA04]. Trends [Ahm08b, Arc04b, BHRR07, Gro08, JJV^{+}09, Mil05, Por09, Zan09, Zdr09]. Trial [HWB^{+}20]. Tricks [GY13]. Tricky [Ksh14]. Trigger [SS19b]. Triggers [Pay19]. Trojan [Fra07a]. Trouble [MOG^{+}20]. True [Cha04b]. TrueCrypt [BM14]. Truly [Kel10, LA10]. Trust [Ano14-39, Ano17t, Bar15, Bel22a, Ber21b, BSSB07, BC21, BDL^{+}14, GA10, GN07b, Heb22, Hor16, Kal12, Les14d, Lyo24, MRO^{+}23, Mic09, MKF21, Pet10, Pfl14c, RR06, SMJM11, Sch13d, Sun16, TMGP13, Trč11, Var09, Wil16a, Wil22]. Trust-Enhanced [Var09]. Trusted [Ano17z, BGMP08, Cur06, EKA14, Fel03, JSS20, Kau10a, KSL^{+}20, LL23, Pot09, Sch07a, SG20, Thi06, TCJ^{+}19, IS05, OR05]. Trusting [BDL^{+}14, Sur16, Wil22]. Trustworthiness [Mic19, Sch08a]. Trustworthy [EH12, HMK12, HSS11, LRW^{+}23, LL11, Mic22, SZ05, Sch23, SS05, SS22, SE07, Vie11b, YI13, ZHC^{+}20]. Trustworthy-Systems [YI13]. Truth [Ano22-74, Cox11, NROS24, RCT06]. Try [YK16]. Try-CybSI [YK16]. Trying [Gee08b, Sch16b, YK16]. TSA [GB12]. Turbulent [JP11]. Turing [BGMP19, Smi05b]. Turn [Tsa07]. turns [Don05b]. TV [Les05]. Tweeting [SBM^{+}13]. Twenty [LH23, BC22]. Twenty-First [BC22]. Twenty-Year [LH23]. Twice [Les04c]. Twin [DP20, Mat23]. Twins [EEA^{+}23]. Twitsper [SBM^{+}13]. Twitter [PRC17]. Two [Ahm08b, Av18, AS09, DeF11, DMB^{+}22, GD13, Hec03, Rya03]. Two-Factor [DeF11]. Type [GC08d]. Types [CRK^{+}13, Eps13]. Typing [PKW04]. Tyranny [CXG^{+}18].

U [SZ06]. U-Key [SZ06]. U.S. [NH23]. Ubiquitous [CM23]. Ugly [Av18, Les03e]. Ultimate [Fra07a, Qua03]. Umbrellas [Hor15]. Unattainable [San14]. Unauthorized [BS13c, TPMT15]. Uncertain [Sty04b]. Unclonable [BH14]. Unconventional [Bel06b]. Underbelly [Her19]. Undergraduate [BP08, OC12, Zat16]. Undermine [vdKHA^{+}20]. Underrepresented [PGP16]. Understanding [AAG15, AA15, BS24, DXS21, EOM09, Fel03, Gor06, GT22, GWS11, KMP20, MBR^{+}21, Mus21, PMA^{+}20, PKBS15, PH18, Sty04a, SCHC24, Ada05]. Undervalued [Sho23]. Undervolting [MOG^{+}20]. Undetectable [CEL^{+}19]. Unequal [SR16]. Unexpected [Blo11, SDKM20]. Unicode [MJF07]. Unintended [EPH12, Ess17, SS19b]. Uninviting [CS09]. Union [ADG23]. Unique [Kau10b]. United [NRH^{+}22, WMDX23]. Universities [HRDR05, Sch13b]. University [HNE^{+}08, Sch06c]. Unknowable [Gee19b]. Unknown [Gif10]. Unknowns [AFO13, Gee19b]. Unleashing [DP20, SCHC24]. Unlimited [Ano18-37]. Unmanned [BKS18]. Unmodified [Bon16]. Unnoticed [Bel18a]. Unobserved [Gee15b]. Unprecedented [NROS24]. Unpredictable [Sag13]. Unreasonable [Gla19]. Unrelated [PMA^{+}20]. Unsafe [Pei22, vO21b, Eps13]. Unseating [vO21b]. Untangling [vO20]. Untapped [CBN11, EKA14]. Until [Koh23a, Sch03c]. Unusable [Zur23]. Unveiling [MB19]. unwanted [PB05]. Unwavering [Ano17-84]. Unwinding [FD10]. Unwittingly [Ano17w]. Update [GP12, Gee21a]. Updated [KCC07]. Updates [Ano21-89, Ano22-87, CDD13]. Upend [Sch17a]. URL [KCC07]. USA [Fra07b, Ten16]. Usability [CB14, DD08, GG05, JW11, Kap07, KNBV17, LB04, SSH^{+}16, SS16, SC04, Yee04]. Usable [BPB^{+}04, BDGS04, CPS^{+}16, CG04, GLS23, GS16, Mal23, NSSS08, RCV^{+}12, RS19, TP11, TGC16, Zur22, ZH23]. Usage [Cal03a, PHS^{+}08]. USB [Ano17s]. Use [App11, CP10b, DSKB22, De 24, Don04h, Lan16, Les06, MCW08, MŽ22, Ost20, Pet09, RP09, WC04, WZE18]. Used [Pei19]. Useful [All07, Cra03, GAA04, ZS19]. Useless [Nic20b]. USENIX [EB19, Ano08g, Ano13f]. User [Ano17-66, AEY10, BC14, BHRR07, CDJS^{+}24, DEKM^{+}21, FTCS12, Gro08, Kni17, MR03, OdH12, PKW04, PRC17, PWVT12, Roe17, Sas07, Sch16b, SR16, VVYY11, Lin05]. User-Centric [BHRR07]. User-Customizable [Gro08]. User-Tailored [Kni17]. Users [BKS13, Cra12, RIC17, VVYY11, WDL21]. Uses [Les03b, McG13b]. Ushers [Gro14]. Using [AJV18, Ano17-66, BB17, BRS^{+}21, Ded17, Der13, DGU^{+}12, Fet16, GSB^{+}04, GW07, HCL11, Hec03, HSS11, JMD06, LGO04, LLGJ16, LLWC11, LH07, PEC^{+}14, dQSzL13, RTL09, SCZ^{+}13, SZ05, SFE10, SyPKZ22, SS19b, SBF^{+}15, TJS^{+}22, THS20, TMU23, WHF07, ZRM14]. Utility [Ano17u, CK10, El 20]. Utilization [Ksh14].

v [GM14, Les03c, Ten16]. V2X [KNL22]. Valid [Mas19]. Validate [AALS21]. Validation [FE11, HRTT03]. Validator [WO09]. Valley [MBLT13]. Valuating [HAF05]. Value [And12, ABAGOB10, Gri04, Irv11, Nic20b, Pre15, Sch15]. Variations [SJZG19]. VAULT [BFS^{+}21]. VAULT-Style [BFS^{+}21]. VAX [LJZ12]. VAX/SVS [LJZ12]. Vectors [Wil24]. Vegas [Jon06]. Vehicle [SSJ^{+}23]. Vehicles [HCL04]. Vehicular [ES14]. Veil [Lin09]. Vendor [Les12d]. Venn [MR08]. Venture [Bac13]. Ventured [GC10b]. Venus [Smi05b, TGS20]. Verifiability [RST15]. Verifiable [BCS16, Cha04b, CEC^{+}08, HCRS18, HWB^{+}20, KZZ17, RS17]. Verification [BCDS22, EH12, Ess17, FE11, Kob19, MRO^{+}23, Slo14, TCC^{+}24, Tom16]. Verified [ASH^{+}08]. Verify [HWF^{+}10, Les14d, Pet10, Sun16, WDL21]. Verifying [BFK16, CMT^{+}19, Dod22, MWM13, RC17]. Version [FG08, GSHU08]. Versus [MBP^{+}24, Can14, Sch08a, Wil16b]. via [CCX^{+}20]. Vibrations [SWCS21]. Victims [AA15]. Vidal [Eva15]. Video [AEP^{+}19, GG04, SPH^{+}05]. Video-Game [GG04]. View [AT13, BM14, KJC^{+}12, Ort14, vO22b]. Viewpoint [FWFA23, MTA23, OGM^{+}24, TAGM23]. Views [Hec03, Rya03, SMK^{+}20b]. Vigilantes [Mar13b]. Vigna [McG12b]. Village [Ber19]. Violating [Oeh05]. Violations [AHB04]. Violent [Hut24]. Virgil [McG09f]. Virtual [Arc07a, BS24, CAL^{+}24, Kan09, KS08a, LWLL10, MRF^{+}24, NBH08, WHD^{+}09, PEC^{+}14]. Virtualization [CZL08, CLS07, Coh10c, GG11a, KS08a, KS08b, PvDS08]. Virtues [Tro08]. Virus [Gor06]. Viruses [Pau05, Sch06a, For05]. Vision [May15a, SPH^{+}05]. Visions [DH13]. visit [Hol05]. Vista [Mic08b]. Visual [Bab16]. Vital [San14]. VM [LJZ12, LWLL10]. Voice [KPM^{+}19, LWL^{+}21, SWCS21, WK05, Ker10, MNS08]. Voice-over-IP [Ker10, MNS08]. Voices [Arc06]. VoIP [Alf05]. vol [Ano03b]. Volume [Ano03a, Ano06c]. Vote [BPR^{+}04, BRST17, KH12]. Voter [ASH^{+}08, Cha04b, CEC^{+}08]. Voter-Verifiable [Cha04b, CEC^{+}08]. Voter-Verified [ASH^{+}08]. Votes [YB08]. Voting [Ano16e, Ano16d, ASH^{+}08, Av18, BPR^{+}04, BF07, BCS16, Cha04b, CEC^{+}08, CEL^{+}19, DR04, Eps15a, EP04, GSHU08, GG04, HCRS18, HWB^{+}20, KH12, RR06, RJ08, RST15, SCM08, Sch04c, BW12, Eps13, KZZ17, SY12]. vs [Cal03b, FBG04, GN07a, KO08, KS12b, RCT06, Sch20c]. Vulnerabilities [CP09b, DXS21, Dia11, GWS11, HS17, How09a, Jon07, SPL^{+}24, Vis10, SCCB05]. Vulnerability [BC23b, BRS^{+}21, CCF^{+}20, CP09b, DFJ^{+}20, Hee11, MPM22, Mas24, MCKS10, McK07, McK08, MSR06, PR08, AT05]. vVote [BCS16]. VW [Res16].

W. [McG13g]. Wafaa [McG18f]. Wake [HB22]. Walking [PRS09]. Wall [How09b]. Walls [Bel13b]. Want [Str12]. Wanted [Pfl06]. War [Bel12a, BS13c, Don09a, Fel06, Fid11, Hut24, SPWS14]. Ware [Gel14, MBA12]. Warning [Kup06, Sto04]. Warnings [BLCDK11]. Warranties [Sty03]. Warranty [Tro04]. Warzala [McG13c]. Was [Fid11, Lan15]. Waste [Ark11]. Watch [Ano15-43, Ano20a]. Watchers [KT20]. Watches [Fri03]. Watching [LS08, MP07]. Water [Lan08b, MC12, PRS09]. Watermarking [HCL11]. Watts [Ano22a, Ano23b, Ano23-66]. Way [BSC^{+}08, Les07b, MKG09, RWD03, SRG03]. Ways [El 20, Sno05]. Weak [Cyb06a]. Weakening [BBCL13]. Weakest [Arc03b, BM10, SS06]. Weakness [Eps15a, NSA^{+}23]. Weaknesses [GGBM21, LL23]. Weapon [Lan11]. Weapons [MV22]. Wearable [vZH^{+}19].

Wearables [Fla16, SWCS21, vZH^{+}19]. Weather [Pet07]. Weaver [McG18e]. Web [CCF^{+}20, Ada05, And06a, And06b, AEV^{+}07, CFH05, CP09a, CA06, DeL07, Dre16, Gro08, GSLA15, HMR06, HAO^{+}22, JS07, Lin05, Mei06, MSW09, NKJ^{+}14, PRCC10, Pre15, RBE03, Sch04d, SC07, VdCA07, VE06, Vis10]. Web-Based [NKJ^{+}14, Ada05]. Website [Boy16, Gro12]. Websites [Fly11, RS11]. Week [Ano22-61, Ano23-43, Ano23-54, SD16a, SD16b, SD17]. Weinberger [McL05b]. Welcome [Ano22-88, Koh22e]. Welcomes [Ano03f]. Well [JCM12, Mar13b]. Wenyuan [McG13h]. Were [Pfl06]. Weren’t [Wil21]. West [McG16a, RWD03]. Whack [GY21]. Whack-a-Meltdown [GY21]. Wheat [All07]. Where [Bur06, Eva14, HKM17, JJV^{+}09, RBB19, SPL^{+}24, Wha11a]. wherever [Ano17-82, Ano18-52]. While [CDJS^{+}24]. White [BCGN16, Gee07, LKL^{+}19, Mic10b]. White-Box [BCGN16, Mic10b]. Whitelisting [Der13]. Whitfield [McG15f]. Whither [Gee23b, Sad16]. Who [Bel17b, BF06, Fri03, LM08, MP07, Sty07b, Wil21]. Whole [Eps15b]. Whose [Don04e, Les04c, Les14a]. Wi [GM14, Sty04b, Ohm14]. Wi-Fi [GM14, Sty04b, Ohm14]. Wide [HMR06, ZKH^{+}23]. WiFi [ACL07]. Wild [AEP^{+}19]. Will [EMM06, Fel03, Les03d, MM11, Mos18, Sch17a, Sch21]. Willis [Gel14]. Wimpy [ZYG15]. Win [KMP20, Les10a, Pal04, Som04]. Windows [Mic08b, DBR22, HL03, LH23, Mic08b, O’D08]. Wine [Che17]. Wineskin [Che17]. Winning [GMB12, McL05c]. Wired [Arb04]. Wireless [ACL07, Arb04, BS13a, BPVS04, HNE^{+}08, HP04, MK13, NSLW20, PA03, RC17, RLB^{+}23, SWYP12, Sty04b, Zan09]. Wiretap [Gid06]. Wiretapping [BBCL13, Lan05a, SCCB05]. Wisdom [Bel06b]. Within [SDKM20, JMD06, PS06, San14]. Without [Cal03e, KMP20, Ano17o, Ano17p, BBCL13, KZZ17, Lys07, Mar07, Sad16]. Witty [Don04d, SM04]. Wonderful [Wil24]. Won’t [JJ13, PP06a, Sas15]. Word [Bel13a, Bel13b, Bel18a, Ber21a, Gee13b, Gee14a, Sch23, Sch13d]. Words [DMB^{+}22]. Work [Ano17v, BF08, CIL^{+}21, CCS06, GB09, Hec11, RS11, Sas15, Sur16]. Workarounds [KKK^{+}21]. Worked [HAO^{+}22]. Workers [FOM23, HKMR24, ZH23]. Workforce [BC23b, BS20b, HBT12, NCA19, PMNT12, TA23]. Working [Bac13, BGM^{+}06]. Workings [NKJ^{+}14]. Works [Car09, Hea03d, MPM22, Pfl13a, SVR^{+}19, TCC^{+}24]. Workshop [RRS06, Wei04, McG03b]. Workshops [Ben16b, SI09]. World [Ano13b, Ano15-43, Ano16i, Ano16-44, Ano19-62, AJQA23, BCDS22, BFK16, BPS16, DPW11, DOP^{+}21, Eva14, Gra13, Kau09, KTS18, Lan14a, Lev03a, MS20, SS10, Smi03a, Sol07, Sty04b, Tom16, Val12, Lan13a, Ano08f, HMR06, Sty04b]. Worlds [Lan03]. Worm [CCW03, Lev05, Orm03, SM04, Sto04, CHK^{+}05, AL03, BCJ^{+}05]. Worms [Pau05, Sch06a, For05, SK05]. Worse [Blo12]. Worst [Lev06]. Worst-Case [Lev06]. Worth [Gee12b, Les11d]. Would [HAO^{+}22, MDR^{+}20]. Wrapped [KM16]. Write [Whi03]. Writers [Gor06]. Wrong [FH12, For04c, SS10].

X [For04a]. Xandra [NTMD^{+}18]. XML [GM09, RSM13]. XML-Based [GM09]. XTS [Mar10]. Xu [McG13h].

Y2K [Pay04]. Year [BO23, LRW^{+}23, LH23, Orm03, TEG^{+}19, Lan07b]. Years [Ahm08b, CCF14, DR10, HS17, KH12, MS21a, Pei23a, Sho22, Vie11b, Vie12d]. Yesterday [Cal09]. Yoshi [McG14f]. young [Don05b]. Younger [BS20a]. Yuck [Sas07].

Zatko  [McG16e].  Zero

[Ano17-66, Ber21b, Lev04a, MBC^{+}18]. Zero-Knowledge [MBC^{+}18]. Zero-Permission [Ano17-66]. ZeroConf [Ano17d]. Zigbee [RSWO18]. Zimmermann [McL06]. zoo [Hol05]. Zoombombing [LSBB22]. Zotob [Sch05d].

References

Aguirre:2012:IAS

[AA12] Idoia Aguirre and Sergio Alonso. Improving the automation of security information management: a collaborative approach. IEEE Security & Privacy, 10(1):55–59, January/February 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Arief:2015:UCSb

[AA15] Budi Arief and Mohd Azeem Bin Adzmi. Understanding cybercrime from its stakeholders’ perspectives: Part 2 — defenders and victims. IEEE Security & Privacy, 13(2):84– 88, March/April 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                                    URL http://www.computer.
                                                    org/csdl/mags/sp/2015/02/
                                                    msp2015020084-abs.html.

Acquisti:2013:GSL

[AAB13] Alessandro Acquisti, Idris Adjerid, and Laura Brandimarte. Gone in 15 seconds: The limits of privacy transparency and control. IEEE Security & Privacy, 11(4):72–74, July/ August 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Aikat:2017:RSE

[AAC^{+}17] Jay Aikat, Aditya Akella, Jeffrey S. Chase, Ari Juels, Michael K. Reiter, Thomas Ristenpart, Vyas Sekar, and Michael Swift. Rethinking security in the era of cloud computing. IEEE Security & Privacy, 15(3):60– 69, May/June 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/03/
msp2017030060-abs.html.

Abrams:2019:AIE

[AACG19] M. Abrams, J. Abrams, P. Cullen, and L. Goldstein. Artificial intelligence, ethics, and enhanced data stewardship. IEEE Security & Privacy, 17(2):17–30, March/ April 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Arief:2015:UCSa

[AAG15] Budi Arief, Mohd Azeem Bin Adzmi, and Thomas Gross. Understanding cybercrime from its stakeholders’ perspectives: Part 1 — attackers. IEEE Security & Privacy, 13(1):71–76, January/February 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.

                                        org/csdl/mags/sp/2015/01/
                                        msp2015010071-abs.html.

Ali:2021:WWH

[AALS21] Sameed Ali, Prashant Anantharaman, Zephyr Lucas, and Sean W. Smith. What we have here is failure to validate: Summer of LangSec. IEEE Security & Privacy, 19(3):17–23, May/ June 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Andoh-Baidoo:2010:HIS

[ABAGOB10] Francis Kofi Andoh-Baidoo, Kwasi Amoako-Gyampah, and Kweku-Muata Osei-Bryson. How Internet security breaches harm market value. IEEE Security & Privacy, 8(1):36–42, January/February 2010. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Avgerinos:2018:MCR

[ABD^{+}18] Thanassis Avgerinos, David Brumley, John Davis, Ryan Goulden, Tyler Nighswander, Alex Rebert, and Ned Williamson. The Mayhem cyber reasoning system. IEEE Security & Privacy, 16(2):52– 60, March/April 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/02/
                                        msp2018020052-abs.html.

Aslam:2004:KTI

[ABK^{+}04] Javed Aslam, Sergey Bratus, David Kotz, Ron Peterson, Brett Tofel, and Daniela Rus.

The Kerf Toolkit for intrusion analysis. IEEE Security & Privacy, 2(6):42–52, November/December 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/06/j6042.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/06/j6042.
pdf.

Ardito:2022:PPG

[ABMS22] Luca Ardito, Luca Barbato, Paolo Mori, and Andrea Saracino. Preserving privacy in the globalized smart home: The SIFIS-Home Project. IEEE Security & Privacy, 20(1):33–44, January/February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Archer:2016:MPP

[ABPP16] David W. Archer, Dan Bogdanov, Benny Pinkas, and Pille Pullonen. Maturity and performance of programmable secure computation. IEEE Security & Privacy, 14(5):48–56, September/October 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/05/
msp2016050048-abs.html.

Ahmad:2021:SAT

[AC21] Wajeeha Ahmad and David D.

Clark. A systems approach toward addressing anonymous abuses: Technical and policy considerations. IEEE Security & Privacy, 19(2):38–47, March/April 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Arias-Cabarcos:2015:BIP

[ACAT^{+}15] Patricia Arias-Cabarcos, Florina Almenarez, Ruben Trapero, Daniel Diaz-Sanchez, and Andres Marin. Blended identity: Pervasive IdM for continuous authentication. IEEE Security & Privacy, 13(3):32– 39, May/June 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/sp/2015/03/
                                        msp2015030032-abs.html.

Anthony:2013:SIT

[ACC^{+}13] Denise Anthony, Andrew T. Campbell, Thomas Candon, Andrew Gettinger, David Kotz, Lisa A. Marsch, Andres Molina-Markham, Karen Page, Sean W. Smith, Carl A. Gunter, and M. Eric Johnson. Securing information technology in healthcare. IEEE Security & Privacy, 11(6):25– 33, November/December 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Aime:2007:DWN

[ACL07] Marco Domenico Aime, Giorgio Calandriello, and Antonio Lioy. Dependability in wireless networks: Can we rely on WiFi? IEEE Security & Privacy, 5(1):23–29, January/ February 2007. CODEN ????

ISSN 1540-7993 (print), 1558-4046 (electronic).

Allodi:2020:NNA

[ACPZ20] L. Allodi, T. Chotza, E. Panina, and N. Zannone. The need for new antiphishing measures against spear-phishing attacks. IEEE Security & Privacy, 18 (2):23–34, March/April 2020. ISSN 1558-4046.

Acquisti:2008:IMP

[Acq08] Alessandro Acquisti. Identity management, privacy, and price discrimination. IEEE Security & Privacy, 6(2):46–50, March/April 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Acquisti:2009:NPB

[Acq09] Alessandro Acquisti. Nudging privacy: The behavioral economics of personal information. IEEE Security & Privacy, 7(6):82–85, November/ December 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Acquisti:2013:CPP

[Acq13] Alessandro Acquisti. Complementary perspectives on privacy and security: Economics. IEEE Security & Privacy, 11 (2):93–95, March/April 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Adams:2005:BSW

[Ada05] C. Adams. Building secure Web-based environments: understanding research interrelationships through a construction metaphor. IEEE Security & Privacy, 3(1):74– 77, January/February 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/30310/01392706.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
30310&arnumber=1392706&count=
17&index=13.

Adams:2015:PMD

[Ada15] A. A. Adams. Possessing mobile devices. IEEE Security & Privacy, 13(6):89–95, November/December 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Aharoni:2022:CET

[ADE^{+}22] Ehud Aharoni, Nir Drucker, Gilad Ezov, Hayim Shaul, and Omri Soceanu. Complex encoded tile tensors: Accelerating encrypted analytics. IEEE Security & Privacy, 20(5):35– 43, September/October 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Adamos:2023:AEU

[ADG23] Konstantinos Adamos, Fabio Di Franco, and Athanasios Grammatopoulos. An analysis of European Union cybersecurity higher education programs through the crowdsourced database CyberHEAD. IEEE Security & Privacy, 21 (5):85–94, September/October 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Ali:2021:PCS

[AED^{+}21] Suzan Ali, Mounir Elgharabawy, Quentin Duchaussoy, Mohammad Mannan, and Amr Youssef. Parental controls: Safer Internet solutions or new pitfalls? IEEE Security & Privacy, 19(6):36–46, November/ December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anton:2004:FPP

[AEH^{+}04] Annie I. Antón, Julia B. Earp, Qingfeng He, William Stufflebeam, Davide Bolchini, University of Lugano, and Carlos Jensen. Financial privacy policies and the need for standardization. IEEE Security & Privacy, 2(2):36–45, March/ April 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2004/02/j2036abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/02/j2036.
pdf.

Axenopoulos:2019:FLS

[AEP^{+}19] A. Axenopoulos, V. Eiselein, A. Penta, E. Koblents, E. La Mattina, and P. Daras. A framework for large-scale analysis of video “in the wild” to assist digital forensic examination. IEEE Security & Privacy, 17(1):23–33, January/ February 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anton:2007:HEW

[AEV^{+}07] Annie I. Antón, Julia B. Eart, Matthew W. Vail, Neha Jain, Carrie M. Gheen, and Jack M. Frink. HIPAA’s effect on Web site privacy policies. IEEE Security & Privacy, 5(1):45–52, January/February 2007. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anton:2010:HIU

[AEY10] Annie I. Anton, Julia B. Earp, and Jessica D. Young. How Internet users’ privacy concerns have evolved since 2002. IEEE Security & Privacy, 8(1):21–27, January/February 2010. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Arbaugh:2011:LI

[AF11] William Arbaugh and Deborah A. Frincke. Living with insecurity. IEEE Security & Privacy, 9(6):12–13, November/ December 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Alves-Foss:2016:DCG

[AF16] Jim Alves-Foss. The DARPA Cyber Grand Challenge: A competitor’s perspective, part 2. IEEE Security & Privacy, 14(1):76–81, January/ February 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Alonso-Fernandez:2012:QMB

[AFFOG12] Fernando Alonso-Fernandez, Julian Fierrez, and Javier Ortega-Garcia. Quality measures in biometric systems. IEEE Security & Privacy, 10 (6):52–62, November/December 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Alves-Foss:2013:KU

[AFO13] Jim Alves-Foss and Paul Oman. The known unknowns. IEEE Security & Privacy, 11 (5):77–79, September/October 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Acquisti:2005:PRI

[AG05] A. Acquisti and J. Grossklags.

Privacy and rationality in individual decision making. IEEE Security & Privacy, 3 (1):26–33, January/February 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/30310/01392696.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
30310&arnumber=1392696&count=
17&index=5.

Ayday:2017:IAA

[AH17] Erman Ayday and Mathias Humbert. Inference attacks against kin genomic privacy. IEEE Security & Privacy, 15(5):29–37, September/October 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/05/
                                        msp2017050029-abs.html.

Anton:2004:IJP

[AHB04] Annie I. Antón, Qingfeng He, and David L. Baumer. Inside JetBlue’s privacy policy violations. IEEE Security & Privacy, 2(6):12–18, November/December 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/sp/2004/06/j6012.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2004/06/j6012.
                                        pdf.

Ahmad:2007:CSS

[Ahm07] David Ahmad. The contemporary software security landscape. IEEE Security & Privacy, 5(3):75–77, May/June 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ahmad:2008:CDD

[Ahm08a] Dave Ahmad. The confused deputy and the domain hijacker. IEEE Security & Privacy, 6(1):74–77, January/ February 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ahmad:2008:ATT

[Ahm08b] David Ahmad. Attack trends: Two years of broken crypto: Debian’s dress rehearsal for a global PKI compromise. IEEE Security & Privacy, 6(5):70–73, September/October 2008. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Alemzadeh:2013:ASC

[AIKR13] Homa Alemzadeh, Ravishankar K. Iyer, Zbigniew Kalbarczyk, and Jai Raman. Analysis of safety-critical computer failures in medical devices. IEEE Security & Privacy, 11(4):14–26, July/ August 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ayub:2023:HIC

[AJQA23] Adeen Ayub, Wooyeon Jo, Syed Ali Qasim, and Irfan Ahmed. How are industrial control systems insecure by design? A deeper insight into real-world programmable logic controllers. IEEE Security & Privacy, 21(4):10–19, July/ August 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Albanese:2018:DSB

[AJV18] Massimiliano Albanese, Sushil Jajodia, and Sridhar Venkatesan. Defending from stealthy botnets using moving target defenses. IEEE Security & Privacy, 16(1):92–97, January/February 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/01/
msp2018010092.html.

Ahmed:2013:RTD

[AJW13] Abdulghani Ali Ahmed, Aman Jantan, and Tat-Chee Wan. Real-time detection of intrusive traffic in QoS network domains. IEEE Security & Privacy, 11(6):45–53, November/ December 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Agrawal:2003:MAD

[AK03] Dakshi Agrawal and Dogan Kesdogan. Measuring anonymity: The disclosure attack. IEEE Security & Privacy, 1(6):27–34, November/ December 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/sp/2003/06/j6027abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2003/06/j6027.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2003/06/j6027.
                                        pdf.

AlSabbagh:2015:STF

[AK15] Bilal Al Sabbagh and Stewart Kowalski. A socio-technical framework for threat modeling a software supply chain. IEEE Security & Privacy, 13(4):30– 39, July/August 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/sp/2015/04/
                                        msp2015040030-abs.html.

Arlat:2012:NSD

[AKN12] Jean Arlat, Zbigniew Kalbarczyk, and Takashi Nanya. Nanocomputing: Small devices, large dependability challenges. IEEE Security & Privacy, 10(1):69–72, January/ February 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Arce:2003:ASW

[AL03] Iván Arce and Elias Levy. An analysis of the Slapper Worm. IEEE Security & Privacy, 1(1): 82–87, January/February 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j1082.pdf;   http://www.
computer.org/security/j1082abs.
htm.

Alfonsi:2005:AAV

[Alf05] Benjamin Alfonsi. Alliance addresses VoIP security. IEEE Security & Privacy, 3(4): 8, July/August 2005. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/comp/mags/sp/2005/04/
j4008.pdf.

Aljifri:2003:ITN

[Alj03]  Hassan Aljifri.  IP trace-

back: a new denial-of-service deterrent? IEEE Security & Privacy, 1(3):24–31, May/June 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://dlib.computer.
                                        org/sp/books/sp2003/pdf/
                                        j3024.pdf;   http://www.
                                        computer.org/security/j3024abs.
                                        htm.

Allen:2005:CF

[All05] William H. Allen. Computer forensics. IEEE Security & Privacy, 3(4):59–62, July/August 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Allen:2007:MWC

[All07] William H. Allen. Mixing wheat with the chaff: Creating useful test data for IDS evaluation. IEEE Security & Privacy, 5(4):65–67, July/August 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Arce:2004:GEI

[AM04] Iváan Arce and Gary Mc-Graw. Guest Editors’ introduction: Why attacking systems is a good idea. IEEE Security & Privacy, 2(4):17–19, July/August 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/sp/2004/04/j4017.
                                        htm.

AlSadeh:2012:SND

[AM12] Ahmad AlSa’deh and Christoph Meinel. Secure neighbor discovery: Review, challenges, perspectives, and recommendations. IEEE Security & Privacy, 10(4):26–34, July/ August 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Akhtar:2015:BLD

[AMF15] Zahid Akhtar, Christian Micheloni, and Gian Luca Foresti. Biometric liveness detection: Challenges and research opportunities. IEEE Security & Privacy, 13(5):63–72, September/October 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/csdl/mags/sp/2015/05/
msp2015050063-abs.html.

Amin:2003:NAE

[Ami03] Massoud Amin. North America’s electricity infrastructure: Are we ready for more perfect storms? IEEE Security & Privacy, 1(5):19–25, September/ October 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2003/05/j5019abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/05/j5019.
pdf.

Amin:2005:GEI

[Ami05] M. Amin. Guest Editor’s introduction: Infrastructure security–reliability and dependability of critical systems. IEEE Security & Privacy, 3(3):15–17, May/June 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/31002/01439497.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        31002&arnumber=1439497&count=
                                        20&index=4.

Al-HajBaddar:2019:BAD

[AMM19] S. Al-Haj Baddar, A. Merlo, and M. Migliardi. Behavioralanomaly detection in forensics analysis. IEEE Security & Privacy, 17(1):55–62, January/ February 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Amoroso:2013:EPM

[Amo13] Edward G. Amoroso. From the enterprise perimeter to a mobility-enabled secure cloud. IEEE Security & Privacy, 11 (1):23–31, January/February 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Amo:2016:AGG

[Amo16] Laura Amo. Addressing gender gaps in teens’ cybersecurity engagement and selfefficacy. IEEE Security & Privacy, 14(1):72–75, January/ February 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Abu-Nimeh:2010:PDB

[ANC10] Saeed Abu-Nimeh and Thomas Chen. Proliferation and detection of blog spam. IEEE Security & Privacy, 8(5):42–47, September/October 2010. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Andresen:2003:NDS

[And03a] Scott L. Andresen. New draft to secure cyberspace leaked. IEEE Security & Privacy, 1(1): 13, January/February 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j1013.pdf.

Andresen:2003:NB

[And03b] Scott L. Andresen. News briefs. IEEE Security & Privacy, 1(5):12–13, September/October 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2003/05/j5012.
pdf.

Andresen:2004:NBP

[And04] Scott L. Andresen. News briefs: Policy; privacy; security. IEEE Security & Privacy, 2(1):18–19, January/ February 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2004/01/j1020abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/01/j1020.
pdf.

Anderson:2006:WSP

[And06a] Anne Anderson. Web services policies (abstract). IEEE Security & Privacy, 4(3):84–87, May/June 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Andrews:2006:GEI

[And06b] Mike Andrews. Guest Editor’s introduction: The state of Web security. IEEE Security & Privacy, 4(4):14–15, July/ August 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/sp/2006/04/j4014.pdf.

Anderson:2007:SSS

[And07] Ross Anderson. Software security: State of the art. IEEE Security & Privacy, 5(1):8, January/February 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anderson:2012:MVS

[And12] Paul Anderson. Measuring the value of static-analysis tool deployments. IEEE Security & Privacy, 10(3):40–47, May/ June 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2003:AII

[Ano03a] Anonymous. 2003 annual index IEEE Security & Privacy Volume 1. IEEE Security & Privacy, 1(6):83–88, November/December 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/sp/2003/06/
                                        j6083.pdf;   http://csdl.
                                        computer.org/dl/mags/sp/
                                        2003/06/j6083.htm.

Anonymous:2003:EHV

[Ano03b] Anonymous. Errata: “On the Horizon” (vol. 1, no. 2) and “Interface” (vol. 1, no. 2, p. 9). IEEE Security & Privacy, 1(3):7, May/June 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/

j3008.pdf. See [McG03b, Ano03c].

Anonymous:2003:LEa

[Ano03c] Anonymous. Letters to the Editor. IEEE Security & Privacy, 1(2):7–10, March/April 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j2007.pdf;   http://www.
computer.org/security/j2007abs.

htm. See erratum [Ano03b].

Anonymous:2003:LEb

[Ano03d] Anonymous. Letters to the editor. IEEE Security & Privacy, 1(5):14, September/ October 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2003/05/j5014.pdf.

Anonymous:2003:N

[Ano03e] Anonymous. News. IEEE Security & Privacy, 1(3):8– 13, May/June 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.

                                        org/sp/books/sp2003/pdf/
                                        j3008.pdf.

Anonymous:2003:SPW

[Ano03f] Anonymous. Security and Privacy welcomes new Editorial Board members. IEEE Security & Privacy, 1(5):6–7, September/October 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/sp/2003/05/
                                        j5006.pdf.

Anonymous:2004:AI

[Ano04a]  Anonymous.  2004 annual

index. IEEE Security & Privacy, 2(6):80–85, November/December 2004. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/sp/2004/06/
                                        j6080.pdf;   http://csdl.
                                        computer.org/dl/mags/sp/
                                        2004/06/j6080.htm.

Anonymous:2004:LE

[Ano04b] Anonymous. Letters to the Editor. IEEE Security & Privacy, 2(5):6–7, September/October 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/sp/2004/05/
                                        j5006.pdf;   http://csdl.
                                        computer.org/dl/mags/sp/
                                        2004/05/j5006.htm.

Anonymous:2004:PCP

[Ano04c] Anonymous. Protecting consumers’ private health information. IEEE Security & Privacy, 2(5):12, September/October 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/comp/mags/sp/2004/05/
j5012.pdf;   http://csdl.
computer.org/dl/mags/sp/
2004/05/j5012.htm.

Anonymous:2004:RT

[Ano04d] Anonymous. Reviewer thanks.

IEEE Security & Privacy, 2 (1):6, January/February 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/comp/mags/sp/2004/01/
j1006.pdf.

Anonymous:2005:AI

[Ano05] Anonymous. 2005 annual index. IEEE Security & Privacy, 3(6):90–95, November/ December 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2006:EC

[Ano06a] Anonymous. 2006 editorial calendar. IEEE Security & Privacy, 4(1):61, January/ February 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2006:ICS

[Ano06b]  Anonymous.  IEEE Com-

puter Society Distance Learning Campus. IEEE Security & Privacy, 4(6):35, November/ December 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2006:ISPb

[Ano06c] Anonymous. IEEE Security & Privacy 2006 annual index, volume 4. IEEE Security & Privacy, 4(6):90–95, November/ December 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2006:ISPa

[Ano06d] Anonymous. IEEE Security & Privacy 2006 editorial calendar. IEEE Security & Privacy, 4(2):92, March/April 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2006:LE

[Ano06e] Anonymous. Letters to the editor. IEEE Security & Privacy, 4(5):5, September/ October 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/sp/2006/05/j5005.pdf.

Anonymous:2006:M

[Ano06f] Anonymous. Masthead. IEEE Security & Privacy, 4(6): 4, November/December 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2006:RC

[Ano06g] Anonymous. RSA(R) Conference 2007. IEEE Security & Privacy, 4(6):14, November/ December 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2006:STS

[Ano06h] Anonymous. Special thanks to S&P’s reviewers. IEEE Security & Privacy, 4(1):7–8, January/February 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2007:I

[Ano07a] Anonymous. Interface. IEEE Security & Privacy, 5(5):5, September/October 2007. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/comp/mags/sp/2007/05/
msp2007050005.pdf.

Anonymous:2007:STS

[Ano07b] Anonymous. Special thanks to S&P’s reviewers. IEEE Security & Privacy, 5(1):6–7, January/February 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2008:MA

[Ano08a] Anonymous. 2008 membership advertisement. IEEE Security & Privacy, 6(4):71–73, July/ August 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2008:AI

[Ano08b] Anonymous. Annual index. IEEE Security & Privacy, 6(6): 0, November/December 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2008:CPO

[Ano08c] Anonymous. Call for papers: Online Gaming Security. IEEE Security & Privacy, 6(4):57, July/August 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2008:CNS

[Ano08d] Anonymous. Corporate network security advertisement. IEEE Security & Privacy, 6(4): c3, July/August 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2008:FC

[Ano08e]  Anonymous.  Front cover.

IEEE Security & Privacy, 6(4): c1, July/August 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2008:ISW

[Ano08f] Anonymous. IT Security World advertisement. IEEE Security & Privacy, 6(4):c2, July/ August 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2008:USS

[Ano08g] Anonymous. Usenix Security Symposium advertisement. IEEE Security & Privacy, 6(4): c4, July/August 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2009:AI

[Ano09a] Anonymous. Annual index.

IEEE Security & Privacy, 7(6): 1, November/December 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2009:ILE

[Ano09b] Anonymous. Interface: Letters to the editor. IEEE Security & Privacy, 7(4):6–7, July/ August 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2009:STS

[Ano09c] Anonymous. Special thanks:

Special thanks to S&P’s peer reviewers. IEEE Security & Privacy, 7(2):5–6, March/April 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2011:RT

[Ano11] Anonymous. 2010 reviewer thanks. IEEE Security & Privacy, 9(1):5–6, January/ February 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2013:FC

[Ano13a]  Anonymous.  Front cover.

IEEE Security & Privacy, 11 (6):c1, November/December 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2013:IWC

[Ano13b] Anonymous. InfoSec World Conference 2014 trade advertisement. IEEE Security & Privacy, 11(6):c3, November/ December 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2013:MSH

[Ano13c] Anonymous. Magazine subscribe [house advertisement]. IEEE Security & Privacy, 11 (6):c4, November/December 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2013:M

[Ano13d] Anonymous. Masthead. IEEE Security & Privacy, 11(6): 4, November/December 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2013:TC

[Ano13e] Anonymous. Table of contents. IEEE Security & Privacy, 11 (6):1–2, November/December 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2013:UHA

[Ano13f] Anonymous. Usenix [house advertisement]. IEEE Security & Privacy, 11(6):c2, November/ December 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2014:A

[Ano14a]  Anonymous.  [Advertise-

ment]. IEEE Security & Privacy, 12(6):c4, November/December 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/06/
msp20140600c4.pdf.

Anonymous:2014:CSA

[Ano14b]  Anonymous.  Co3 systems

advertisement. IEEE Security & Privacy, 12(5):c2, September/October 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/05/
msp20140500c2.pdf.

Anonymous:2014:FYJb

[Ano14c] Anonymous. Focus on your job search [advertisement]. IEEE Security & Privacy, 12(6): c2, November/December 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/06/
msp20140600c2.pdf.

Anonymous:2014:FYJa

[Ano14d] Anonymous. Focus on your job search [house advertisement]. IEEE Security & Privacy, 12 (3):c2, May/June 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2014:FCa

[Ano14e]  Anonymous.  Front cover.

IEEE Security & Privacy, 12 (1):c1, January/February 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2014:FCb

[Ano14f]  Anonymous.  Front cover.

IEEE Security & Privacy, 12 (2):c1, March/April 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2014:FCc

[Ano14g] Anonymous. [Front cover]. IEEE Security & Privacy, 12 (4):c1, July/August 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/sp/2014/04/
                                        msp20140400c1.pdf.

Anonymous:2014:FCd

[Ano14h] Anonymous. [Front cover]. IEEE Security & Privacy, 12 (6):c1, November/December 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

                                        //www.computer.org/csdl/
                                        mags/sp/2014/06/msp20140600c1.
                                        pdf.

Anonymous:2014:ICS

[Ano14i] Anonymous. IEEE Computer Society [advertisement]. IEEE Security & Privacy, 12(6): 67, November/December 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/06/
msp2014060067.pdf.

Anonymous:2014:ISPa

[Ano14j] Anonymous. IEEE Security & Privacy [advertisement]. IEEE Security & Privacy, 12(3):c3, May/June 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2014:ISPb

[Ano14k] Anonymous. IEEE Security & Privacy [advertisement]. IEEE Security & Privacy, 12(4): c3, July/August 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/04/
msp20140400c3.pdf.

Anonymous:2014:ISPc

[Ano14l] Anonymous. IEEE Security & Privacy [advertisement]. IEEE Security & Privacy, 12(5):c3, September/October 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/05/
msp20140500c3.pdf.

Anonymous:2014:ISPd

[Ano14m] Anonymous. IEEE security & privacy [advertisement]. IEEE Security & Privacy, 12(6): c3, November/December 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.

                                        org/csdl/mags/sp/2014/06/
                                        msp20140600c3.pdf.

Anonymous:2014:IA

[Ano14n]  Anonymous.  Intelect [ad-

vertisement]. IEEE Security & Privacy, 12(5):94, September/October 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/sp/2014/05/
                                        msp2014050094.pdf.

Anonymous:2014:JBH

[Ano14o] Anonymous. Jobs board [house advertisement]. IEEE Security & Privacy, 12(2):27, March/ April 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2014:MSH

[Ano14p] Anonymous. Magazine subscribe [house advertisement]. IEEE Security & Privacy, 12 (2):c2, March/April 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2014:Ma

[Ano14q] Anonymous. Masthead. IEEE Security & Privacy, 12(1):5, January/February 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2014:Mb

[Ano14r] Anonymous. Masthead. IEEE Security & Privacy, 12(2):5, March/April 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2014:Mc

[Ano14s] Anonymous. [Masthead]. IEEE Security & Privacy, 12(3):5, May/June 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2014:Md

[Ano14t] Anonymous. [Masthead]. IEEE Security & Privacy, 12(4): 7, July/August 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/04/
msp2014040007.pdf.

Anonymous:2014:Me

[Ano14u] Anonymous. [Masthead]. IEEE Security & Privacy, 12(5):5, September/October 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/05/
msp2014050005.pdf.

Anonymous:2014:Mf

[Ano14v] Anonymous. [Masthead]. IEEE Security & Privacy, 12(6): 6, November/December 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/06/
msp2014060006.pdf.

Anonymous:2014:MMH

[Ano14w]  Anonymous.  Membership

matters [house advertisement]. IEEE Security & Privacy, 12 (2):6, March/April 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2014:RT

[Ano14x] Anonymous. Reviewer thanks. IEEE Security & Privacy, 12 (6):7–8, November/December 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//www.computer.org/csdl/
mags/sp/2014/06/msp2014060007.
html.

Anonymous:2014:RSB

[Ano14y] Anonymous. Rock stars of big data analytics [advertisement]. IEEE Security & Privacy, 12(5):c4, September/October 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/05/
msp20140500c4.pdf.

Anonymous:2014:RSCb

[Ano14z] Anonymous. Rock stars of cybersecurity [advertisement]. IEEE Security & Privacy, 12 (4):c4, July/August 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/04/
msp20140400c4.pdf.

Anonymous:2014:RSCa

[Ano14-27] Anonymous. Rock stars of cybersecurity [house advertisement]. IEEE Security & Privacy, 12(3):c4, May/June 2014.

CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2014:RSMa

[Ano14-28] Anonymous. Rock stars of mobile cloud [house advertisement]. IEEE Security & Privacy, 12(1):c2, January/ February 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2014:RSMb

[Ano14-29] Anonymous. Rock stars of mobile cloud [house advertisement]. IEEE Security & Privacy, 12(2):c4, March/April 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2014:SISb

[Ano14-30] Anonymous. Seeking IEEE Security & Security Editor in Chief. IEEE Security & Privacy, 12(4):c2, July/August 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/04/
msp20140400c2.pdf.

Anonymous:2014:SES

[Ano14-31] Anonymous. Software experts summit [house advertisement]. IEEE Security & Privacy, 12 (2):c3, March/April 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2014:SISa

[Ano14-32] Anonymous. Subscribe to IEEE Security & Privacy [house advertisement]. IEEE Security & Privacy, 12(1):c3, January/February 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2014:TCa

[Ano14-33] Anonymous. Table of contents. IEEE Security & Privacy, 12(1):1–2, January/February 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2014:TCb

[Ano14-34] Anonymous. Table of contents. IEEE Security & Privacy, 12(2):1–2, March/April 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2014:TCc

[Ano14-35] Anonymous. Table of contents. IEEE Security & Privacy, 12 (3):1–2, May/June 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2014:TCd

[Ano14-36] Anonymous. Table of contents. IEEE Security & Privacy, 12 (4):1–2, July/August 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/sp/2014/04/
                                        msp2014040001.pdf.

Anonymous:2014:TCe

[Ano14-37] Anonymous. Table of contents. IEEE Security & Privacy, 12(5):1–2, September/October 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/05/
msp2014050001.pdf.

Anonymous:2014:TCf

[Ano14-38] Anonymous. Table of contents. IEEE Security & Privacy, 12(6):1–2, November/December 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/06/
msp2014060001.pdf.

Anonymous:2014:TAB

[Ano14-39] Anonymous. Taking action to build trust in security. IEEE Security & Privacy, 12(2):3–4, March/April 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2015:RMDa

[Ano15a] Anonymous. 2016 Richard E.

Merwin Distinguished Service Award house advertisement. IEEE Security & Privacy, 13 (4):c2, July/August 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/04/
msp20150400c2.pdf.

Anonymous:2015:RMDb

[Ano15b] Anonymous. 2016 Richard E. Merwin Distinguished Service Award house advertisement. IEEE Security & Privacy, 13(5):11, September/October 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/csdl/mags/sp/2015/05/
msp2015050011.pdf.

Anonymous:2015:AIC

[Ano15c] Anonymous. 39th Annual International Computers, Software & Applications Conference house advertisement. IEEE Security & Privacy, 13 (2):11, March/April 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/02/
msp2015020011.pdf.

Anonymous:2015:ABS

[Ano15d] Anonymous. 3rd Annual Best Scientific Cybersecurity Paper Competition [house advertisement]. IEEE Security & Privacy, 13(1):c4, January/February 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/01/
msp20150100c4.pdf.

Anonymous:2015:CNH

[Ano15e] Anonymous. Call for nominees house advertisement. IEEE Security & Privacy, 13(5):73, September/October 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/csdl/mags/sp/2015/05/
msp2015050073.pdf.

Anonymous:2015:CPH

[Ano15f] Anonymous. Call for papers house advertisement. IEEE Security & Privacy, 13(6): 51, November/December 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2015:CSA

[Ano15g] Anonymous. Call for standards award nominations house advertisement. IEEE Security & Privacy, 13(5):62, September/October 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/csdl/mags/sp/2015/05/
msp2015050062.pdf.

Anonymous:2015:FYJa

[Ano15h] Anonymous. Focus on your job search house advertisement. IEEE Security & Privacy, 13 (2):47, March/April 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/02/
msp2015020047.pdf.

Anonymous:2015:FYJb

[Ano15i] Anonymous. Focus on your job search house advertisement. IEEE Security & Privacy, 13 (3):87, May/June 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/sp/2015/03/
                                        msp2015030087.pdf.

Anonymous:2015:FYJc

[Ano15j] Anonymous. Focus on your job search house advertisement. IEEE Security & Privacy, 13(5):7, September/October 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://csdl.computer.
                                        org/csdl/mags/sp/2015/05/
                                        msp2015050007.pdf.

Anonymous:2015:FYJd

[Ano15k] Anonymous. Focus on your job search house advertisement. IEEE Security & Privacy, 13 (6):c3, November/December 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2015:FCa

[Ano15l]  Anonymous.  Front cover.

IEEE Security & Privacy, 13 (1):c1, January/February 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/sp/2015/01/
                                        msp20150100c1.pdf.

Anonymous:2015:FCb

[Ano15m]  Anonymous.  Front cover.

IEEE Security & Privacy, 13 (2):c1, March/April 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/02/
msp20150200c1.pdf.

Anonymous:2015:FCc

[Ano15n]  Anonymous.  Front cover.

IEEE Security & Privacy, 13 (3):c1, May/June 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/03/
msp20150300c1.pdf.

Anonymous:2015:FCd

[Ano15o]  Anonymous.  Front cover.

IEEE Security & Privacy, 13 (4):c1, July/August 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/04/
msp20150400c1.pdf.

Anonymous:2015:FCe

[Ano15p]  Anonymous.  Front cover.

IEEE Security & Privacy, 13(5):c1, September/October 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/csdl/
mags/sp/2015/05/msp20150500c1.
pdf.

Anonymous:2015:FCf

[Ano15q]  Anonymous.  Front cover.

IEEE Security & Privacy, 13 (6):c1, November/December 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2015:GMLa

[Ano15r] Anonymous. Get more, for less! House advertisement. IEEE Security & Privacy, 13 (2):c4, March/April 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/02/
msp20150200c4.pdf.

Anonymous:2015:GMLb

[Ano15s] Anonymous. Get more, for less! House advertisement. IEEE Security & Privacy, 13 (3):c4, May/June 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/03/
msp20150300c4.pdf.

Anonymous:2015:GRY

[Ano15t] Anonymous. Get the recognition you deserve house advertisement. IEEE Security & Privacy, 13(6):29, November/ December 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2015:ICSb

[Ano15u]  Anonymous.  IEEE Com-

puter Society 2015 call for major award nominations house advertisement. IEEE Security & Privacy, 13(5):33, September/October 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/csdl/mags/sp/2015/05/
msp2015050033.pdf.

Anonymous:2015:ICSc

[Ano15v] Anonymous. IEEE Computer Society: Be at the center of it all house advertisement. IEEE Security & Privacy, 13(6): c2, November/December 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2015:ICSa

[Ano15w] Anonymous. IEEE Computer Society house advertisement. IEEE Security & Privacy, 13 (4):c3, July/August 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/04/
msp20150400c3.pdf.

Anonymous:2015:ISPa

[Ano15x] Anonymous. IEEE Security & Privacy house advertisement. IEEE Security & Privacy, 13 (2):c3, March/April 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/02/
msp20150200c3.pdf.

Anonymous:2015:ISPb

[Ano15y] Anonymous. IEEE Security & Privacy house advertisement. IEEE Security & Privacy, 13 (3):c3, May/June 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/03/
msp20150300c3.pdf.

Anonymous:2015:ISPc

[Ano15z] Anonymous. IEEE Security & Privacy house advertisement. IEEE Security & Privacy, 13 (4):c4, July/August 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/04/
msp20150400c4.pdf.

Anonymous:2015:ISPd

[Ano15-27] Anonymous. IEEE Security & Privacy house advertisement. IEEE Security & Privacy, 13(5):c4, September/October 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/csdl/mags/sp/2015/05/
msp20150500c4.pdf.

Anonymous:2015:KYC

[Ano15-28] Anonymous. Keep your career moving forward house advertisement. IEEE Security & Privacy, 13(5):c3, September/October 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/csdl/mags/sp/2015/05/
msp20150500c3.pdf.

Anonymous:2015:MYC

[Ano15-29] Anonymous. Mark your calendars [house advertisement]. IEEE Security & Privacy, 13 (1):c2, January/February 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/01/
msp20150100c2.pdf.

Anonymous:2015:Ma

[Ano15-30] Anonymous. Masthead. IEEE Security & Privacy, 13(1):6, January/February 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/01/
msp2015010006.pdf.

Anonymous:2015:Mb

[Ano15-31] Anonymous. Masthead. IEEE Security & Privacy, 13(2): 6, March/April 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/02/
msp2015020006.pdf.

Anonymous:2015:Mc

[Ano15-32] Anonymous. Masthead. IEEE Security & Privacy, 13(3): 8, May/June 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/03/
msp2015030008.pdf.

Anonymous:2015:Md

[Ano15-33] Anonymous. Masthead. IEEE Security & Privacy, 13(4): 6, July/August 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/04/
msp2015040006.pdf.

Anonymous:2015:Me

[Ano15-34] Anonymous. Masthead. IEEE Security & Privacy, 13(5):6, September/October 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://csdl.computer.
                                        org/csdl/mags/sp/2015/05/
                                        msp2015050006.pdf.

Anonymous:2015:Mf

[Ano15-35] Anonymous. Masthead. IEEE Security & Privacy, 13(6): 11, November/December 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2015:RSCa

[Ano15-36] Anonymous. Rock stars of cyber security [Advertisement]. IEEE Security & Privacy, 13 (2):c2, March/April 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/sp/2015/02/
                                        msp20150200c2.pdf.

Anonymous:2015:RSCb

[Ano15-37] Anonymous. Rock stars of cybersecurity house advertisement. IEEE Security & Privacy, 13(3):c2, May/June 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/sp/2015/03/
                                        msp20150300c2.pdf.

Anonymous:2015:RSCc

[Ano15-38] Anonymous. Rock stars of cybersecurity house advertisement. IEEE Security & Privacy, 13(5):c2, September/October 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/csdl/mags/sp/2015/05/
msp20150500c2.pdf.

Anonymous:2015:SES

[Ano15-39] Anonymous. Software experts summit house advertisement. IEEE Security & Privacy, 13(5):85, September/October 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/csdl/mags/sp/2015/05/
msp2015050085.pdf.

Anonymous:2015:S

[Ano15-40] Anonymous. Sponsor. IEEE Security & Privacy, 13(6): ??, November/December 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2015:SRS

[Ano15-41] Anonymous. Startup rock stars [house advertisement]. IEEE Security & Privacy, 13(1):c3, January/February 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/01/
msp20150100c3.pdf.

Anonymous:2015:TC

[Ano15-42] Anonymous. Table of contents.

IEEE Security & Privacy, 13 (6):1–2, November/December 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2015:WWL

[Ano15-43] Anonymous. Watch the world’s leading experts take multi-core strategies to new heights house advertisement. IEEE Security & Privacy, 13(6):71, November/December 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2016:ABS

[Ano16a] Anonymous. 4th annual best scientific cybersecurity paper competition. IEEE Security & Privacy, 14(1):c4, January/ February 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2016:CN

[Ano16b] Anonymous. Call for nominees. IEEE Security & Privacy, 14 (2):11, March/April 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2016:CNE

[Ano16c] Anonymous. Call for nominees education awards nominations house advertisement. IEEE Security & Privacy, 14 (4):86, July/August 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2016/04/
                                        msp2016040086.pdf.

Anonymous:2016:CPSc

[Ano16d] Anonymous. Call for papers special issue on electronic voting. IEEE Security & Privacy, 14(4):45, July/August 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/04/
msp2016040045.pdf.

Anonymous:2016:CPSb

[Ano16e] Anonymous. Call for papers special issue on electronic voting house advertisement. IEEE Security & Privacy, 14(3):c3, May/June 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2016:CPSa

[Ano16f] Anonymous. Call for papers special issue on genome privacy and security house advertisement. IEEE Security & Privacy, 14(2):65, March/April 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2016:CPSd

[Ano16g] Anonymous. Call for papers special issue on postquantum cryptography. IEEE Security & Privacy, 14(4): 63, July/August 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/04/
msp2016040063.pdf.

Anonymous:2016:CPSe

[Ano16h] Anonymous. Call for papers special issue on postquantum cryptography. IEEE Security & Privacy, 14(5):57, September/October 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/05/
msp2016050057.pdf.

Anonymous:2016:CYI

[Ano16i] Anonymous. Can you invent a better world through technologies? IEEE Security & Privacy, 14(6):c4, November/December 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/06/
msp20160600c4.pdf.

Anonymous:2016:CEA

[Ano16j] Anonymous. Computer entrepreneur award house advertisement. IEEE Security & Privacy, 14(4):85, July/August 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/04/
msp2016040085.pdf.

Anonymous:2016:FCa

[Ano16k]  Anonymous.  Front cover.

IEEE Security & Privacy, 14 (1):c1, January/February 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2016:FCb

[Ano16l]  Anonymous.  Front cover.

IEEE Security & Privacy, 14 (2):c1, March/April 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2016:FCc

[Ano16m]  Anonymous.  Front cover.

IEEE Security & Privacy, 14 (3):c1, May/June 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2016:FCd

[Ano16n]  Anonymous.  Front cover.

IEEE Security & Privacy, 14 (4):c1, July/August 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/04/
msp20160400c1.pdf.

Anonymous:2016:FCe

[Ano16o]  Anonymous.  Front cover.

IEEE Security & Privacy, 14(5):c1, September/October 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL https:

//www.computer.org/csdl/
mags/sp/2016/05/msp20160500c1.
pdf.

Anonymous:2016:FCf

[Ano16p]  Anonymous.  Front cover.

IEEE Security & Privacy, 14 (6):c1, November/December 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL https:

//www.computer.org/csdl/
mags/sp/2016/06/msp20160600c1.
pdf.

Anonymous:2016:GF

[Ano16q] Anonymous. Got flaws? IEEE Security & Privacy, 14(5):87, September/October 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/05/
msp2016050087.pdf.

Anonymous:2016:GFHa

[Ano16r] Anonymous. Got flaws? House advertisement. IEEE Security & Privacy, 14(2):c2, March/ April 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2016:GFHb

[Ano16s] Anonymous. Got flaws? House advertisement. IEEE Security & Privacy, 14(3):c2, May/June 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2016:GFHc

[Ano16t]  Anonymous.  Got flaws?

House advertisement. IEEE Security & Privacy, 14(4): c3, July/August 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/04/
msp20160400c3.pdf.

Anonymous:2016:ICSb

[Ano16u] Anonymous. IEEE Computer Society 2016 call for major award nominations house advertisement. IEEE Security & Privacy, 14(2):47, March/April 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2016:ICSf

[Ano16v]  Anonymous.  IEEE Com-

puter Society 2016 call for major award nominations house advertisement. IEEE Security & Privacy, 14(4):81, July/August 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/04/
msp2016040081.pdf.

Anonymous:2016:ICSa

[Ano16w] Anonymous. IEEE Computer Society: Be at the center of it all house advertisement. IEEE Security & Privacy, 14(1):c3, January/February 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2016:ICSd

[Ano16x] Anonymous. IEEE Computer Society: Be at the center of it all house advertisement. IEEE Security & Privacy, 14(2):c3, March/April 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2016:ICSe

[Ano16y] Anonymous. IEEE Computer Society: Be at the center of it all house advertisement. IEEE Security & Privacy, 14 (4):33, July/August 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/04/
msp2016040033.pdf.

Anonymous:2016:ICSc

[Ano16z]  Anonymous.  IEEE Com-

puter Society Richard E. Merwin Student Leadership Scholarship house advertisement. IEEE Security & Privacy, 14 (2):72, March/April 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2016:ICD

[Ano16-27] Anonymous. IEEE cybersecurity development SecDev 2016. IEEE Security & Privacy, 14(5):c2, September/October 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/05/
msp20160500c2.pdf.

Anonymous:2016:ISPe

[Ano16-28] Anonymous. IEEE Security & Privacy. IEEE Security & Privacy, 14(5):c4, September/October 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.

org/csdl/mags/sp/2016/05/
msp20160500c4.pdf.

Anonymous:2016:ISPb

[Ano16-29] Anonymous. IEEE Security & Privacy house advertisement. IEEE Security & Privacy, 14 (2):c4, March/April 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2016:ISPc

[Ano16-30] Anonymous. IEEE Security & Privacy house advertisement. IEEE Security & Privacy, 14 (3):c4, May/June 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2016:ISPd

[Ano16-31] Anonymous. IEEE Security & Privacy house advertisement. IEEE Security & Privacy, 14 (4):c4, July/August 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/04/
msp20160400c4.pdf.

Anonymous:2016:ISPa

[Ano16-32] Anonymous. IEEE Security & Privacy qmags subscription house advertisement. IEEE Security & Privacy, 14(1):c2, January/February 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2016:Ma

[Ano16-33] Anonymous. Masthead. IEEE Security & Privacy, 14(1):6, January/February 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2016:Mb

[Ano16-34] Anonymous. Masthead. IEEE Security & Privacy, 14(2):7, March/April 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2016:Mc

[Ano16-35] Anonymous. Masthead. IEEE Security & Privacy, 14(3):6, May/June 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2016:Md

[Ano16-36] Anonymous. Masthead. IEEE Security & Privacy, 14(4): 6, July/August 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/04/
msp2016040006.pdf.

Anonymous:2016:Me

[Ano16-37] Anonymous. Masthead. IEEE Security & Privacy, 14(5):6, September/October 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/05/
msp2016050006.pdf.

Anonymous:2016:Mf

[Ano16-38] Anonymous. Masthead. IEEE Security & Privacy, 14(6): 6, November/December 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/06/
msp2016060006.pdf.

Anonymous:2016:NMOa

[Ano16-39] Anonymous. New membership options for a better fit. IEEE Security & Privacy, 14(5):c3, September/October 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/05/
msp20160500c3.pdf.

Anonymous:2016:NMOb

[Ano16-40] Anonymous. New membership options for a better fit. IEEE Security & Privacy, 14(6): c2, November/December 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/06/
msp20160600c2.pdf.

Anonymous:2016:RSC

[Ano16-41] Anonymous. Rock stars of cybersecurity house advertisement. IEEE Security & Privacy, 14(4):c2, July/August 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/04/
msp20160400c2.pdf.

Anonymous:2016:RSP

[Ano16-42] Anonymous. Rock stars of pervasive, predictive analytics. IEEE Security & Privacy, 14(5):47, September/October 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2016/05/
                                        msp2016050047.pdf.

Anonymous:2016:RSR

[Ano16-43] Anonymous. Rock stars of risk-based security house advertisement. IEEE Security & Privacy, 14(2):29, March/April 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2016:SIR

[Ano16-44] Anonymous. Special issue on real-world cryptography call for papers house advertisement. IEEE Security & Privacy, 14(1):62, January/ February 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2016:TCa

[Ano16-45] Anonymous. Table of contents. IEEE Security & Privacy, 14(1):1–2, January/February 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2016:TCb

[Ano16-46] Anonymous. Table of contents. IEEE Security & Privacy, 14(2):1–2, March/April 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2016:TCc

[Ano16-47] Anonymous. Table of contents.

IEEE Security & Privacy, 14 (3):1–2, May/June 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2016:TCd

[Ano16-48] Anonymous. Table of contents.

IEEE Security & Privacy, 14 (4):1–2, July/August 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/04/
msp2016040001.pdf.

Anonymous:2016:TCe

[Ano16-49] Anonymous. Table of contents. IEEE Security & Privacy, 14(5):1–2, September/October 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/05/
msp2016050001.pdf.

Anonymous:2016:TCf

[Ano16-50] Anonymous. Table of contents. IEEE Security & Privacy, 14(6):1–2, November/December 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/06/
msp2016060001.pdf.

Anonymous:2016:T

[Ano16-51] Anonymous. TechIgnite. IEEE Security & Privacy, 14(6): c3, November/December 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/06/
msp20160600c3.pdf.

Anonymous:2017:BRR

[Ano17a] Anonymous. 2017 B. Ramakrishna Rau Award call for nominations house advertisement. IEEE Security & Privacy, 15 (2):61, March/April 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/02/
msp2017020061.pdf.

Anonymous:2017:AYC

[Ano17b] Anonymous. Achieve your career goals with the fit that’s right for you. IEEE Security & Privacy, 15(3): 11, May/June 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/03/
msp2017030011.pdf.

Anonymous:2017:AIC

[Ano17c] Anonymous. AI industrial complex: The challenge of AI ethics. IEEE Security & Privacy, 15(5):3–5, September/October 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/05/
msp2017050003.html.

Anonymous:2017:AZH

[Ano17d] Anonymous. Apple ZeroConf holes: How hackers can steal iPhone photos. IEEE Security & Privacy, 15(2):42– 49, March/April 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/02/
msp2017020042-abs.html.

Anonymous:2017:AIA

[Ano17e] Anonymous. Authorship integrity and attacks. IEEE Security & Privacy, 15(4):3– 5, July/August 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/04/
msp2017040003.html.

Anonymous:2017:BSN

[Ano17f] Anonymous. Behind the scenes at NSA advertisement. IEEE Security & Privacy, 15(2): c2, March/April 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/02/
msp20170200c2.pdf.

Anonymous:2017:ISPe

[Ano17g]  Anonymous.  IEEE Secu-

rity & Privacy. IEEE Security & Privacy, 15(5):c4, September/October 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.

                                        org/csdl/mags/sp/2017/05/
                                        msp20170500c4.pdf.

Anonymous:2017:BFA

[Ano17h] Anonymous. Botnet fingerprinting: Anomaly detection in SMTP conversations. IEEE Security & Privacy, 15(6):25– 32, November/December 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/06/
                                        msp2017060025.html.

Anonymous:2017:CNEb

[Ano17i] Anonymous. Call for nominees: Education awards nominations. IEEE Security & Privacy, 15(3):51, May/June 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/03/
                                        msp2017030051.pdf.

Anonymous:2017:CNEc

[Ano17j] Anonymous. Call for nominees education awards nominations. IEEE Security & Privacy, 15(5):13, September/October 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/05/
                                        msp2017050013.pdf.

Anonymous:2017:CNEa

[Ano17k] Anonymous. Call for nominees education awards nominations house advertisement. IEEE Security & Privacy, 15(2): 101, March/April 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/02/
msp2017020101.pdf.

Anonymous:2017:CPSc

[Ano17l] Anonymous. Call for papers: Special issue on AI ethics: The privacy challenge. IEEE Security & Privacy, 15 (3):39, May/June 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/03/
msp2017030039.pdf.

Anonymous:2017:CPSe

[Ano17m] Anonymous. Call for papers special issue on AI ethics: The privacy challenge. IEEE Security & Privacy, 15(4): c2, July/August 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/04/
msp20170400c2.pdf.

Anonymous:2017:CPSb

[Ano17n] Anonymous. Call for papers: Special issue on digital forensics. IEEE Security & Privacy, 15(1):41, January/February 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/01/
msp2017010041.pdf.

Anonymous:2017:CPSa

[Ano17o] Anonymous. Call for papers: Special issue on hacking without humans. IEEE Security & Privacy, 15(1):21, January/February 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/01/
                                        msp2017010021.pdf.

Anonymous:2017:CPSd

[Ano17p] Anonymous. Call for papers: Special issue on hacking without humans. IEEE Security & Privacy, 15(3): c2, May/June 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/03/
                                        msp20170300c2.pdf.

Anonymous:2017:CPSf

[Ano17q] Anonymous. Call for papers special issue on security and privacy research in Brazil. IEEE Security & Privacy, 15(5):c2, September/October 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/05/
                                        msp20170500c2.pdf.

Anonymous:2017:CPSg

[Ano17r] Anonymous. Call for papers special issue on security and privacy research in Brazil. IEEE Security & Privacy, 15(6):c2, November/December 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/06/
msp20170600c2.pdf.

Anonymous:2017:DUD

[Ano17s]  Anonymous.  The danger

of USB drives. IEEE Security & Privacy, 15(2):62– 69, March/April 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/02/
msp2017020062-abs.html.

Anonymous:2017:DDD

[Ano17t] Anonymous. Dawn of the dead domain: Measuring the exploitation of residual trust in domains. IEEE Security & Privacy, 15(2):70– 77, March/April 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/02/
msp2017020070-abs.html.

Anonymous:2017:DEU

[Ano17u] Anonymous. Decision and experienced utility: Computational applications in privacy decision making. IEEE Security & Privacy, 15(6):68– 72, November/December 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.

                                        org/csdl/mags/sp/2017/06/
                                        msp2017060068-abs.html.

Anonymous:2017:DIS

[Ano17v]  Anonymous.  Does indus-

try self-regulation of consumer data privacy work? IEEE Security & Privacy, 15(2):92– 95, March/April 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/02/
                                        msp2017020092-abs.html.

Anonymous:2017:DOC

[Ano17w] Anonymous. Does the online card payment landscape unwittingly facilitate fraud? IEEE Security & Privacy, 15(2):78– 86, March/April 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/02/
                                        msp2017020078-abs.html.

Anonymous:2017:EIS

[Ano17x] Anonymous. Ethics in information security. IEEE Security & Privacy, 15(3):3– 4, May/June 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/03/
                                        msp2017030003.html.

Anonymous:2017:EES

[Ano17y] Anonymous. An exploration of the effects of sensory stimuli on the completion of security tasks. IEEE Security & Privacy, 15(6):52–60, November/December 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/06/
msp2017060052-abs.html.

Anonymous:2017:FSC

[Ano17z] Anonymous. Faster secure cloud computations with a trusted proxy. IEEE Security & Privacy, 15(6):61–67, November/December 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/06/
msp2017060061-abs.html.

Anonymous:2017:FYJa

[Ano17-27] Anonymous. Focus on your job search. IEEE Security & Privacy, 15(1):86, January/February 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/01/
msp2017010086.pdf.

Anonymous:2017:FYJc

[Ano17-28] Anonymous. Focus on your job search. IEEE Security & Privacy, 15(3):c3, May/June 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/03/
msp20170300c3.pdf.

Anonymous:2017:FYJb

[Ano17-29] Anonymous. Focus on your job search house advertisement. IEEE Security & Privacy, 15 (2):31, March/April 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/02/
                                        msp2017020031.pdf.

Anonymous:2017:FCa

[Ano17-30] Anonymous. Front cover. IEEE Security & Privacy, 15 (1):c1, January/February 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/01/
                                        msp20170100c1.pdf.

Anonymous:2017:FCb

[Ano17-31] Anonymous. Front cover. IEEE Security & Privacy, 15 (2):c1, March/April 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/02/
                                        msp20170200c1.pdf.

Anonymous:2017:FCc

[Ano17-32] Anonymous. Front cover. IEEE Security & Privacy, 15 (3):c1, May/June 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/03/
                                        msp20170300c1.pdf.

Anonymous:2017:FCd

[Ano17-33] Anonymous. Front cover.

IEEE Security & Privacy, 15 (4):c1, July/August 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/04/
msp20170400c1.pdf.

Anonymous:2017:FCe

[Ano17-34] Anonymous. Front cover.

IEEE Security & Privacy, 15(5):c1, September/October 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL https:

//www.computer.org/csdl/
mags/sp/2017/05/msp20170500c1.
pdf.

Anonymous:2017:FCf

[Ano17-35] Anonymous. Front cover.

IEEE Security & Privacy, 15 (6):c1, November/December 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL https:

//www.computer.org/csdl/
mags/sp/2017/06/msp20170600c1.
pdf.

Anonymous:2017:FDF

[Ano17-36] Anonymous. The future of digital forensics: Challenges and the road ahead. IEEE Security & Privacy, 15(6):12– 17, November/December 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/06/
msp2017060012-abs.html.

Anonymous:2017:GF

[Ano17-37] Anonymous. Got flaws? IEEE Security & Privacy, 15(6): c4, November/December 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/06/
                                        msp20170600c4.pdf.

Anonymous:2017:HBC

[Ano17-38] Anonymous. How businesses can speed up international cybercrime investigation. IEEE Security & Privacy, 15(2):102– 106, March/April 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/02/
                                        msp2017020102-abs.html.

Anonymous:2017:HIR

[Ano17-39] Anonymous. How Internet resources might be helping you develop faster but less securely. IEEE Security & Privacy, 15 (2):50–60, March/April 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/02/
                                        msp2017020050-abs.html.

Anonymous:2017:ICSc

[Ano17-40] Anonymous. IEEE Computer Society. IEEE Security & Privacy, 15(3):77, May/June 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.

org/csdl/mags/sp/2017/03/
msp2017030077.pdf.

Anonymous:2017:ICSb

[Ano17-41] Anonymous. IEEE Computer Society 2017 call for major award nominations. IEEE Security & Privacy, 15(3): 59, May/June 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/03/
msp2017030059.pdf.

Anonymous:2017:ICSa

[Ano17-42] Anonymous. IEEE Computer Society 2017 call for major award nominations house advertisement. IEEE Security & Privacy, 15(2):107, March/April 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/02/
msp2017020107.pdf.

Anonymous:2017:ICSd

[Ano17-43] Anonymous. IEEE Computer Society Harlan D. Mills Award. IEEE Security & Privacy, 15 (3):89, May/June 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/03/
msp2017030089.pdf.

Anonymous:2017:ISPa

[Ano17-44] Anonymous. IEEE Security & Privacy. IEEE Security & Privacy, 15(1):c4, January/February 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/01/
                                        msp20170100c4.pdf.

Anonymous:2017:ISPb

[Ano17-45] Anonymous. IEEE Security & Privacy. IEEE Security & Privacy, 15(2):c4, March/April 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/02/
                                        msp20170200c4.pdf.

Anonymous:2017:ISPc

[Ano17-46] Anonymous. IEEE Security & Privacy. IEEE Security & Privacy, 15(3):c4, May/June 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/03/
                                        msp20170300c4.pdf.

Anonymous:2017:ISPd

[Ano17-47] Anonymous. IEEE Security & Privacy. IEEE Security & Privacy, 15(4):c4, July/August 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/04/
                                        msp20170400c4.pdf.

Anonymous:2017:LMI

[Ano17-48] Anonymous. The last mile for IoT privacy. IEEE Security & Privacy, 15(6):73– 76, November/December 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/06/
msp2017060073-abs.html.

Anonymous:2017:LBT

[Ano17-49] Anonymous. Looking for the BEST tech job for you? [advertisement]. IEEE Security & Privacy, 15(4):11, July/August 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/04/
msp2017040011.pdf.

Anonymous:2017:Ma

[Ano17-50] Anonymous. Masthead. IEEE Security & Privacy, 15(1):5, January/February 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/01/
msp2017010005.pdf.

Anonymous:2017:Mb

[Ano17-51] Anonymous. Masthead. IEEE Security & Privacy, 15(2): 7, March/April 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/02/
msp2017020007.pdf.

Anonymous:2017:Mc

[Ano17-52] Anonymous. Masthead. IEEE Security & Privacy, 15(3):

5, May/June 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/03/
msp2017030005.pdf.

Anonymous:2017:Me

[Ano17-53] Anonymous. Masthead. IEEE Security & Privacy, 15(4): 6, July/August 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/04/
msp2017040006.pdf.

Anonymous:2017:Mf

[Ano17-54] Anonymous. Masthead. IEEE Security & Privacy, 15(5):6, September/October 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/05/
msp2017050006.pdf.

Anonymous:2017:Mg

[Ano17-55] Anonymous. Masthead. IEEE Security & Privacy, 15(6): 6, November/December 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/06/
msp2017060006.pdf.

Anonymous:2017:MFA

[Ano17-56] Anonymous. Mobile forensics: Advances, challenges, and research opportunities. IEEE Security & Privacy, 15(6):42– 51, November/December 2017.

CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/06/
msp2017060042-abs.html.

Anonymous:2017:Md

[Ano17-57] Anonymous. myCS. IEEE Security & Privacy, 15(3): 88, May/June 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/03/
msp2017030088.pdf.

Anonymous:2017:MHA

[Ano17-58] Anonymous. myCS house advertisement. IEEE Security & Privacy, 15(2):87, March/April 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/02/
msp2017020087.pdf.

Anonymous:2017:NFM

[Ano17-59] Anonymous. NAND flash memory forensic analysis and the growing challenge of bit errors. IEEE Security & Privacy, 15(6):82–87, November/December 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/06/
msp2017060082-abs.html.

Anonymous:2017:NMOa

[Ano17-60] Anonymous. New membership options for a better fit. IEEE Security & Privacy, 15(1):c3, January/February 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/01/
msp20170100c3.pdf.

Anonymous:2017:NMOc

[Ano17-61] Anonymous. New membership options for a better fit. IEEE Security & Privacy, 15 (3):10, May/June 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/03/
msp2017030010.pdf.

Anonymous:2017:NMOe

[Ano17-62] Anonymous. New membership options for a better fit. IEEE Security & Privacy, 15(5):c3, September/October 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/05/
msp20170500c3.pdf.

Anonymous:2017:NMOd

[Ano17-63] Anonymous. New membership options for a better fit [advertisement]. IEEE Security & Privacy, 15(4): c3, July/August 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/04/
msp20170400c3.pdf.

Anonymous:2017:NMOb

[Ano17-64] Anonymous. New membership options for a better fit. house advertisement. IEEE Security & Privacy, 15(2): c3, March/April 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/02/
msp20170200c3.pdf.

Anonymous:2017:SHA

[Ano17-65] Anonymous. Not so secure house advertisement. IEEE Security & Privacy, 15(2): 2, March/April 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/02/
msp2017020002.pdf.

Anonymous:2017:PUT

[Ano17-66] Anonymous. The perils of user tracking using zero-permission mobile apps. IEEE Security & Privacy, 15(2):32– 41, March/April 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/02/
msp2017020032-abs.html.

Anonymous:2017:PPS

[Ano17-67] Anonymous. Prepose: Privacy, security, and reliability for gesture-based programming. IEEE Security & Privacy, 15 (2):14–23, March/April 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/02/
                                        msp2017020014-abs.html.

Anonymous:2017:PPF

[Ano17-68] Anonymous. PROFORMA: Proactive forensics with message analytics. IEEE Security & Privacy, 15(6):33– 41, November/December 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/06/
                                        msp2017060033-abs.html.

Anonymous:2017:PLC

[Ano17-69] Anonymous. Programmable logic controller forensics. IEEE Security & Privacy, 15(6):18– 24, November/December 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/06/
                                        msp2017060018-abs.html.

Anonymous:2017:RT

[Ano17-70] Anonymous. Reviewer thanks. IEEE Security & Privacy, 15(1):6–7, January/February 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL https:

                                        //www.computer.org/csdl/
                                        mags/sp/2017/01/msp2017010006.
                                        html.

Anonymous:2017:SBC

[Ano17-71] Anonymous. Securing binary code. IEEE Security & Privacy, 15(6):77–81, November/December 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/06/
msp2017060077-abs.html.

Anonymous:2017:SAC

[Ano17-72] Anonymous. Security advice that can be followed. IEEE Security & Privacy, 15(6):3– 5, November/December 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/06/
msp2017060003.html.

Anonymous:2017:SCO

[Ano17-73] Anonymous. Security challenges and opportunities of software-defined networking. IEEE Security & Privacy, 15 (2):96–100, March/April 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/02/
msp2017020096-abs.html.

Anonymous:2017:SIP

[Ano17-74] Anonymous. Security implications of permission models in smart-home application frameworks. IEEE Security & Privacy, 15(2):24– 30, March/April 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/02/
msp2017020024-abs.html.

Anonymous:2017:SBS

[Ano17-75] Anonymous. #SP18 is back in San Francisco! IEEE Security & Privacy, 15(6):c3, November/December 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/06/
                                        msp20170600c3.pdf.

Anonymous:2017:TCa

[Ano17-76] Anonymous. Table of contents. IEEE Security & Privacy, 15(1):1–2, January/ February 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/01/
                                        msp2017010001.pdf.

Anonymous:2017:TCb

[Ano17-77] Anonymous. Table of contents. IEEE Security & Privacy, 15 (2):3–4, March/April 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/02/
                                        msp2017020003.pdf.

Anonymous:2017:TCc

[Ano17-78] Anonymous. Table of contents. IEEE Security & Privacy, 15 (3):1–2, May/June 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/03/
                                        msp2017030001.pdf.

Anonymous:2017:TCd

[Ano17-79] Anonymous. Table of contents.

IEEE Security & Privacy, 15 (4):1–2, July/August 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/04/
msp2017040001.pdf.

Anonymous:2017:TCe

[Ano17-80] Anonymous. Table of contents. IEEE Security & Privacy, 15(5):1–2, September/October 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/05/
msp2017050001.pdf.

Anonymous:2017:TCf

[Ano17-81] Anonymous. Table of contents. IEEE Security & Privacy, 15(6):1–2, November/December 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/06/
msp2017060001.pdf.

Anonymous:2017:TCL

[Ano17-82] Anonymous. Take the CS Library wherever you go! IEEE Security & Privacy, 15(5):75, September/October 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/05/
msp2017050075.pdf.

Anonymous:2017:T

[Ano17-83] Anonymous. TechIgnite. IEEE Security & Privacy, 15(1):c2, January/February 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/01/
                                        msp20170100c2.pdf.

Anonymous:2017:UMU

[Ano17-84] Anonymous. Unwavering mission unwavering commitment advertisement. IEEE Security & Privacy, 15(2): 1, March/April 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/02/
                                        msp2017020001.pdf.

Anonymous:2018:ISPe

[Ano18a] Anonymous. IEEE Security & Privacy. IEEE Security & Privacy, 16(5):c4, September/ October 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2018:ISPd

[Ano18b] Anonymous. IEEE Security & Privacy special issue call for papers. IEEE Security & Privacy, 16(5):c2, September/ October 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2018:ITB

[Ano18c] Anonymous. IEEE Transactions on Big Data. IEEE Security & Privacy, 16(5): 81, September/October 2018.

ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2018:ITSa

[Ano18d] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Security & Privacy, 16(1):63, January/ February 2018. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/01/
msp2018010063.pdf.

Anonymous:2018:C

[Ano18e]  Anonymous.  COMPSAC

2018. IEEE Security & Privacy, 16(1):43, January/ February 2018. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/01/
msp2018010043.pdf.

Anonymous:2018:CSE

[Ano18f] Anonymous. Computing in science & engineering. IEEE Security & Privacy, 16(3): 95, May/June 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/03/
msp2018030095.pdf.

Anonymous:2018:CI

[Ano18g] Anonymous. Connect on interface. IEEE Security & Privacy, 16(3):c2, May/June 2018. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/03/
                                        msp20180300c2.pdf.

Anonymous:2018:CPY

[Ano18h] Anonymous. Cross-pollinate your ideas. IEEE Security & Privacy, 16(1):23, January/February 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/01/
                                        msp2018010023.pdf.

Anonymous:2018:EDS

[Ano18i] Anonymous. Encouraging diversity in security and privacy research. IEEE Security & Privacy, 16(4):3– 5, July/August 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/04/
                                        msp2018040003.html.

Anonymous:2018:FCa

[Ano18j]  Anonymous.  Front cover.

IEEE Security & Privacy, 16 (1):c1, January/February 2018. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/01/
                                        msp20180100c1.pdf.

Anonymous:2018:FCb

[Ano18k]  Anonymous.  Front cover.

IEEE Security & Privacy, 16 (2):c1, March/April 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/02/
msp20180200c1.pdf.

Anonymous:2018:FCc

[Ano18l]  Anonymous.  Front cover.

IEEE Security & Privacy, 16 (3):c1, May/June 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/03/
msp20180300c1.pdf.

Anonymous:2018:FCd

[Ano18m]  Anonymous.  Front cover.

IEEE Security & Privacy, 16 (4):c1, July/August 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/04/
msp20180400c1.pdf.

Anonymous:2018:FCe

[Ano18n]  Anonymous.  Front cover.

IEEE Security & Privacy, 16(5):c1, September/October 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2018:FCf

[Ano18o]  Anonymous.  Front cover.

IEEE Security & Privacy, 16 (6):C1, November/December 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2018:HMT

[Ano18p] Anonymous. How many is too many candidates? IEEE Security & Privacy, 16(3):

3–5, May/June 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/03/
msp2018030003.html.

Anonymous:2018:ICSb

[Ano18q]  Anonymous.  IEEE Com-

puter Society. IEEE Security & Privacy, 16(1):c3, January/February 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/01/
msp20180100c3.pdf.

Anonymous:2018:ICSc

[Ano18r]  Anonymous.  IEEE Com-

puter Society. IEEE Security & Privacy, 16(2):9, March/April 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/02/
msp2018020009.pdf.

Anonymous:2018:ICSd

[Ano18s]  Anonymous.  IEEE Com-

puter Society. IEEE Security & Privacy, 16(3):55, May/June 2018. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/03/
msp2018030055.pdf.

Anonymous:2018:ICSf

[Ano18t] Anonymous. IEEE Computer Society. IEEE Security & Privacy, 16(5):37, September/ October 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2018:ICSa

[Ano18u]  Anonymous.  IEEE Com-

puter Society: Be at the center of it all. IEEE Security & Privacy, 16(1):101, January/February 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/01/
msp2018010101.pdf.

Anonymous:2018:ICSe

[Ano18v] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 16(4): 95, July/August 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/04/
msp2018040095.pdf.

Anonymous:2018:ICSg

[Ano18w] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 16(6): 77, November/December 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2018:ILC

[Ano18x] Anonymous. IEEE letters of the Computer Society. IEEE Security & Privacy, 16(5): c3, September/October 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2018:ISPa

[Ano18y]  Anonymous.  IEEE Secu-

rity & Privacy. IEEE Security & Privacy, 16(2):c4, March/April 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/02/
                                        msp20180200c4.pdf.

Anonymous:2018:ISPb

[Ano18z] Anonymous. IEEE Security & Privacy. IEEE Security & Privacy, 16(3):c4, May/June 2018. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/03/
                                        msp20180300c4.pdf.

Anonymous:2018:ISPc

[Ano18-27] Anonymous. IEEE Security & Privacy. IEEE Security & Privacy, 16(4):c2, July/August 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/04/
                                        msp20180400c2.pdf.

Anonymous:2018:IBA

[Ano18-28] Anonymous. Impact a broader audience. IEEE Security & Privacy, 16(5):71, September/ October 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2018:ITSb

[Ano18-29] Anonymous. Internet of Things security: Is anything new?

IEEE Security & Privacy, 16 (5):3–5, September/October 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2018:INE

[Ano18-30] Anonymous. Introduction from the new EIC. IEEE Security & Privacy, 16(2):3– 4, March/April 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/02/
msp2018020003.html.

Anonymous:2018:Ma

[Ano18-31] Anonymous. Masthead. IEEE Security & Privacy, 16(1):5, January/February 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/01/
msp2018010005.pdf.

Anonymous:2018:Mc

[Ano18-32] Anonymous. Masthead. IEEE Security & Privacy, 16(2): 5, March/April 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/02/
msp2018020005.pdf.

Anonymous:2018:Md

[Ano18-33] Anonymous. Masthead. IEEE Security & Privacy, 16(3): 11, May/June 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/03/
                                        msp2018030011.pdf.

Anonymous:2018:Me

[Ano18-34] Anonymous. Masthead. IEEE Security & Privacy, 16(4): 6, July/August 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/04/
                                        msp2018040006.pdf.

Anonymous:2018:Mf

[Ano18-35] Anonymous. [Masthead]. IEEE Security & Privacy, 16(5): 6, September/October 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2018:Mb

[Ano18-36] Anonymous. myCS. IEEE Security & Privacy, 16(1):90, January/February 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/01/
                                        msp2018010090.pdf.

Anonymous:2018:OMU

[Ano18-37] Anonymous. One membership. Unlimited knowledge. IEEE Security & Privacy, 16 (3):c3, May/June 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/03/
                                        msp20180300c3.pdf.

Anonymous:2018:PC

[Ano18-38] Anonymous. Prepare to connect. IEEE Security & Privacy, 16(1):91, January/ February 2018. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/01/
msp2018010091.pdf.

Anonymous:2018:RT

[Ano18-39] Anonymous. Reviewer thanks.

IEEE Security & Privacy, 16(1):3–4, January/February 2018. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL https:

//www.computer.org/csdl/
mags/sp/2018/01/msp2018010003.
html.

Anonymous:2018:SCS

[Ano18-40] Anonymous. Seymour Cray, Sidney Fernbach & Ken Kennedy awards. IEEE Security & Privacy, 16(2):c2, March/April 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/02/
msp20180200c2.pdf.

Anonymous:2018:SGKa

[Ano18-41] Anonymous. Share the gift of knowledge: give your favorite student a membership to the IEEE Computer Society! IEEE Security & Privacy, 16 (2):c3, March/April 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/02/
                                        msp20180200c3.pdf.

Anonymous:2018:SGKb

[Ano18-42] Anonymous. Share the gift of knowledge: Give your favorite student a membership to the IEEE Computer Society! IEEE Security & Privacy, 16 (4):c3, July/August 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/04/
                                        msp20180400c3.pdf.

Anonymous:2018:SBS

[Ano18-43] Anonymous. SP18 is back in San Francisco! IEEE Security & Privacy, 16(1):c2, January/February 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/01/
                                        msp20180100c2.pdf.

Anonymous:2018:SL

[Ano18-44] Anonymous. Staff list. IEEE Security & Privacy, 16(6): 3, November/December 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2018:SC

[Ano18-45] Anonymous. Stay connected. IEEE Security & Privacy, 16 (4):c4, July/August 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.

org/csdl/mags/sp/2018/04/
msp20180400c4.pdf.

Anonymous:2018:TCa

[Ano18-46] Anonymous. Table of contents. IEEE Security & Privacy, 16(1):1–2, January/ February 2018. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/01/
msp2018010001.pdf.

Anonymous:2018:TCb

[Ano18-47] Anonymous. Table of contents.

IEEE Security & Privacy, 16 (2):1–2, March/April 2018. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/02/
msp2018020001.pdf.

Anonymous:2018:TCc

[Ano18-48] Anonymous. Table of contents.

IEEE Security & Privacy, 16 (3):1–2, May/June 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/03/
msp2018030001.pdf.

Anonymous:2018:TCd

[Ano18-49] Anonymous. Table of contents.

IEEE Security & Privacy, 16 (4):1–2, July/August 2018. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/04/
msp2018040001.pdf.

Anonymous:2018:TCe

[Ano18-50] Anonymous. Table of contents. IEEE Security & Privacy, 16(5):1–2, September/ October 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2018:TCf

[Ano18-51] Anonymous. Table of contents. IEEE Security & Privacy, 16 (6):1–2, November/December 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2018:TCL

[Ano18-52] Anonymous. Take the CS Library wherever you go! IEEE Security & Privacy, 16(1):c4, January/February 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/01/
                                        msp20180100c4.pdf.

Anonymous:2019:ICSj

[Ano19a] Anonymous. 2019 IEEE Computer Society election. IEEE Security & Privacy, 17(5): 9, September/October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ICSl

[Ano19b] Anonymous. 2019 IEEE Computer Society election. IEEE Security & Privacy, 17(5): 95, September/October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:SIS

[Ano19c] Anonymous. 40th Symposium on IEEE Security Privacy. IEEE Security & Privacy, 17(2):16, March/April 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:IESa

[Ano19d] Anonymous. 4th IEEE European Symposium on Security and Privacy. IEEE Security & Privacy, 17(1):C2, January/ February 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:IESb

[Ano19e] Anonymous. 4th IEEE European Symposium on Security Privacy. IEEE Security & Privacy, 17(2):C2, March/ April 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:BRR

[Ano19f] Anonymous. B. Ramkrishna Rau Award. IEEE Security & Privacy, 17(1):C3, January/ February 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:CEM

[Ano19g] Anonymous. Computing Edge magazine. IEEE Security & Privacy, 17(6):C2, November/ December 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ISPb

[Ano19h] Anonymous. IEEE & Security Privacy subscribe. IEEE Security & Privacy, 17(2):58, March/April 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:IAHa

[Ano19i] Anonymous. IEEE Annals of the History of Computing. IEEE Security & Privacy, 17(2):79, March/April 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:IAHb

[Ano19j] Anonymous. IEEE Annals of the History of Computing. IEEE Security & Privacy, 17(4):83, July/August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ICSb

[Ano19k] Anonymous. IEEE CG&A subscribe. IEEE Security & Privacy, 17(1):54, January/ February 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ICGb

[Ano19l]  Anonymous.  IEEE Com-

puter Graphics and Applications Magazine. IEEE Security & Privacy, 17(6):36, November/December 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ICGa

[Ano19m] Anonymous. IEEE Computer Graphics Applications magazine. IEEE Security & Privacy, 17(4):87, July/August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ICEa

[Ano19n] Anonymous. IEEE Computing Edge. IEEE Security & Privacy, 17(4):C2, July/August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ICEb

[Ano19o] Anonymous. IEEE Computing Edge. IEEE Security & Privacy, 17(5):C4, September/ October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:IIS

[Ano19p] Anonymous. IEEE Intelligent Systems magazine. IEEE Security & Privacy, 17(4):99, July/ August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:IIC

[Ano19q] Anonymous. IEEE Internet Computing magazine. IEEE Security & Privacy, 17(4):97, July/August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:IIPa

[Ano19r] Anonymous. IEEE IT Professionals call for articles. IEEE Security & Privacy, 17(4):93, July/August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:IIPb

[Ano19s] Anonymous. IEEE IT Professionals call for papers. IEEE Security & Privacy, 17(5): 21, September/October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:IMM

[Ano19t] Anonymous. IEEE Multimedia magazine. IEEE Security & Privacy, 17(4):99, July/August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:IPCb

[Ano19u] Anonymous. IEEE Pervasive Computing call for articles. IEEE Security & Privacy, 17(1):77, January/ February 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:IPCc

[Ano19v] Anonymous. IEEE Pervasive Computing call for articles. IEEE Security & Privacy, 17(4):17, July/August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:IPCd

[Ano19w] Anonymous. IEEE Pervasive Computing call for papers. IEEE Security & Privacy, 17(5):37, September/October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ISPc

[Ano19x]  Anonymous.  IEEE Secu-

rity & Privacy magazine subscribe. IEEE Security & Privacy, 17(4):52, July/August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ISPa

[Ano19y] Anonymous. IEEE Security & Privacy subscribe. IEEE Security & Privacy, 17(1):33, January/February 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ISPd

[Ano19z] Anonymous. IEEE Security & Privacy subscribe. IEEE Security & Privacy, 17(5): 46, September/October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ISM

[Ano19-27] Anonymous. IEEE Software Magazine call for articles. IEEE Security & Privacy, 17(5):83, September/October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ITBa

[Ano19-28] Anonymous. IEEE Transactions on Big Data. IEEE Security & Privacy, 17(1):8, January/February 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ITBb

[Ano19-29] Anonymous. IEEE Transactions on Big Data. IEEE Security & Privacy, 17(2):72, March/April 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ITBc

[Ano19-30] Anonymous. IEEE Transactions on Big Data. IEEE Security & Privacy, 17(4):70, July/ August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ITBd

[Ano19-31] Anonymous. IEEE Transactions on Big Data. IEEE Security & Privacy, 17(5): 57, September/October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ITSa

[Ano19-32] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Security & Privacy, 17(1):22, January/ February 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ITSb

[Ano19-33] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Security & Privacy, 17(3):95, May/June 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ITSc

[Ano19-34] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Security & Privacy, 17(4):77, July/August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ITSd

[Ano19-35] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Security & Privacy, 17(5):89, September/October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:IPCa

[Ano19-36] Anonymous. IT Professional call for articles. IEEE Security & Privacy, 17(1):71, January/ February 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:SPM

[Ano19-37] Anonymous. Security & Privacy magazine. IEEE Security & Privacy, 17(3):94, May/June 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:TBD

[Ano19-38] Anonymous. Transactions on Big Data. IEEE Security & Privacy, 17(6):6, November/ December 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:TSC

[Ano19-39] Anonymous. Transactions on Sustainable Computing. IEEE Security & Privacy, 17(6): 30, November/December 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:FCa

[Ano19-40] Anonymous. Front cover.

IEEE Security & Privacy, 17(1):C1, January/February 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:FCb

[Ano19-41] Anonymous. Front cover. IEEE Security & Privacy, 17(2):C1, March/April 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:FCc

[Ano19-42] Anonymous. Front cover. IEEE Security & Privacy, 17 (3):C1, May/June 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:FCd

[Ano19-43] Anonymous. [Front cover]. IEEE Security & Privacy, 17(4):C1, July/August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:FCe

[Ano19-44] Anonymous. Front cover. IEEE Security & Privacy, 17(5):C1, September/October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:FCf

[Ano19-45] Anonymous. [Front cover]. IEEE Security & Privacy, 17 (6):C1, November/December 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ICSa

[Ano19-46] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 17(1):9, January/February 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ICSd

[Ano19-47] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 17(2):C3, March/April 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ICSf

[Ano19-48] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 17(3):C3, May/June 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ICSh

[Ano19-49] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 17(4):C3, July/August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ICSm

[Ano19-50] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 17(5): C3, September/October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ICSn

[Ano19-51] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 17(6): C3, November/December 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ICSc

[Ano19-52] Anonymous. IEEE Computer Society Jobs Boards. IEEE Security & Privacy, 17(1):C4, January/February 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ICSe

[Ano19-53] Anonymous. IEEE Computer Society Jobs Boards. IEEE Security & Privacy, 17(2):C4, March/April 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ICSi

[Ano19-54] Anonymous. IEEE Computer Society Jobs Boards. IEEE Security & Privacy, 17(4):C4, July/August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ICSg

[Ano19-55] Anonymous. IEEE Computer Society social networking. IEEE Security & Privacy, 17(4):9, July/August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ICSk

[Ano19-56] Anonymous. IEEE Computer Society social networking. IEEE Security & Privacy, 17(5):94, September/October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ID

[Ano19-57] Anonymous. IEEE Data-Port. IEEE Security & Privacy, 17(5):73, September/ October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:IJBa

[Ano19-58] Anonymous. IEEE Jobs Board.

IEEE Security & Privacy, 17 (3):C4, May/June 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:IJBb

[Ano19-59] Anonymous. IEEE Jobs Board.

IEEE Security & Privacy, 17(5):C2, September/October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:IJBc

[Ano19-60] Anonymous. IEEE Jobs Board.

IEEE Security & Privacy, 17 (6):C4, November/December 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:ILC

[Ano19-61] Anonymous. IEEE letters of the Computer Society. IEEE Security & Privacy, 17(4):98, July/August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:IWC

[Ano19-62] Anonymous. IEEE World Congress on Services 2019. IEEE Security & Privacy, 17 (3):C2, May/June 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:LCS

[Ano19-63] Anonymous. Letters of the Computer Society. IEEE Security & Privacy, 17(3):89, May/ June 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:Ma

[Ano19-64] Anonymous. Masthead. IEEE Security & Privacy, 17(1):3, January/February 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:Mb

[Ano19-65] Anonymous. Masthead. IEEE Security & Privacy, 17(2):3, March/April 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:Mc

[Ano19-66] Anonymous. Masthead. IEEE Security & Privacy, 17(3):3, May/June 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:Md

[Ano19-67] Anonymous. Masthead. IEEE Security & Privacy, 17(4):3, July/August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:Me

[Ano19-68] Anonymous. Masthead. IEEE Security & Privacy, 17(5): 3, September/October 2019.

ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:Mf

[Ano19-69] Anonymous. Masthead. IEEE Security & Privacy, 17(6): 3, November/December 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:RSa

[Ano19-70] Anonymous. Reliability Society. IEEE Security & Privacy, 17(1):6, January/ February 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:RSb

[Ano19-71] Anonymous. Reliability Society. IEEE Security & Privacy, 17(2):7, March/April 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:RSc

[Ano19-72] Anonymous. Reliability Society. IEEE Security & Privacy, 17(3):6, May/June 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:RSd

[Ano19-73] Anonymous. Reliability Society. IEEE Security & Privacy, 17(4):6, July/August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:RSe

[Ano19-74] Anonymous. Reliability Society. IEEE Security & Privacy, 17(5):6, September/October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:RSf

[Ano19-75] Anonymous. Reliability Society. IEEE Security & Privacy, 17(6):5, November/December 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:SN

[Ano19-76] Anonymous. Social networking. IEEE Security & Privacy, 17(3):7, May/June 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:SSP

[Ano19-77] Anonymous. Subscribe to Security & Privacy magazine. IEEE Security & Privacy, 17 (3):88, May/June 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:TCa

[Ano19-78] Anonymous. Table of contents. IEEE Security & Privacy, 17(1):1–2, January/February 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:TCb

[Ano19-79] Anonymous. Table of contents. IEEE Security & Privacy, 17(2):1–2, March/April 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:TCc

[Ano19-80] Anonymous. Table of contents. IEEE Security & Privacy, 17 (3):1–2, May/June 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:TCd

[Ano19-81] Anonymous. Table of contents. IEEE Security & Privacy, 17(4):1–2, July/August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:TCe

[Ano19-82] Anonymous. Table of contents. IEEE Security & Privacy, 17(5):1–2, September/ October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2019:TCf

[Ano19-83] Anonymous. Table of contents.

IEEE Security & Privacy, 17 (6):1–2, November/December 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:AWC

[Ano20a] Anonymous. AI’s 10 to watch call for nominations. IEEE Security & Privacy, 18(4):55, July/August 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:CEM

[Ano20b] Anonymous. Computing Edge magazine. IEEE Security & Privacy, 18(1):11, January/ February 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:IAH

[Ano20c] Anonymous. IEEE Annals of the History of Computing. IEEE Security & Privacy, 18(5):17, September/October 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:ICGa

[Ano20d] Anonymous. IEEE Computer Graphics & Applications. IEEE Security & Privacy, 18(4):37, July/August 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:ICGb

[Ano20e] Anonymous. IEEE Computer Graphics & Applications. IEEE Security & Privacy, 18(5): 37, September/October 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:ICEa

[Ano20f] Anonymous. IEEE Computing Edge. IEEE Security & Privacy, 18(4):C4, July/August 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:ICEc

[Ano20g] Anonymous. IEEE Computing Edge. IEEE Security & Privacy, 18(6):C4, November/ December 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:ICEb

[Ano20h] Anonymous. IEEE Computing Edge magazine. IEEE Security & Privacy, 18(5):C4, September/October 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:IIPa

[Ano20i] Anonymous. IEEE IT Professional call for articles. IEEE Security & Privacy, 18(4):19, July/August 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:IIPb

[Ano20j] Anonymous. IEEE IT Professional call for articles. IEEE Security & Privacy, 18(5): 56, September/October 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:IIPc

[Ano20k] Anonymous. IEEE IT Professional call for articles. IEEE Security & Privacy, 18(6): 63, November/December 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:IOA

[Ano20l] Anonymous. IEEE Open Access journal of the Computer Society. IEEE Security & Privacy, 18(6):C2, November/ December 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:IPCa

[Ano20m] Anonymous. IEEE Pervasive Computing call for articles. IEEE Security & Privacy, 18(4):66, July/August 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:IPCb

[Ano20n] Anonymous. IEEE Pervasive Computing call for articles. IEEE Security & Privacy, 18(5):73, September/October 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:ISP

[Ano20o] Anonymous. IEEE Security & Privacy [advertisement]. IEEE Security & Privacy, 18(4):70, July/August 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:ITC

[Ano20p] Anonymous. IEEE Transactions on Computers. IEEE Security & Privacy, 18(1):7, January/February 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:ITS

[Ano20q] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Security & Privacy, 18(5):7, September/ October 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:C

[Ano20r] Anonymous. Compsac 2020. IEEE Security & Privacy, 18(1):C2, January/February 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:CST

[Ano20s] Anonymous. Computer Society technology predictions. IEEE Security & Privacy, 18(1):10, January/February 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:Ea

[Ano20t] Anonymous. Errata. IEEE Security & Privacy, 18(2):55, March/April 2020. ISSN 1558-4046.

Anonymous:2020:Eb

[Ano20u] Anonymous. Errata. IEEE Security & Privacy, 18(3):61, May/June 2020. ISSN 1558-4046.

Anonymous:2020:FCa

[Ano20v]  Anonymous.  Front cover.

IEEE Security & Privacy, 18(1):C1, January/February 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:FCb

[Ano20w]  Anonymous.  Front cover.

IEEE Security & Privacy, 18(4):C1, July/August 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:FCc

[Ano20x]  Anonymous.  Front cover.

IEEE Security & Privacy, 18(5):C1, September/October 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:FCd

[Ano20y]  Anonymous.  Front cover.

IEEE Security & Privacy, 18 (6):C1, November/December 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:HDS

[Ano20z] Anonymous. HOST 2020: 6–9 Dec. 2020, San Jose, CA. IEEE Security & Privacy, 18(4):71, July/August 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:ICSe

[Ano20-27] Anonymous. IEEE Computer Society call for papers. IEEE Security & Privacy, 18(5): 46, September/October 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:ICSg

[Ano20-28] Anonymous. IEEE Computer Society call for papers. IEEE Security & Privacy, 18(6): 22, November/December 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:ICSa

[Ano20-29] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 18(1):C3, January/February 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:ICSb

[Ano20-30] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 18(2):C3, March/April 2020. ISSN 1558-4046.

Anonymous:2020:ICSd

[Ano20-31] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 18(4):C3, July/August 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:ICSf

[Ano20-32] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 18(5): C3, September/October 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:ICSh

[Ano20-33] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 18(6): C3, November/December 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:ICSc

[Ano20-34] Anonymous. IEEE Computer Society member benefits. IEEE Security & Privacy, 18(4):C2, July/August 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:IJB

[Ano20-35] Anonymous. IEEE Jobs Board.

IEEE Security & Privacy, 18(1):C4, January/February 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:Ma

[Ano20-36] Anonymous. Masthead. IEEE Security & Privacy, 18(1):3, January/February 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:Mb

[Ano20-37] Anonymous. Masthead. IEEE Security & Privacy, 18(2):3, March/April 2020. ISSN 1558-4046.

Anonymous:2020:Mc

[Ano20-38] Anonymous. Masthead. IEEE Security & Privacy, 18(3):3, May/June 2020. ISSN 1558-4046.

Anonymous:2020:Md

[Ano20-39] Anonymous. Masthead. IEEE Security & Privacy, 18(4):3, July/August 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:Me

[Ano20-40] Anonymous. Masthead. IEEE Security & Privacy, 18(5): 2, September/October 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:Mf

[Ano20-41] Anonymous. Masthead. IEEE Security & Privacy, 18(6): 3, November/December 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:RSa

[Ano20-42] Anonymous. Reliability Society. IEEE Security & Privacy, 18(1):6, January/ February 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:RSb

[Ano20-43] Anonymous. Reliability Society. IEEE Security & Privacy, 18(2):6, March/April 2020. ISSN 1558-4046.

Anonymous:2020:RSc

[Ano20-44] Anonymous. Reliability Society. IEEE Security & Privacy, 18(3):5, May/June 2020. ISSN 1558-4046.

Anonymous:2020:RSd

[Ano20-45] Anonymous. Reliability society. IEEE Security & Privacy, 18(4):5, July/August 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:RSe

[Ano20-46] Anonymous. Reliability society. IEEE Security & Privacy, 18(5):4, September/October 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:RSf

[Ano20-47] Anonymous. Reliability society. IEEE Security & Privacy, 18(6):5, November/December 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:STI

[Ano20-48] Anonymous. Submit today:

IEEE Transactions on Big Data. IEEE Security & Privacy, 18(4):75, July/August 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:TCa

[Ano20-49] Anonymous. Table of contents.

IEEE Security & Privacy, 18(1):1–2, January/February 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:TCb

[Ano20-50] Anonymous. Table of contents. IEEE Security & Privacy, 18(2):1–2, March/April 2020. ISSN 1558-4046.

Anonymous:2020:TCc

[Ano20-51] Anonymous. Table of contents. IEEE Security & Privacy, 18 (3):1–2, May/June 2020. ISSN 1558-4046.

Anonymous:2020:TCd

[Ano20-52] Anonymous. Table of contents. IEEE Security & Privacy, 18(4):1–2, July/August 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:TCe

[Ano20-53] Anonymous. Table of contents. IEEE Security & Privacy, 18 (5):C2–1, September/October 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2020:TCf

[Ano20-54] Anonymous. Table of contents. IEEE Security & Privacy, 18 (6):1–2, November/December 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ISS

[Ano21a] Anonymous. 42nd IEEE Symposium on Security & Privacy. IEEE Security & Privacy, 19(2):13, March/April 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:IAHb

[Ano21b] Anonymous. IEEE Annals of the History Computing. IEEE Security & Privacy, 19(2):28, March/April 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:IAHa

[Ano21c] Anonymous. IEEE Annals of the History of Computing. IEEE Security & Privacy, 19(1):23, January/February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:IAHd

[Ano21d] Anonymous. IEEE Annals of the History of Computing. IEEE Security & Privacy, 19(4):69, July/August 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:IAHe

[Ano21e] Anonymous. IEEE Annals of the History of Computing. IEEE Security & Privacy, 19(5):15, September/October 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:IAHf

[Ano21f] Anonymous. IEEE Annals of the History of Computing. IEEE Security & Privacy, 19(6):7, November/December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:IAHc

[Ano21g] Anonymous. IEEE Annals of the History of Computing [advertisement]. IEEE Security & Privacy, 19(3):7, May/June 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICGb

[Ano21h] Anonymous. IEEE Computer Graphics & Applications. IEEE Security & Privacy, 19(2):47, March/April 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICGc

[Ano21i] Anonymous. IEEE Computer Graphics & Applications. IEEE Security & Privacy, 19(6): 14, November/December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICGa

[Ano21j]  Anonymous.  IEEE Com-

puter Graphics and Applications. IEEE Security & Privacy, 19(1):33, January/ February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICEa

[Ano21k] Anonymous. IEEE Computing Edge. IEEE Security & Privacy, 19(1):C2, January/ February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICEb

[Ano21l] Anonymous. IEEE Computing Edge. IEEE Security & Privacy, 19(2):C2, March/April 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICEd

[Ano21m] Anonymous. IEEE Computing Edge. IEEE Security & Privacy, 19(4):C2, July/August 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICEe

[Ano21n] Anonymous. IEEE Computing Edge. IEEE Security & Privacy, 19(5):71, September/ October 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICEf

[Ano21o] Anonymous. IEEE Computing Edge. IEEE Security & Privacy, 19(6):C2, November/ December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICEc

[Ano21p] Anonymous. IEEE Computing Edge [advertisement]. IEEE Security & Privacy, 19(3):C2, May/June 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICSn

[Ano21q] Anonymous. IEEE Computing in Science & Engineering. IEEE Security & Privacy, 19(5):25, September/October 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICSj

[Ano21r] Anonymous. IEEE Computing in Science and Engineering. IEEE Security & Privacy, 19(4):64, July/August 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICSs

[Ano21s] Anonymous. IEEE Computing in Science and Engineering. IEEE Security & Privacy, 19(6):92, November/December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:IIPa

[Ano21t] Anonymous. IEEE IT Professional call for articles. IEEE Security & Privacy, 19(1):64, January/February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:IIPb

[Ano21u] Anonymous. IEEE IT Professional call for articles. IEEE Security & Privacy, 19(2):69, March/April 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:IIPc

[Ano21v] Anonymous. IEEE IT Professional call for articles. IEEE Security & Privacy, 19(5): 45, September/October 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:IIPd

[Ano21w] Anonymous. IEEE IT Professional call for articles. IEEE Security & Privacy, 19 (6):102, November/December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:IOJa

[Ano21x] Anonymous. IEEE Open Journal of the Computer Society. IEEE Security & Privacy, 19(1):89, January/February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:IOJb

[Ano21y] Anonymous. IEEE Open Journal of the Computer Society. IEEE Security & Privacy, 19 (3):57, May/June 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:IOJc

[Ano21z] Anonymous. IEEE Open Journal of the Computer Society. IEEE Security & Privacy, 19(4):75, July/August 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:IOJd

[Ano21-27] Anonymous. IEEE Open Journal of the Computer Society. IEEE Security & Privacy, 19(5):77, September/October 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:IOJe

[Ano21-28] Anonymous. IEEE Open Journal of the Computer Society. IEEE Security & Privacy, 19 (6):103, November/December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:IPCa

[Ano21-29] Anonymous. IEEE Pervasive Computing call for articles. IEEE Security & Privacy, 19(1):73, January/ February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:IPCb

[Ano21-30] Anonymous. IEEE Pervasive Computing call for articles. IEEE Security & Privacy, 19(2):73, March/April 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:IPCc

[Ano21-31] Anonymous. IEEE Pervasive Computing call for articles. IEEE Security & Privacy, 19 (6):110, November/December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ISP

[Ano21-32] Anonymous. IEEE Security and Privacy subscription. IEEE Security & Privacy, 19 (6):118, November/December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ITBa

[Ano21-33] Anonymous. IEEE Transactions on Big Data. IEEE Security & Privacy, 19(1):80, January/February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ITBb

[Ano21-34] Anonymous. IEEE Transactions on Big Data. IEEE Security & Privacy, 19(6):123, November/December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ITC

[Ano21-35] Anonymous. IEEE Transactions on Computers call for papers. IEEE Security & Privacy, 19(2):78, March/April 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ITSa

[Ano21-36] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Security & Privacy, 19(1):88, January/ February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ITSb

[Ano21-37] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Security & Privacy, 19(2):79, March/April 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ITSc

[Ano21-38] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Security & Privacy, 19(5):95, September/October 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:DES

[Ano21-39] Anonymous. A discussion of election security, cryptography, and exceptional access with Michael Alan Specter. IEEE Security & Privacy, 19(6):15– 22, November/December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:E

[Ano21-40] Anonymous. Errata. IEEE Security & Privacy, 19(1):103, January/February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ETA

[Ano21-41] Anonymous. Erratum [for “A Taxonomy of Attacks on Federated Learning”]. IEEE Security & Privacy, 19(3):62, May/June 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:FCa

[Ano21-42] Anonymous. Front cover. IEEE Security & Privacy, 19(1):C1, January/February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:FCb

[Ano21-43] Anonymous. Front cover.

IEEE Security & Privacy, 19(2):C1, March/April 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:FCc

[Ano21-44] Anonymous. [Front cover].

IEEE Security & Privacy, 19 (3):C1, May/June 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:FCd

[Ano21-45] Anonymous. [Front cover].

IEEE Security & Privacy, 19(4):C1, July/August 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:FCe

[Ano21-46] Anonymous. Front cover.

IEEE Security & Privacy, 19(5):C1, September/October 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:FCf

[Ano21-47] Anonymous. Front cover.

IEEE Security & Privacy, 19 (6):C1, November/December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICSf

[Ano21-48] Anonymous. IEEE Computer Society call for papers. IEEE Security & Privacy, 19(3):63, May/June 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICSi

[Ano21-49] Anonymous. IEEE Computer Society call for papers. IEEE Security & Privacy, 19(4):53, July/August 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICSo

[Ano21-50] Anonymous. IEEE Computer Society call for papers. IEEE Security & Privacy, 19(5): 89, September/October 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICSt

[Ano21-51] Anonymous. IEEE Computer Society call for papers. IEEE Security & Privacy, 19(6): 97, November/December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICSr

[Ano21-52] Anonymous. IEEE Computer Society Diversity and Integrity Fund. IEEE Security & Privacy, 19(6):25, November/ December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICSm

[Ano21-53] Anonymous. IEEE Computer Society elections. IEEE Security & Privacy, 19(5): C2, September/October 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICSb

[Ano21-54] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 19(1):C3, January/February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICSd

[Ano21-55] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 19(2):C3, March/April 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICSg

[Ano21-56] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 19(3):C3, May/June 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICSk

[Ano21-57] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 19(4):C3, July/August 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICSp

[Ano21-58] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 19(5): C3, September/October 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICSv

[Ano21-59] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 19(6): C3, November/December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICSc

[Ano21-60] Anonymous. IEEE Computer Society member benefits. IEEE Security & Privacy, 19(1):C4, January/February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICSe

[Ano21-61] Anonymous. IEEE Computer Society member benefits. IEEE Security & Privacy, 19(2):C4, March/April 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICSh

[Ano21-62] Anonymous. IEEE Computer Society member benefits. IEEE Security & Privacy, 19(3):C4, May/June 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICSl

[Ano21-63] Anonymous. IEEE Computer Society member benefits. IEEE Security & Privacy, 19(4):C4, July/August 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICSq

[Ano21-64] Anonymous. IEEE Computer Society member benefits. IEEE Security & Privacy, 19(5):C4, September/October 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICSu

[Ano21-65] Anonymous. IEEE Computer Society member benefits. IEEE Security & Privacy, 19 (6):111, November/December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:ICSa

[Ano21-66] Anonymous. IEEE Computing Society call for papers. IEEE Security & Privacy, 19(1):43, January/February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:IJB

[Ano21-67] Anonymous. IEEE Job Boards.

IEEE Security & Privacy, 19 (6):C4, November/December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:IJ

[Ano21-68] Anonymous. IEEE JobBoards.

IEEE Security & Privacy, 19 (3):50, May/June 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:Ma

[Ano21-69] Anonymous. Masthead. IEEE Security & Privacy, 19(1):3, January/February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:Mb

[Ano21-70] Anonymous. Masthead. IEEE Security & Privacy, 19(2):3, March/April 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:Mc

[Ano21-71] Anonymous. Masthead. IEEE Security & Privacy, 19(3):3, May/June 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:Md

[Ano21-72] Anonymous. Masthead. IEEE Security & Privacy, 19(4):3, July/August 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:Me

[Ano21-73] Anonymous. Masthead. IEEE Security & Privacy, 19(5): 3, September/October 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:Mf

[Ano21-74] Anonymous. Masthead. IEEE Security & Privacy, 19(6): 3, November/December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:RSa

[Ano21-75] Anonymous. Reliability society. IEEE Security & Privacy, 19(1):7, January/ February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:RSb

[Ano21-76] Anonymous. Reliability Society. IEEE Security & Privacy, 19(2):5, March/April 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:RSc

[Ano21-77] Anonymous. Reliability Society. IEEE Security & Privacy, 19(3):5, May/June 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:RSd

[Ano21-78] Anonymous. Reliability society. IEEE Security & Privacy, 19(4):5, July/August 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:RSe

[Ano21-79] Anonymous. Reliability society. IEEE Security & Privacy, 19(5):5, September/October 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:RSf

[Ano21-80] Anonymous. Reliability Society. IEEE Security & Privacy, 19(6):5, November/December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:SIS

[Ano21-81] Anonymous. Special issue on security and privacy issues of home globalization. IEEE Security & Privacy, 19(1):11, January/February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:TCa

[Ano21-82] Anonymous. Table of contents. IEEE Security & Privacy, 19(1):1–2, January/February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:TCb

[Ano21-83] Anonymous. Table of contents. IEEE Security & Privacy, 19(2):1–2, March/April 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:TCc

[Ano21-84] Anonymous. Table of contents. IEEE Security & Privacy, 19 (3):1–2, May/June 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:TCd

[Ano21-85] Anonymous. Table of contents. IEEE Security & Privacy, 19(4):1–2, July/August 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:TCe

[Ano21-86] Anonymous. Table of contents. IEEE Security & Privacy, 19(5):1–2, September/ October 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:TCf

[Ano21-87] Anonymous. Table of contents. IEEE Security & Privacy, 19 (6):1–2, November/December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:TFS

[Ano21-88] Anonymous. Toward fail safety for security decisions. IEEE Security & Privacy, 19(6):4– 7, November/December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2021:UIS

[Ano21-89] Anonymous. Updates from IEEE Security & Privacy’s Editor in Chief Sean Peisert. IEEE Security & Privacy, 19(5):8, September/October 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ICSl

[Ano22a] Anonymous. 2023 IEEE Computer Society Nomination for Watts S. Humphrey Award. IEEE Security & Privacy, 20(4):70, July/August 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:CE

[Ano22b] Anonymous. Computing Edge.

IEEE Security & Privacy, 20 (3):C2, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:IAH

[Ano22c] Anonymous. IEEE Annals of the History of Computing. IEEE Security & Privacy, 20(2):48, March/April 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:IC

[Ano22d]  Anonymous.  IEEE CGA.

IEEE Security & Privacy, 20 (3):21, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ICGa

[Ano22e]  Anonymous.  IEEE Com-

puter Graphics and Applications. IEEE Security & Privacy, 20(1):95, January/ February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ICGb

[Ano22f]  Anonymous.  IEEE Com-

puter Graphics and Applications. IEEE Security & Privacy, 20(2):95, March/April 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ICEa

[Ano22g] Anonymous. IEEE Computing Edge. IEEE Security & Privacy, 20(1):C2, January/ February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ICEb

[Ano22h] Anonymous. IEEE Computing Edge. IEEE Security & Privacy, 20(2):C2, March/April 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ICEc

[Ano22i] Anonymous. IEEE Computing Edge. IEEE Security & Privacy, 20(4):C2, July/August 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ICSa

[Ano22j]  Anonymous.  IEEE Com-

puting in Science and Engineering. IEEE Security & Privacy, 20(1):22, January/ February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ICSk

[Ano22k] Anonymous. IEEE Computing in Science and Engineering. IEEE Security & Privacy, 20(4):60, July/August 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ICSr

[Ano22l] Anonymous. IEEE Computing in Science and Engineering. IEEE Security & Privacy, 20(6):82, November/December 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:IIPa

[Ano22m] Anonymous. IEEE IT Professional call for articles. IEEE Security & Privacy, 20(1):71, January/February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:IIPb

[Ano22n] Anonymous. IEEE IT Professional call for articles. IEEE Security & Privacy, 20(6): 72, November/December 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:IOJa

[Ano22o] Anonymous. IEEE Open Journal of the Computer Society. IEEE Security & Privacy, 20(1):81, January/February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:IOJc

[Ano22p] Anonymous. IEEE Open Journal of the Computer Society. IEEE Security & Privacy, 20(6):73, November/December 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:IOJb

[Ano22q] Anonymous. IEEE Open Journal of the Computing Society. IEEE Security & Privacy, 20(5):100, September/October 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:IPCb

[Ano22r] Anonymous. IEEE Pervasive Computing. IEEE Security & Privacy, 20(3):70, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:IPCa

[Ano22s] Anonymous. IEEE Pervasive Computing call for articles. IEEE Security & Privacy, 20(1):86, January/ February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:IPCc

[Ano22t] Anonymous. IEEE Pervasive Computing call for articles. IEEE Security & Privacy, 20(5):76, September/October 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:IPCd

[Ano22u] Anonymous. IEEE Pervasive Computing call for papers. IEEE Security & Privacy, 20(6):52, November/December 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ISPa

[Ano22v] Anonymous. IEEE Security & Privacy subscribe. IEEE Security & Privacy, 20(4):71, July/ August 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ISPd

[Ano22w] Anonymous. IEEE Security and Privacy magazine subscribe. IEEE Security & Privacy, 20(6):42, November/ December 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ISPc

[Ano22x] Anonymous. IEEE Security and Privacy Over the Rainbow podcast. IEEE Security & Privacy, 20(5):107, September/October 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ISPb

[Ano22y] Anonymous. IEEE Security and Privacy subscribe. IEEE Security & Privacy, 20(5): 34, September/October 2022.

ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ITB

[Ano22z] Anonymous. IEEE Transactions on Big Data. IEEE Security & Privacy, 20(1):112, January/February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ITC

[Ano22-27] Anonymous. IEEE Transactions on Computers call for papers. IEEE Security & Privacy, 20(1):118, January/ February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ITSa

[Ano22-28] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Security & Privacy, 20(1):119, January/ February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:FCa

[Ano22-29] Anonymous. [Front cover]. IEEE Security & Privacy, 20(1):C1, January/February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:FCb

[Ano22-30] Anonymous. Front cover. IEEE Security & Privacy, 20(2):C1, March/April 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:FCc

[Ano22-31] Anonymous. Front cover.

IEEE Security & Privacy, 20 (3):C1, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:FCd

[Ano22-32] Anonymous. Front cover.

IEEE Security & Privacy, 20(4):C1, July/August 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:FCe

[Ano22-33] Anonymous. Front cover.

IEEE Security & Privacy, 20(5):C1, September/October 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:FCf

[Ano22-34] Anonymous. Front cover.

IEEE Security & Privacy, 20 (6):C1, November/December 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ICSc

[Ano22-35] Anonymous. IEEE Computer Society call for papers. IEEE Security & Privacy, 20(1):61, January/February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ICSf

[Ano22-36] Anonymous. IEEE Computer Society call for papers. IEEE Security & Privacy, 20(2):107, March/April 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ICSj

[Ano22-37] Anonymous. IEEE Computer Society call for papers. IEEE Security & Privacy, 20(4):50, July/August 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ICSn

[Ano22-38] Anonymous. IEEE Computer Society call for papers. IEEE Security & Privacy, 20(5): 10, September/October 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ICSq

[Ano22-39] Anonymous. IEEE Computer Society call for papers. IEEE Security & Privacy, 20(6): 26, November/December 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ICSe

[Ano22-40] Anonymous. IEEE Computer Society call for special issue proposals. IEEE Security & Privacy, 20(2):75, March/April 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ICSp

[Ano22-41] Anonymous. IEEE Computer Society diversity and inclusion. IEEE Security & Privacy, 20(6):C2, November/ December 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ICSi

[Ano22-42] Anonymous. IEEE Computer Society election. IEEE Security & Privacy, 20(4):22, July/ August 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ICSd

[Ano22-43] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 20(1):C3, January/February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ICSg

[Ano22-44] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 20(2):C3, March/April 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ICSh

[Ano22-45] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 20(3):C3, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ICSm

[Ano22-46] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 20(4):C3, July/August 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ICSo

[Ano22-47] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 20(5): C3, September/October 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ICSs

[Ano22-48] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 20(6): C3, November/December 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ICSb

[Ano22-49] Anonymous. IEEE Computer Society member benefits. IEEE Security & Privacy, 20(1):32, January/February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ICC

[Ano22-50] Anonymous. IEEE CS CFP filler. IEEE Security & Privacy, 20(3):23, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ICI

[Ano22-51] Anonymous. IEEE CS IT Professional filler. IEEE Security & Privacy, 20(3):32, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ICJ

[Ano22-52] Anonymous. IEEE CS Job Boards. IEEE Security & Privacy, 20(3):C4, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:IDI

[Ano22-53] Anonymous. IEEE Diversity and Inclusion Fund. IEEE Security & Privacy, 20(1):101, January/February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:IJBb

[Ano22-54] Anonymous. IEEE Job Boards.

IEEE Security & Privacy, 20(4):C4, July/August 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:IJBa

[Ano22-55] Anonymous. IEEE Jobs Board.

IEEE Security & Privacy, 20(2):C4, March/April 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:IJBc

[Ano22-56] Anonymous. IEEE Jobs Board.

IEEE Security & Privacy, 20(5):C4, September/October 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:IJBd

[Ano22-57] Anonymous. IEEE Jobs Board.

IEEE Security & Privacy, 20 (6):C4, November/December 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:IJ

[Ano22-58] Anonymous. IEEE JobSite.

IEEE Security & Privacy, 20(1):C4, January/February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:IOC

[Ano22-59] Anonymous. IEEE OJ Computer Society filler. IEEE Security & Privacy, 20(3):101, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:IOA

[Ano22-60] Anonymous. IEEE Open Access of the Computer Society. IEEE Security & Privacy, 20(2):87, March/April 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:IQW

[Ano22-61] Anonymous. IEEE Quantum Week 2022. IEEE Security & Privacy, 20(5):C2, September/ October 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ISR

[Ano22-62] Anonymous. IEEE SP over the rainbow filler2. IEEE Security & Privacy, 20(3):42, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ITF

[Ano22-63] Anonymous. IEEE tbd filler. IEEE Security & Privacy, 20 (3):90, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ITG

[Ano22-64] Anonymous. IEEE TC generic CFP filler. IEEE Security & Privacy, 20(3):100, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:ITSb

[Ano22-65] Anonymous. IEEE Trans Sustainable Comp filler. IEEE Security & Privacy, 20(3):107, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:IIS

[Ano22-66] Anonymous. Introducing IEEE Security & Privacy’s “Off by one” column. IEEE Security & Privacy, 20(3):9, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:Ma

[Ano22-67] Anonymous. Masthead. IEEE Security & Privacy, 20(1):3, January/February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:Mb

[Ano22-68] Anonymous. Masthead. IEEE Security & Privacy, 20(2):3, March/April 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:Mc

[Ano22-69] Anonymous. Masthead. IEEE Security & Privacy, 20(3):3, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:Md

[Ano22-70] Anonymous. Masthead. IEEE Security & Privacy, 20(4):3, July/August 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:Me

[Ano22-71] Anonymous. Masthead. IEEE Security & Privacy, 20(5): 3, September/October 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:Mf

[Ano22-72] Anonymous. Masthead. IEEE Security & Privacy, 20(6): 3, November/December 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:RP

[Ano22-73] Anonymous. Over the Rainbow podcast. IEEE Security & Privacy, 20(6):83, November/ December 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:PGT

[Ano22-74] Anonymous. Pseudo groundtruth generators and largescale studies. IEEE Security & Privacy, 20(2):4–7, March/ April 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:RSa

[Ano22-75] Anonymous. Reliability Society. IEEE Security & Privacy, 20(1):5, January/ February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:RSb

[Ano22-76] Anonymous. Reliability Society. IEEE Security & Privacy, 20(2):5, March/April 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:RSc

[Ano22-77] Anonymous. Reliability society. IEEE Security & Privacy, 20(3):5, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:RSd

[Ano22-78] Anonymous. Reliability society. IEEE Security & Privacy, 20(4):5, July/August 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:RSe

[Ano22-79] Anonymous. Reliability society. IEEE Security & Privacy, 20(5):5, September/October 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:RSf

[Ano22-80] Anonymous. Reliability society. IEEE Security & Privacy, 20(6):7, November/December 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:TCa

[Ano22-81] Anonymous. Table of contents.

IEEE Security & Privacy, 20(1):1–2, January/February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:TCb

[Ano22-82] Anonymous. Table of contents. IEEE Security & Privacy, 20(2):1–2, March/April 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:TCc

[Ano22-83] Anonymous. Table of contents. IEEE Security & Privacy, 20 (3):1–2, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:TCd

[Ano22-84] Anonymous. Table of contents. IEEE Security & Privacy, 20(4):1–2, July/August 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:TCe

[Ano22-85] Anonymous. Table of contents. IEEE Security & Privacy, 20(5):1–2, September/ October 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:TCf

[Ano22-86] Anonymous. Table of contents. IEEE Security & Privacy, 20 (6):1–2, November/December 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:UIS

[Ano22-87] Anonymous. Updates from IEEE Security & Privacy’s Editor-in-Chief Sean Peisert: And now for something completely different. IEEE Security & Privacy, 20(3):11, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2022:WLO

[Ano22-88] Anonymous. Welcome to the latest off by one column. IEEE Security & Privacy, 20(5): 15, September/October 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

xx-90315775

[Ano23a] Anonymous. 0. IEEE Security & Privacy, 21(6):95, November/December 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ICSf

[Ano23b]  Anonymous.  2024 IEEE

Computer Society Watts S. Humphery Software Quality Award call for nominations. IEEE Security & Privacy, 21(2):47, March/April 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:CE

[Ano23c] Anonymous. Computing Edge.

IEEE Security & Privacy, 21 (6):C4, November/December 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:IAF

[Ano23d] Anonymous. IEEE Annals filler. IEEE Security & Privacy, 21(6):94, November/ December 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ICEa

[Ano23e] Anonymous. IEEE Computing Edge. IEEE Security & Privacy, 21(1):C4, January/ February 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ICEb

[Ano23f] Anonymous. IEEE Computing Edge. IEEE Security & Privacy, 21(2):C4, March/April 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ICEd

[Ano23g] Anonymous. IEEE Computing Edge. IEEE Security & Privacy, 21(5):C4, September/ October 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ISPb

[Ano23h] Anonymous. IEEE Security & Privacy subscribe filler. IEEE Security & Privacy, 21(6): 7, November/December 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ISPa

[Ano23i] Anonymous. IEEE Security and Privacy subscribe. IEEE Security & Privacy, 21(1):90, January/February 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ISMa

[Ano23j]  Anonymous.  IEEE Soft-

ware magazine call for articles. IEEE Security & Privacy, 21(2):89, March/April 2023.

ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ISMb

[Ano23k]  Anonymous.  IEEE Soft-

ware magazine call for articles. IEEE Security & Privacy, 21 (4):8, July/August 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:FCa

[Ano23l]  Anonymous.  Front cover.

IEEE Security & Privacy, 21(1):C1, January/February 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:FCb

[Ano23m]  Anonymous.  Front cover.

IEEE Security & Privacy, 21(2):C1, March/April 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:FCc

[Ano23n]  Anonymous.  Front cover.

IEEE Security & Privacy, 21 (3):C1, May/June 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:FCd

[Ano23o]  Anonymous.  Front cover.

IEEE Security & Privacy, 21(4):C1, July/August 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:FCe

[Ano23p]  Anonymous.  Front cover.

IEEE Security & Privacy, 21(5):C1, September/October 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:FCf

[Ano23q]  Anonymous.  Front cover.

IEEE Security & Privacy, 21 (6):C1, November/December 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ICI

[Ano23r] Anonymous. IEEE CAI 2023: IEEE Conference on Artificial Intelligence. IEEE Security & Privacy, 21(2):17, March/April 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ICC

[Ano23s] Anonymous. IEEE career center. IEEE Security & Privacy, 21(5):67, September/October 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ICSg

[Ano23t] Anonymous. IEEE Computer Society. IEEE Security & Privacy, 21(2):C3, March/April 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ICSi

[Ano23u] Anonymous. IEEE Computer Society. IEEE Security & Privacy, 21(3):C3, May/June 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ICSa

[Ano23v] Anonymous. IEEE Computer Society benefits. IEEE Security & Privacy, 21(1):C2, January/ February 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ICSj

[Ano23w] Anonymous. IEEE Computer Society Computing Edge. IEEE Security & Privacy, 21(3):C4, May/June 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ICSc

[Ano23x] Anonymous. IEEE Computer Society call for papers. IEEE Security & Privacy, 21(1):36, January/February 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ICSm

[Ano23y] Anonymous. IEEE Computer Society call for papers. IEEE Security & Privacy, 21(5): 17, September/October 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ICSr

[Ano23z] Anonymous. IEEE Computer Society Career Center filler. IEEE Security & Privacy, 21 (6):11, November/December 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ICSs

[Ano23-27] Anonymous. IEEE Computer Society CFP filler. IEEE Security & Privacy, 21(6): 23, November/December 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ICSn

[Ano23-28] Anonymous. IEEE Computer Society diversity and inclusion. IEEE Security & Privacy, 21(5):78, September/October 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ICSp

[Ano23-29] Anonymous. IEEE Computer Society diversity filler. IEEE Security & Privacy, 21(6): C2, November/December 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ICSq

[Ano23-30] Anonymous. IEEE Computer Society has you covered. IEEE Security & Privacy, 21(6):10, November/December 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ICSd

[Ano23-31] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 21(1):C3, January/February 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ICSl

[Ano23-32] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 21(4):C3, July/August 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ICSu

[Ano23-33] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 21(6): C3, November/December 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ICSe

[Ano23-34] Anonymous. IEEE Computer Society member benefits. IEEE Security & Privacy, 21(2):C2, March/April 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ICSh

[Ano23-35] Anonymous. IEEE Computer Society member benefits. IEEE Security & Privacy, 21(3):C2, May/June 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ICSk

[Ano23-36] Anonymous. IEEE Computer Society member benefits. IEEE Security & Privacy, 21(4):83, July/August 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ICSo

[Ano23-37] Anonymous. IEEE Computer Society member benefits. IEEE Security & Privacy, 21(5):C3, September/October 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ICSb

[Ano23-38] Anonymous. IEEE Computer Society publications seek 2024 Editors in Chief. IEEE Security & Privacy, 21(1):7, January/February 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ICSt

[Ano23-39] Anonymous. IEEE Computer Society seeks applicants for Editors in Chief. IEEE Security & Privacy, 21(6):89, November/ December 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:ICEc

[Ano23-40] Anonymous. IEEE computing edge. IEEE Security & Privacy, 21(4):C4, July/August 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:IJBa

[Ano23-41] Anonymous. IEEE Job Boards. IEEE Security & Privacy, 21(1):19, January/February 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:IJBb

[Ano23-42] Anonymous. IEEE Job Boards. IEEE Security & Privacy, 21(2):83, March/April 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:IQW

[Ano23-43] Anonymous. IEEE quantum week. IEEE Security & Privacy, 21(3):36, May/June 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:IRS

[Ano23-44] Anonymous. IEEE Reliability Society. IEEE Security & Privacy, 21(5):77, September/ October 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:Ma

[Ano23-45] Anonymous. Masthead. IEEE Security & Privacy, 21(1):3, January/February 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:Mb

[Ano23-46] Anonymous. Masthead. IEEE Security & Privacy, 21(2):3, March/April 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:Mc

[Ano23-47] Anonymous. Masthead. IEEE Security & Privacy, 21(3):3, May/June 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:Md

[Ano23-48] Anonymous. Masthead. IEEE Security & Privacy, 21(4):3, July/August 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:Me

[Ano23-49] Anonymous. Masthead. IEEE Security & Privacy, 21(5):

3, September/October 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:Mf

[Ano23-50] Anonymous. Masthead. IEEE Security & Privacy, 21(6): 3, November/December 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:RPa

[Ano23-51] Anonymous. Over the rainbow podcast. IEEE Security & Privacy, 21(2):42, March/April 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:RPb

[Ano23-52] Anonymous. Over the rainbow podcast. IEEE Security & Privacy, 21(3):15, May/June 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:RPc

[Ano23-53] Anonymous. Over the rainbow podcast. IEEE Security & Privacy, 21(4):99, July/August 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:QW

[Ano23-54] Anonymous. Quantum week 2023. IEEE Security & Privacy, 21(4):C2, July/August 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:RSa

[Ano23-55] Anonymous. Reliability society. IEEE Security & Privacy, 21(1):5, January/ February 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:RSb

[Ano23-56] Anonymous. Reliability society. IEEE Security & Privacy, 21(2):5, March/April 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:RSc

[Ano23-57] Anonymous. Reliability society. IEEE Security & Privacy, 21(3):5, May/June 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:RSd

[Ano23-58] Anonymous. Reliability society. IEEE Security & Privacy, 21(4):5, July/August 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:RSe

[Ano23-59] Anonymous. Reliability society. IEEE Security & Privacy, 21(6):5, November/December 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:TCa

[Ano23-60] Anonymous. Table of contents.

IEEE Security & Privacy, 21(1):1–2, January/February 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:TCb

[Ano23-61] Anonymous. Table of contents. IEEE Security & Privacy, 21(2):1–2, March/April 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:TCc

[Ano23-62] Anonymous. Table of contents. IEEE Security & Privacy, 21 (3):1–2, May/June 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:TCd

[Ano23-63] Anonymous. Table of contents. IEEE Security & Privacy, 21(4):1–2, July/August 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:TCe

[Ano23-64] Anonymous. Table of contents. IEEE Security & Privacy, 21(5):1–2, September/ October 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:TCf

[Ano23-65] Anonymous. Table of contents. IEEE Security & Privacy, 21 (6):1–2, November/December 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2023:WHS

[Ano23-66] Anonymous. Watts S. Humphrey Software Quality Award nominations. IEEE Security & Privacy, 21(3):26, May/June 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:CSE

[Ano24a] Anonymous. Computing in Science & Engineering. IEEE Security & Privacy, 22(2):68, March/April 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:IAC

[Ano24b] Anonymous. IEEE Annals in Computing. IEEE Security & Privacy, 22(1):75, January/ February 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:IAH

[Ano24c] Anonymous. IEEE Annals of the History of Computing. IEEE Security & Privacy, 22(2):48, March/April 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:ICG

[Ano24d]  Anonymous.  IEEE Com-

puter Graphics and Applications. IEEE Security & Privacy, 22(2):59, March/April 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:ICSm

[Ano24e] Anonymous. IEEE Computers In Science & Engineering. IEEE Security & Privacy, 22 (3):61, May/June 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:ICEa

[Ano24f] Anonymous. IEEE Computing Edge. IEEE Security & Privacy, 22(1):C4, January/ February 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:ICEb

[Ano24g] Anonymous. IEEE Computing Edge. IEEE Security & Privacy, 22(2):C4, March/April 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:ICEc

[Ano24h] Anonymous. IEEE Computing Edge. IEEE Security & Privacy, 22(3):C4, May/June 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:IPCb

[Ano24i] Anonymous. IEEE Pervasive Computing. IEEE Security & Privacy, 22(3):67, May/June 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:IPCa

[Ano24j] Anonymous. IEEE Pervasive Computing — call for articles. IEEE Security & Privacy, 22(2):95, March/April 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:ITB

[Ano24k] Anonymous. IEEE Transactions on Big Data. IEEE Security & Privacy, 22(2):74, March/April 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:ITC

[Ano24l] Anonymous. IEEE Transactions on Computers. IEEE Security & Privacy, 22(2):7, March/April 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:ITS

[Ano24m] Anonymous. IEEE Transactions on Sustainable Computing. IEEE Security & Privacy, 22(2):80, March/April 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:FCa

[Ano24n]  Anonymous.  Front cover.

IEEE Security & Privacy, 22(1):C1, January/February 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:FCb

[Ano24o]  Anonymous.  Front cover.

IEEE Security & Privacy, 22(2):C1, March/April 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:FCc

[Ano24p]  Anonymous.  Front cover.

IEEE Security & Privacy, 22 (3):C1, May/June 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:ICCa

[Ano24q] Anonymous. IEEE Career Center. IEEE Security & Privacy, 22(1):C3, January/ February 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:ICCb

[Ano24r] Anonymous. IEEE Career Center. IEEE Security & Privacy, 22(3):C3, May/June 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:ICSf

[Ano24s]  Anonymous.  IEEE Com-

puter Society — call for papers. IEEE Security & Privacy, 22(2):37, March/April 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:ICSb

[Ano24t] Anonymous. IEEE Computer Society benefits. IEEE Security & Privacy, 22(1):9, January/ February 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:ICSl

[Ano24u] Anonymous. IEEE Computer Society benefits. IEEE Security & Privacy, 22(3):11, May/June 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:ICSc

[Ano24v] Anonymous. IEEE Computer Society call for papers. IEEE Security & Privacy, 22(1):42, January/February 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:ICSk

[Ano24w] Anonymous. IEEE Computer Society call for papers. IEEE Security & Privacy, 22(3):10, May/June 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:ICSi

[Ano24x] Anonymous. IEEE Computer Society Career Center. IEEE Security & Privacy, 22(2):C3, March/April 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:ICSa

[Ano24y] Anonymous. IEEE Computer Society diversity and inclusion fund. IEEE Security & Privacy, 22(1):C2, January/ February 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:ICSe

[Ano24z] Anonymous. IEEE Computer Society Diversity and Inclusion Fund. IEEE Security & Privacy, 22(2):C2, March/April 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:ICSj

[Ano24-27] Anonymous. IEEE Computer Society Diversity and Inclusion Fund. IEEE Security & Privacy, 22(3):C2, May/June 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:ICSg

[Ano24-28] Anonymous. IEEE Computer Society has you covered! IEEE Security & Privacy, 22(2):85, March/April 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:ICSd

[Ano24-29] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 22(1):73, January/February 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:ICSh

[Ano24-30] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 22(2):93, March/April 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:ICSn

[Ano24-31] Anonymous. IEEE Computer Society information. IEEE Security & Privacy, 22(3):87, May/June 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:IRP

[Ano24-32] Anonymous. IEEE Over the Rainbow podcast. IEEE Security & Privacy, 22(1):62, January/February 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:IRSa

[Ano24-33] Anonymous. IEEE Reliability Society. IEEE Security & Privacy, 22(1):5, January/ February 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:IRSb

[Ano24-34] Anonymous. IEEE Reliability Society. IEEE Security & Privacy, 22(2):5, March/April 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:Ma

[Ano24-35] Anonymous. Masthead. IEEE Security & Privacy, 22(1):3, January/February 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:Mb

[Ano24-36] Anonymous. Masthead. IEEE Security & Privacy, 22(2):3, March/April 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:Mc

[Ano24-37] Anonymous. Masthead. IEEE Security & Privacy, 22(3):3, May/June 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:PSE

[Ano24-38] Anonymous. Publications seek 2025 Editors in Chief. IEEE Security & Privacy, 22(1):63, January/February 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:RS

[Ano24-39] Anonymous. Reliability society. IEEE Security & Privacy, 22(3):5, May/June 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:TCa

[Ano24-40] Anonymous. Table of contents.

IEEE Security & Privacy, 22(1):1–2, January/February 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:TCb

[Ano24-41] Anonymous. Table of contents. IEEE Security & Privacy, 22(2):1–2, March/April 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anonymous:2024:TCc

[Ano24-42] Anonymous. Table of contents. IEEE Security & Privacy, 22 (3):1–2, May/June 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Apvrille:2005:SSD

[AP05] Axelle Apvrille and Makan Pourzandi. Secure software development by example. IEEE Security & Privacy, 3(4):10–17, July/August 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Applegate:2011:CPH

[App11] Scott D. Applegate. Cybermilitias and political hackers: Use of irregular forces in cyberwarfare. IEEE Security & Privacy, 9(5):16–22, September/ October 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Arp:2023:LLM

[AQP^{+}23] Daniel Arp, Erwin Quiring, Feargus Pendlebury, Alexander Warnecke, Fabio Pierazzi, Christian Wressnegger, Lorenzo Cavallaro, and Konrad Rieck. Lessons learned on machine learning for computer security. IEEE Security & Privacy, 21(5):72–77, September/ October 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Archer:2015:CDP

[AR15] David W. Archer and Kurt Rohloff. Computing with data privacy: Steps toward realization. IEEE Security & Privacy, 13(1):22–29, January/February 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/01/
msp2015010022-abs.html.

Ahmed:2018:PIT

[AR18] Irfan Ahmed and Vassil Roussev. Peer instruction teaching methodology for cybersecurity education. IEEE Security & Privacy, 16(4):88– 91, July/August 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/04/
msp2018040088-abs.html.

Arbuckle:2019:FSR

[AR19] L. Arbuckle and F. Ritchie.

The five safes of risk-based anonymization. IEEE Security & Privacy, 17(5):84– 89, September/October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Aiello:2024:SAS

[AR24] Samuel Aiello and Bhaskar P. Rimal. Secure access service edge convergence: Recent progress and open issues. IEEE Security & Privacy, 22(2):8–16, March/April 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Arbaugh:2004:GEI

[Arb04] William A. Arbaugh. Guest Editor’s introduction: Wired on wireless. IEEE Security & Privacy, 2(3):26–27, May/June 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/sp/2004/03/j3026.
                                        htm.

Arce:2003:RG

[Arc03a] Iván Arce. The rise of the gadgets. IEEE Security & Privacy, 1(5):78–81, September/ October 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/sp/2003/05/j5078abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2003/05/j5078.
                                        pdf.

Arce:2003:WLR

[Arc03b]  Iván Arce.  The weakest

link revisited. IEEE Security & Privacy, 1(2):72– 76, March/April 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j2072.pdf;   http://www.
computer.org/security/j2072abs.
htm.

Arce:2004:KC

[Arc04a]  Ivaán Arce.  The kernel

craze. IEEE Security & Privacy, 2(3):79–81, May/ June 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/03/j3079.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/03/j3079.
pdf.

Arce:2004:ATM

[Arc04b] Iván Arce. Attack trends:

More bang for the bug: An account of 2003’s attack trends. IEEE Security & Privacy, 2 (1):66–68, January/February 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2004/01/j1066abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/01/j1066.
pdf.

Arce:2004:SG

[Arc04c] Iván Arce. The shellcode generation. IEEE Security & Privacy, 2(5):72–76, September/October 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/05/j5072.

                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2004/05/j5072.
                                        pdf.

Arce:2005:BP

[Arc05a] I. Arce. Bad peripherals. IEEE Security & Privacy, 3 (1):70–73, January/February 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/30310/01392705.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        30310&arnumber=1392705&count=
                                        17&index=12.

Arce:2005:LB

[Arc05b] Iván Arce. The land of the blind. IEEE Security & Privacy, 3(4):63–67, July/August 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Arce:2006:VHV

[Arc06] Ivan Arce. Voices, I hear voices. IEEE Security & Privacy, 4(4): 80–83, July/August 2006. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Arce:2007:GVM

[Arc07a] Iván Arce. Ghost in the virtual machine. IEEE Security & Privacy, 5(4):68–71, July/ August 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Arce:2007:SPY

[Arc07b] Iván Arce. A surprise party (on your computer)? IEEE Security & Privacy, 5(2):15–16, March/April 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Al-Rubaie:2019:PPM

[ARC19] M. Al-Rubaie and J. M.

Chang. Privacy-preserving machine learning: Threats and solutions. IEEE Security & Privacy, 17(2):49–58, March/ April 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Arkin:2011:NWC

[Ark11] Brad Arkin. Never waste a crisis. IEEE Security & Privacy, 9(3):82–85, May/June 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Amin:2012:DTM

[ARvD12] Rohan M. Amin, Julie J. C. H.

Ryan, and Johan Rene van Dorp. Detecting targeted malicious email. IEEE Security & Privacy, 10(3):64–71, May/ June 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anderson:2005:GEI

[AS05] R. Anderson and B. Schneier.

Guest Editors’ introduction: Economics of information security. IEEE Security & Privacy, 3(1):12–13, January/February 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/30310/01392693.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
30310&arnumber=1392693&count=
17&index=2.

Avramopoulos:2009:PDR

[AS09] Ioannis Avramopoulos and Martin Suchara. Protecting the DNS from routing attacks: Two alternative anycast implementations. IEEE Security & Privacy, 7(5):14–20, September/October 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Alexander:2011:DT

[AS11] James Alexander and Jonathan Smith. Disinformation: a taxonomy. IEEE Security & Privacy, 9(1):58–63, January/ February 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Anderson:2013:MSN

[AS13] Jonathan Anderson and Frank Stajano. Must social networking conflict with privacy? IEEE Security & Privacy, 11 (3):51–60, May/June 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Avgerou:2015:PAD

[AS15] A. D. Avgerou and Y. C. Stamatiou. Privacy awareness diffusion in social networks. IEEE Security & Privacy, 13(6):44– 50, November/December 2015.

CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Alt:2022:BPC

[AS22] Florian Alt and Stefan Schneegass. Beyond passwords: Challenges and opportunities of future authentication. IEEE Security & Privacy, 20(1):82–86, January/February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Anthony:2015:BBI

[ASC15] Denise Anthony, Timothy Stablein, and Emily K. Carian. Big Brother in the information age: Concerns about government information gathering over time. IEEE Security & Privacy, 13(4):12– 19, July/August 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/04/
msp2015040012-abs.html.

Amit:2021:SHM

[ASE21] G. Amit, A. Shabtai, and Y. Elovici. A self-healing mechanism for Internet of Things devices. IEEE Security & Privacy, 19(1):44–53, January/ February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Ansari:2008:EEV

[ASH^{+}08] Nirwan Ansari, Pitipatana Sakarindr, Ehsan Haghani, Chao Zhang, Aridaman K. Jain, and Yun Q. Shi. Evaluating electronic voting systems equipped with voter-verified paper records. IEEE Security & Privacy, 6(3):30–39, May/ June 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Aciicmez:2007:MAC

[ASK07] Onur Aciiçmez, Jean-Pierre Seifert, and Çetin Kaya Koç. Micro-architectural cryptanalysis. IEEE Security & Privacy, 5(4):62–64, July/August 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Arkin:2005:SPT

[ASM05] B. Arkin, S. Stender, and G. McGraw. Software penetration testing. IEEE Security & Privacy, 3(1):84– 87, January/February 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/30310/01392709.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
30310&arnumber=1392709&count=
17&index=15.

Arora:2005:ESV

[AT05] A. Arora and R. Telang. Economics of software vulnerability disclosure. IEEE Security & Privacy, 3(1):20– 25, January/February 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/30310/01392695.pdf;

http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
30310&arnumber=1392695&count=
17&index=4.

Amoroso:2013:VCS

[AT13] Edward G. Amoroso and Hugh Thompson. A view from the C-Suite. IEEE Security & Privacy, 11(1):11–12, January/ February 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Aura:2006:WYS

[Aur06]  Tuomas Aura.  Why you

shouldn’t study security. IEEE Security & Privacy, 4(3):74–76, May/June 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Aranha:2018:GBU

[Av18] D. F. Aranha and J. van de Graaf. The good, the bad, and the ugly: Two decades of e-voting in Brazil. IEEE Security & Privacy, 16(6):22– 30, November/December 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Asghari:2015:EFB

[AvEB15] Hadi Asghari, Michel J. G.

van Eeten, and Johannes M. Bauer. Economics of fighting botnets: Lessons from a decade of mitigation. IEEE Security & Privacy, 13(5):16–23, September/October 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/csdl/mags/sp/2015/05/
msp2015050016-abs.html.

Andrews:2004:CS

[AW04] Mike Andrews and James A. Whittaker. Computer security. IEEE Security & Privacy, 2(5):68–71, September/October 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/05/j5068.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/05/j5068.
pdf.

Altman:2018:HRF

[AWV18] Micah Altman, Alexandra Wood, and Effy Vayena. A harm-reduction framework for algorithmic fairness. IEEE Security & Privacy, 16(3):34– 45, May/June 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/03/
msp2018030034-abs.html.

Ahmed:2020:COC

[AZ20] C. M. Ahmed and J. Zhou. Challenges and opportunities in cyberphysical systems security: a physics-based perspective. IEEE Security & Privacy, 18(6):14–22, November/ December 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Barcellos:2018:RSP

[BA18] M. Barcellos and D. F. Aranha.

Research in security and privacy in Brazil. IEEE Security & Privacy, 16(6):14– 21, November/December 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Baase:2003:GFS

[Baa03] Sara Baase. A Gift of Fire: Social, Legal, and Ethical Issues in Computing and the Internet. Pearson Education, Upper Saddle River, NJ, USA, second edition, 2003. ISBN 0-13-008215-5. xv + 464 pp. LCCN QA76.9.C66 B3 2003.

Babaguchi:2016:EPC

[Bab16] Noboru Babaguchi. Evaluating protection capability for visual privacy information. IEEE Security & Privacy, 14(1):55–61, January/February 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bace:2013:PME

[Bac13] Rebecca Bace. Pain management for entrepreneurs: Working with venture capital. IEEE Security & Privacy, 11(4):78– 81, July/August 2013. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bailey:2012:MMM

[Bai12a] Don A. Bailey. Moving 2 mishap: M2M’s impact on privacy and safety. IEEE Security & Privacy, 10(1):84–87, January/February 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Baize:2012:DSP

[Bai12b] Eric Baize. Developing secure products in the age of advanced persistent threats. IEEE Security & Privacy, 10(3):88–92, May/June 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Barthe:2015:HAC

[Bar15] Gilles Barthe. High-assurance cryptography: Cryptographic software we can trust. IEEE Security & Privacy, 13(5):86– 89, September/October 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/csdl/mags/sp/2015/05/
msp2015050086-abs.html.

Bayuk:2011:SSE

[Bay11] Jennifer L. Bayuk. Systems security engineering. IEEE Security & Privacy, 9(2):72–74, March/April 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Blanton:2017:ISE

[BB17] Marina Blanton and Fattaneh Bayatbabolghani. Improving the security and efficiency of private genomic computation using server aid. IEEE Security & Privacy, 15(5):20–28, September/October 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/05/
msp2017050020-abs.html.

Bellovin:2008:IAT

[BBB^{+}08] Steven M. Bellovin, Terry V.

Benzel, Bob Blakley, Dorothy E. Denning, Whitfield Diffie, Jeremy Epstein, and Paulo Veríssimo. Information Assurance Technology Forecast 2008. IEEE Security & Privacy, 6(1): 16–23, January/February 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Balliu:2023:CPS

[BBB^{+}23] Musard Balliu, Benoit Baudry, Sofia Bobadilla, Mathias Ekstedt, Martin Monperrus, Javier Ron, Aman Sharma, Gabriel Skoglund, César Soto-Valero, and Martin Wittlinger. Challenges of producing software bill of materials for Java. IEEE Security & Privacy, 21(6):12– 23, November/December 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Beguin:2019:CSO

[BBC^{+}19] E. Beguin, S. Besnard, A. Cros, B. Joannes, O. Leclerc-Istria, A. Noel, N. Roels, F. Taleb, J. Thongphan, E. Alata, and V. Nicomette. Computersecurity-oriented escape room. IEEE Security & Privacy, 17 (4):78–83, July/August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Bellovin:2013:GBW

[BBCL13] Steven M. Bellovin, Matt Blaze, Sandy Clark, and Susan Landau. Going bright: Wiretapping without weakening communications infrastructure. IEEE Security & Privacy, 11(1):62–72, January/ February 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bellovin:2008:RCS

[BBD^{+}08] Steven M. Bellovin, Matt Blaze, Whitfield Diffie, Susan Landau, Peter G. Neumann, and Jennifer Rexford. Risking communications security: Potential hazards of the Protect America Act. IEEE Security & Privacy, 6(1):24–33, January/ February 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bellatti:2017:DHD

[BBL^{+}17] Jacob Bellatti, Andrew Brunner, Joseph Lewis, Prasad Annadata, Wisam Eltarjaman, Rinku Dewri, and Ramakrishna Thurimella. Driving habits data: Location privacy implications and solutions. IEEE Security & Privacy, 15(1):12–20, January/February 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/01/
msp2017010012-abs.html.

Balliu:2019:SIA

[BBS19] M. Balliu, I. Bastys, and A. Sabelfeld. Securing IoT apps. IEEE Security & Privacy, 17(5):22–29, September/ October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Bonver:2008:BSD

[BC08] Edward Bonver and Michael Cohen. Building security in: Developing and retaining a security testing mindset. IEEE Security & Privacy, 6(5):82–85, September/October 2008. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bellovin:2009:CTS

[BC09] Steven M. Bellovin and Daniel G.

Conway. Clear text: Security as a systems property. IEEE Security & Privacy, 7(5):88, September/October 2009. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Balebako:2014:IAP

[BC14] Rebecca Balebako and Lorrie Cranor. Improving app privacy: Nudging app developers to protect user privacy. IEEE Security & Privacy, 12(4):55– 58, July/August 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/04/
msp2014040055-abs.html.

Blakley:2021:DPH

[BC21] Bob Blakley and Lorrie Cranor. A discussion of public health, trust, and privacy with Susan Landau. IEEE Security & Privacy, 19(5):9– 15, September/October 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Blakley:2022:HAT

[BC22] Bob Blakley and Lorrie Cranor. High assurance in the Twenty-First Century with Roger Schell. IEEE Security & Privacy, 20(3):12–21, May/ June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Blakley:2023:AEC

[BC23a] Bob Blakley and Lorrie Cranor. A 20th anniversary episode chat with SP editors George Cybenko, Carl Landwehr, Shari Lawrence Pfleeger, and Sean Peisert. IEEE Security & Privacy, 21 (2):9–16, March/April 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Blakley:2023:KMV

[BC23b] Bob Blakley and Lorrie Cranor. Katie Moussouris: Vulnerability disclosure and security workforce development. IEEE Security & Privacy, 21(1):11– 18, January/February 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Bellandi:2015:TEA

[BCD^{+}15] V. Bellandi, S. Cimato, E. Damiani, G. Gianini, and A. Zilli. Toward economicaware risk assessment on the cloud. IEEE Security & Privacy, 13(6):30–37, November/ December 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Basin:2022:TVL

[BCDS22] David Basin, Cas Cremers, Jannik Dreier, and Ralf Sasse. Tamarin: Verification of large-scale, real-world, cryptographic protocols. IEEE Security & Privacy, 20(3):24–32, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Balfanz:2012:FA

[BCE^{+}12] Dirk Balfanz, Richard Chow, Ori Eisen, Markus Jakobsson, Steve Kirsch, Scott Matsumoto, Jesus Molina, and Paul van Oorschot. The future of authentication. IEEE Security & Privacy, 10(1):22–27, January/February 2012. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Black:2006:SAI

[BCG06] John Black, Martin Cochran, and Ryan Gardner. A security analysis of the Internet Chess Club. IEEE Security & Privacy, 4(1):46– 52, January/February 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/33481/01588825.pdf;
 http://ieeexplore.ieee.

                                        org/xpls/abs_all.jsp?isnumber=
                                        33481&arnumber=1588825.

Bier:2009:RRI

[BCG^{+}09] Eric Bier, Richard Chow, Philippe Golle, Tracy Holloway King, and Jessica Staddon. The rules of redaction: Identify, protect, review (and repeat). IEEE Security & Privacy, 7(6):46–53, November/ December 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Beunardeau:2016:WBC

[BCGN16] Marc Beunardeau, Aisling Connolly, Remi Geraud, and David Naccache. White-box cryptography: Security in an insecure environment. IEEE Security & Privacy, 14(5):88– 92, September/October 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2016/05/
                                        msp2016050088-abs.html.

Bailey:2005:BWT

[BCJ^{+}05] Michael Bailey, Evan Cooke, Farnam Jahanian, David Watson, and Jose Nazario. The Blaster Worm: Then and now. IEEE Security & Privacy, 3(4): 26–31, July/August 2005. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bono:2009:RAS

[BCLM09] Stephen Bono, Dan Caselden, Gabriel Landau, and Charlie Miller. Reducing the attack surface in massively multiplayer online role-playing games. IEEE Security & Privacy, 7(3):13–19, May/June 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ball:2003:PPE

[BCM03] Edward Ball, David W. Chadwick, and Darren Mundy. Patient privacy in electronic prescription transfer. IEEE Security & Privacy, 1(2):77– 80, March/April 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j2077.pdf;   http://www.
computer.org/security/j2077abs.
htm.

Basin:2015:ISC

[BCM^{+}15a] David Basin, Cas Cremers, Kunihiko Miyazaki, Sasa Radomirovic, and Dai Watanabe. Improving the security of cryptographic protocol standards. IEEE Security & Privacy, 13(3):24– 31, May/June 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/03/
msp2015030024-abs.html.

Bond:2015:PEP

[BCM^{+}15b] Mike Bond, Marios O. Choudary, Steven J. Murdoch, Sergei Skorobogatov, and Ross Anderson. Be prepared: The EMV preplay attack. IEEE Security & Privacy, 13(2):56– 64, March/April 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/sp/2015/02/
                                        msp2015020056-abs.html.

Burton:2016:VVE

[BCS16] Craig Burton, Chris Culnane, and Steve Schneider. vVote: Verifiable electronic voting in practice. IEEE Security & Privacy, 14(4):64– 73, July/August 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2016/04/
                                        msp2016040064-abs.html.

Berson:2011:C

[BD11] Thomas A. Berson and Dorothy E. Denning. Cyberwarfare. IEEE Security & Privacy, 9(5):13–15, September/October 2011. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bryant:2018:HAC

[BD18] Timothy Bryant and Shaun Davenport. A honeybug for automated cyber reasoning systems. IEEE Security & Privacy, 16(2):32– 36, March/April 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/02/
                                        msp2018020032-abs.html.

Balfanz:2004:SUS

[BDGS04] Dirk Balfanz, Glenn Durfee, Rebecca E. Grinter, and D. K. Smetters. In search of usable security: Five lessons from the field. IEEE Security & Privacy, 2(5):19–24, September/October 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/05/j5019.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/05/j5019.
pdf.

Bailey:2012:MR

[BDKM12] Michael Bailey, David Dittrich, Erin Kenneally, and Doug Maughan. The Menlo Report. IEEE Security & Privacy, 10 (2):71–75, March/April 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bratus:2014:BPB

[BDL^{+}14] Sergey Bratus, Trey Darley, Michael Locasto, Meredith L. Patterson, Rebecca “bx” Shapiro, and Anna Shubina. Beyond planted bugs in ‘trusting trust’: The inputprocessing frontier. IEEE Security & Privacy, 12(1):83–87, January/February 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bradley:2017:GSL

[BDT17] Tatiana Bradley, Xuhua Ding, and Gene Tsudik. Genomic security (lest we forget). IEEE Security & Privacy, 15(5):38– 46, September/October 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/05/
                                        msp2017050038-abs.html.

Breidenbach:2019:HFP

[BDTJ19] L. Breidenbach, P. Daian, F. Tramer, and A. Juels. The Hydra framework for principled, automated bug bounties. IEEE Security & Privacy, 17 (4):53–61, July/August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Bellovin:2005:SPE

[Bel05] S. Bellovin. Security and privacy: Enemies or allies? IEEE Security & Privacy, 3(3):92, May/June 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/31002/01439512.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        31002&arnumber=1439512&count=
                                        20&index=19.

Bellia:2006:FAE

[Bel06a] Patricia L. Bellia. The Fourth Amendment and emerging communications technologies. IEEE Security & Privacy, 4(3): 20–28, May/June 2006. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bellovin:2006:UW

[Bel06b] Steve Bellovin. Unconventional wisdom. IEEE Security & Privacy, 4(1):88, January/February 2006. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/33481/01588836.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
33481&arnumber=1588836.

Bellovin:2006:BSI

[Bel06c] Steven M. Bellovin. On the brittleness of software and the infeasibility of security metrics. IEEE Security & Privacy, 4(4): 96, July/August 2006. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bellovin:2007:DCC

[Bel07a] Steve Bellovin. DRM, complexity, and correctness. IEEE Security & Privacy, 5(1):80, January/February 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bellovin:2007:SC

[Bel07b] Steve Bellovin. Seers and craftspeople. IEEE Security & Privacy, 5(5):88, September/October 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://bell.computer.org/
dlcomments/.

Bellovin:2008:SC

[Bel08a]  Steve Bellovin.  Security

by checklist. IEEE Security & Privacy, 6(2):88, March/April 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

                                        http://bell.computer.org/
                                        dlcomments/.

Bellovin:2008:CTP

[Bel08b] Steven M. Bellovin. Clear text: The puzzle of privacy. IEEE Security & Privacy, 6(5):88, September/October 2008. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bellovin:2009:CTG

[Bel09] Steven M. Bellovin. Clear text: The government and cybersecurity. IEEE Security & Privacy, 7(2):96, March/April 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bellovin:2010:IS

[Bel10a] Steven M. Bellovin. Identity and security. IEEE Security & Privacy, 8(2):88, March/April 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bellovin:2010:PR

[Bel10b] Steven M. Bellovin. Perceptions and reality. IEEE Security & Privacy, 8(5):88, 87, September/October 2010. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bellovin:2011:CBS

[Bel11a] Steven M. Bellovin. Clouds from both sides. IEEE Security & Privacy, 9(3):88, May/ June 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bellovin:2011:ST

[Bel11b] Steven M. Bellovin. Security think. IEEE Security & Privacy, 9(6):88, November/ December 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bellovin:2012:FLW

[Bel12a] Steven M. Bellovin. Fighting the last war. IEEE Security & Privacy, 10(3):96, May/June 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bellovin:2012:MCI

[Bel12b]  Steven M. Bellovin.  The

Major Cyberincident Investigations Board. IEEE Security & Privacy, 10(6):96, November/ December 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bellovin:2013:LWM

[Bel13a] Steven M. Bellovin. Last word: Military cybersomethings. IEEE Security & Privacy, 11(3):88, May/June 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bellovin:2013:LWW

[Bel13b] Steven M. Bellovin. The last word: Walls and gates. IEEE Security & Privacy, 11(6): 88, November/December 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bellovin:2014:MPH

[Bel14a]  Steven M. Bellovin.  By

any means possible: How intelligence agencies have gotten their data. IEEE Security & Privacy, 12(4):80– 84, July/August 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/04/
msp2014040080-abs.html.

Bellovin:2014:DS

[Bel14b]  Steven M. Bellovin.  Dr.

Strangecode. IEEE Security & Privacy, 12(3):88, May/June 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bellovin:2014:WSC

[Bel14c] Steven M. Bellovin. What should crypto look like? IEEE Security & Privacy, 12(6):108, November/December 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/06/
msp2014060108.html.

Bellovin:2015:KK

[Bel15a] Steven M. Bellovin. The key to the key. IEEE Security & Privacy, 13(6):96, November/ December 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bellovin:2015:WRC

[Bel15b] Steven M. Bellovin. What a real cybersecurity bill should address. IEEE Security & Privacy, 13(3):92, May/June 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//www.computer.org/csdl/
mags/sp/2015/03/msp2015030092.
html.

Bellovin:2016:AS

[Bel16a] Steven M. Bellovin. Attack surfaces. IEEE Security & Privacy, 14(3):88, May/June 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bellovin:2016:EEE

[Bel16b] Steven M. Bellovin. Easy email encryption. IEEE Security & Privacy, 14(6):96, November/December 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/06/
msp2016060096.html.

Bellovin:2017:JI

[Bel17a] Steven M. Bellovin. Jurisdiction and the Internet. IEEE Security & Privacy, 15(3): 96, May/June 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/03/
                                        msp2017030096.html.

Bellovin:2017:WY

[Bel17b] Steven M. Bellovin. Who are you? IEEE Security & Privacy, 15(6):88, November/December 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/06/
                                        msp2017060088.html.

Bellovin:2018:UCL

[Bel18a] S. M. Bellovin. Unnoticed consent [last word]. IEEE Security & Privacy, 16(6):80– 79, November/December 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Bellovin:2018:TNC

[Bel18b] Steven M. Bellovin. Toward a national cybersecurity policy. IEEE Security & Privacy, 16 (3):108, May/June 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/03/
                                        msp2018030108.html.

Bellovin:2019:LI

[Bel19] S. M. Bellovin. Layered insecurity. IEEE Security & Privacy, 17(3):96–95, May/June 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Bellovin:2020:PP

[Bel20a] S. M. Bellovin. Policies on privacy. IEEE Security & Privacy, 18(2):76, March/April 2020. ISSN 1558-4046.

Bellovin:2020:SPS

[Bel20b] S. M. Bellovin. Security, privacy, and scale. IEEE Security & Privacy, 18(6):63– 64, November/December 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Bellovin:2021:LLH

[Bel21] Steven M. Bellovin. The law and lawful hacking. IEEE Security & Privacy, 19(4):76, July/August 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Bellovin:2022:OST

[Bel22a] Steven M. Bellovin. Open source and trust. IEEE Security & Privacy, 20(2):107–108, March/April 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Bellovin:2022:WDW

[Bel22b] Steven M. Bellovin. What do we owe? IEEE Security & Privacy, 20(6):83–84, November/ December 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Bellovin:2023:CLL

[Bel23] Steven M. Bellovin. Is cybersecurity liability a liability? IEEE Security & Privacy, 21 (4):99–100, July/August 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Bellovin:2024:DA

[Bel24] Steven M. Bellovin. Degenerative AI? IEEE Security & Privacy, 22(3):88, May/June 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Benzel:2014:SMC

[Ben14] Terry Benzel. A symposium, a magazine, and a community [guest editorial]. IEEE Security & Privacy, 12(3):13–14, May/June 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Benzel:2015:ESL

[Ben15a] Terry Benzel. An enduring symposium for leading research in security and privacy. IEEE Security & Privacy, 13(2):12– 13, March/April 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/02/
msp2015020012.html.

Benzel:2015:SPC

[Ben15b] Terry Benzel. A strategic plan for cybersecurity research and development. IEEE Security & Privacy, 13(4):3– 5, July/August 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/04/
msp2015040003.html.

Benzel:2016:GCC

[Ben16a] Terry Benzel. The growth of a conference, a community, and an industry. IEEE Security & Privacy, 14(4):3– 5, July/August 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/04/
msp2016040003.html.

Benzel:2016:ISP

[Ben16b] Terry Benzel. The IEEE Security and Privacy symposium workshops. IEEE Security & Privacy, 14(2):12–14, March/ April 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Benzel:2017:SPI

[Ben17] Terry Benzel. Selected papers from the 2016 IEEE Symposium on Security and Privacy. IEEE Security & Privacy, 15 (2):11–13, March/April 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/02/
msp2017020011.html.

Benzel:2021:RIP

[Ben21] Terry Benzel. Research and industry partnerships in cybersecurity and privacy research: New frontiers or fueling the tech sector? IEEE Security & Privacy, 19(5):4–7, September/October 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Benzel:2023:SPR

[Ben23] Terry Benzel. Security and privacy research artifacts: Are we making progress? IEEE Security & Privacy, 21(1):4–6, January/February 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Bertino:2019:ITV

[Ber19] E. Bertino. It takes a village to secure cellular networks. IEEE Security & Privacy, 17(5):96– 95, September/October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Berger:2020:ACT

[Ber20a] K. M. Berger. Addressing cyber threats in biology. IEEE Security & Privacy, 18(3):58– 61, May/June 2020. ISSN 1558-4046.

Bertino:2020:QDT

[Ber20b] E. Bertino. The quest for data transparency. IEEE Security & Privacy, 18(3):67–68, May/ June 2020. ISSN 1558-4046.

Bertino:2021:AAI

[Ber21a] E. Bertino. Attacks on artificial intelligence [last word]. IEEE Security & Privacy, 19(1):103– 104, January/February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Bertino:2021:ZTA

[Ber21b] Elisa Bertino. Zero trust architecture: Does it help? IEEE Security & Privacy, 19(5):95– 96, September/October 2021.

ISSN 1540-7993 (print), 1558-4046 (electronic).

Bertino:2022:PPS

[Ber22] Elisa Bertino. The persistent problem of software insecurity. IEEE Security & Privacy, 20 (3):107–108, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Bertino:2023:PEI

[Ber23] Elisa Bertino. Privacy in the era of 5G, IoT, Big Data, and machine learning. IEEE Security & Privacy, 21(1):91–92, January/February 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Bezzi:2024:LLM

[Bez24] Michele Bezzi. Large language models and security. IEEE Security & Privacy, 22(2):60–68, March/April 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Bishop:2004:TRP

[BF04] Matt Bishop and Deb Frincke.

Teaching robust programming. IEEE Security & Privacy, 2(2):54–57, March/April 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2004/02/j2054abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/02/j2054.
pdf.

Bishop:2005:HEL

[BF05a] Matt Bishop and Deborah Frincke. A human endeavor: Lessons from Shakespeare and beyond. IEEE Security & Privacy, 3(4):49–51, July/August 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bishop:2005:TSP

[BF05b] Matt Bishop and Deborah A. Frincke. Teaching secure programming. IEEE Security & Privacy, 3(5):54–56, September/October 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bishop:2006:WOY

[BF06] Matt Bishop and Deborah A. Frincke. Who owns your computer? IEEE Security & Privacy, 4(2):61–63, March/April 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bishop:2007:ALO

[BF07] Matt Bishop and Deborah A. Frincke. Achieving learning objectives through E-voting case studies. IEEE Security & Privacy, 5(1):53–56, January/ February 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bishop:2008:EIA

[BF08] Matt Bishop and Deborah A. Frincke. Education: Information assurance education: a work in progress. IEEE Security & Privacy, 6(5):54–57, September/October 2008. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Batcheller:2017:BSB

[BFC^{+}17] Archer Batcheller, Summer Craze Fowler, Robert Cunningham, Dinara Doyle, Trent Jaeger, and Ulf Lindqvist. Building on the success of building security in. IEEE Security & Privacy, 15(4):85– 87, July/August 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/04/
msp2017040085.html.

Bhargavan:2016:MVP

[BFK16] Karthikeyan Bhargavan, Cedric Fournet, and Markulf Kohlweiss. miTLS: Verifying protocol implementations against realworld attacks. IEEE Security & Privacy, 14(6):18– 25, November/December 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/06/
msp2016060018-abs.html.

Benaloh:2021:VSR

[BFS^{+}21] Josh Benaloh, Kammi Foote, Philip B. Stark, Vanessa Teague, and Dan S. Wallach. VAULT-style risk-limiting audits and the Inyo County Pilot. IEEE Security & Privacy, 19(4):8–18, July/August 2021.

ISSN 1540-7993 (print), 1558-4046 (electronic).

Breaux:2013:WES

[BG13] Travis Breaux and David Gordon. What engineers should know about US security and privacy law. IEEE Security & Privacy, 11(3):72–76, May/ June 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Barreto:2014:CSP

[BGC^{+}14] Carlos Barreto, Jairo Giraldo, Alvaro A. Cardenas, Eduardo Mojica-Nava, and Nicanor Quijano. Control systems for the power grid and their resiliency to attacks. IEEE Security & Privacy, 12(6):15– 23, November/December 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/sp/2014/06/
                                        msp2014060015-abs.html.

Boudot:2022:SAI

[BGG^{+}22] Fabrice Boudot, Pierrick Gaudry, Aurore Guillevic, Nadia Heninger, Emmanuel Thomé, and Paul Zimmermann. The state of the art in integer factoring and breaking public-key cryptography. IEEE Security & Privacy, 20(2):80–86, March/ April 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Burbano:2023:OAR

[BGL^{+}23] Luis Burbano, Kunal Garg, Santiago J. Leudo, Alvaro A.

Cardenas, and Ricardo G. Sanfelice. Online attack recovery in cyberphysical systems. IEEE Security & Privacy, 21 (4):20–28, July/August 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Bruno:2015:APC

[BGLEP15] Clementina Bruno, Luca Guidi, Azahara Lorite-Espejo, and Daniela Pestonesi. Assessing a potential cyberattack on the Italian electric system. IEEE Security & Privacy, 13(5):42– 51, September/October 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/csdl/mags/sp/2015/05/
msp2015050042-abs.html.

Bloomfield:2006:IWG

[BGM^{+}06] Robin E. Bloomfield, Sofia Guerra, Ann Miller, Marcelo Masera, and Charles B. Weinstock. International working group on assurance cases (for security). IEEE Security & Privacy, 4(3):66–68, May/June 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Balfe:2008:TCC

[BGMP08] Shane Balfe, Eimear Gallery, Chris J. Mitchell, and Kenneth G. Paterson. Trusted computing: Challenges for trusted computing. IEEE Security & Privacy, 6(6):60– 66, November/December 2008.

CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Boneh:2019:HRT

[BGMP19] D. Boneh, A. J. Grotto, P. Mc-Daniel, and N. Papernot. How relevant is the Turing Test in the age of sophisbots? IEEE Security & Privacy, 17(6):64– 71, November/December 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Bauer:2014:PUF

[BH14] Todd Bauer and Jason Hamlet. Physical unclonable functions: A primer. IEEE Security & Privacy, 12(6):97–101, November/December 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/sp/2014/06/
                                        msp2014060097-abs.html.

Bates:2019:CDP

[BH19] A. Bates and W. U. Hassan. Can data provenance put an end to the data breach? IEEE Security & Privacy, 17(4):88– 93, July/August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Benzel:2022:SPI

[BH22] Terry Benzel and Thorsten Holz. Selected papers from the 2021 IEEE Symposium on Security and Privacy. IEEE Security & Privacy, 20(2):8–9, March/April 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Barth:2019:TES

[BHJM19] S. Barth, P. Hartel, M. Junger, and L. Montoya. Teaching empirical social-science research to cybersecurity students: The case of “Thinking like a thief”. IEEE Security & Privacy, 17(3):8–16, May/June 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Backendal:2024:EEE

[BHP24] Matilda Backendal, Miro Haller, and Kenny Paterson. End-to-end encrypted cloud storage. IEEE Security & Privacy, 22(2):69–74, March/ April 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Bramhall:2007:UCI

[BHRR07] Pete Bramhall, Marit Hansen, Kai Rannenberg, and Thomas Roessler. User-centric identity management: New trends in standardization and regulation. IEEE Security & Privacy, 5(4):84–87, July/August 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bilar:2010:DST

[Bil10] Daniel Bilar. Degradation and subversion through subsystem attacks. IEEE Security & Privacy, 8(4):70–73, July/August 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Binns:2018:WCP

[Bin18] Reuben Binns. What can political philosophy teach us about algorithmic fairness? IEEE Security & Privacy, 16(3):73– 80, May/June 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/03/
msp2018030073-abs.html.

Bishop:2003:WCS

[Bis03] Matt Bishop. What is computer security? IEEE Security & Privacy, 1(1):67–69, January/February 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j1067.pdf;   http://www.
computer.org/security/j1067abs.
htm.

Bishop:2007:APT

[Bis07] Matt Bishop. About penetration testing. IEEE Security & Privacy, 5(6):84–87, November/December 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bishop:2010:CSP

[Bis10a] Matt Bishop. A clinic for “Secure” programming. IEEE Security & Privacy, 8(2):54–56, March/April 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bishop:2010:TTT

[Bis10b]  Matt Bishop.  Technology,

training, and transformation. IEEE Security & Privacy, 8 (5):72–75, September/October 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bishop:2011:TSS

[Bis11] Matt Bishop. Teaching security stealthily. IEEE Security & Privacy, 9(2):69–71, March/April 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bisogni:2016:WNE

[Bis16] Fabio Bisogni. What’s new in the economics of cybersecurity? IEEE Security & Privacy, 14(3):11–13, May/June 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bugeja:2022:ESH

[BJD22] Joseph Bugeja, Andreas Jacobsson, and Paul Davidsson. The ethical smart home: Perspectives and guidelines. IEEE Security & Privacy, 20(1):72–80, January/February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Brouchier:2009:CCT

[BKMN09] Julien Brouchier, Tom Kean, Carol Marsh, and David Naccache. Crypto corner: Temperature attacks. IEEE Security & Privacy, 7(2):79–82, March/April 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Blythe:2013:CSG

[BKS13] Jim Blythe, Ross Koppel, and Sean W. Smith. Circumvention of security: Good users do bad things. IEEE Security & Privacy, 11(5):80–83, September/ October 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Blank:2018:PAR

[BKS18] Peter Blank, Sabrina Kirrane, and Sarah Spiekermann. Privacy-aware restricted areas for unmanned aerial systems. IEEE Security & Privacy, 16 (2):70–79, March/April 2018. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/02/
msp2018020070-abs.html.

Benzel:2013:CGDa

[BL13a] Terry V. Benzel and Steve Lipner. Crossing the great divide: Transferring security technology from research to the market. IEEE Security & Privacy, 11(2):12–13, March/ April 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bloomfield:2013:SCS

[BL13b] Robin Bloomfield and Jay Lala. Safety-critical systems: The next generation. IEEE Security & Privacy, 11(4):11–13, July/August 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Blaze:2003:RAM

[Bla03] Matt Blaze. Rights amplification in master-keyed mechanical locks. IEEE Security & Privacy, 1(2):24– 32, March/April 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j2024.pdf;   http://www.
computer.org/security/j2024abs.
htm.

Blaze:2009:SST

[Bla09] Matt Blaze. Secure systems:

Taking surveillance out of the shadows. IEEE Security & Privacy, 7(5):75–77, September/ October 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bravo-Lillo:2011:BGC

[BLCDK11] Cristian Bravo-Lillo, Lorrie Faith Cranor, Julie Downs, and Saranga Komanduri. Bridging the gap in computer security warnings: a mental model approach. IEEE Security & Privacy, 9(2):18–26, March/ April 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bashir:2015:CCH

[BLG^{+}15] Masooda Bashir, April Lambert, Boyi Guo, Nasir Memon, and Tzipora Halevi. Cybersecurity competitions: The human angle. IEEE Security & Privacy, 13(5):74–79, September/October 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://csdl.computer.
                                        org/csdl/mags/sp/2015/05/
                                        msp2015050074-abs.html.

Buchmann:2017:PCS

[BLM17] Johannes Buchmann, Kristin Lauter, and Michele Mosca. Postquantum cryptography — state of the art. IEEE Security & Privacy, 15(4):12– 13, July/August 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/04/
                                        msp2017040012.html.

Buchmann:2018:PCP

[BLM18] J. Buchmann, K. Lauter, and M. Mosca. Postquantum cryptography, part 2. IEEE Security & Privacy, 16(5):12– 13, September/October 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Bloomfield:2011:RU

[Blo11] Robin E. Bloomfield. Resilient to the unexpected. IEEE Security & Privacy, 9(3):3–4, May/ June 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bloomfield:2012:TGW

[Blo12] Robin Bloomfield. Are things getting worse? IEEE Security & Privacy, 10(5):3, September/October 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bloomfield:2013:OA

[Blo13] Robin Bloomfield. Open assurance. IEEE Security & Privacy, 11(5):3–4, September/ October 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bloomfield:2015:ARD

[Blo15] Robin E. Bloomfield. Autonomy, robotics, and dependability. IEEE Security & Privacy, 13(5):3–5, September/October 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/csdl/mags/sp/2015/05/
msp2015050003.html.

Bartolini:2019:DPR

[BLR19] C. Bartolini, G. Lenzini, and L. Robaldo. The DAta Protection REgulation COmpliance model. IEEE Security & Privacy, 17(6):37–45, November/ December 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Blumenthal:2010:HSC

[Blu10] Marjory S. Blumenthal. Hide and seek in the cloud. IEEE Security & Privacy, 8(2):57–58, March/April 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Barnum:2005:KSS

[BM05] S. Barnum and G. McGraw. Knowledge for software security. IEEE Security & Privacy, 3(2):74–78, March/April 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/30742/01423967.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
30742&arnumber=1423967&count=
15&index=13.

Bohme:2010:IWL

[BM10] Rainer Bohme and Tyler Moore. The iterated weakest link. IEEE Security & Privacy, 8(1):53–55, January/February 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bau:2011:SMA

[BM11] Jason Bau and John C. Mitchell. Security modeling and analysis. IEEE Security & Privacy, 9(3):18–25, May/ June 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Broz:2014:TDF

[BM14] Milan Broz and Vashek Matyas. The TrueCrypt on-disk format

— an independent view. IEEE Security & Privacy, 12(3):74– 77, May/June 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bruschi:2007:CNS

[BMM07] Danilo Bruschi, Lorenzo Martignoni, and Mattia Monga. Code normalization for selfmutating malware. IEEE Security & Privacy, 5(2):46–54, March/April 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Butler:2010:NSA

[BMMM10] Kevin Butler, Steve McLaughlin, Thomas Moyer, and Patrick McDaniel. New security architectures based on emerging disk functionality. IEEE Security & Privacy, 8 (5):34–41, September/October 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bratus:2008:WDS

[BMS08] Sergey Bratus, Chris Masone, and Sean W. Smith. Why do street-smart people do stupid things online? IEEE Security & Privacy, 6(3):71–74, May/June 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bhatt:2014:ORS

[BMZ14] Sandeep Bhatt, Pratyusa K.

Manadhata, and Loai Zomlot. The operational role of security information and event management systems. IEEE Security & Privacy, 12(5):35–41, September/October 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/sp/2014/05/
                                        msp2014050035-abs.html.

Brandle:2008:PCS

[BN08] Markus Brändle and Martin Naedele. Process control security: Security for process control systems: An overview. IEEE Security & Privacy, 6(6): 24–29, November/December 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

BenNetanel:2021:DSD

[BNSE21] R. Ben Netanel, B. Nassi, A. Shamir, and Y. Elovici. Detecting spying drones. IEEE Security & Privacy, 19(1):65– 73, January/February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Bonebrake:2014:AGT

[BO14] Chris Bonebrake and Lori Ross O’Neil. Attacks on GPS time reliability. IEEE Security & Privacy, 12(3):82–84, May/June 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Baracaldo:2022:MLS

[BO22] Nathalie Baracaldo and Alina Oprea. Machine learning security and privacy. IEEE Security & Privacy, 20(5):11– 13, September/October 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Benzel:2023:ISP

[BO23] Terry Benzel and Hilarie Orman. IEEE Security and Privacy symposium in the year 2003. IEEE Security & Privacy, 21(2):43–47, March/ April 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Bonver:2008:STI

[Bon08] Edward Bonver. Security testing of internal tools. IEEE Security & Privacy, 6(1):81–83, January/February 2008. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Boneh:2016:SDA

[Bon16] Dan Boneh. Stickler: Defending against malicious content distribution networks in an unmodified browser. IEEE Security & Privacy, 14(2):22–28, March/April 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Borg:2005:ECC

[Bor05]  Scott Borg.  Economically

complex cyberattacks. IEEE Security & Privacy, 3(6): 64–67, November/December 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/33104/01556539.pdf;
http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
33104&arnumber=1556539.

Benzel:2013:CGDb

[BOR^{+}13a] Terry V. Benzel, Eric O’Brien, Robert Rodriguez, William Arbaugh, and John Sebes. Crossing the Great Divide: From research to market. IEEE Security & Privacy, 11(2):42–46, March/April 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Borgesius:2013:BTE

[Bor13b] Frederik Zuiderveen Borgesius. Behavioral targeting: A European legal perspective. IEEE Security & Privacy, 11(1):82– 85, January/February 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Borgesius:2015:ICW

[Bor15] Frederik Zuiderveen Borgesius. Informed consent: We can do better to defend privacy. IEEE Security & Privacy, 13(2):103– 107, March/April 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/02/
msp2015020103-abs.html.

Bos:2016:BRA

[Bos16] Herbert Bos. Binary rejuvenation: Applications and challenges. IEEE Security & Privacy, 14(1):68–71, January/ February 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Boyce:2016:BOT

[Boy16] Griffin Boyce. Bake in .onion for tear-free and stronger Website authentication. IEEE Security & Privacy, 14(2):15–21, March/April 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Baer:2007:CIS

[BP07] Walter S. Baer and Andrew Parkinson. Cyberinsurance in IT security management. IEEE Security & Privacy, 5(3):50–56, May/June 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Batten:2008:TDF

[BP08] Lynn Batten and Lei Pan.

Teaching digital forensics to undergraduate students. IEEE Security & Privacy, 6(3):54–56, May/June 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bishop:2012:SE

[BP12] Matt Bishop and Sean Peisert.

Security and elections. IEEE Security & Privacy, 10(5):64– 67, September/October 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Benzel:2018:SPI

[BP18] Terry Benzel and Sean Peisert. Selected papers from the 2017 IEEE Symposium on Security and Privacy. IEEE Security & Privacy, 16(1):10–11, January/February 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/01/
msp2018010010.html.

Bakken:2004:DOA

[BPB^{+}04] David E. Bakken, Rupa Parameswaran, Douglas M. Blough, Andy A. Franz, and Ty J. Palmer. Data obfuscation: Anonymity and desensitization of usable data sets. IEEE Security & Privacy, 2(6):34–41, November/ December 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/06/j6034.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/06/j6034.
pdf.

Bodden:2024:EST

[BPFG24] Eric Bodden, Jens Pottebaum, Markus Fockel, and Iris Gräßler. Evaluating security through isolation and defense in depth. IEEE Security & Privacy, 22(1):69–72, January/ February 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Bannet:2004:VHV

[BPR^{+}04] Jonathan Bannet, David W. Price, Algis Rudys, Justin Singer, and Dan S. Wallach. E-voting: Hack-a-vote: Security issues with electronic voting systems. IEEE Security & Privacy, 2(1):32–37, January/ February 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2004/01/j1032abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/01/j1032.
pdf.

Boneh:2016:BCR

[BPS16] Dan Boneh, Kenny Paterson, and Nigel P. Smart. Building a community of real-world cryptographers. IEEE Security & Privacy, 14(6):7–9, November/December 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/06/
msp2016060007.html.

Branch:2004:AWL

[BPVS04] Joel W. Branch, Nick L.

Petroni, Jr., Leendert Van Doorn, and David Safford. Autonomic 802.11 wireless LAN security auditing. IEEE Security & Privacy, 2(3):56–65, May/June 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/03/j3056.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/03/j3056.
pdf.

Bradner:2006:EEE

[Bra06] Scott Bradner. The end of endto-end security? IEEE Security & Privacy, 4(2):76–79, March/April 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bratus:2007:WHL

[Bra07] Sergey Bratus. What hackers learn that the rest of us don’t: Notes on hacker curriculum. IEEE Security & Privacy, 5(4):72–75, July/August 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Brennan:2012:AIF

[Bre12] Michael Brennan. Academic impact at the Federal Trade Commission. IEEE Security & Privacy, 10(6):78–82, November/December 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Boechat:2021:VRC

[BRS^{+}21] François Boechat, Gabriel Ribas, Lucas Senos, Miguel Bicudo, Mateus Schulz Nogueira, Leandro Pfleger de Aguiar, and Daniel Sadoc Menasche. Is vulnerability report confidence redundant? Pitfalls using temporal risk scores. IEEE Security & Privacy, 19(4):44–53, July/ August 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Benaloh:2017:VC

[BRST17] Josh Benaloh, Peter Y. A. Ryan, Steve Schneider, and Vanessa Teague. A vote of confidence? IEEE Security & Privacy, 15(3):12– 13, May/June 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/03/
msp2017030012.html.

Birman:2009:IMM

[BS09] Kenneth P. Birman and Fred B. Schneider. IT monoculture: The monoculture risk put into context. IEEE Security & Privacy, 7(1):14–17, January/ February 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bikos:2013:LSS

[BS13a] Anastasios N. Bikos and Nicolas Sklavos. LTE/SAE security issues on 4G wireless networks. IEEE Security & Privacy, 11 (2):55–62, March/April 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Birrell:2013:FIM

[BS13b] Eleanor Birrell and Fred B.

Schneider. Federated identity management systems: A privacy-based characterization. IEEE Security & Privacy, 11 (5):36–48, September/October 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bratus:2013:AWU

[BS13c] Sergey Bratus and Anna Shubina. Avoiding a war on unauthorized computation. IEEE Security & Privacy, 11(2):83– 88, March/April 2013. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Beuhring:2014:BBC

[BS14] Aaron Beuhring and Kyle Salous. Beyond blacklisting: Cyberdefense in the era of advanced persistent threats. IEEE Security & Privacy, 12(5):90–93, September/October 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/05/
msp2014050090-abs.html.

Benzel:2020:IES

[BS20a] T. Benzel and F. Stajano. IEEE Euro S&P: The younger sibling across the pond following in Oakland’s footsteps. IEEE Security & Privacy, 18 (3):6–7, May/June 2020. ISSN 1558-4046.

Borum:2020:PAC

[BS20b] R. Borum and R. Sanders. Preparing America’s cyber intelligence workforce. IEEE Security & Privacy, 18(5):67– 73, September/October 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Brehm:2024:UPV

[BS24] Karoline Brehm and Yan Shvartzshnaider. Understanding privacy in virtual reality classrooms: a contextual integrity perspective. IEEE Security & Privacy, 22(1):53–62, January/February 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Bessani:2008:CWC

[BSC^{+}08] Alysson Neves Bessani, Paulo Sousa, Miguel Correia, Nuno Ferreira Neves, and Paulo Veríssimo. The crutial way of critical infrastructure protection. IEEE Security & Privacy, 6(6):44– 51, November/December 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bowen:2009:DHN

[BSH^{+}09] Brian Bowen, Malek Ben Salem, Shlomo Hershkop, Angelos Keromytis, and Salvatore Stolfo. Designing host and network sensors to mitigate the insider threat. IEEE Security & Privacy, 7(6):22–29, November/December 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bejtlich:2011:DID

[BSP11] Richard Bejtlich, John Steven, and Gunnar Peterson. Directions in incident detection and response. IEEE Security & Privacy, 9(1):91–92, January/ February 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Barnum:2010:RRE

[BSS10] Sean Barnum, Shankar Sastry, and John A. Stankovic. Roundtable: Reliability of embedded and cyber-physical systems. IEEE Security & Privacy, 8(5):27–32, September/ October 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bhargav-Spantzel:2007:TNI

[BSSB07] Abhilasha Bhargav-Spantzel, Anna C. Squicciarini, and Elisa Bertino. Trust negotiation in identity management. IEEE Security & Privacy, 5(2):55–63, March/April 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Burr:2003:SAE

[Bur03]  William E. Burr.  Select-

ing the Advanced Encryption Standard. IEEE Security & Privacy, 1(2):43– 52, March/April 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://dlib.computer.
                                        org/sp/books/sp2003/pdf/
                                        j2043.pdf;   http://www.
                                        computer.org/security/j2043abs.
                                        htm.

Burr:2006:CHS

[Bur06] William E. Burr. Cryptographic hash standards: Where do we go from here? IEEE Security & Privacy, 4(2):88–91, March/April 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Burr:2008:NHC

[Bur08] William E. Burr. A new hash competition. IEEE Security & Privacy, 6(3):60–62, May/June 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Butin:2017:HBS

[But17] Denis Butin. Hash-based signatures: State of play. IEEE Security & Privacy, 15(4):37– 43, July/August 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/04/
msp2017040037-abs.html.

Brooks:2006:KAN

[BV06] Richard R. Brooks and Christopher Vutsinas. Kafka in the academy: a note on ethics in IA education. IEEE Security & Privacy, 4(4):50–53, July/ August 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Barrera:2011:SSI

[BV11a] David Barrera and Paul Van Oorschot. Secure software installation on smartphones. IEEE Security & Privacy, 9(3): 42–48, May/June 2011. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Beyah:2011:RAP

[BV11b] Raheem Beyah and Aravind Venkataraman. Rogue-access-point detection: Challenges, solutions, and future directions. IEEE Security & Privacy, 9(5):56–61, September/ October 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Boureanu:2015:CDB

[BV15] Ioana Boureanu and Serge Vaudenay. Challenges in distance bounding. IEEE Security & Privacy, 13(1):41–48, January/February 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/01/
msp2015010041-abs.html.

Bringhenti:2022:TCP

[BVB22] Daniele Bringhenti, Fulvio Valenza, and Cataldo Basile. Toward cybersecurity personalization in smart homes. IEEE Security & Privacy, 20(1):45– 53, January/February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Baker:2007:ISU

[BW07] Wade H. Baker and Linda Wallace. Is information security under control?: Investigating quality in information security management. IEEE Security & Privacy, 5(1):36–44, January/ February 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Bruening:2010:DTN

[BW10] Paula J. Bruening and K. Krasnow Waterman. Data tagging for new information governance models. IEEE Security & Privacy, 8(5):64–68, September/ October 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Buckland:2012:FVA

[BW12] Richard Buckland and Roland Wen. The future of E-voting in Australia. IEEE Security & Privacy, 10(5):25–32, September/October 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Byers:2004:ILC

[Bye04]  Simon Byers.  Informa-

tion leakage caused by hidden data in published documents. IEEE Security & Privacy, 2(2):23–27, March/ April 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2004/02/j2023abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/02/j2023.
pdf.

Bhattacharya:2014:LMS

[BYG^{+}14] Prabir Bhattacharya, Li Yang, Minzhe Guo, Kai Qian, and Ming Yang. Learning mobile security with Labware. IEEE Security & Privacy, 12(1):69– 72, January/February 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Conti:2005:FCD

[CA05] Gregory Conti and Mustaque Ahamad. A framework for countering denial-of-information attacks. IEEE Security & Privacy, 3(6): 50–56, November/December 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/33104/01556536.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        33104&arnumber=1556536.

Curphey:2006:WAS

[CA06] Mark Curphey and Rudolph Araujo. Web application security assessment tools. IEEE Security & Privacy, 4(4):32–41, July/August 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Chess:2011:SSP

[CA11] Brian Chess and Brad Arkin. Software security in practice. IEEE Security & Privacy, 9(2): 89–92, March/April 2011. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Cofer:2022:CSE

[CAB^{+}22] Darren Cofer, Isaac Amundson, Junaid Babar, David Hardin, Konrad Slind, Perry Alexander, John Hatcliff, Robby, Gerwin Klein, Corey Lewis, Eric Mercer, and John Shackleton. Cyberassured systems engineering at scale. IEEE Security & Privacy, 20 (3):52–64, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Caloyannides:2003:KOC

[Cal03a] Michael Caloyannides. Keeping offline computer usage private. IEEE Security & Privacy, 1(5):93–95, September/ October 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2003/05/j5093abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/05/j5093.
pdf.

Caloyannides:2003:PVI

[Cal03b] Michael Caloyannides. Privacy vs. information technology. IEEE Security & Privacy, 1(1):100–103, January/February 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j1100.pdf;   http://www.
computer.org/security/j1100abs.
htm.

Caloyannides:2003:DER

[Cal03c] Michael A. Caloyannides. Digital ‘evidence’ and reasonable doubt. IEEE Security & Privacy, 1(6):89–91, November/ December 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2003/06/j6089abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/06/j6089.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/06/j6089.
pdf.

Caloyannides:2003:ESC

[Cal03d] Michael A. Caloyannides. Engineering or sloganeering? the counterattack on privacy. IEEE Security & Privacy, 1(2): 84–87, March/April 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://dlib.computer.
                                        org/sp/books/sp2003/pdf/
                                        j2084.pdf;   http://www.
                                        computer.org/security/j2084abs.
                                        htm.

Caloyannides:2003:SCF

[Cal03e] Michael A. Caloyannides. Society cannot function without privacy. IEEE Security & Privacy, 1(3):84–86, May/June 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://dlib.computer.
                                        org/sp/books/sp2003/pdf/
                                        j3084.pdf.

Caloyannides:2004:ESC

[Cal04a] Michael Caloyannides. Enhancing security: Not for the conformist. IEEE Security & Privacy, 2(6):88, 86– 87, November/December 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/sp/
                                        2004/06/j6088.htm;   http:
                                        //csdl.computer.org/dl/mags/
                                        sp/2004/06/j6088.pdf.

Caloyannides:2004:CCF

[Cal04b] Michael A. Caloyannides. The cost of convenience: a Faustian deal. IEEE Security & Privacy, 2(2):84–87, March/ April 2004. CODEN ????

ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2004/02/j2084abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/02/j2084.
pdf.

Caloyannides:2004:PRC

[Cal04c] Michael A. Caloyannides. Is privacy really constraining security or is this a red herring? IEEE Security & Privacy, 2(4):86–87, July/ August 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/04/j4086.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/04/j4086.
pdf.

Caloyannides:2004:PMO

[Cal04d] Michael A. Caloyannides. Privacy matters: Online monitoring: Security or social control? IEEE Security & Privacy, 2(1):81–83, January/ February 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2004/01/j1081abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/01/j1081.
pdf.

Caloyannides:2004:SC

[Cal04e] Michael A. Caloyannides. Security or cosmetology? IEEE Security & Privacy, 2(3):93– 95, May/June 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/sp/2004/03/j3093.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2004/03/j3093.
                                        pdf.

Caloyannides:2004:SPT

[Cal04f] Michael A. Caloyannides. Speech privacy technophobes need not apply. IEEE Security & Privacy, 2(5):86– 87, September/October 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/sp/
                                        2004/05/j5086.htm;   http:
                                        //csdl.computer.org/dl/mags/
                                        sp/2004/05/j5086.pdf.

Caloyannides:2009:FY

[Cal09] Michael A. Caloyannides. Forensics is so “Yesterday”. IEEE Security & Privacy, 7(2): 18–25, March/April 2009. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Calo:2013:TSM

[Cal13] Ryan Calo. Tiny salespeople: Mediated transactions and the Internet of Things. IEEE Security & Privacy, 11(5):70–72, September/October 2013. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Cayir:2024:ASP

[CAL^{+}24] Derin Cayir, Abbas Acar, Riccardo Lazzeretti, Marco Angelini, Mauro Conti, and Selcuk Uluagac. Augmenting security and privacy in the virtual realm: an analysis of extended reality devices. IEEE Security & Privacy, 22(1):10– 23, January/February 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Camp:2003:AD

[Cam03a] Jean Camp. Access denied.

IEEE Security & Privacy, 1 (5):82–85, September/October 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2003/05/j5082abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/05/j5082.
pdf.

Campbell:2003:CSI

[Cam03b] Robert J. Campbell. Crime scene investigators: The next generation. IEEE Security & Privacy, 1(5):15, September/October 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/comp/mags/sp/2003/05/
j5015.pdf;   http://csdl.
computer.org/comp/mags/sp/
2003/05/j5015abs.htm.

Campbell:2005:DSD

[Cam05]  Philip L. Campbell.  The

denial-of-service dance. IEEE Security & Privacy, 3(6): 34–40, November/December 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/33104/01556534.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        33104&arnumber=1556534.

Camp:2010:IMM

[Cam10] Jean Camp. Identity management’s misaligned incentives. IEEE Security & Privacy, 8(6):90–94, November/ December 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Candea:2014:TVP

[Can14] George Candea. The testsversus-proofs conundrum. IEEE Security & Privacy, 12(1):65– 68, January/February 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Carvalho:2008:SMA

[Car08] Marco Carvalho. Security in mobile ad hoc networks. IEEE Security & Privacy, 6(2):72–75, March/April 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Carrier:2009:DFW

[Car09] Brian D. Carrier. Digital forensics works. IEEE Security & Privacy, 7(2):26–29, March/ April 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Cate:2009:PIS

[Cat09] Fred H. Cate. Privacy interests: Security, privacy, and the role of law. IEEE Security & Privacy, 7(5):60–63, September/ October 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Cate:2010:GAP

[Cat10a] Fred H. Cate. Government access to private-sector data. IEEE Security & Privacy, 8(6): 68–71, November/December 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Cate:2010:LNC

[Cat10b] Fred H. Cate. The limits of notice and choice. IEEE Security & Privacy, 8(2):59–62, March/April 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Cate:2011:TCP

[Cat11] Fred H. Cate. A transatlantic convergence on privacy? IEEE Security & Privacy, 9(1):76–79, January/February 2011. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Cranor:2014:BTU

[CB14] Lorrie Faith Cranor and Norbou Buchler. Better together: Usability and security go hand in hand. IEEE Security & Privacy, 12(6):89–93, November/December 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/06/
msp2014060089-abs.html.

Crain:2015:BSE

[CB15] J. Adam Crain and Sergey Bratus. Bolt-on security extensions for industrial control system protocols: A case study of DNP3 SAv5. IEEE Security & Privacy, 13(3):74– 79, May/June 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/sp/2015/03/
                                        msp2015030074-abs.html.

Chess:2012:SAMb

[CBE^{+}12] Brian Chess, Kris Britton, Chris Eng, Bill Pugh, Lakshmikanth Raghavan, and Jacob West. Static analysis in motion. IEEE Security & Privacy, 10(3):53–56, May/June 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Conti:2011:HCT

[CBN11] Gregory Conti, Thomas Babbitt, and John Nelson. Hacking competitions and their untapped potential for security education. IEEE Security & Privacy, 9(3):56–59, May/June 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Carlin:2020:YCM

[CBOS20] D. Carlin, J. Burgess, P. O’Kane, and S. Sezer. You could be mine(d): The rise of cryptojacking. IEEE Security & Privacy, 18(2):16–22, March/ April 2020. ISSN 1558-4046.

Cachin:2010:EKS

[CC10] Christian Cachin and Jan Camenisch. Encrypting keys securely. IEEE Security & Privacy, 8(4):66–69, July/August 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Conti:2011:EKM

[CC11] Gregory Conti and James Caroland. Embracing the Kobayashi Maru: Why you should teach your students to cheat. IEEE Security & Privacy, 9(4):48–51, July/August 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Clark:2010:NPI

[CCF^{+}10] Paul C. Clark, Glenn R. Cook, Edward L. Fisher, John D. Fulp, Valerie Linhoff, and Cynthia E. Irvine. New pathways in identity management. IEEE Security & Privacy, 8(6):64– 67, November/December 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Cybenko:2014:ISP

[CCF14] George Cybenko and Kathy Clark-Fisher. IEEE Security & Privacy: The early years. IEEE Security & Privacy, 12(3):18– 19, May/June 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Calzavara:2020:MLW

[CCF^{+}20] S. Calzavara, M. Conti, R. Focardi, A. Rabitti, and G. Tolomei. Machine learning for web vulnerability detection: The case of cross-site request forgery. IEEE Security & Privacy, 18(3):8–16, May/June 2020. ISSN 1558-4046.

Cunningham:2021:RRM

[CCLS21] Robert Cunningham, Anita D. Carleton, Tom Longstaff, and Forrest J. Shull. A research road map for building secure and resilient software-intensive systems. IEEE Security & Privacy, 19(6):8–14, November/ December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Callegati:2009:BTM

[CCR09] Franco Callegati, Walter Cerroni, and Marco Ramilli. Basic training: Man-in-the-middle attack to the HTTPS protocol. IEEE Security & Privacy, 7(1): 78–81, January/February 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Chandra:2006:PTW

[CCS06] Pravir Chandra, Brian Chess, and John Steven. Putting the tools to work: How to succeed with source code analysis. IEEE Security & Privacy, 4(3): 80–83, May/June 2006. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Capek:2003:MCE

[CCW03] Peter G. Capek, David M.

Chess, and Steve R. White. Merry Christma: An early network worm. IEEE Security & Privacy, 1(5):26– 34, September/October 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2003/05/j5026abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/05/j5026.
pdf.

Chen:2020:SSI

[CCX^{+}20] G. Chen, S. Chen, Y. Xiao, Y. Zhang, Z. Lin, and T. Lai. SgxPectre: Stealing Intel secrets from SGX enclaves via speculative execution. IEEE Security & Privacy, 18(3):28– 37, May/June 2020. ISSN 1558-4046.

Coppens:2013:PYS

[CDD13] Bart Coppens, Bjorn De Sutter, and Koen De Bosschere. Protecting your software updates. IEEE Security & Privacy, 11(2):47–54, March/ April 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Cavoukian:2012:PFC

[CDF^{+}12] Ann Cavoukian, Alan Davidson, Ed Felton, Marit Hansen, Susan Landau, and Anna Slomovic. Privacy: Front and center. IEEE Security & Privacy, 10(5):10–15, September/ October 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Carvalho:2014:H

[CDFW14] Marco Carvalho, Jared De-Mott, Richard Ford, and David A. Wheeler. Heartbleed 101. IEEE Security & Privacy, 12(4):63– 67, July/August 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/sp/2014/04/
                                        msp2014040063-abs.html.

Corbett:2024:SBP

[CDJS^{+}24] Matthew Corbett, Brendan David-John, Jiacheng Shang, Y. Charlie Hu, and Bo Ji. Securing bystander privacy in mixed reality while protecting the user experience. IEEE Security & Privacy, 22(1):33–42, January/February 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Chakrabarti:2008:GCS

[CDS08] Anirban Chakrabarti, Anish Damodaran, and Shubhashis Sengupta. Grid computing security: a taxonomy. IEEE Security & Privacy, 6(1):44–51, January/February 2008. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Chaum:2008:SEE

[CEC^{+}08] David Chaum, Aleks Essex, Richard Carback, Jeremy Clark, Stefan Popoveniuc, Alan Sherman, and Poorvi Vora. Scantegrity: End-to-end voter-verifiable opticalscan voting. IEEE Security & Privacy, 6(3):40–46, May/June 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Calero:2010:TMT

[CEK^{+}10] Jose M. Alcaraz Calero, Nigel Edwards, Johannes Kirschnick, Lawrence Wilcock, and Mike Wray. Toward a multi-tenancy authorization system for cloud services. IEEE Security & Privacy, 8(6):48–55, November/ December 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Celik:2016:MLA

[Cel16] Z. Berkay Celik. Machine learning in adversarial settings. IEEE Security & Privacy, 14(3):68–72, May/June 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Culnane:2019:KKR

[CEL^{+}19] C. Culnane, A. Essex, S. J.

Lewis, O. Pereira, and V. Teague. Knights and knaves run elections: Internet voting and undetectable electoral fraud. IEEE Security & Privacy, 17 (4):62–70, July/August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Carvalho:2014:MTD

[CF14] Marco Carvalho and Richard Ford. Moving-target defenses for computer networks. IEEE Security & Privacy, 12(2):73– 76, March/April 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Carminati:2005:EPI

[CFH05] Barbara Carminati, Elena Ferrari, and Patrick C. K. Hung. Exploring privacy issues in Web services discovery agencies. IEEE Security & Privacy, 3(5):14–21, September/ October 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Cioranesco:2013:CCC

[CFN13] Jean-Michel Cioranesco, Houda Ferradi, and David Naccache. Crypto corner: Communicating covertly through CPU monitoring. IEEE Security & Privacy, 11(6):71–73, November/ December 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Cranor:2004:GEI

[CG04] Lorrie Faith Cranor and Simson Garfinkel. Guest Editors’ introduction: Secure or usable? IEEE Security & Privacy, 2(5):16–18, September/October 2004. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/sp/2004/05/

j5016.pdf;   http://csdl.
computer.org/dl/mags/sp/
2004/05/j5016.htm.

Carvalho:2019:IMC

[CGC19] R. Carvalho, M. Goldsmith, and S. Creese. Investigating malware campaigns with semantic technologies. IEEE Security & Privacy, 17(1):43–54, January/February 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Cunningham:2016:ISP

[CGL^{+}16] Robert Cunningham, Pamela Gupta, Ulf Lindqvist, Stelios Sidiroglou-Douskos, and Michael Hicks. IEEE SecDev 2016: Prioritizing secure development. IEEE Security & Privacy, 14(4):82– 84, July/August 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/04/
msp2016040082.html.

Cho:2022:SDN

[CGR^{+}22] Minsu Cho, Zahra Ghodsi, Brandon Reagen, Siddharth Garg, and Chinmay Hegde. Sphynx: a deep neural network design for private inference. IEEE Security & Privacy, 20(5):22–34, September/ October 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Chamales:2004:HCR

[Cha04a] George Chamales. The Honeywall CD-ROM. IEEE Security & Privacy, 2(2):77–79, March/ April 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2004/02/j2077abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/02/j2077.
pdf.

Chaum:2004:VSB

[Cha04b]  David Chaum.  E-voting:

Secret-ballot receipts: True voter-verifiable elections. IEEE Security & Privacy, 2(1):38– 47, January/February 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2004/01/j1038abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/01/j1038.
pdf.

Charney:2009:HEO

[Cha09] Scott Charney. On the horizon: The evolution of online identity. IEEE Security & Privacy, 7(5):56–59, September/ October 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Charney:2012:CDA

[Cha12] Scott Charney. Collective defense: Applying the publichealth model to the Internet. IEEE Security & Privacy, 10 (2):54–59, March/April 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Cheung:2006:DSA

[Che06] Steven Cheung. Denial of service against the Domain Name System. IEEE Security & Privacy, 4(1):40– 45, January/February 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/33481/01588824.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
33481&arnumber=1588824.

Cheung:2011:SCI

[Che11] Steven Cheung. Securing collaborative intrusion detection systems. IEEE Security & Privacy, 9(6):36–42, November/ December 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Chen:2017:CSQ

[Che17] Lidong Chen. Cryptography standards in quantum time: New wine in an old wineskin? IEEE Security & Privacy, 15 (4):51–57, July/August 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/04/
msp2017040051-abs.html.

Cai:2005:CIW

[CHK^{+}05] Min Cai, Kai Hwang, Yu-Kwong Kwok, Shanshan Song, and Yu Chen. Collaborative Internet worm containment. IEEE Security & Privacy, 3(3):25–33, May/June 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/31002/01439499.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        31002&arnumber=1439499&count=
                                        20&index=6.

Camp:2020:TSI

[CHK^{+}20] J. Camp, R. Henry, T. Kohno, S. Mare, S. Myers, S. Patel, and J. Streiff. Toward a secure Internet of Things: Directions for research. IEEE Security & Privacy, 18(4):28–37, July/August 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Caulkins:2007:OIS

[CHMO07] Jonathan Caulkins, Eric D. Hough, Nancy R. Mead, and Hassan Osman. Optimizing investments in security countermeasures: a practical tool for fixed budgets. IEEE Security & Privacy, 5(5):57–60, September/October 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Chu:2008:BTI

[Chu08] Chengyun Chu. Basic training: Introduction to Microsoft .NET security. IEEE Security & Privacy, 6(6):73–78, November/December 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Campanile:2021:PRS

[CIL^{+}21] L. Campanile, M. Iacono, A. H.

Levis, F. Marulli, and M. Mastroianni. Privacy regulations, smart roads, blockchain, and liability insurance: Putting technologies to work. IEEE Security & Privacy, 19(1):34–43, January/February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Claffy:2010:DPU

[CK10] Kc Claffy and Erin Kenneally.

Dialing privacy and utility: a proposed data-sharing framework to advance Internet research. IEEE Security & Privacy, 8(4):31–39, July/August 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Charlet:2020:FCP

[CK20] K. Charlet and H. King. The future of cybersecurity policy. IEEE Security & Privacy, 18(1):8–10, January/February 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Cifuentes:2012:TPD

[CKL^{+}12] Cristina Cifuentes, Nathan Keynes, Lian Li, Nathan Hawes, and Manuel Valdiviezo. Transitioning Parfait into a development tool. IEEE Security & Privacy, 10(3):16–23, May/ June 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Cavallaro:2023:MLM

[CKPP23] Lorenzo Cavallaro, Johannes Kinder, Feargus Pendlebury, and Fabio Pierazzi. Are machine learning models for malware detection ready for prime time? IEEE Security & Privacy, 21(2):53–56, March/ April 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Chandramouli:2007:ISS

[CL07] Ramaswamy Chandramouli and Philip Lee. Infrastructure standards for smart ID card deployment. IEEE Security & Privacy, 5(2):92–96, March/April 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Cybenko:2012:SAM

[CL12] George Cybenko and Carl E. Landwehr. Security analytics and measurements. IEEE Security & Privacy, 10(3):5–8, May/June 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Caine:2013:SPH

[CL13] Kelly Caine and Michael Lesk. Security and privacy in health IT [Guest editors’ introduction]. IEEE Security & Privacy, 11(6):10–11, November/ December 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Camenisch:2012:EIN

[CLN12] Jan Camenisch, Anja Lehmann, and Gregory Neven. Electronic identities need private credentials. IEEE Security & Privacy, 10(1):80–83, January/ February 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Carpenter:2007:HVA

[CLS07] Matthew Carpenter, Tom Liston, and Ed Skoudis. Hiding virtualization from attackers and malware. IEEE Security & Privacy, 5(3):62–65, May/June 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Chess:2004:SAS

[CM04] Brian Chess and Gary Mc-Graw. Static analysis for security. IEEE Security & Privacy, 2(6):76–79, November/December 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/06/j6076.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/06/j6076.
pdf.

Caviglione:2022:NMM

[CM22] Luca Caviglione and Wojciech Mazurczyk. Never mind the malware, here’s the stegomalware. IEEE Security & Privacy, 20(5):101–106, September/October 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Ciclosi:2023:DPO

[CM23] Francesco Ciclosi and Fabio Massacci. The Data Protection Officer: a ubiquitous role that no one really knows. IEEE Security & Privacy, 21(1):66–77, January/February 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Cardenas:2013:SSB

[CMR13] Alvaro A. Cardenas, Pratyusa K. Manadhata, and Sreeranga P. Rajan. Systems security: Big data analytics for security. IEEE Security & Privacy, 11(6):74–76, November/ December 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Caputo:2009:DIT

[CMS09] Deanna Caputo, Marcus Maloof, and Gregory Stephens. Detecting insider theft of trade secrets. IEEE Security & Privacy, 7(6):14–21, November/ December 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Celik:2019:VIT

[CMT^{+}19] Z. B. Celik, P. McDaniel, G. Tan, L. Babun, and A. S. Uluagac. Verifying Internet of Things safety and security in physical spaces. IEEE Security & Privacy, 17(5):30– 37, September/October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Caloyannides:2009:GEI

[CMV09] Michael A. Caloyannides, Nasir Memon, and Wietse Venema. Guest Editors’ introduction: Digital forensics. IEEE Security & Privacy, 7(2):16–17, March/April 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Carettoni:2007:SBM

[CMZ07] Luca Carettoni, Claudio Merloni, and Stefano Zanero. Studying Bluetooth malware propagation: The BlueBag project. IEEE Security & Privacy, 5(2):17–25, March/April 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Cotroneo:2013:FIS

[CN13] Domenico Cotroneo and Roberto Natella. Fault injection for software certification. IEEE Security & Privacy, 11(4):38– 45, July/August 2013. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Cohen:2010:ACS

[Coh10a] Fred Cohen. Automated control system security. IEEE Security & Privacy, 8(5):62–63, September/October 2010. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Cohen:2010:SG

[Coh10b] Fred Cohen. The smarter grid.

IEEE Security & Privacy, 8(1): 60–63, January/February 2010.

CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Cohen:2010:VS

[Coh10c] Fred Cohen. The virtualization solution. IEEE Security & Privacy, 8(3):60–63, May/June 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Connolly:2018:FE

[Con18] Aisling Connolly. Freedom of encryption. IEEE Security & Privacy, 16(1):102–103, January/February 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/01/
msp2018010102.html.

Coron:2006:WC

[Cor06] Jean-Sebastien Coron. What is cryptography? IEEE Security & Privacy, 4(1):70– 73, January/February 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/33481/01588831.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
33481&arnumber=1588831.

Cowan:2003:SSO

[Cow03] Crispin Cowan. Software security for open-source systems. IEEE Security & Privacy, 1(1): 38–45, January/February 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j1038.pdf;   http://www.
computer.org/security/j1038abs.
htm.

Cowan:2023:OCS

[Cow23] Crispin Cowan. Open and closed software security redux. IEEE Security & Privacy, 21 (2):18–23, March/April 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Cox:2011:TC

[Cox11] Landon P. Cox. Truth in crowdsourcing. IEEE Security & Privacy, 9(5):74–76, September/October 2011. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Chuvakin:2009:BSL

[CP09a] Anton Chuvakin and Gunnar Peterson. Building security in: Logging in the age of Web services. IEEE Security & Privacy, 7(3):82–85, May/June 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Cukier:2009:VRP

[CP09b] Michel Cukier and Susmit Panjwani. Vulnerability remediation: Prioritizing vulnerability remediation by determining attacker-targeted vulnerabilities. IEEE Security & Privacy, 7(1):42–48, January/ February 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Chuvakin:2010:HDA

[CP10a] Anton Chuvakin and Gunnar Peterson. How to do application logging right. IEEE Security & Privacy, 8(4):82–85, July/August 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Cook:2010:SDS

[CP10b] Ian P. Cook and Shari Lawrence Pfleeger. Security decision support challenges in data collection and use. IEEE Security & Privacy, 8(3):28–35, May/June 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Caulfield:2015:ISP

[CP15] Tristan Caulfield and David Pym. Improving security policy decisions with models. IEEE Security & Privacy, 13(5):34–41, September/October 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://csdl.computer.
                                        org/csdl/mags/sp/2015/05/
                                        msp2015050034-abs.html.

Ceschin:2018:NSA

[CPC^{+}18] F. Ceschin, F. Pinage, M. Castilho, D. Menotti, L. S. Oliveira, and A. Gregio. The need for speed: An analysis of Brazilian malware classifiers. IEEE Security & Privacy, 16(6):31– 41, November/December 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Caputo:2014:GSP

[CPFJ14] Deanna D. Caputo, Shari Lawrence Pfleeger, Jesse D. Freeman, and M. Eric Johnson. Going spear phishing: Exploring embedded training and awareness. IEEE Security & Privacy, 12 (1):28–38, January/February 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Caputo:2016:BUS

[CPS^{+}16] Deanna D. Caputo, Shari Lawrence Pfleeger, M. Angela Sasse, Paul Ammann, Jeff Offutt, and Lin Deng. Barriers to usable security? Three organizational case studies. IEEE Security & Privacy, 14(5):22–32, September/October 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/05/
msp2016050022-abs.html.

Chandramouli:2006:CSD

[CR06] Ramaswamy Chandramouli and Scott Rose. Challenges in securing the Domain Name System. IEEE Security & Privacy, 4(1):84– 87, January/February 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/33481/01588835.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
33481&arnumber=1588835.

Callegati:2009:FL

[CR09a] Franco Callegati and Marco Ramilli. Frightened by links. IEEE Security & Privacy, 7(6): 72–76, November/December 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Chandramouli:2009:OIS

[CR09b] Ramaswamy Chandramouli and Scott Rose. Open issues in secure DNS deployment. IEEE Security & Privacy, 7(5):29–35, September/October 2009. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Crowther:2020:BCI

[CR20] K. G. Crowther and B. Rust. Built-in cybersecurity: Insights into product security for cyberphysical systems at a large company. IEEE Security & Privacy, 18(5):74–79, September/October 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Cranor:2003:PMP

[Cra03] Lorrie Faith Cranor. P3P: Making privacy policies more useful. IEEE Security & Privacy, 1(6):50–55, November/ December 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/sp/2003/06/j6050abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2003/06/j6050.
                                        htm;   http://csdl.computer.

org/dl/mags/sp/2003/06/j6050.
pdf.

Cranor:2012:CUC

[Cra12] Lorrie Faith Cranor. Can users control online behavioral advertising effectively? IEEE Security & Privacy, 10(2):93–96, March/April 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Chaboya:2006:NID

[CRBM06] David J. Chaboya, Richard A.

Raines, Rusty O. Baldwin, and Barry E. Mullins. Network intrusion detection: Automated and manual methods prone to attack and evasion. IEEE Security & Privacy, 4(6):36– 43, November/December 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Chignell:2013:NPT

[CRK^{+}13] Mark Chignell, Mahsa Rouzbahman, Ryan Kealey, Reza Samavi, Erin Yu, and Tammy Sieminowski. Nonconfidential patient types in emergency clinical decision support. IEEE Security & Privacy, 11(6):12– 18, November/December 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Conti:2009:PIM

[CS09] Gregory Conti and Edward Sobiesk. Privacy interests: Malicious interfaces and personalization’s uninviting future. IEEE Security & Privacy, 7(3):

64–67, May/June 2009. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Cranor:2013:SPE

[CS13] Lorrie Faith Cranor and Norman Sadeh. A shortage of privacy engineers. IEEE Security & Privacy, 11(2):77–79, March/April 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Clark:2014:DAC

[CSSvdH14] Kas Clark, Don Stikvoort, Eelco Stofbergen, and Elly van den Heuvel. A Dutch approach to cybersecurity through participation. IEEE Security & Privacy, 12(5):27– 34, September/October 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/05/
msp2014050027-abs.html.

Chen:2014:OPP

[CSZ^{+}14] Tiffani R. Chen, Daniel B. Shore, Stephen J. Zaccaro, Reeshad S. Dalal, Lois E. Tetrick, and Aiva K. Gorab. An organizational psychology perspective to examining Computer Security Incident Response Teams. IEEE Security & Privacy, 12(5):61–67, September/October 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/05/
msp2014050061-abs.html.

Chabanne:2011:SPE

[CT11] Herve Chabanne and Mehdi Tibouchi. Securing E-passports with elliptic curves. IEEE Security & Privacy, 9(2):75–78, March/April 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Corbett:2014:CIJ

[CUCD14] Cherita Corbett, Jason Uher, Jarriel Cook, and Angela Dalton. Countering intelligent jamming with full protocol stack agility. IEEE Security & Privacy, 12(2):44–50, March/ April 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Curran:2006:CSS

[Cur06] Charles D. Curran. Combating spam, spyware, and other desktop intrusions: Legal considerations in operating trusted intermediary technologies. IEEE Security & Privacy, 4(3):45–51, May/June 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Chillarege:2010:GEI

[CV10] Ram Chillarege and Jeffrey Voas. Guest Editors’ introduction: Reliability of embedded and cyber-physical systems. IEEE Security & Privacy, 8(5):12–13, September/ October 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Chan:2008:ETC

[CW08a] Yuen-Yan Chan and Victor K. Wei. Education: Teaching for conceptual change in security awareness. IEEE Security & Privacy, 6(6):67–69, November/December 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Coyne:2008:RII

[CW08b] Ed Coyne and Tim Weil. An RBAC implementation and interoperability standard: The INCITS cyber security 1.1 model. IEEE Security & Privacy, 6(1):84–87, January/ February 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Chan:2009:ETC

[CW09] Yuen-Yan Chan and Victor K. Wei. Education: Teaching for conceptual change in security awareness: a case study in higher education. IEEE Security & Privacy, 7(1):68–71, January/February 2009. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Chess:2012:SAMa

[CW12] Brian Chess and Chris Wysopal. Software assurance for the masses. IEEE Security & Privacy, 10(3):14–15, May/June 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Caviglione:2022:SPI

[CWVM22] Luca Caviglione, Steffen Wendzel, Simon Vrhovec, and Aleksandra Mileva. Security and privacy issues of home globalization. IEEE Security & Privacy, 20(1):10–11, January/ February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Chen:2018:TMI

[CXG^{+}18] Lin Chen, Lei Xu, Zhimin Gao, Yang Lu, and Weidong Shi. Tyranny of the majority: On the (Im)possibility of correctness of smart contracts. IEEE Security & Privacy, 16(4):30– 37, July/August 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/04/
msp2018040030-abs.html.

Cybenko:2003:CNA

[Cyb03a] George Cybenko. A critical need, an ambitious mission, a new magazine. IEEE Security & Privacy, 1(1):5–9, January/February 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/security/j1005.pdf.

Cybenko:2003:ESS

[Cyb03b] George Cybenko. From the Editor: Sapphire/Slammer redux. IEEE Security & Privacy, 1 (2):6, March/April 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j2006.pdf;   http://www.
computer.org/security/j2006abs.
htm.

Cybenko:2003:EPI

[Cyb03c] George Cybenko. From the Editors: Privacy is the issue. IEEE Security & Privacy, 1(6):5–7, November/ December 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/comp/mags/sp/2003/06/
j6005.pdf;   http://csdl.
computer.org/comp/mags/sp/
2003/06/j6005abs.htm;   http:
//csdl.computer.org/dl/mags/
sp/2003/06/j6005.htm.

Cybenko:2004:EDB

[Cyb04a] George Cybenko. From the editors: Don’t bring a knife to a gunfight. IEEE Security & Privacy, 2(2):5, March/ April 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2004/02/j2005.pdf.

Cybenko:2004:SA

[Cyb04b] George Cybenko. Security alchemy. IEEE Security & Privacy, 2(6):5, November/December 2004. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/comp/mags/sp/2004/06/
j6005.pdf;   http://csdl.

computer.org/dl/mags/sp/
2004/06/j6005.htm.

Cybenko:2005:OEM

[Cyb05] George Cybenko. The oneeyed man is king. IEEE Security & Privacy, 3(4):4–5, July/ August 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2005/04/j4004.pdf.

Cybenko:2006:WLS

[Cyb06a] George Cybenko. Weak links, strong ties. IEEE Security & Privacy, 4(6):3, November/ December 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Cybenko:2006:WJC

[Cyb06b]  George Cybenko.  Why

Johnny can’t evaluate security risk. IEEE Security & Privacy, 4(1):5, January/February 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/33481/01588814.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
33481&arnumber=1588814.

Ciocarlie:2023:SCI

[CZ23] Gabriela F. Ciocarlie and Jianying Zhou. Securing critical infrastructure across cyber and physical dimensions. IEEE Security & Privacy, 21(4):9, July/August 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Carbone:2008:TV

[CZL08] Martim Carbone, Diego Zamboni, and Wenke Lee. Taming virtualization. IEEE Security & Privacy, 6(1):65–67, January/February 2008. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Danezis:2005:ERC

[DA05] G. Danezis and R. Anderson. The economics of resisting censorship. IEEE Security & Privacy, 3(1):45– 50, January/February 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/30310/01392699.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        30310&arnumber=1392699&count=
                                        17&index=8.

delAlamo:2016:PES

[dA16] Jose M. del Alamo. Privacy engineering: Shaping an emerging field of research and practice. IEEE Security & Privacy, 14(2):40–46, March/ April 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

deAndrade:2013:ABS

[dAMM13] Norberto Nuno Gomes de Andrade, Aaron Martin, and Shara Monteleone. “All the better to see you with, my dear”: Facial recognition and privacy in online social networks. IEEE Security & Privacy, 11(3):21–28, May/June 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Dark:2014:ACE

[Dar14] Melissa Dark. Advancing cybersecurity education. IEEE Security & Privacy, 12(6):79– 83, November/December 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/06/
msp2014060079-abs.html.

Dark:2015:TAC

[Dar15] Melissa Dark. Thinking about cybersecurity. IEEE Security & Privacy, 13(1):61–65, January/February 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/01/
msp2015010061-abs.html.

Datta:2010:LMS

[Dat10] Anupam Datta. Logical methods in security and privacy. IEEE Security & Privacy, 8(6): 86–89, November/December 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Dittrich:2011:BAC

[DBD11] David Dittrich, Michael Bailey, and Sven Dietrich. Building an active computer security ethics community. IEEE Security & Privacy, 9(4):32–40, July/August 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Dube:2008:HRE

[DBR^{+}08] Thomas E. Dube, Bobby D. Birrer, Richard A. Raines, Rusty O. Baldwin, Barry E. Mullins, Robert W. Bennington, and Christopher E. Reuter. Hindering reverse engineering: Thinking outside the box. IEEE Security & Privacy, 6(2):58–65, March/April 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Demetrio:2022:PAM

[DBR22] Luca Demetrio, Battista Biggio, and Fabio Roli. Practical attacks on machine learning: a case study on adversarial Windows malware. IEEE Security & Privacy, 20(5):77– 85, September/October 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Duran:2009:BSI

[DCC^{+}09] Felicia Duran, Stephen H. Conrad, Gregory N. Conrad, David P. Duggan, and Edward Bruce Held. Building a system for insider security. IEEE Security & Privacy, 7(6): 30–38, November/December 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Davis:2003:TSD

[DD03] Jim Davis and Melissa Dark. Teaching students to design secure systems. IEEE Security & Privacy, 1(2):56– 58, March/April 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j2056.pdf;   http://www.
computer.org/security/j2056abs.
htm.

Dhamija:2008:SFI

[DD08] Rachna Dhamija and Lisa Dusseault. The seven flaws of identity management: Usability and security challenges. IEEE Security & Privacy, 6(2): 24–29, March/April 2008. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Dessiatnikoff:2012:PAO

[DDAN12] Anthony Dessiatnikoff, Yves Deswarte, Eric Alata, and Vincent Nicomette. Potential attacks on onboard aerospace systems. IEEE Security & Privacy, 10(4):71–74, July/ August 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

DeCristofaro:2014:GPR

[De 14] Emiliano De Cristofaro. Genomic privacy and the rise of a new research community. IEEE Security & Privacy, 12(2):80– 83, March/April 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

DeCristofaro:2021:COP

[De 21] Emiliano De Cristofaro. A critical overview of privacy in machine learning. IEEE Security & Privacy, 19(4):19–27, July/ August 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

DeCristofaro:2024:SDM

[De 24] Emiliano De Cristofaro. Synthetic data: Methods, use cases, and risks. IEEE Security & Privacy, 22(3):62–67, May/ June 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Dedeke:2017:CFA

[Ded17] Adenekan Dedeke. Cybersecurity framework adoption: Using capability levels for implementation tiers and profiles. IEEE Security & Privacy, 15(5):47–54, September/October 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/05/
                                        msp2017050047-abs.html.

DeFigueiredo:2011:CMT

[DeF11] Dimitri DeFigueiredo. The case for mobile two-factor authentication. IEEE Security & Privacy, 9(5):81–85, September/October 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Das:2021:PPE

[DEKM^{+}21] Sauvik Das, W. Keith Edwards, DeBrae Kennedy-Mayo, Peter Swire, and Yuxi Wu. Privacy for the people? Exploring collective action as a mechanism to shift power to consumers in end-user privacy. IEEE Security & Privacy, 19 (5):66–70, September/October 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

DeLooze:2007:PWS

[DeL07] Lori DeLooze. Providing Web service security in a federated environment. IEEE Security & Privacy, 5(1):73–75, January/ February 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

DeLong:2014:ACJ

[DeL14] John DeLong. Aligning the compasses: A journey through compliance and technology. IEEE Security & Privacy, 12 (4):85–89, July/August 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/04/
msp2014040085-abs.html.

DeMott:2015:BE

[DeM15]  Jared DeMott.  Bypassing

EMET 4.1. IEEE Security & Privacy, 13(4):66– 72, July/August 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/04/
msp2015040066-abs.html.

Dern:2003:PC

[Der03]  Daniel P. Dern.  Privacy

concerns. IEEE Security & Privacy, 1(2):11–13, March/ April 2003. CODEN ????

ISSN 1540-7993 (print), 1558-4046 (electronic). URL

                                        http://dlib.computer.org/
                                        sp/books/sp2003/pdf/j2011.
                                        pdf;   http://www.computer.
                                        org/security/j2011abs.htm.

Dery:2013:UWC

[Der13] Susan Dery. Using whitelisting to combat malware attacks at Fannie Mae. IEEE Security & Privacy, 11(4):90–92, July/ August 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Datta:2011:AMC

[DFG^{+}11] Anupam Datta, Jason Franklin, Deepak Garg, Limin Jia, and Dilsun Kaynar. On adversary models and compositional security. IEEE Security & Privacy, 9(3):26–32, May/June 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Dessouky:2020:GCC

[DFJ^{+}20] G. Dessouky, T. Frassetto, P. Jauernig, A.-R. Sadeghi, and E. Stapf. With great complexity comes great vulnerability: From stand-alone fixes to reconfigurable security. IEEE Security & Privacy, 18(5):57– 66, September/October 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Drinan:2005:NBb

[DFK05] Heather Drinan, Nancy Fontaine, and Brent Kesler. News briefs. IEEE Security & Privacy, 3 (6):7–8, November/December 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Drinan:2006:NB

[DFK06] Heather Drinan, Nancy Fontaine, and Brent Kesler. News briefs. IEEE Security & Privacy, 4 (1):14–16, January/February 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/33481/01588819.pdf;
http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
33481&arnumber=1588819.

Dunlop:2012:BMB

[DGU^{+}12] Matthew Dunlop, Stephen Groat, William Urbanski, Randy Marchany, and Joseph Tront. The blind man’s bluff approach to security using IPv6. IEEE Security & Privacy, 10(4):35–43, July/ August 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Dewey-Hagborg:2013:SVP

[DH13] Heather Dewey-Hagborg. Stranger visions: A provocation. IEEE Security & Privacy, 11(6):69– 70, November/December 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Dhillon:2011:DDT

[Dhi11] Danny Dhillon. Developerdriven threat modeling: Lessons learned in the trenches. IEEE Security & Privacy, 9(4):41–47, July/August 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Dechesne:2014:EIS

[DHP14] Francien Dechesne, Dina Hadziosmanovic, and Wolter Pieters. Experimenting with incentives: Security in pilots for future grids. IEEE Security & Privacy, 12(6):59–66, November/December 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/06/
msp2014060059-abs.html.

Davis:2004:PPS

[DHR^{+}04] Noopur Davis, Watts Humphrey, Samuel T. Redwine, Jr., Gerlinde Zibulski, and Gary Mc-Graw. Processes for producing secure software: Summary of US National Cybersecurity Summit Subgroup Report. IEEE Security & Privacy, 2(3):18–25, May/ June 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/03/j3018.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/03/j3018.
pdf.

Diamant:2011:RSA

[Dia11] John Diamant. Resilient security architecture: a complementary approach to reducing vulnerabilities. IEEE Security & Privacy, 9(4):80–84, July/ August 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Diffie:2006:CAS

[Dif06] Whitfield Diffie. Chattering about SIGINT. IEEE Security & Privacy, 4(1):9, January/February 2006. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/33481/01588817.pdf;
http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
33481&arnumber=1588817.

Dimitriadis:2007:IMC

[Dim07] Christos K. Dimitriadis. Improving mobile core network security with honeynets. IEEE Security & Privacy, 5(4):40–47, July/August 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Drinan:2005:NBa

[DK05] Heather Drinan and Brent Kesler. News briefs. IEEE Security & Privacy, 3(5):8–10, September/October 2005. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/comp/mags/sp/2005/05/
j5008.pdf.

DeCristofaro:2010:SLI

[DK10] Emiliano De Cristofaro and Jihye Kim. Some like it private: Sharing confidential information based on oblivious authorization. IEEE Security & Privacy, 8(4):18–24, July/ August 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Dragoni:2021:WPS

[DLMS21] N. Dragoni, A. Lluch Lafuente, F. Massacci, and A. Schlichtkrull. Are we preparing students to build security in? A survey of European cybersecurity in higher education programs [education]. IEEE Security & Privacy, 19(1):81–88, January/ February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Dark:2015:ETP

[DM15a] Melissa Dark and Jelena Mirkovic. Evaluation theory and practice applied to cybersecurity education. IEEE Security & Privacy, 13(2):75– 80, March/April 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/02/
msp2015020075-abs.html.

Dimitrakakis:2015:DBP

[DM15b] Christos Dimitrakakis and Aikaterini Mitrokotsa. Distancebounding protocols: Are you close enough? IEEE Security & Privacy, 13(4):47– 51, July/August 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.

org/csdl/mags/sp/2015/04/
msp2015040047-abs.html.

Drimer:2009:FTP

[DMA09] Saar Drimer, Steven J. Murdoch, and Ross Anderson. Failures of tamper-proofing in PIN entry devices. IEEE Security & Privacy, 7(6):39–45, November/December 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Dykstra:2022:ABT

[DMB^{+}22] Josiah Dykstra, Jamie Met, Nicole Backert, Rebecca Mattie, and Douglas Hough. Action bias and the two most dangerous words in cybersecurity incident response: An argument for more measured incident response. IEEE Security & Privacy, 20(3):102–106, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Deutschbein:2022:THS

[DMR^{+}22] Calvin Deutschbein, Andres Meza, Francesco Restuccia, Matthew Gregoire, Ryan Kastner, and Cynthia Sturton. Toward hardware security property generation at scale. IEEE Security & Privacy, 20(3):43– 51, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Dingledine:2007:DLL

[DMS07] Roger Dingledine, Nick Mathewson, and Paul Syverson. Deploying low-latency anonymity:

Design challenges and social factors. IEEE Security & Privacy, 5(5):83–87, September/ October 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

DeGramatica:2015:IIE

[DMS^{+}15] Martina De Gramatica, Fabio Massacci, Woohyun Shim, Alessandra Tedeschi, and Julian Williams. IT interdependence and the economic fairness of cybersecurity regulations for civil aviation. IEEE Security & Privacy, 13(5):52– 61, September/October 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/csdl/mags/sp/2015/05/
msp2015050052-abs.html.

Dushin:2007:HMC

[DN07] Fred Dushin and Eric Newcomer. Handling multiple credentials in a heterogeneous SOA environment. IEEE Security & Privacy, 5(5):80–82, September/October 2007. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Divakaran:2022:PDL

[DO22] Dinil Mon Divakaran and Adam Oest. Phishing detection leveraging machine learning and deep learning: a review. IEEE Security & Privacy, 20(5):86–95, September/ October 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Dodds:2022:FVI

[Dod22] Mike Dodds. Formally verifying industry cryptography. IEEE Security & Privacy, 20 (3):65–70, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

DAmico:2013:BBA

[DOL13] Anita D’Amico, Brianne O’Brien, and Mark Larkin. Building a bridge across the transition chasm. IEEE Security & Privacy, 11(2):24–33, March/ April 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Donner:2003:ABM

[Don03a] Marc Donner. AI bites man?

IEEE Security & Privacy, 1(1): 63–66, January/February 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j1063.pdf;   http://www.
computer.org/security/j1063abs.
htm.

Donner:2003:HR

[Don03b] Marc Donner. Hey, robot!

IEEE Security & Privacy, 1(3): 51–55, May/June 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j3051.pdf.

Donner:2003:PAN

[Don03c] Marc Donner. Post-apocalypse now. IEEE Security & Privacy, 1(2):53–55, March/April 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j2053.pdf;   http://www.
computer.org/security/j2053abs.
htm.

Donner:2003:TSO

[Don03d] Marc Donner. Toward a security ontology. IEEE Security & Privacy, 1(3):6– 7, May/June 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j3006.pdf.

Donner:2004:BTG

[Don04a] Marc Donner. Biblio tech: Die Gedanken Sind Frei. IEEE Security & Privacy, 2 (1):53–55, January/February 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2004/01/j1053abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/01/j1053.
pdf.

Donner:2004:CC

[Don04b]  Marc Donner.  Cult clas-

sics. IEEE Security & Privacy, 2(3):66–68, May/ June 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/03/j3066.

htm;   http://csdl.computer.
org/dl/mags/sp/2004/03/j3066.
pdf.

Donner:2004:DEM

[Don04c]  Marc Donner.  Deus est

machina. IEEE Security & Privacy, 2(4):51–53, July/ August 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/04/j4051.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/04/j4051.
pdf.

Donner:2004:EWL

[Don04d] Marc Donner. From the editors: a witty lesson. IEEE Security & Privacy, 2(4):5, July/August 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/04/j4005.
htm.

Donner:2004:EWD

[Don04e]  Marc Donner.  From the

editors: Whose data are these, anyway? IEEE Security & Privacy, 2(3):5–6, May/June 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/03/j3005.
htm.

Donner:2004:HBS

[Don04f] Marc Donner. Hacking the best-seller list. IEEE Security & Privacy, 2(2):51–53, March/ April 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2004/02/j2051abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/02/j2051.
pdf.

Donner:2004:JG

[Don04g] Marc Donner. Jennifer Government. IEEE Security & Privacy, 2(5):57–59, September/October 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/05/j5057.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/05/j5057.
pdf.

Donner:2004:UFL

[Don04h] Marc Donner. Use the force, Luke! IEEE Security & Privacy, 2(6):53–55, November/December 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/06/j6053.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/06/j6053.
pdf.

Donner:2005:WN

[Don05a] M. Donner. What’s in a name? IEEE Security & Privacy, 3(2):4–5, March/April 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/30742/01423949.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
30742&arnumber=1423949&count=
15&index=0.

Donner:2005:YGF

[Don05b] M. Donner. A young geek’s fancy turns to… science fiction? [book recommendations]. IEEE Security & Privacy, 3(3):58–60, May/June 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/31002/01439503.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
31002&arnumber=1439503&count=
20&index=10.

Donner:2005:TAN

[Don05c] Marc Donner. There ain’t no inside, there ain’t no outside. … IEEE Security & Privacy, 3(5):4–5, September/ October 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2005/05/j5004.pdf.

Donner:2006:ID

[Don06a] Marc Donner. The impending debate. IEEE Security & Privacy, 4(2):4–5, March/ April 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2006/02/j2004.pdf.

Donner:2006:ITO

[Don06b]  Marc Donner.  Insecurity

through obscurity. IEEE Security & Privacy, 4(5):4, September/October 2006. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/sp/2006/05/
                                        j5004.pdf.

Donner:2007:CE

[Don07] Marc Donner. Cyberassault on Estonia. IEEE Security & Privacy, 5(4):4, July/August 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Donner:2008:CLB

[Don08a] Marc Donner. Charge of the light brigade. IEEE Security & Privacy, 6(1):5, January/February 2008. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/sp/2008/01/
                                        msp2008010005.pdf.

Donner:2008:LEI

[Don08b] Marc Donner. Lessons from electrification for identification. IEEE Security & Privacy, 6(2):3, March/April 2008. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/sp/2008/02/
                                        msp2008020003.pdf.

Donner:2009:BTW

[Don09a] Marc Donner. Biblio tech:

War stories. IEEE Security & Privacy, 7(3):60–63, May/June 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Donner:2009:ENM

[Don09b] Marc Donner. From the editors: New models for old. IEEE Security & Privacy, 7(4):3–4, July/August 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Donner:2009:ERE

[Don09c] Marc Donner. From the Editors: Reading (with) the enemy. IEEE Security & Privacy, 7(1):3, January/February 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Donner:2010:IB

[Don10a] Marc Donner. International blues. IEEE Security & Privacy, 8(2):3, March/April 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Donner:2010:PC

[Don10b] Marc Donner. Phagocytes in cyberspace. IEEE Security & Privacy, 8(5):3–4, September/ October 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Donner:2011:IC

[Don11a] Marc Donner. The invisible computers. IEEE Security & Privacy, 9(6):3, November/ December 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Donner:2011:PSL

[Don11b] Marc Donner. Privacy and the system life cycle. IEEE Security & Privacy, 9(2):3, March/ April 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Dettmer:2021:LLD

[DOP^{+}21] Amy Dettmer, Hamed Okhravi, Kevin Perry, Nabil Schear, Richard Shay, Mary Ellen Zurko, and Paula Donovan. Lessons learned from designing a security architecture for realworld government agencies. IEEE Security & Privacy, 19 (4):28–36, July/August 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Ding:2017:CSM

[DP17] Jintai Ding and Albrecht Petzoldt. Current state of multivariate cryptography. IEEE Security & Privacy, 15(4):28– 36, July/August 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/04/
msp2017040028-abs.html.

Dunphy:2018:FLI

[DP18] Paul Dunphy and Fabien A. P.

Petitcolas. A first look at identity management schemes on the blockchain. IEEE Security & Privacy, 16(4):20– 29, July/August 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/04/
msp2018040020-abs.html.

Dietz:2020:UDT

[DP20] M. Dietz and G. Pernul. Unleashing the digital twin’s potential for ICS security. IEEE Security & Privacy, 18(4):20– 27, July/August 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Dellios:2015:ISC

[DPP15] Kleanthis Dellios, Dimitrios Papanikas, and Despina Polemi. Information security compliance over intelligent transport systems: Is IT possible? IEEE Security & Privacy, 13(3):9– 15, May/June 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/03/
msp2015030009-abs.html.

Degabriele:2011:PSR

[DPW11] Jean Paul Degabriele, Kenny Paterson, and Gaven J. Watson. Provable security in the real world. IEEE Security & Privacy, 9(3):33–41, May/June 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Qiu:2013:IDR

[dQSzL13] Wei dong Qiu, Qian Su, and Bo zhong Liu. iOS data recovery using low-level NAND images. IEEE Security & Privacy, 11(5):49–55, September/ October 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Dill:2004:GEI

[DR04] David L. Dill and Aviel D. Rubin. Guest Editors’ introduction: E-voting security. IEEE Security & Privacy, 2(1):22–23, January/February 2004. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/comp/mags/sp/2004/01/
j1022.pdf;   http://csdl.
computer.org/comp/mags/sp/
2004/01/j1022abs.htm.

Dodge:2005:TEU

[DR05] R. Dodge and D. Ragsdale. Technology education at the US Military Academy. IEEE Security & Privacy, 3(2):49– 53, March/April 2005. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/30742/01423961.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
30742&arnumber=1423961&count=
15&index=7.

Dempsey:2006:GEI

[DR06] James X. Dempsey and Ira Rubinstein. Guest Editors’ introduction: Lawyers and technologists—joined at the hip? IEEE Security & Privacy, 4(3):15–19, May/June 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/comp/mags/sp/2006/03/
j3015.pdf.

Daemen:2010:FYA

[DR10] Joan Daemen and Vincent Rijmen. The first 10 years of advanced encryption. IEEE Security & Privacy, 8(6):72– 74, November/December 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Dressler:2016:CWS

[Dre16] Falko Dressler. Cleaning up Web 2.0’s security mess — at least partly. IEEE Security & Privacy, 14(2):48–57, March/ April 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Dixon:2016:NTO

[DRS16] Lucas Dixon, Thomas Ristenpart, and Thomas Shrimpton. Network traffic obfuscation and automated Internet censorship. IEEE Security & Privacy, 14(6):43– 53, November/December 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.

                                        org/csdl/mags/sp/2016/06/
                                        msp2016060043-abs.html.

Dan:2012:CPS

[DSEB12] Gyorgy Dan, Henrik Sandberg, Mathias Ekstedt, and Gunnar Bjorkman. Challenges in power system information security. IEEE Security & Privacy, 10 (4):62–70, July/August 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

DarvishRouani:2019:SML

[DSJK19] B. Darvish Rouani, M. Samragh, T. Javidi, and F. Koushanfar. Safe machine learning and defeating adversarial attacks. IEEE Security & Privacy, 17 (2):31–38, March/April 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Daffalla:2022:DTU

[DSKB22] Alaa Daffalla, Lucy Simko, Tadayoshi Kohno, and Alexandru G. Bardas. Defensive technology use during the 2018–2019 Sudanese revolution. IEEE Security & Privacy, 20(2):40–48, March/ April 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Du:2011:SHL

[Du11] Wenliang Du. SEED: Handson lab exercises for computer security education. IEEE Security & Privacy, 9(5):70–73, September/October 2011. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Dunning:2010:TBB

[Dun10] John Paul Dunning. Taming the blue beast: a survey of Bluetooth based threats. IEEE Security & Privacy, 8(2):20–27, March/April 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Doumanidis:2023:DIC

[DXR^{+}23] Constantine Doumanidis, Yongyu Xie, Prashant H. N. Rajput, Ryan Pickren, Burak Sahin, Saman Zonouz, and Michail Maniatakos. Dissecting the industrial control systems software supply chain. IEEE Security & Privacy, 21(4):39–50, July/August 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Deng:2021:UIP

[DXS21] Shuwen Deng, Wenjie Xiong, and Jakub Szefer. Understanding the insecurity of processor caches due to cache timingbased vulnerabilities. IEEE Security & Privacy, 19(3):42–49, May/June 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Eagle:2013:CSC

[Eag13] Chris Eagle. Computer security competitions: Expanding educational outcomes. IEEE Security & Privacy, 11(4):69– 71, July/August 2013. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Enck:2019:SPU

[EB19] W. Enck and T. Benzel. Selected papers from the 2018 USENIX Security Symposium. IEEE Security & Privacy, 17 (4):7–8, July/August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

English:2007:MAC

[ECG^{+}07] Jennifer English, David Coe, Rhonda Gaede, David Hyde, and Jeffrey Kulick. MEMSassisted cryptography for CPI protection. IEEE Security & Privacy, 5(4):14–21, July/ August 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Edelman:2010:LCA

[Ede10] Benjamin Edelman. Least-cost avoiders in online fraud and abuse. IEEE Security & Privacy, 8(4):78–81, July/August 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Edelman:2014:APS

[Ede14] Benjamin Edelman. Accountable? The problems and solutions of online ad optimization. IEEE Security & Privacy, 12(6):102–107, November/December 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/sp/2014/06/
                                        msp2014060102-abs.html.

Abou-Tair:2022:DSS

[eDIATK22] Dhiah el Diehn I. Abou-Tair and Ala Khalifeh. Distributed self-sovereign-based access control system. IEEE Security & Privacy, 20(6):35–42, November/December 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Eckhart:2023:SED

[EEA^{+}23] Matthias Eckhart, Andreas Ekelhart, David Allison, Magnus Almgren, Katharina Ceesay-Seitz, Helge Janicke, Simin Nadjm-Tehrani, Awais Rashid, and Mark Yampolskiy. Securityenhancing digital twins: Characteristics, indicators, and future perspectives. IEEE Security & Privacy, 21(6):64– 75, November/December 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

ElBansarkhani:2018:PSD

[EGB18] Rachid El Bansarkhani, Matthias Geihs, and Johannes Buchmann. PQChain: Strategic design decisions for distributed ledger technologies against future threats. IEEE Security & Privacy, 16(4):57– 65, July/August 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/04/
msp2018040057-abs.html.

Essex:2012:HTE

[EH12] Aleksander Essex and Urs Hengartner. Hover: Trustworthy elections with hash-only verification. IEEE Security & Privacy, 10(5):18–24, September/ October 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ellison:2013:PPH

[EH13] Nicole B. Ellison and Jeffrey T. Hancock. Profile as promise: Honest and deceptive signals in online dating. IEEE Security & Privacy, 11(5):84–88, September/October 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Evans:2004:RBS

[EHK^{+}04] Shelby Evans, David Heinbuch, Elizabeth Kyule, John Piorkowski, and James Wallner. Risk-based systems security engineering: Stopping attacks with intention. IEEE Security & Privacy, 2(6):59–62, November/December 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/06/j6059.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/06/j6059.
pdf.

Egan:2006:BBF

[EJ06] Erin Egan and Tim Jucovy. Building a better filter: How to create a safer Internet and avoid the litigation trap. IEEE Security & Privacy, 4(3):37–44, May/June 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

ElEmam:2009:PIP

[EK09] Khaled El Emam and Patricia Kosseim. Privacy interests: Privacy interests in prescription data, part 2: Patient privacy. IEEE Security & Privacy, 7(2):75–78, March/April 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ekberg:2014:UPT

[EKA14] Jan-Erik Ekberg, Kari Kostiainen, and N. Asokan. The untapped potential of trusted execution environments on mobile devices. IEEE Security & Privacy, 12(4):29– 37, July/August 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/04/
msp2014040029-abs.html.

ElEmam:2008:HIH

[El 08] Khaled El Emam. Heuristics for de-identifying health data. IEEE Security & Privacy, 6(4): 58–61, July/August 2008. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

ElEmam:2010:RBI

[El 10] Khaled El Emam. Risk-based de-identification of health data. IEEE Security & Privacy, 8(3): 64–67, May/June 2010. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

ElEmam:2020:SWE

[El 20] K. El Emam. Seven ways to evaluate the utility of synthetic data. IEEE Security & Privacy, 18(4):56–59, July/ August 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Elliott:2004:QC

[Ell04]  Chip Elliott.  Quantum

cryptography. IEEE Security & Privacy, 2(4):57–61, July/August 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/04/j4057.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/04/j4057.
pdf.

Elliott:2011:DSC

[Ell11] David Elliott. Deterring strategic cyberattack. IEEE Security & Privacy, 9(5):36–40, September/October 2011. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ellis:2014:RCI

[Ell14] Ryan Ellis. Regulating cybersecurity: Institutional learning or a lesson in futility? IEEE Security & Privacy, 12(6):48– 54, November/December 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/06/
msp2014060048-abs.html.

Epstein:2006:SSS

[EMM06] Jeremy Epstein, Scott Matsumoto, and Gary McGraw. Software security and SOA: Danger, Will Robinson! IEEE Security & Privacy, 4(1):80– 83, January/February 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/33481/01588834.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
33481&arnumber=1588834.

Emami-Naeini:2022:ISP

[ENDAC22] Pardis Emami-Naeini, Janarth Dheenadhayalan, Yuvraj Agarwal, and Lorrie Faith Cranor. An informative security and privacy nutrition label for Internet of Things devices. IEEE Security & Privacy, 20(2):31– 39, March/April 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Enck:2009:FUA

[EOM09] William Enck, Machigar Ongtang, and Patrick McDaniel. Focus: Understanding Android security. IEEE Security & Privacy, 7(1):50–57, January/ February 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Evans:2004:VES

[EP04] David Evans and Nathanael Paul. E-voting: Election security: Perception and reality. IEEE Security & Privacy, 2(1):24–31, January/ February 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/

                                        mags/sp/2004/01/j1024abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2004/01/j1024.
                                        pdf.

Endicott-Popovsky:2012:UCD

[EPH12] Barbara Endicott-Popovsky and Donald J. Horowitz. Unintended consequences: Digital evidence in our legal system. IEEE Security & Privacy, 10 (2):80–83, March/April 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Epstein:2008:SLL

[Eps08]  Jeremy Epstein.  Security

lessons learned from Société Générale. IEEE Security & Privacy, 6(3):80–82, May/June 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Epstein:2012:CWT

[Eps12a] Jeremy Epstein. Can we be too careful? IEEE Security & Privacy, 10(2):3–5, March/ April 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Epstein:2012:RSP

[Eps12b] Jeremy Epstein. Reflecting on some past predictions. IEEE Security & Privacy, 10(2):7–10, March/April 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Epstein:2013:EAT

[Eps13] Jeremy Epstein. From the Editors: Are all types of Internet voting unsafe? IEEE Security & Privacy, 11(3):3–4, May/June 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Epstein:2014:POE

[Eps14] Jeremy Epstein. Phishing our employees. IEEE Security & Privacy, 12(3):3–4, May/June 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Epstein:2015:WDV

[Eps15a] Jeremy Epstein. Weakness in depth: A voting machine’s demise. IEEE Security & Privacy, 13(3):55– 58, May/June 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/03/
msp2015030055-abs.html.

Epstein:2015:WLT

[Eps15b] Jeremy Epstein. The whole is less than the sum of the parts. IEEE Security & Privacy, 13 (2):3–5, March/April 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/02/
msp2015020003.html.

Epstein:2016:RNP

[Eps16] Jeremy Epstein. Reflections of an NSF Program Officer. IEEE Security & Privacy, 14(2):3–6, March/April 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Epstein:2017:PCD

[Eps17] Jeremy Epstein. Privacy is context dependent. IEEE Security & Privacy, 15(2):5– 6, March/April 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/02/
                                        msp2017020005.html.

Epstein:2019:HRC

[Eps19] J. Epstein. A horse’s rump and cybersecurity. IEEE Security & Privacy, 17(4):4–6, July/August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Evans:2011:GEI

[ES11] David Evans and Sal Stolfo. Guest Editors’ introduction: The science of security. IEEE Security & Privacy, 9(3):16–17, May/June 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Eckhoff:2014:DBD

[ES14] David Eckhoff and Christoph Sommer. Driving for big data? Privacy concerns in vehicular networking. IEEE Security & Privacy, 12(1):77–79, January/ February 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Essex:2017:DDU

[Ess17] Aleksander Essex. Detecting the detectable: Unintended consequences of cryptographic election verification. IEEE Security & Privacy, 15(3):30– 38, May/June 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/03/
msp2017030030-abs.html.

Edwards:2018:EAR

[EV18] Lilian Edwards and Michael Veale. Enslaving the algorithm: From a “Right to an explanation” to a “Right to better decisions”? IEEE Security & Privacy, 16(3):46– 54, May/June 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/03/
msp2018030046-abs.html.

Evans:2014:WWM

[Eva14] Katrine Evans. Where in the world is my information?: Giving people access to their data. IEEE Security & Privacy, 12(5):78–81, September/October 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/05/
msp2014050078-abs.html.

Evans:2015:VHR

[Eva15] Katrine Evans. Vidal–Hall and risk management for privacy breaches. IEEE Security & Privacy, 13(5):80–84, September/October 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://csdl.computer.
                                        org/csdl/mags/sp/2015/05/
                                        msp2015050080-abs.html.

Enck:2022:TFC

[EW22] William Enck and Laurie Williams. Top five challenges in software supply chain security: Observations from 30 industry and government organizations. IEEE Security & Privacy, 20 (2):96–100, March/April 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Ford:2007:HS

[FA07a] Richard Ford and William H. Allen. How not to be seen. IEEE Security & Privacy, 5(1): 67–69, January/February 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ford:2007:HSI

[FA07b] Richard Ford and William H. Allen. How not to be seen II: The defenders fight back. IEEE Security & Privacy, 5(6):65– 68, November/December 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ford:2009:MSG

[FA09] Richard Ford and William H. Allen. Malware shall greatly increase …. IEEE Security & Privacy, 7(6):69–71, November/December 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Farahmand:2020:QIC

[Far20]  F. Farahmand.  Quantita-

tive issues in cyberinsurance: Lessons from behavioral economics, counterfactuals, and causal inference. IEEE Security & Privacy, 18(2):8–15, March/ April 2020. ISSN 1558-4046.

Farahmand:2021:ICA

[Far21] Fariborz Farahmand. Integrating cybersecurity and artificial intelligence research in engineering and computer science education. IEEE Security & Privacy, 19(6): 104–110, November/December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Frincke:2004:ADP

[FB04a] Deborah Frincke and Matt Bishop. Academic degrees and professional certification. IEEE Security & Privacy, 2(6):56–58, November/ December 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/06/j6056.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/06/j6056.
pdf.

Frincke:2004:BS

[FB04b] Deborah Frincke and Matt Bishop. Back to school. IEEE Security & Privacy, 2(4):54–56, July/August 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/04/j4054.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/04/j4054.
pdf.

Frincke:2004:JSE

[FB04c] Deborah Frincke and Matt Bishop. Joining the security education community. IEEE Security & Privacy, 2(5):61– 63, September/October 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/sp/
2004/05/j5061.htm;   http:
//csdl.computer.org/dl/mags/
sp/2004/05/j5061.pdf.

Frincke:2004:GCK

[FB04d] Deborah A. Frincke and Matt Bishop. Guarding the castle keep: Teaching with the fortress metaphor. IEEE Security & Privacy, 2(3):69–72, May/June 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/03/j3069.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/03/j3069.
pdf.

Ferguson-Boucher:2011:CCR

[FB11] Kirsten Ferguson-Boucher. Cloud computing: a records and information management perspective. IEEE Security & Privacy, 9(6):63–66, November/ December 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Fischmann:2004:LEE

[FBG04] Matthias Fischmann, Matthias Bauer, and Simson Garfinkel. Letters to the editor: EBIA vs. PKI. IEEE Security & Privacy, 2(2):6–7, March/April 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/comp/mags/sp/2004/02/
j2006.pdf.

Ford:2014:PM

[FC14] Richard Ford and Marco Carvalho. Protecting me. IEEE Security & Privacy, 12(1):80– 82, January/February 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Furon:2010:TPC

[FD10] Teddy Furon and Gwenael Doerr. Tracing pirated content on the Internet: Unwinding Ariadne’s thread. IEEE Security & Privacy, 8(5):69–71, September/October 2010. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Fuglerud:2011:SIA

[FD11] Kristin Fuglerud and Øystein Dale. Secure and inclusive authentication with a talking mobile one-time-password client. IEEE Security & Privacy, 9(2): 27–34, March/April 2011. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Fenz:2011:VVE

[FE11] Stefan Fenz and Andreas Ekelhart. Verification, validation, and evaluation in information security risk management. IEEE Security & Privacy, 9(2):58–65, March/April 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Felten:2003:UTC

[Fel03]  Edward W. Felten.  Un-

derstanding trusted computing: Will its benefits outweigh its drawbacks? IEEE Security & Privacy, 1(3):60– 62, May/June 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://dlib.computer.
                                        org/sp/books/sp2003/pdf/
                                        j3060.pdf.

Felker:2006:IWG

[Fel06] Mikhael Felker. Internet war games: Power of the masses. IEEE Security & Privacy, 4 (2):7, March/April 2006. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/sp/2006/02/
                                        j2007.pdf.

Fetzer:2016:BCA

[Fet16] Christof Fetzer. Building critical applications using microservices. IEEE Security & Privacy, 14(6):86–89, November/December 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/06/
msp2016060086-abs.html.

Ford:2010:BBB

[FF10] Richard Ford and Deborah A.

Frincke. Building a better boot camp. IEEE Security & Privacy, 8(1):68–71, January/ February 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Frankel:2008:IPV

[FG08] Sheila Frankel and David Green. Internet Protocol Version 6. IEEE Security & Privacy, 6(3):83–86, May/June 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Fischer:2022:NSD

[FG22] Felix Fischer and Jens Grossklags.

Nudging software developers toward secure code. IEEE Security & Privacy, 20(2):76–79, March/April 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Felten:2006:DRM

[FH06] Edward W. Felten and J. Alex Halderman. Digital rights management, spyware, and security. IEEE Security & Privacy, 4(1):18–23, January/February 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/33481/01588821.pdf;

                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        33481&arnumber=1588821.

Florencio:2012:EWK

[FH12] Dinei Florencio and Cormac Herley. Is everything we know about password stealing wrong? IEEE Security & Privacy, 10(6):63–69, November/ December 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Fink:2014:DMA

[FHMF14] Glenn A. Fink, Jereme N. Haack, A. David McKinnon, and Errin W. Fulp. Defense on the move: Ant-based cyber defense. IEEE Security & Privacy, 12(2):36–43, March/ April 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Fidler:2011:WSA

[Fid11] David P. Fidler. Was Stuxnet an act of war? decoding a cyberattack. IEEE Security & Privacy, 9(4):56–59, July/ August 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Figueroa:2009:NBa

[Fig09a] James Figueroa. News briefs. IEEE Security & Privacy, 7(3): 6–7, May/June 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Figueroa:2009:NBb

[Fig09b] James Figueroa. News briefs.

IEEE Security & Privacy, 7(4): 8–10, July/August 2009. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Figueroa:2009:NBc

[Fig09c] James Figueroa. News briefs.

IEEE Security & Privacy, 7 (5):9–10, September/October 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Figueroa:2009:NBd

[Fig09d] James Figueroa. News briefs.

IEEE Security & Privacy, 7 (6):8–9, November/December 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Figueroa:2010:NBa

[Fig10a] James Figueroa. News briefs.

IEEE Security & Privacy, 8(1): 6–7, January/February 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Figueroa:2010:NBb

[Fig10b] James Figueroa. News briefs.

IEEE Security & Privacy, 8(2): 8–10, March/April 2010. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ferraiolo:2007:RSR

[FKS07] David Ferraiolo, Rick Kuhn, and Ravi Sandhu. RBAC standard rationale: Comments on “A Critique of the ANSI Standard on Role-Based Access Control”. IEEE Security & Privacy, 5(6):51–53, November/December 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Flammer:2016:GWB

[Fla16] Ivo Flammer. Genteel wearables: Bystander-centered design. IEEE Security & Privacy, 14(5):73–79, September/October 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/05/
msp2016050073-abs.html.

Fasano:2019:RLB

[FLDGB19] A. Fasano, T. Leek, B. Dolan-Gavitt, and J. Bundt. The Rode0day to less-buggy programs. IEEE Security & Privacy, 17(6):84–88, November/ December 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Fly:2011:DFW

[Fly11] Robert Fly. Detecting fraud on Websites. IEEE Security & Privacy, 9(6):80–85, November/December 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Figueroa:2008:NSC

[FO08] James Figueroa and Brandi Ortega. News: Shaking up the cybersecurity landscape. IEEE Security & Privacy, 6(6):18– 21, November/December 2008.

CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Fazelnia:2023:SAI

[FOM23] Mohamad Fazelnia, Ahmet Okutan, and Mehdi Mirakhorli. Supporting artificial intelligence/machine learning security workers through an adversarial techniques, tools, and common knowledge framework. IEEE Security & Privacy, 21(1):37–48, January/ February 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Forbes:2004:BRN

[For04a] Scott Forbes. Book reviews: A .NET Gold Mine: .NET Security and Cryptography, by Peter Thorsteinson and G. Gnana Arun Ganesh (Prentice-Hall 2004, ISBN 0-13-100851-X). IEEE Security & Privacy, 2(4): 10, July/August 2004. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/sp/
2004/04/j4010.htm.

Forbes:2004:PLR

[For04b] Scott Forbes. Privacy law resource for students and professionals. IEEE Security & Privacy, 2(6):7, November/December 2004. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/comp/mags/sp/2004/06/
j6007.pdf;   http://csdl.

                                        computer.org/dl/mags/sp/
                                        2004/06/j6007.htm.

Ford:2004:WS

[For04c] Richard Ford. The wrong stuff? IEEE Security & Privacy, 2(3):86–89, May/ June 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/sp/2004/03/j3086.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2004/03/j3086.
                                        pdf.

Ford:2005:MMR

[For05] R. Ford. Malcode mysteries revealed [computer viruses and worms]. IEEE Security & Privacy, 3(3):72–75, May/June 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/31002/01439507.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        31002&arnumber=1439507&count=
                                        20&index=14.

Fichtinger:2012:DSS

[FPP12] Barbara Fichtinger, Frances Paulisch, and Peter Panholzer. Driving secure software development experience in a diverse product environment. IEEE Security & Privacy, 10(2):97– 101, March/April 2012. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Franz:2007:CUT

[Fra07a] Michael Franz. Containing the ultimate Trojan horse. IEEE Security & Privacy, 5(4):52–56, July/August 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Fraser:2007:CRU

[Fra07b] David Fraser. The Canadian response to the USA Patriot Act. IEEE Security & Privacy, 5(5):66–68, September/ October 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Fabro:2010:NGL

[FRA10] Mark Fabro, Tim Roxey, and Michael Assante. No grid left behind. IEEE Security & Privacy, 8(1):72–76, January/ February 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Franz:2018:MMP

[Fra18] Michael Franz. Making multivariant programming practical and inexpensive. IEEE Security & Privacy, 16(3):90– 94, May/June 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/03/
msp2018030090-abs.html.

Fernandes:2017:ITS

[FREP17] Earlence Fernandes, Amir Rahmati, Kevin Eykholt, and Atul Prakash. Internet of Things security research: A rehash of old ideas or new intellectual challenges? IEEE Security & Privacy, 15(4):79– 84, July/August 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/04/
msp2017040079-abs.html.

Frincke:2003:WWS

[Fri03] Deborah Frincke. Who watches the security educators? IEEE Security & Privacy, 1(3):56– 58, May/June 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j3056.pdf.

Furman:2012:BCT

[FTCS12] Susanne M. Furman, Mary Frances Theofanos, Yee-Yin Choong, and Brian Stanton. Basing cybersecurity training on user perceptions. IEEE Security & Privacy, 10(2):40–49, March/ April 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Franken:2019:ECP

[FVJ19] G. Franken, T. Van Goethem, and W. Joosen. Exposing cookie policy flaws through an extensive evaluation of browsers and their extensions. IEEE Security & Privacy, 17 (4):25–34, July/August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Felici:2015:WNE

[FWBC15] Massimo Felici, Nick Wainwright, Fabio Bisogni, and Simona Cavallini. What’s new in the economics of cybersecurity?: Observational and empirical studies. IEEE Security & Privacy, 13(5):12–15, September/October 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/csdl/mags/sp/2015/05/
msp2015050012.html.

Fourne:2023:VHF

[FWFA23] Marcel Fourné, Dominik Wermke, Sascha Fahl, and Yasemin Acar. A viewpoint on human factors in software supply chain security: a research agenda. IEEE Security & Privacy, 21(6):59–63, November/ December 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Goth:2003:Nb

[GA03] Greg Goth and Benjamin J. Alfonsi. News. IEEE Security & Privacy, 1(6):8–13, November/ December 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2003/06/j6008abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/06/j6008.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/06/j6008.
pdf.

Goth:2004:Nb

[GA04a] Greg Goth and Benjamin Alfonsi. News. IEEE Security & Privacy, 2(6):8–9, November/December 2004. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/sp/2004/06/
                                        j6008.pdf;   http://csdl.
                                        computer.org/dl/mags/sp/
                                        2004/06/j6008.htm.

Goth:2004:NRC

[GA04b] Greg Goth and Scott L. Andresen. News: Richard Clarke talks cybersecurity and JELL-O. IEEE Security & Privacy, 2(3):11–15, May/June 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/sp/2004/03/j3011.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2004/03/j3011.
                                        pdf.

Ghosh:2010:GEI

[GA10] Anup Ghosh and Ivan Arce. Guest Editors’ introduction: In cloud computing we trust — but should we? IEEE Security & Privacy, 8(6):14–16, November/December 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2012:SYG

[GA12] Daniel E. Geer and Jerry Archer. Stand your ground. IEEE Security & Privacy, 10 (4):96, July/August 2012. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Goth:2004:NHU

[GAA04] Greg Goth, Benjamin J. Alfonsi, and Scott L. Andresen. News: How useful are attack trend resources? IEEE Security & Privacy, 2(2):9–11, March/April 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/02/j2009.
pdf.

Garfinkel:2003:EBI

[Gar03] Simson L. Garfinkel. Emailbased identification and authentication: An alternative to PKI? IEEE Security & Privacy, 1(6):20–26, November/ December 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2003/06/j6020abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/06/j6020.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/06/j6020.
pdf.

Garfinkel:2008:SFF

[Gar08] Simson Garfinkel. Sharp figures, fuzzy purpose. IEEE Security & Privacy, 6(2): 5, March/April 2008. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.

                                        org/comp/mags/sp/2008/02/
                                        msp2008020005.pdf.

Garber:2011:NBa

[Gar11a] Lee Garber. News briefs. IEEE Security & Privacy, 9(1):15–17, January/February 2011. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Garber:2011:NBb

[Gar11b] Lee Garber. News briefs. IEEE Security & Privacy, 9(2):5–7, March/April 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Garber:2011:NBc

[Gar11c] Lee Garber. News briefs. IEEE Security & Privacy, 9(3):6–8, May/June 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Garber:2011:NBd

[Gar11d] Lee Garber. News briefs. IEEE Security & Privacy, 9(4):6–8, July/August 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Garber:2011:NBe

[Gar11e] Lee Garber. News briefs. IEEE Security & Privacy, 9(5):6–8, September/October 2011. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Garber:2011:NBf

[Gar11f] Lee Garber. News briefs. IEEE Security & Privacy, 9(6):9– 11, November/December 2011.

CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Garber:2012:SPPa

[Gar12a] Lee Garber. Security, privacy, and policy roundup. IEEE Security & Privacy, 10(1):15–17, January/February 2012. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Garber:2012:SPPb

[Gar12b] Lee Garber. Security, privacy, and policy roundup. IEEE Security & Privacy, 10(2):15–17, March/April 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Garber:2012:SPPc

[Gar12c] Lee Garber. Security, privacy, and policy roundup. IEEE Security & Privacy, 10(3):12–13, May/June 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Garber:2012:SPPd

[Gar12d] Lee Garber. Security, privacy, and policy roundup. IEEE Security & Privacy, 10(4):16–18, July/August 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Garber:2012:SPPe

[Gar12e] Lee Garber. Security, privacy, and policy roundup. IEEE Security & Privacy, 10(5):8–9, September/October 2012. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Garber:2012:SPPf

[Gar12f] Lee Garber. Security, privacy, policy, and dependability roundup. IEEE Security & Privacy, 10(6):6–8, November/ December 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Garber:2013:NBSa

[Gar13a] Lee Garber. News briefs: Security, privacy, policy, and dependability roundup. IEEE Security & Privacy, 11(5):6–7, September/October 2013. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Garber:2013:NBSb

[Gar13b] Lee Garber. News briefs: Security, privacy, policy, and dependability roundup. IEEE Security & Privacy, 11(6):8– 9, November/December 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Garber:2013:SPPa

[Gar13c] Lee Garber. Security, privacy, policy, and dependability roundup. IEEE Security & Privacy, 11(1):6–7, January/ February 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Garber:2013:SPPb

[Gar13d] Lee Garber. Security, privacy, policy, and dependability roundup. IEEE Security & Privacy, 11(2):6–7, March/ April 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Garber:2013:SPPc

[Gar13e] Lee Garber. Security, privacy, policy, and dependability roundup. IEEE Security & Privacy, 11(3):6–7, May/June 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Garber:2013:SPPd

[Gar13f] Lee Garber. Security, privacy, policy, and dependability roundup. IEEE Security & Privacy, 11(4):6–7, July/ August 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Garber:2014:SPPa

[Gar14a] Lee Garber. Security, privacy, policy, and dependability roundup. IEEE Security & Privacy, 12(1):9–10, January/ February 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Garber:2014:SPPb

[Gar14b] Lee Garber. Security, privacy, policy, and dependability roundup. IEEE Security & Privacy, 12(2):11–13, March/ April 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Garber:2014:SPPc

[Gar14c] Lee Garber. Security, privacy, policy, and dependability roundup. IEEE Security & Privacy, 12(3):6–8, May/June 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Garber:2014:SPPd

[Gar14d] Lee Garber. Security, privacy, policy, and dependability roundup. IEEE Security & Privacy, 12(4):8– 10, July/August 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/04/
msp2014040008.html.

Garfinkel:2014:LSI

[Gar14e] Simson L. Garfinkel. Leaking sensitive information in complex document files–and how to prevent it. IEEE Security & Privacy, 12(1):20–27, January/ February 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Gattiker:2005:E

[Gat05] Urs E. Gattiker. EICAR 2005. IEEE Security & Privacy, 3(4): 45–48, July/August 2005. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2003:M

[GAW03] Daniel E. Geer, Jr., Dave Aucsmith, and James A. Whittaker. Monoculture. IEEE Security & Privacy, 1(6): 14–17, November/December 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2003/06/j6014abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/06/j6014.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/06/j6014.
pdf.

Goldstein:2009:DCS

[GB09] Adam Goldstein and David Bucciero. The Dartmouth Cyber Security Initiative: Faculty, staff, and students work together. IEEE Security & Privacy, 7(6):57–59, November/ December 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2012:YST

[GB12] Daniel E. Geer and Bob Blakley. Are you smarter than the TSA? (hint: No). IEEE Security & Privacy, 10(4):94–95, July/August 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Gardner:2009:BSP

[GBK09] Ryan W. Gardner, Matt Bishop, and Tadayoshi Kohno. Building security in: Are patched machines really fixed? IEEE Security & Privacy, 7 (5):82–85, September/October 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geiger:2006:SSD

[GC06] Matthew Geiger and Lorrie Faith Cranor. Scrubbing stubborn data: An evaluation of counter-forensic privacy tools. IEEE Security & Privacy, 4(5):16–25, September/ October 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2008:WWG

[GC08a] Dan Geer and Daniel Conway. What we got for Christmas. IEEE Security & Privacy, 6(1): 88, January/February 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://bell.computer.
                                        org/dlcomments/.

Geer:2008:BIM

[GC08b] Daniel E. Geer, Jr. and Daniel G. Conway. Beware the IDs of March. IEEE Security & Privacy, 6(2):87, March/April 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2008:GMS

[GC08c] Daniel E. Geer, Jr. and Daniel G. Conway. For good measure: Security is a subset of reliability. IEEE Security & Privacy, 6(6):86–87, November/December 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2008:GMT

[GC08d] Daniel E. Geer, Jr. and Daniel G. Conway. For good measure: Type II reverse engineering. IEEE Security & Privacy, 6(5):86–87, September/ October 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2008:SA

[GC08e] Daniel E. Geer, Jr. and Daniel G. Conway. Strong attractors. IEEE Security & Privacy, 6(4):78–79, July/August 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2009:GMD

[GC09a] Daniel E. Geer, Jr. and Daniel G. Conway. For good measure: a doubt of the benefit. IEEE Security & Privacy, 7(3):86–87, May/June 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2009:GMH

[GC09b] Daniel E. Geer, Jr. and Daniel G. Conway. For good measure: Hard data is good to find. IEEE Security & Privacy, 7(2):94–95, March/April 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2009:GMR

[GC09c] Daniel E. Geer, Jr. and Daniel G. Conway. For good measure: Risk concentration. IEEE Security & Privacy, 7 (5):86–87, September/October 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2009:GMO

[GC09d] Daniel E. Geer, Jr. and Daniel G. Conway. For good measure: The owned price index. IEEE Security & Privacy, 7(1):86–87, January/February 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2009:PGP

[GC09e] Daniel E. Geer, Jr. and Daniel G. Conway. Patch grief with proverbs. IEEE Security & Privacy, 7(6):86–87, November/December 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2010:LSH

[GC10a] Daniel E. Geer, Jr. and Daniel G. Conway. A life is short, a half-life is forever. IEEE Security & Privacy, 8(1): 86–87, January/February 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2010:NVN

[GC10b] Daniel E. Geer, Jr. and Daniel G. Conway. Nothing ventured, nothing gained. IEEE Security & Privacy, 8(2): 86–87, March/April 2010. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2013:TTC

[GC13] Daniel E. Geer, Jr. and Daniel G. Conway. The times, they are a changin’. IEEE Security & Privacy, 11(1):94–95, January/February 2013. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Giechaskiel:2018:WCC

[GCR18] Ilias Giechaskiel, Cas Cremers, and Kasper B. Rasmussen. When the crypto in cryptocurrencies breaks: Bitcoin security under broken primitives. IEEE Security & Privacy, 16 (4):46–56, July/August 2018. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/04/
msp2018040046-abs.html.

Grimes:2022:PBS

[GCY^{+}22] Howard D. Grimes, Gabriela F.

Ciocarlie, Bo Yu, Duminda Wijesekera, Greg Shannon, Wayne Austad, Charles Fracchia, Dongyan Xu, Thomas R. Kurfess, Lisa Strama, Michael Mylrea, and Bill Reid. PURE biomanufacturing: Secure, pandemic-adaptive biomanufacturing. IEEE Security & Privacy, 20(6):53–65, November/December 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Garfinkel:2011:ODS

[GD11] Simson L. Garfinkel and George Dinolt. Operations with degraded security. IEEE Security & Privacy, 9(6):43– 48, November/December 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Gurses:2013:TTP

[GD13] Seda Gurses and Claudia Diaz. Two tales of privacy in online social networks. IEEE Security & Privacy, 11(3):29–37, May/June 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2017:MDD

[GD17] Daniel E. Geer and Richard Danzig. Mutual dependence demands mutual sharing. IEEE Security & Privacy, 15(1):87–88, January/ February 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/01/
                                        msp2017010087.html.

Goodman:2018:PPF

[GD18] Peter Goodman and Artem Dinaburg. The past, present, and future of Cyberdyne. IEEE Security & Privacy, 16(2):61– 69, March/April 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/02/
                                        msp2018020061-abs.html.

Geer:2005:WPS

[Gee05a] Dan Geer. When is a product a security product? IEEE Security & Privacy, 3(5):80, September/October 2005. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://csdl.computer.

org/comp/mags/sp/2005/05/
j5080.pdf.

Geer:2005:PSP

[Gee05b]  D. Geer, Jr.  The prob-

lem statement is the problem. IEEE Security & Privacy, 3(2):80, March/April 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/30742/01423969.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
30742&arnumber=1423969&count=
15&index=14.

Geer:2006:CA

[Gee06a] Daniel E. Geer, Jr. Convergence (abstract). IEEE Security & Privacy, 4(3):88, May/ June 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2006:EE

[Gee06b] Daniel E. Geer, Jr. Evidently evidentiary. IEEE Security & Privacy, 4(6):96, November/ December 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2007:EBW

[Gee07] Daniel E. Geer. The end of black and white. IEEE Security & Privacy, 5(4):88, July/ August 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2008:CTC

[Gee08a] Daniel E. Geer, Jr. Clear text: Complexity is the enemy. IEEE Security & Privacy, 6(6): 88, November/December 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2008:LAT

[Gee08b] Daniel E. Geer, Jr. Learn by analogy or die trying. IEEE Security & Privacy, 6(3):88, 87, May/June 2008. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://bell.computer.
                                        org/dlcomments/.

Geer:2009:CDE

[Gee09a] Daniel E. Geer, Jr. Clear-Text: Digital endosymbiosis. IEEE Security & Privacy, 7(3): 88, May/June 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2009:DDS

[Gee09b] Daniel E. Geer, Jr. Deskilling digital security. IEEE Security & Privacy, 7(6):88, November/ December 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2010:F

[Gee10a] Daniel E. Geer, Jr. Fratricide. IEEE Security & Privacy, 8(3): 88, 87, May/June 2010. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2010:IC

[Gee10b] Daniel E. Geer, Jr. An index of cybersecurity. IEEE Security & Privacy, 8(6):96, 95, November/December 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2010:TR

[Gee10c] Daniel E. Geer, Jr. A time to rethink. IEEE Security & Privacy, 8(4):86–87, July/August 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2011:ASI

[Gee11a] Daniel E. Geer, Jr. Attack surface inflation. IEEE Security & Privacy, 9(4):85–86, July/ August 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2011:CC

[Gee11b] Daniel E. Geer, Jr. Correlation is not causation. IEEE Security & Privacy, 9(2):93–94, March/April 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2011:DRT

[Gee11c] Daniel E. Geer, Jr. Does a rising tide lift all boats? IEEE Security & Privacy, 9(1):93–94, January/February 2011. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2011:ER

[Gee11d] Daniel E. Geer, Jr. Eisenhower revisited. IEEE Security & Privacy, 9(4):88, 87, July/ August 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2011:NM

[Gee11e] Daniel E. Geer, Jr. New measures. IEEE Security & Privacy, 9(3):86–87, May/June 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2011:SBB

[Gee11f] Daniel E. Geer, Jr. Small is beautiful, big is inevitable. IEEE Security & Privacy, 9(6): 86–87, November/December 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2011:TC

[Gee11g] Daniel E. Geer, Jr. A time for choosing. IEEE Security & Privacy, 9(1):96, 95, January/ February 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2012:ML

[Gee12a] Daniel E. Geer, Jr. More or less. IEEE Security & Privacy, 10(1):96, January/ February 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2012:NWH

[Gee12b] Daniel E. Geer, Jr. Numbers worth having. IEEE Security & Privacy, 10(2):102–103, March/April 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2012:PL

[Gee12c] Daniel E. Geer, Jr. Power.

law. IEEE Security & Privacy, 10(1):94–95, January/ February 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2012:RA

[Gee12d] Daniel E. Geer, Jr. Risk aversion. IEEE Security & Privacy, 10(5):86–87, September/ October 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2013:A

[Gee13a] Daniel E. Geer. On abandonment. IEEE Security & Privacy, 11(4):96, July/August 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2013:LWI

[Gee13b] Daniel E. Geer, Jr. Last word:

Identity as privacy. IEEE Security & Privacy, 11(1):96, January/February 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2014:LWP

[Gee14a] Daniel E. Geer. Last word: Polarization. IEEE Security & Privacy, 12(1):88, January/ February 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2014:PDG

[Gee14b] Daniel E. Geer. Personal data and government surveillance. IEEE Security & Privacy, 12 (4):90–96, July/August 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/04/
msp2014040090.html.

Geer:2015:CM

[Gee15a] Daniel E. Geer. Children of the magenta. IEEE Security & Privacy, 13(5):104, September/October 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/csdl/mags/sp/2015/05/
msp2015050104.html.

Geer:2015:RU

[Gee15b] Daniel E. Geer. The right to be unobserved. IEEE Security & Privacy, 13(4): 88, July/August 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/04/
msp2015040088.html.

Geer:2015:LMS

[Gee15c] Daniel E. Geer, Jr. Less is more: Saving the Internet from itself. IEEE Security & Privacy, 13(1):80, January/February 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/01/
msp2015010080.html.

Geer:2016:PP

[Gee16a] Daniel E. Geer. Privacy’s paradigm. IEEE Security & Privacy, 14(4):87– 88, July/August 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/04/
msp2016040087.html.

Geer:2016:P

[Gee16b] Daniel E. Geer. Provenance.

IEEE Security & Privacy, 14 (2):88, March/April 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2017:A

[Gee17] Daniel E. Geer. Attribution.

IEEE Security & Privacy, 15 (4):88, July/August 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/04/
msp2017040088.html.

Geer:2018:TP

[Gee18a]  Daniel E. Geer.  Trad-

ing places. IEEE Security & Privacy, 16(1):104, January/February 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/01/
                                        msp2018010104.html.

Geer:2018:YWY

[Gee18b] Daniel E. Geer. You are what you eat. IEEE Security & Privacy, 16(4):96, July/August 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/04/
                                        msp2018040096.html.

Geer:2019:O

[Gee19a] D. E. Geer. Ownership. IEEE Security & Privacy, 17(4):100– 99, July/August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2019:UU

[Gee19b] D. E. Geer, Jr. Unknowable unknowns. IEEE Security & Privacy, 17(2):80–79, March/ April 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2020:STB

[Gee20] D. E. Geer. Security theater, the beat goes on. IEEE Security & Privacy, 18(4):75–76, July/August 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2021:AUC

[Gee21a] Daniel E. Geer. Auto-update considered harmful. IEEE Security & Privacy, 19(2):79–80, March/April 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2021:C

[Gee21b]  Daniel E. Geer.  Conver-

gence. IEEE Security & Privacy, 19(6):123–124, November/December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2022:I

[Gee22] Daniel E. Geer. Identity. IEEE Security & Privacy, 20(4):71– 72, July/August 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2023:CO

[Gee23a] Daniel E. Geer. Convergence:

Ongoing. IEEE Security & Privacy, 21(5):95–96, September/ October 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2023:WS

[Gee23b] Daniel E. Geer. Whither software? IEEE Security & Privacy, 21(2):89–90, March/ April 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Gellman:2014:WWL

[Gel14] Robert Gellman. Willis Ware’s lasting contribution to privacy: Fair information practices. IEEE Security & Privacy, 12 (4):51–54, July/August 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/04/
msp2014040051-abs.html.

Gennaro:2006:RC

[Gen06] Rosario Gennaro. Randomness in cryptography. IEEE Security & Privacy, 4(2):64–67, March/April 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Gu:2018:FCP

[GFJ^{+}18] Q. Gu, D. Formby, S. Ji, H. Cam, and R. Beyah. Fingerprinting for cyber-physical system security: Device physics matters too. IEEE Security & Privacy, 16(5):49–59, September/October 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Goth:2004:NVS

[GG04] Greg Goth and Pam Frost Gorder. News: E-voting security: The electoral dialect gets hot; balancing video-game piracy issues. IEEE Security & Privacy, 2(1):14–17, January/ February 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2004/01/j1014abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/01/j1014.
pdf.

Gutmann:2005:SU

[GG05] Peter Gutmann and Ian Grigg.

Security usability. IEEE Security & Privacy, 3(4):56–58, July/August 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Greamo:2011:SVM

[GG11a] Chris Greamo and Anup Ghosh. Sandboxing and virtualization: Modern tools for combating malware. IEEE Security & Privacy, 9(2):79–82, March/April 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Grigg:2011:CCN

[GG11b] Ian Grigg and Peter Gutmann.

The curse of cryptographic numerology. IEEE Security & Privacy, 9(3):70–72, May/June 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Gueye:2021:DRS

[GGBM21] Assane Gueye, Carlos E. C.

Galhardo, Irena Bojanova, and Peter Mell. A decade of reoccurring software weaknesses. IEEE Security & Privacy, 19(6):74–82, November/ December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Geva:2014:BDD

[GHG14] Moti Geva, Amir Herzberg, and Yehoshua Gev. Bandwidth distributed denial of service:

Attacks and defenses. IEEE Security & Privacy, 12(1):54– 61, January/February 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Gligor:2006:IAT

[GHK^{+}06] Virgil D. Gligor, Tom Haigh, Dick Kemmerer, Carl Landwehr, Steve Lipner, and John McLean. Information assurance technology forecast 2005. IEEE Security & Privacy, 4 (1):62–69, January/February 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/33481/01588830.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
33481&arnumber=1588830.

Grosse:2010:CCR

[GHR^{+}10] Eric Grosse, John Howie, James Ransome, Jim Reavis, and Steve Schmidt. Cloud computing roundtable. IEEE Security & Privacy, 8(6):17– 23, November/December 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Gilad:2014:PHI

[GHS14] Yossi Gilad, Amir Herzberg, and Haya Shulman. Offpath hacking: The illusion of challenge–response authentication. IEEE Security & Privacy, 12(5):68–77, September/October 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/05/
msp2014050068-abs.html.

Gidari:2006:DRW

[Gid06] Albert Gidari. Designing the right wiretap solution: Setting standards under CALEA. IEEE Security & Privacy, 4(3): 29–36, May/June 2006. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Giffin:2010:NMB

[Gif10] Jon Giffin. The next malware battleground: Recovery after unknown infection. IEEE Security & Privacy, 8(3):74–76, May/June 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Garfinkel:2005:RPO

[GJP05] S. L. Garfinkel, A. Juels, and R. Pappu. RFID privacy: an overview of problems and proposed solutions. IEEE Security & Privacy, 3(3):34–43, May/ June 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/31002/01439500.pdf;
http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
31002&arnumber=1439500&count=
20&index=7.

Geer:2011:WBE

[GK11] Daniel E. Geer, Jr. and Peter Kuper. When $80 billion is not enough. IEEE Security & Privacy, 9(5):86–87, September/ October 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2014:IMH

[GK14] Daniel E. Geer and Poul-Henning Kamp. Inviting more heartbleed. IEEE Security & Privacy, 12(4):46– 50, July/August 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/04/
msp2014040046-abs.html.

Gervais:2014:BDC

[GKCC14] Arthur Gervais, Ghassan O. Karame, Vedran Capkun, and Srdjan Capkun. Is Bitcoin a decentralized currency? IEEE Security & Privacy, 12(3):54– 60, May/June 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Gruteser:2004:PPC

[GL04] Marco Gruteser and Xuan Liu. Protecting privacy in continuous location-tracking applications. IEEE Security & Privacy, 2(2):28–34, March/ April 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2004/02/j2028abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/02/j2028.
pdf.

Geer:2012:PI

[GL12] Daniel E. Geer, Jr. and Daniel B. Larremore. Progress is infectious. IEEE Security & Privacy, 10(6):94–95, November/December 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Goldrich:2014:MTG

[GL14] Luanne Goldrich and Carl E.

Landwehr. Moving target [Guest Editors’ introduction]. IEEE Security & Privacy, 12 (2):14–15, March/April 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Gladyshev:2019:CCU

[Gla19] P. Gladyshev. Cybercrime as a consequence of unreasonable expectations. IEEE Security & Privacy, 17(4):84–87, July/ August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Gunter:2011:EBA

[GLM11] Carl A. Gunter, David M.

Liebovitz, and Bradley Malin. Experience-based access management: a life-cycle framework for identity and access management systems. IEEE Security & Privacy, 9(5):48–55, September/October 2011. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Gorski:2023:ELU

[GLS23] Peter Leo Gorski, Luigi Lo Iacono, and Matthew Smith.

Eight lightweight usable security principles for developers. IEEE Security & Privacy, 21(1):20–26, January/ February 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Goth:2003:Na

[GM03] Greg Goth and Sandra Kay Miller. News. IEEE Security & Privacy, 1(5):8–11, September/October 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2003/05/j5008.
pdf.

Garfinkel:2009:NXB

[GM09] Simson L. Garfinkel and James J. Migletz. New XMLbased files implications for forensics. IEEE Security & Privacy, 7(2):38–44, March/April 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ghosh:2012:LDG

[GM12] Anup Ghosh and Gary Mc-Graw. Lost decade or golden era: Computer security since 9/11. IEEE Security & Privacy, 10(1):6–10, January/ February 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Garfinkel:2014:CWS

[GM14] Simson L. Garfinkel and Michael McCarrin. Can we sniff wi-fi?: Implications of Joffe v. Google. IEEE Security & Privacy, 12(4):22– 28, July/August 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/04/
msp2014040022-abs.html.

Gavas:2012:WCO

[GMB12] Efstratios Gavas, Nasir Memon, and Douglas Britton. Winning cybersecurity one challenge at a time. IEEE Security & Privacy, 10(4):75–79, July/ August 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Greitzer:2008:CIC

[GMC^{+}08] Frank L. Greitzer, Andrew P.

Moore, Dawn M. Cappelli, Dee H. Andrews, Lynn A. Carroll, and Thomas D. Hull. Combating the insider cyber threat. IEEE Security & Privacy, 6(1):61–64, January/ February 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Gratzer:2006:CLE

[GN06] Vanessa Gratzer and David Naccache. Cryptography, law enforcement, and mobile communications. IEEE Security & Privacy, 4(6):67–70, November/December 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Gratzer:2007:AVQ

[GN07a] Vanessa Gratzer and David Naccache. Alien vs. Quine. IEEE Security & Privacy, 5(2): 26–31, March/April 2007. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Gratzer:2007:TNS

[GN07b] Vanessa Gratzer and David Naccache. Trust on a nationwide scale. IEEE Security & Privacy, 5(5):69–71, September/October 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Gutmann:2005:WHC

[GNP05] P. Gutmann, D. Naccache, and C. C. Palmer. When hashes collide [applied cryptography]. IEEE Security & Privacy, 3(3):68–71, May/June 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/31002/01439506.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        31002&arnumber=1439506&count=
                                        20&index=13.

Garcia:2012:ERP

[GOPB12] Sergio Sanchez Garcia, Ana Gomez Oliva, and Emilia Perez-Belleboni. Is Europe ready for a pan-European identity management system? IEEE Security & Privacy, 10(4):44–49, July/August 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Gordon:2006:UAV

[Gor06] Sarah Gordon. Understanding the adversary: Virus writers and beyond. IEEE Security & Privacy, 4(5):67–70, September/October 2006. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Goth:2004:Na

[Got04] Greg Goth. News. IEEE Security & Privacy, 2(5):8–11, September/October 2004. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/comp/mags/sp/2004/05/
j5008.pdf;   http://csdl.
computer.org/dl/mags/sp/
2004/05/j5008.htm.

Geer:2012:IU

[GP12] Daniel E. Geer, Jr. and Mukul Pareek. ICS update. IEEE Security & Privacy, 10(3):93–95, May/June 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geer:2019:FD

[GP19] D. E. Geer and D. Peterson.

Failure as design. IEEE Security & Privacy, 17(6):90– 89, November/December 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Gondree:2013:STP

[GPD13] Mark Gondree, Zachary N. J.

Peterson, and Tamara Denning. Security through play. IEEE Security & Privacy, 11 (3):64–67, May/June 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Gorbenko:2013:TOI

[GR13] Anatoliy Gorbenko and Alexander Romanovsky. Time-outing Internet services. IEEE Security & Privacy, 11(2):68–71, March/April 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Gradon:2013:CSI

[Gra13] Kacper Gradon. Crime science and the Internet battlefield: Securing the analog world from digital crime. IEEE Security & Privacy, 11(5):93–95, September/October 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Gradon:2023:ESP

[Gra23] Kacper T. Gradon. Electric sheep on the pastures of disinformation and targeted phishing campaigns: The security implications of Chat-GPT. IEEE Security & Privacy, 21(3):58–61, May/June 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Green:2013:TC

[Gre13] Matthew Green. The threat in the cloud. IEEE Security & Privacy, 11(1):86–89, January/ February 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Grimaila:2004:EMB

[Gri04] Michael Russell Grimaila. Education: Maximizing business information security’s educational value. IEEE Security & Privacy, 2(1):56–60, January/ February 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2004/01/j1056abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/01/j1056.
pdf.

Grossman:2008:ATF

[Gro08] Jeremiah Grossman. Attack trends: Five user-customizable Web site security features. IEEE Security & Privacy, 6(6): 79–81, November/December 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Grossman:2012:SWS

[Gro12] Jeremiah Grossman. The state of Website security. IEEE Security & Privacy, 10(4):91–93, July/August 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Grossman:2014:EUN

[Gro14] Wendy M. Grossman. “Emergency” ushers in a new era in British communications surveillance. IEEE Security & Privacy, 12(6):84–88, November/December 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/sp/2014/06/
                                        msp2014060084-abs.html.

Grotto:2020:DCA

[Gro20] A. Grotto. Deconstructing cyber attribution: A proposed framework and lexicon. IEEE Security & Privacy, 18(1):12– 20, January/February 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Grumbach:2016:CDI

[Gru16] Stephane Grumbach. Chasing data in the intermediation era: Economy and security at stake. IEEE Security & Privacy, 14(3):22–31, May/June 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Garfinkel:2003:RDP

[GS03] Simson L. Garfinkel and Abhi Shelat. Remembrance of data passed: a study of disk sanitization practices. IEEE Security & Privacy, 1(1):17–27, January/February 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://dlib.computer.
                                        org/sp/books/sp2003/pdf/
                                        j1017.pdf;   http://www.
                                        computer.org/security/j1017abs.
                                        htm.

Garfinkel:2006:GEI

[GS06] Simson Garfinkel and Michael D. Smith. Guest Editors’ introduction: Data surveillance. IEEE Security & Privacy, 4(6):

15–17, November/December 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Geiselmann:2007:SPH

[GS07] Willi Geiselmann and Rainer Steinwandt. Special-purpose hardware in cryptanalysis: The case of 1,024-bit RSA. IEEE Security & Privacy, 5(1):63–66, January/February 2007. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Green:2016:DEN

[GS16] Matthew Green and Matthew Smith. Developers are not the enemy!: The need for usable security APIs. IEEE Security & Privacy, 14(5):40–46, September/October 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/05/
msp2016050040-abs.html.

Guida:2004:DUP

[GSB^{+}04] Richard Guida, Robert Stahl, Thomas Bunt, Gary Secrest, and Joseph Moorcones. Deploying and using public key technology: Lessons learned in real life. IEEE Security & Privacy, 2(4):67–71, July/August 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/04/j4067.
htm;   http://csdl.computer.

                                        org/dl/mags/sp/2004/04/j4067.
                                        pdf.

Goirizelaia:2008:OSV

[GSHU08] Iñaki Goirizelaia, Ted Selker, Maider Huarte, and Juanjo Unzilla. An optical scan E-voting system based on N-version programming. IEEE Security & Privacy, 6(3):47–53, May/June 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Gugelmann:2015:CCB

[GSLA15] David Gugelmann, Pascal Studerus, Vincent Lenders, and Bernhard Ager. Can content-based data loss prevention solutions prevent data leakage in Web traffic? IEEE Security & Privacy, 13(4):52– 59, July/August 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/sp/2015/04/
                                        msp2015040052-abs.html.

Gouert:2022:DMU

[GT22] Charles Gouert and Nektarios Georgios Tsoutsos. Dirty metadata: Understanding a threat to online privacy. IEEE Security & Privacy, 20(6):27– 34, November/December 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Gagnon:2007:SPT

[GTG07] Michael N. Gagnon, Stephen Taylor, and Anup K. Ghosh. Software protection through anti-debugging. IEEE Security & Privacy, 5(3):82–84, May/ June 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Grosse:2013:AS

[GU13] Eric Grosse and Mayank Upadhyay. Authentication at scale. IEEE Security & Privacy, 11(1):15–22, January/ February 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://www.computer.org/
cms/Computer.org/ComputingNow/
pdfs/AuthenticationAtScale.
pdf.

Gueron:2016:MEG

[Gue16] Shay Gueron. Memory encryption for general-purpose processors. IEEE Security & Privacy, 14(6):54–62, November/December 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/06/
msp2016060054-abs.html.

Guido:2011:CSI

[Gui11] Dan Guido. A case study of intelligence-driven defense. IEEE Security & Privacy, 9(6): 67–70, November/December 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Gupta:2017:MLP

[Gup17] Chetan Gupta. The market’s law of privacy: Case studies in privacy and security adoption. IEEE Security & Privacy, 15(3):78–83, May/June 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/03/
                                        msp2017030078-abs.html.

Gupta:2007:UAG

[GW07] Suvajit Gupta and Joel Winstead. Using attack graphs to design systems. IEEE Security & Privacy, 5(4):80–83, July/ August 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Grobauer:2011:UCC

[GWS11] Bernd Grobauer, Tobias Walloschek, and Elmar Stocker. Understanding cloud computing vulnerabilities. IEEE Security & Privacy, 9(2):50–57, March/April 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Goh:2013:TOT

[GY13] Weihan Goh and Chai Kiat Yeo. Teaching an old TPM new tricks: Repurposing for identity-based signatures. IEEE Security & Privacy, 11 (5):28–35, September/October 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Genkin:2021:WMM

[GY21] D. Genkin and Y. Yarom. Whack-a-Meltdown: Microarchitectural security games [systems attacks and defenses]. IEEE Security & Privacy, 19 (1):95–98, January/February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Huberman:2005:VP

[HAF05] Bernardo A. Huberman, Eytan Adar, and Leslie R. Fine. Valuating privacy. IEEE Security & Privacy, 3(5):22–25, September/October 2005. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Hagen:2009:EHR

[Hag09] Janne Hagen. Education: Human relationships: a neverending security education challenge? IEEE Security & Privacy, 7(4):65–67, July/August 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Halderman:2010:SSC

[Hal10]  J. Alex Halderman.  To

strengthen security, change developers’ incentives. IEEE Security & Privacy, 8(2):79–82, March/April 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Handschuh:2012:HASa

[Han12a] Helena Handschuh. Hardwareanchored security based on SRAM PUFs, Part 1. IEEE Security & Privacy, 10(3):80–83, May/June 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Handschuh:2012:HASb

[Han12b] Helena Handschuh. Hardwareanchored security based on SRAM PUFs, Part 2. IEEE Security & Privacy, 10(4):80– 81, July/August 2012. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Huaman:2022:TWD

[HAO^{+}22] Nicolas Huaman, Sabrina Amft, Marten Oltrogge, Yasemin Acar, and Sascha Fahl. They would do better if they worked together: Interaction problems between password managers and the Web. IEEE Security & Privacy, 20(2):49–60, March/ April 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Haverkort:2013:DSS

[Hav13] Boudewijn R. Haverkort. The dependable systems-of-systems design challenge. IEEE Security & Privacy, 11(5):62–65, September/October 2013. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Hayes:2013:NSA

[Hay13] Adrian Hayes. Network service authentication timing attacks. IEEE Security & Privacy, 11 (2):80–82, March/April 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Huck:2022:WDF

[HB22] Jan Huck and Frank Breitinger. Wake up digital forensics community and help combat ransomware. IEEE Security & Privacy, 20(4):61–70, July/August 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Hay:2009:LAP

[HBN09] Brian Hay, Matt Bishop, and Kara Nance. Live analysis: Progress and challenges. IEEE Security & Privacy, 7(2):30–37, March/April 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Hoffman:2012:HBC

[HBT12] Lance J. Hoffman, Diana L.

Burley, and Costis Toregas. Holistically building the cybersecurity workforce. IEEE Security & Privacy, 10(2):33–39, March/April 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Hughes:2009:SDS

[HCC09] Gordon F. Hughes, Tom Coughlin, and Daniel M. Commins. Secure data sanitization: Disposal of disk and tape data by secure sanitization. IEEE Security & Privacy, 7(4):29–34, July/August 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Hubaux:2004:SPS

[HCL04] Jean-Pierre Hubaux, Srdjan Capkun, and Jun Luo. The security and privacy of smart vehicles. IEEE Security & Privacy, 2(3):49–55, May/June 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/03/j3049.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/03/j3049.
pdf.

Han:2011:TDE

[HCL11] ShuiHua Han, Chao-Hsien Chu, and Zongwei Luo. Tamper detection in the EPC network using digital watermarking. IEEE Security & Privacy, 9(5):62–69, September/ October 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Hao:2018:VCV

[HCRS18] Feng Hao, Dylan Clarke, Brian Randell, and Siamak F. Shahandashti. Verifiable classroom voting in practice. IEEE Security & Privacy, 16(1):72–81, January/February 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/01/
msp2018010072.html.

Hughes:2008:CNP

[HDB08] Gordon Hughes, Sophie Dawson, and Tim Brookes. Considering new privacy laws in Australia. IEEE Security & Privacy, 6(3):57–59, May/June 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Hearn:2003:IPC

[Hea03a]  Jim Hearn.  International

participation: The continuing march toward security and privacy. IEEE Security & Privacy, 1(1):79–81, January/February 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j1079.pdf;   http://www.
computer.org/security/j1079abs.
htm.

Hearn:2003:MF

[Hea03b] Jim Hearn. Moving forward?

IEEE Security & Privacy, 1(2): 70–71, March/April 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j2070.pdf;   http://www.
computer.org/security/j2070abs.
htm.

Hearn:2003:SD

[Hea03c] Jim Hearn. Slow dancing.

IEEE Security & Privacy, 1(3): 67–68, May/June 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j3067.pdf.

Hearn:2003:WW

[Hea03d] Jim Hearn. What works?

IEEE Security & Privacy, 1 (5):76–77, September/October 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2003/05/j5076abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/05/j5076.
pdf.

Hearn:2004:GPD

[Hea04] Jim Hearn. Global perspectives: Does the common criteria paradigm have a future? IEEE Security & Privacy, 2(1):64–65, January/ February 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2004/01/j1064abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/01/j1064.
pdf.

Hebert:2022:TMI

[Heb22] Cédric Hebert. Trust me, I’m a liar. IEEE Security & Privacy, 20(6):79–82, November/ December 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Heckman:2003:TVS

[Hec03] Carey Heckman. Two views on security software liability: Using the right legal tools. IEEE Security & Privacy, 1(1):73–75, January/February 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j1073.pdf;   http://www.
computer.org/security/j1073abs.
htm.

Heckle:2011:SDH

[Hec11] Rosa R. Heckle. Security dilemma: Healthcare clinicians at work. IEEE Security & Privacy, 9(6):14–19, November/ December 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Heelan:2011:VDS

[Hee11] Sean Heelan. Vulnerability detection systems: Think cyborg, not robot. IEEE Security & Privacy, 9(3):74–77, May/June 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Heikkila:2007:ESC

[Hei07] Faith M. Heikkila. Encryption: Security considerations for portable media devices. IEEE Security & Privacy, 5(4): 22–27, July/August 2007. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Heimes:2016:GIB

[Hei16] Rita Heimes. Global InfoSec and breach standards. IEEE Security & Privacy, 14(5):68– 72, September/October 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/05/
msp2016050068-abs.html.

Henning:2009:BTP

[Hen09] Ronda Henning. Basic training: Predictable surprises. IEEE Security & Privacy, 7(4):

74–76, July/August 2009. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Herzog:2006:APA

[Her06] Jonathan Herzog. Applying protocol analysis to security device interfaces. IEEE Security & Privacy, 4(4):84–87, July/August 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Herley:2013:WDT

[Her13] Cormac Herley. When does targeting make sense for an attacker? IEEE Security & Privacy, 11(2):89–92, March/ April 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Herley:2014:MA

[Her14] Cormac Herley. More is not the answer. IEEE Security & Privacy, 12(1):14–19, January/ February 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Herardian:2019:SUC

[Her19] R. Herardian. The soft underbelly of cloud security. IEEE Security & Privacy, 17(3):90– 93, May/June 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Hermann:2022:WAE

[Her22] Ben Hermann. What has artifact evaluation ever done for us? IEEE Security & Privacy, 20(5):96–99, September/ October 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Hernandez:2019:TAF

[HFT^{+}19] G. Hernandez, F. Fowze, D. J.

Tang, T. Yavuz, P. Traynor, and K. R. B. Butler. Toward automated firmware analysis in the IoT era. IEEE Security & Privacy, 17(5):38–46, September/October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Henry:2018:BAP

[HHK18] Ryan Henry, Amir Herzberg, and Aniket Kate. Blockchain access privacy: Challenges and directions. IEEE Security & Privacy, 16(4):38– 45, July/August 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/04/
msp2018040038-abs.html.

Harding:2008:DSA

[HJK08] Patrick Harding, Leif Johansson, and Nate Klingenstein. Dynamic Security Assertion Markup Language: Simplifying single sign-on. IEEE Security & Privacy, 6(2):83–85, March/April 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Harauz:2009:IAD

[HK09] John Harauz and Lori M. Kaufman. It all depends: a new era of presidential security: The President and his BlackBerry.

IEEE Security & Privacy, 7(2): 67–70, March/April 2009. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Hubaux:2017:GDP

[HKM17] Jean-Pierre Hubaux, Stefan Katzenbeisser, and Bradley Malin. Genomic data privacy and security: Where we stand and where we are heading. IEEE Security & Privacy, 15(5):10–12, September/October 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/05/
msp2017050010.html.

Hamilton:2024:SDI

[HKMR24] Vaughn Hamilton, Gabriel Kaptchuk, Allison McDonald, and Elissa M. Redmiles. Safer digital intimacy for sex workers and beyond: a technical research agenda. IEEE Security & Privacy, 22(2):17–28, March/April 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Hole:2009:APR

[HKN^{+}09] Kjell J. Hole, André N. Klingsheim, Lars-Helge Netland, Yngve Espelid, Thomas TjØstheim, and VebjØrn Moen. Assessing PKI: Risk assessment of a National Security Infrastructure. IEEE Security & Privacy, 7(1):34–41, January/February 2009. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Hiltgen:2006:SIB

[HKW06] Alain Hiltgen, Thorsten Kramp, and Thomas Weigold. Secure Internet banking authentication. IEEE Security & Privacy, 4(2):21–29, March/April 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Howard:2003:IWS

[HL03] Michael Howard and Steve Lipner. Inside the Windows security push. IEEE Security & Privacy, 1(1):57–61, January/February 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j1057.pdf;   http://www.
computer.org/security/j1057abs.
htm.

Hargittai:2013:NSE

[HL13] Eszter Hargittai and Eden Litt.

New strategies for employment? Internet skills and online privacy practices during people’s job search. IEEE Security & Privacy, 11(3):38–45, May/June 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Haney:2021:CAF

[HLJ21] Julie Haney, Wayne Lutters, and Jody Jacobs. Cybersecurity advocates: Force multipliers in security behavior change. IEEE Security & Privacy, 19 (4):54–59, July/August 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Herzberg:2021:SMA

[HLS^{+}21] Amir Herzberg, Hemi Leibowitz, Kent Seamons, Elham Vaziripour, Justin Wu, and Daniel Zappala. Secure messaging authentication ceremonies are broken. IEEE Security & Privacy, 19(2):29–37, March/April 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Herzberg:2012:TJA

[HM12] Amir Herzberg and Ronen Margulies. Training Johnny to authenticate (safely). IEEE Security & Privacy, 10(1):37–45, January/February 2012. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Hilburn:2013:BSR

[HM13] Thomas B. Hilburn and Nancy R. Mead. Building security in: A road to competency. IEEE Security & Privacy, 11(5):89–92, September/ October 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Heymann:2022:SSP

[HM22] Elisa R. Heymann and Barton P. Miller. Software security for the people: Free and open resources for software security training. IEEE Security & Privacy, 20(2):88–95, March/April 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Hope:2004:MAC

[HMA04] Paco Hope, Gary McGraw, and Annie I. Antón. Misuse and abuse cases: Getting past the positive. IEEE Security & Privacy, 2(3):90– 92, May/June 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/03/j3090.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/03/j3090.
pdf.

Heiser:2012:TTS

[HMK12] Gernot Heiser, Toby Murray, and Gerwin Klein. It’s time for trustworthy systems. IEEE Security & Privacy, 10(2):67–70, March/April 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Hole:2007:LNA

[HMKT07] Kjell J. Hole, Vebjørn Moen, André N. Klingsheim, and Knut M. Tande. Lessons from the Norwegian ATM system. IEEE Security & Privacy, 5(6):25–31, November/ December 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Hohenberger:2015:OAL

[HMPS15] Susan Hohenberger, Steven Myers, Rafael Pass, and Abhi Shelat. An overview of AN-ONIZE: A large-scale anonymous survey system. IEEE Security & Privacy, 13(2):22– 29, March/April 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/02/
msp2015020022-abs.html.

Holz:2006:NTA

[HMR06] Thorsten Holz, Simon Marechal, and Frédéric Raynal. New threats and attacks on the World Wide Web. IEEE Security & Privacy, 4(2):72–75, March/April 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Hole:2006:CSO

[HMT06] Kjell J. Hole, Vebjørn Moen, and Thomas Tjøstheim. Case study: Online banking security. IEEE Security & Privacy, 4(2): 14–20, March/April 2006. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Hole:2010:TRA

[HN10] Kjell Hole and Lars-Helge Netland. Toward risk assessment of large-impact and rare events. IEEE Security & Privacy, 8(3): 21–27, May/June 2010. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Hole:2008:OWN

[HNE^{+}08] Kjell J. Hole, Lars-Helge Netland, Yngve Espelid, André N. Klingsheim, Hallvar Helleseth, and Jan B. Henriksen. Open wireless networks on university campuses. IEEE Security & Privacy, 6(4):14–20, July/ August 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Holz:2005:SVB

[Hol05] T. Holz. A short visit to the bot zoo [malicious bots software]. IEEE Security & Privacy, 3(3):76–79, May/June 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/31002/01439508.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
31002&arnumber=1439508&count=
20&index=15.

Hole:2015:DRI

[Hol15a] Kjell Jorgen Hole. Diversity reduces the impact of malware. IEEE Security & Privacy, 13 (3):48–54, May/June 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/03/
msp2015030048-abs.html.

Hole:2015:TAF

[Hol15b] Kjell Jorgen Hole. Toward anti-fragility: A malwarehalting technique. IEEE Security & Privacy, 13(4):40– 46, July/August 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.

                                        org/csdl/mags/sp/2015/04/
                                        msp2015040040-abs.html.

Hoofnagle:2016:AFT

[Hoo16] Chris Jay Hoofnagle. Assessing the Federal trade Commission’s privacy assessments. IEEE Security & Privacy, 14(2):58–64, March/April 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Horne:2014:HL

[Hor14a] Bill Horne. Humans in the loop. IEEE Security & Privacy, 12(1):3–4, January/February 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Horne:2014:CSI

[Hor14b] Bill Horne. On Computer Security Incident Response Teams. IEEE Security & Privacy, 12(5):13–15, September/October 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/sp/2014/05/
                                        msp2014050013.html.

Horne:2015:UO

[Hor15] Bill Horne. Umbrellas and octopuses. IEEE Security & Privacy, 13(1):3–5, January/February 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/sp/2015/01/
                                        msp2015010003.html.

Horne:2016:TMT

[Hor16] Bill Horne. Trust me. trust me not. IEEE Security & Privacy, 14(3):3–5, May/June 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Horsman:2019:CPE

[Hor19]  G. Horsman.  A call for

the prohibition of encryption: Panacea or problem? IEEE Security & Privacy, 17(2):59–66, March/April 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Horowitz:2020:CRC

[Hor20] B. M. Horowitz. Cyberattackresilient cyberphysical systems. IEEE Security & Privacy, 18 (1):55–60, January/February 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Howard:2004:BMS

[How04] Michael Howard. Building more secure software with improved development processes. IEEE Security & Privacy, 2(6):63–65, November/ December 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/06/j6063.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/06/j6063.
pdf.

Howard:2006:PPS

[How06] Michael Howard. A process for performing security code reviews. IEEE Security & Privacy, 4(4):74–79, July/August 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Howard:2008:BSE

[How08] Michael Howard. Becoming a security expert. IEEE Security & Privacy, 6(1):71–73, January/February 2008. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Howard:2009:BTI

[How09a] Michael Howard. Basic training: Improving software security by eliminating the CWE top 25 vulnerabilities. IEEE Security & Privacy, 7(3):68–71, May/June 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Howard:2009:BTM

[How09b] Michael Howard. Basic training: Managing the security wall of data. IEEE Security & Privacy, 7(5):66–68, September/October 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Hu:2004:SSW

[HP04] Yih-Chun Hu and Adrian Perrig. A survey of secure wireless ad hoc routing. IEEE Security & Privacy, 2(3):28– 39, May/June 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/03/j3028.

htm;   http://csdl.computer.
org/dl/mags/sp/2004/03/j3028.
pdf.

Harnik:2010:SCC

[HPSP10] Danny Harnik, Benny Pinkas, and Alexandra Shulman-Peleg. Side channels in cloud services: Deduplication in cloud storage. IEEE Security & Privacy, 8(6):40–47, November/ December 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Hofstede:2018:FBC

[HPSR18] Rick Hofstede, Aiko Pras, Anna Sperotto, and Gabi Dreo Rodosek. Flow-based compromise detection: Lessons learned. IEEE Security & Privacy, 16(1):82–89, January/February 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/01/
msp2018010082.html.

Hariri:2003:IAF

[HQD^{+}03] Salim Hariri, Guangzhi Qu, Tushneem Dharmagadda, Modukuri Ramkishore, and Cauligi S. Raghavendra. Impact analysis of faults and attacks in large-scale networks. IEEE Security & Privacy, 1(5): 49–54, September/October 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2003/05/j5049abs.

                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2003/05/j5049.
                                        pdf.

Hoffman:2005:ENC

[HRDR05] Lance J. Hoffman, Tim Rosenberg, Ronald Dodge, and Daniel Ragsdale. Exploring a national cybersecurity exercise for universities. IEEE Security & Privacy, 3(5):27–33, September/October 2005. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Hernandez-Ramos:2020:TDD

[HRGK^{+}20] J. L. Hernandez-Ramos, D. Geneiatakis, I. Kounelis, G. Steri, and I. Nai Fovino. Toward a datadriven society: A technological perspective on the development of cybersecurity and data-protection policies. IEEE Security & Privacy, 18(1):28– 38, January/February 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Hussain:2021:FSF

[HRK21] Siam U. Hussain, M. Sadegh Riazi, and Farinaz Koushanfar. The fusion of secure function evaluation and logic synthesis. IEEE Security & Privacy, 19 (2):48–55, March/April 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Hernandez-Ramos:2021:SPI

[HRMS^{+}21a] J. L. Hernandez-Ramos, J. A.

Martinez, V. Savarino, M. Angelini, V. Napolitano, A. F.

Skarmeta, and G. Baldini. Security and privacy in Internet of Things-enabled smart cities: Challenges and future directions. IEEE Security & Privacy, 19(1):12–23, January/ February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Hernandez-Ramos:2021:CSC

[HRMS21b] J. L. Hernandez-Ramos, S. N.

Matheu, and A. Skarmeta. The challenges of software cybersecurity certification [building security in]. IEEE Security & Privacy, 19(1):99–102, January/February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Haines:2003:VSA

[HRTT03] Joshua Haines, Dorene Kewley Ryder, Laura Tinnel, and Stephen Taylor. Validation of sensor alert correlators. IEEE Security & Privacy, 1(1):46–56, January/February 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j1046.pdf;   http://www.
computer.org/security/j1046abs.
htm.

Homaei:2017:SYS

[HS17] Hossein Homaei and Hamid Reza Shahriari. Seven years of software vulnerabilities: The ebb and flow. IEEE Security & Privacy, 15(1):58–65, January/February 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/01/
                                        msp2017010058-abs.html.

Hansen:2008:PIM

[HSC08] Marit Hansen, Ari Schwartz, and Alissa Cooper. Privacy and identity management. IEEE Security & Privacy, 6(2):38–45, March/April 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Hively:2011:TST

[HSS11] Lee Hively, Frederick Sheldon, and Anna Cinzia Squicciarini. Toward scalable trustworthy computing using the human-physiology-immunity metaphor. IEEE Security & Privacy, 9(4):14–23, July/August 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Hwang:2006:SES

[HSTV06] David D. Hwang, Patrick Schaumont, Kris Tiri, and Ingrid Verbauwhede. Securing embedded systems. IEEE Security & Privacy, 4(2):40–49, March/April 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Hu:2016:MSA

[Hu16] Hongxin Hu. Mules, seals, and attacking tools: Analyzing 12 online marketplaces. IEEE Security & Privacy, 14(3):32–43, May/June 2016. CODEN ????

ISSN 1540-7993 (print), 1558-4046 (electronic).

Hutchings:2024:AOD

[Hut24] Alice Hutchings. The amplification of online deviancy through the language of violent crime, war, and aggression. IEEE Security & Privacy, 22 (2):81–84, March/April 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Howard:2020:BCF

[HV20] J. P. Howard and M. E.

Vachino. Blockchain compliance with federal cryptographic information-processing standards. IEEE Security & Privacy, 18(1):65–70, January/ February 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Herley:2012:RAA

[HvO12] Cormac Herley and Paul van Oorschot. A research agenda acknowledging the persistence of passwords. IEEE Security & Privacy, 10(1):28–36, January/ February 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Herley:2018:SSC

[HvO18] Cormac Herley and P. C. van Oorschot. Science of security: Combining theory and measurement to reflect the observable. IEEE Security & Privacy, 16(1):12–22, January/February 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/01/
                                        msp2018010012.html.

Harrison:2013:ADC

[HW13] Keith Harrison and Gregory B. White. Anonymous and distributed community cyberincident detection. IEEE Security & Privacy, 11(5):20–27, September/October 2013. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Hao:2020:EEV

[HWB^{+}20] F. Hao, S. Wang, S. Bag, R. Procter, S. F. Shahandashti, M. Mehrnezhad, E. Toreini, R. Metere, and L. Y. J. Liu. End-to-end verifiable e-voting trial for polling station voting. IEEE Security & Privacy, 18 (6):6–13, November/December 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Hansson:2010:AMV

[HWF^{+}10] Jorgen Hansson, Lutz Wrage, Peter H. Feiler, John Morley, Bruce Lewis, and Jerome Hugues. Architectural modeling to verify security and nonfunctional behavior. IEEE Security & Privacy, 8(1):43–49, January/February 2010. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ishitani:2003:MBA

[IAM03] Lucila Ishitani, Virgilio Almeida, and Wagner Meira, Jr. Masks: Bringing anonymity and personalization together. IEEE Security & Privacy, 1(3):18– 23, May/June 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j3018.pdf;   http://www.
computer.org/security/j3018abs.
htm.

Editor-in-Chief:2015:ISP

[iC15] Editor in Chief. IEEE Security & Privacy house advertisement. IEEE Security & Privacy, 13(6):c4, November/ December 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Iyer:2007:TAA

[IKP^{+}07] Ravishankar K. Iyer, Zbigniew Kalbarczyk, Karthik Pattabiraman, William Healey, Wen-Mei W. Hwu, Peter Klemperer, and Reza Farivar. Toward application-aware security and reliability. IEEE Security & Privacy, 5(1):57–62, January/ February 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

IntVeld:2007:DSA

[In 07] Sophie In ’t Veld. Data sharing across the Atlantic. IEEE Security & Privacy, 5(4):58–61, July/August 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Irvine:2010:ESS

[IN10] Cynthia Irvine and Thuy D.

Nguyen. Educating the systems security engineer’s apprentice. IEEE Security & Privacy, 8(4):58–61, July/August 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Irvine:2010:CCN

[IP10] Cynthia E. Irvine and Charles C. Palmer. Call in the Cyber National Guard! IEEE Security & Privacy, 8(1):56–59, January/ February 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Irvine:2011:GEI

[IR11] Cynthia Irvine and J. R. Rao. Guest Editors’ introduction: Engineering secure systems. IEEE Security & Privacy, 9(1): 18–21, January/February 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Irvine:2003:TCS

[Irv03] Cynthia E. Irvine. Teaching constructive security. IEEE Security & Privacy, 1(6): 59–61, November/December 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/sp/2003/06/j6059abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2003/06/j6059.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2003/06/j6059.
                                        pdf.

Irvine:2011:VCF

[Irv11] Cynthia Irvine. The value of capture-the-flag exercises in education: An interview with Chris Eagle. IEEE Security & Privacy, 9(6):58–60, November/December 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Irvine:2013:CP

[Irv13] Cynthia Irvine. A cyberoperations program. IEEE Security & Privacy, 11(5):66–69, September/October 2013. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Iliev:2005:PCP

[IS05] A. Iliev and S. W. Smith.

Protecting client privacy with trusted computing at the server. IEEE Security & Privacy, 3(2):20–28, March/April 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/30742/01423957.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
30742&arnumber=1423957&count=
15&index=4.

Irvine:2005:CGI

[ITA05] C. E. Irvine, M. F. Thompson, and K. Allen. Cyber-CIEGE: gaming for information assurance. IEEE Security & Privacy, 3(3):61–64, May/ June 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/31002/01439504.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        31002&arnumber=1439504&count=
                                        20&index=11.

Iverson:2005:FC

[Ive05] R. Iverson. A framework to consider. IEEE Security & Privacy, 3(2):14, March/April 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/30742/01423954.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        30742&arnumber=1423954&count=
                                        15&index=2.

Jaeger:2023:BS

[Jae23] Trent Jaeger. On bridges and software. IEEE Security & Privacy, 21(3):4–5, May/June 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Jakobsson:2019:RTL

[Jak19] M. Jakobsson. The rising threat of launchpad attacks. IEEE Security & Privacy, 17 (5):68–72, September/October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Jakobsson:2020:PP

[Jak20] M. Jakobsson. Permissions and privacy. IEEE Security & Privacy, 18(2):46–55, March/ April 2020. ISSN 1558-4046.

Jutla:2005:SAO

[JB05] D. N. Jutla and P. Bodorik.

Sociotechnical architecture for online privacy. IEEE Security & Privacy, 3(2):29–39, March/ April 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/30742/01423958.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
30742&arnumber=1423958&count=
15&index=5.

Jain:2015:ISG

[JBZ^{+}15] Bhushan Jain, Mirza Basim Baig, Dongli Zhang, Donald E. Porter, and Radu Sion. Introspections on the semantic gap. IEEE Security & Privacy, 13 (2):48–55, March/April 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/02/
msp2015020048-abs.html.

Jakobsson:2012:AWD

[JCM12] Markus Jakobsson, Richard Chow, and Jesus Molina. Authentication — are we doing well enough? IEEE Security & Privacy, 10(1):19–21, January/ February 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Jere:2021:TAF

[JFK21] Malhar S. Jere, Tyler Farnan, and Farinaz Koushanfar. A taxonomy of attacks on federated learning. IEEE Security & Privacy, 19(2):20–28, March/ April 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Johnson:2007:EIS

[JG07] M. Eric Johnson and Eric Goetz. Embedding information security into the organization. IEEE Security & Privacy, 5(3): 16–24, May/June 2007. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Johnson:2009:IRS

[JGP09] M. Eric Johnson, Eric Goetz, and Shari Lawrence Pfleeger. Information risk: Security through information risk management. IEEE Security & Privacy, 7(3):45–52, May/June 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Jensen:2013:FIM

[JJ13] Jostein Jensen and Martin Gilje Jaatun. Federated Identity Management — we built it; why won’t they come? IEEE Security & Privacy, 11 (2):34–41, March/April 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Jakobsson:2008:WHP

[JJF08] Markus Jakobsson, Nathaniel Johnson, and Peter Finn. Why and how to perform fraud experiments. IEEE Security & Privacy, 6(2):66–68, March/ April 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Jaatun:2009:ATF

[JJV^{+}09] Martin Gilje Jaatun, Jostein Jensen, Håvard Vegge, Finn Michael Halvorsen, and Rune Walsø Nergård. Attack trends: Fools download where angels fear to tread. IEEE Security & Privacy, 7(2):83–86, March/April 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Jones:2018:AEA

[JKE18] Meg Leta Jones, Ellen Kaufman, and Elizabeth Edenberg. AI and the ethics of automating consent. IEEE Security & Privacy, 16(3):64– 72, May/June 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/03/
msp2018030064.html.

Jaeger:2023:IEH

[JKMS23] Trent Jaeger, Brent ByungHoon Kang, Nele Mentens, and Cynthia Sturton. Impact of emerging hardware on security and privacy. IEEE Security & Privacy, 21(3):6–7, May/June 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Jalali:2019:ITP

[JKSM19] M. S. Jalali, J. P. Kaiser, M. Siegel, and S. Madnick. The Internet of Things promises new benefits and risks: A systematic analysis of adoption dynamics of IoT products.

IEEE Security & Privacy, 17 (2):39–48, March/April 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Jordan:2018:QCS

[JL18] S. P. Jordan and Y. Liu. Quantum cryptanalysis: Shor, Grover, and beyond. IEEE Security & Privacy, 16(5):14– 21, September/October 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Jacoby:2006:UBC

[JMD06] Grant A. Jacoby, Randy Marchany, and Nathaniel J. Davis IV. Using battery constraints within mobile hosts to improve network security. IEEE Security & Privacy, 4 (5):40–49, September/October 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Jonas:2006:TFI

[Jon06] Jeff Jonas. Threat and fraud intelligence, Las Vegas style. IEEE Security & Privacy, 4(6): 28–34, November/December 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Jones:2007:ESV

[Jon07] Jeffrey R. Jones. Estimating software vulnerabilities. IEEE Security & Privacy, 5(4):28–32, July/August 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Johnson:2011:AIR

[JP11] M. Eric Johnson and Shari Pfleeger. Addressing information risk in turbulent times. IEEE Security & Privacy, 9(1): 49–57, January/February 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Juels:2014:HEE

[JR14] Ari Juels and Thomas Ristenpart. Honey encryption: Encryption beyond the bruteforce barrier. IEEE Security & Privacy, 12(4):59– 62, July/August 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/04/
msp2014040059-abs.html.

Jakobsson:2007:WCP

[JS07] Markus Jakobsson and Sid Stamm. Web camouflage: Protecting your clients from browser-sniffing attacks. IEEE Security & Privacy, 5(6):16– 24, November/December 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Jaeger:2010:OCC

[JS10] Trent Jaeger and Joshua Schiffman. Outlook: Cloudy with a chance of security challenges and improvements. IEEE Security & Privacy, 8(1):77–80, January/February 2010. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Jauernig:2020:TEE

[JSS20] P. Jauernig, A. Sadeghi, and E. Stapf. Trusted execution environments: Properties, applications, and challenges. IEEE Security & Privacy, 18(2):56– 60, March/April 2020. ISSN 1558-4046.

Just:2004:DEC

[Jus04] Mike Just. Designing and evaluating challenge-question systems. IEEE Security & Privacy, 2(5):32–39, September/October 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/05/j5032.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/05/j5032.
pdf.

Johnston:2004:OIS

[JW04] David Johnston and Jesse Walker. Overview of IEEE 802.16 security. IEEE Security & Privacy, 2(3):40–48, May/June 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/03/j3040.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/03/j3040.
pdf.

Johnson:2011:UFH

[JW11] M. Eric Johnson and Nicholas D. Willey. Usability failures and healthcare data hemorrhages. IEEE Security & Privacy, 9(2):

35–42, March/April 2011. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Jiwnani:2004:SMN

[JZ04] Kanta Jiwnani and Marvin Zelkowitz. Susceptibility matrix: a new aid to software auditing. IEEE Security & Privacy, 2(2):16–21, March/ April 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2004/02/j2016abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/02/j2016.
pdf.

Kott:2013:PCC

[KA13] Alexander Kott and Curtis Arnold. The promises and challenges of continuous monitoring and risk scoring. IEEE Security & Privacy, 11(1):90–93, January/February 2013. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kolhar:2017:CDA

[KAAEa17] Manjur Kolhar, Mosleh M.

Abu-Alhaj, and Saied M. Abd El-atty. Cloud data auditing techniques with a focus on privacy and security. IEEE Security & Privacy, 15(1):42–51, January/February 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/01/
msp2017010042-abs.html.

Kallberg:2012:CCM

[Kal12]  Jan Kallberg.  The com-

mon criteria meets Realpolitik: Trust, alliances, and potential betrayal. IEEE Security & Privacy, 10(4):50–53, July/ August 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kampanakis:2014:SAT

[Kam14]  Panos Kampanakis.  Secu-

rity automation and threat information-sharing options. IEEE Security & Privacy, 12 (5):42–51, September/October 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

                                        //www.computer.org/csdl/
                                        mags/sp/2014/05/msp2014050042-
                                        abs.html.

Kane:2009:VJL

[Kan09] Sean F. Kane. Virtual judgment: Legal implications of online gaming. IEEE Security & Privacy, 7(3):23–28, May/June 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kapadia:2007:CSU

[Kap07] Apu Kapadia. A case (study) for usability in secure email communication. IEEE Security & Privacy, 5(2):80–84, March/ April 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kaufman:2009:IAD

[Kau09] Lori M. Kaufman. It all depends: Data security in the world of cloud computing. IEEE Security & Privacy, 7(4): 61–64, July/August 2009. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kaufman:2010:CTE

[Kau10a] Lori M. Kaufman. Can a trusted environment provide security? IEEE Security & Privacy, 8(1):50–52, January/ February 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kaufman:2010:CPC

[Kau10b] Lori M. Kaufman. Can publiccloud security meet its unique challenges? IEEE Security & Privacy, 8(4):55–57, July/ August 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kaufman:2011:HPI

[Kau11] Lori M. Kaufman. How private is the Internet? IEEE Security & Privacy, 9(1):73–75, January/February 2011. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kiely:2006:SSM

[KB06] Laree Kiely and Terry V. Benzel. Systemic security management. IEEE Security & Privacy, 4(6):74–77, November/ December 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Karame:2018:BSP

[KC18] Ghassan Karame and Srdjan Capkun. Blockchain security and privacy. IEEE Security & Privacy, 16(4):11– 12, July/August 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/04/
msp2018040011.html.

Kim:2007:SFD

[KCC07] Jangbok Kim, Kihyun Chung, and Kyunghee Choi. Spam filtering with dynamically updated URL statistics. IEEE Security & Privacy, 5(4):33–39, July/August 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kohno:2022:BOC

[KCL^{+}22] Tadayoshi Kohno, Camille Cobb, Ada Lerner, Michelle Lin, and Adam Shostack. The buffet overflow café. IEEE Security & Privacy, 20(4):4–7, July/August 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Krishnan:2023:WSA

[KCL^{+}23] Padmanabhan Krishnan, Cristina Cifuentes, Li Li, Tegawendé F. Bissyandé, and Jacques Klein. Why is static application security testing hard to learn? IEEE Security & Privacy, 21 (5):68–72, September/October 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Kesler:2006:NBb

[KD06a] Brent Kesler and Heather Drinan. News briefs. IEEE Security & Privacy, 4(3):6–10, May/June 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2006/03/j3006.pdf.

Kesler:2006:NBc

[KD06b] Brent Kesler and Heather Drinan. News briefs. IEEE Security & Privacy, 4(4):6–8, July/ August 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2006/04/j4006.pdf.

Kostiainen:2020:DSC

[KDC20] K. Kostiainen, A. Dhar, and S. Capkun. Dedicated security chips in the age of secure enclaves. IEEE Security & Privacy, 18(5):38–46, September/ October 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Kesler:2006:NBa

[KDF06] Brent Kesler, Heather Drinan, and Nancy Fontaine. News briefs. IEEE Security & Privacy, 4(2):8–13, March/ April 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2006/02/j2008.pdf.

Kosseim:2009:PIP

[KE09] Patricia Kosseim and Khaled El Emam. Privacy interests:

Privacy interests in prescription data, Part I: Prescriber privacy. IEEE Security & Privacy, 7(1):72–76, January/ February 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Keikkila:2005:SE

[Kei05] Faith M. Keikkila. Secure-World Expo 2005. IEEE Security & Privacy, 3(6): 57–60, November/December 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/33104/01556537.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
33104&arnumber=1556537.

Kellerman:2010:CTP

[Kel10]  Tom Kellerman.  Cyber-

threat proliferation: Today’s truly pervasive global epidemic. IEEE Security & Privacy, 8(3):70–73, May/June 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Keromytis:2009:RIS

[Ker09] Angelos D. Keromytis. Randomized instruction sets and runtime environments past research and future directions. IEEE Security & Privacy, 7(1): 18–25, January/February 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Keromytis:2010:VIS

[Ker10] Angelos D. Keromytis. Voiceover-IP security: Research and practice. IEEE Security & Privacy, 8(2):76–78, March/April 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Koss:2013:ACM

[KF13] Ian Koss and Richard Ford.

Authorship is continuous: Managing code plagiarism. IEEE Security & Privacy, 11 (2):72–74, March/April 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kalaimannan:2017:SDL

[KG17] Ezhil Kalaimannan and Jatinder N. D. Gupta. The security development lifecycle in the context of accreditation policies and standards. IEEE Security & Privacy, 15(1):52–57, January/February 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/01/
msp2017010052-abs.html.

King:2012:EVS

[KH12] Merle S. King and Brian Hancock. Electronic voting security 10 years after the Help America Vote Act. IEEE Security & Privacy, 10(5):50–52, September/October 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Khurana:2010:SGS

[KHLF10] Himanshu Khurana, Mark Hadley, Ning Lu, and Deborah A. Frincke. Smart-grid security issues. IEEE Security & Privacy, 8(1):81–85, January/ February 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kwon:2015:PPD

[KJ15] Juhee Kwon and M. Eric Johnson. Protecting patient data — the economic perspective of healthcare security. IEEE Security & Privacy, 13(5):90–95, September/October 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/csdl/mags/sp/2015/05/
msp2015050090-abs.html.

Kwon:2012:ECP

[KJC^{+}12] Mischel Kwon, Michael J. Jacobs, David Cullinane, Christopher G. Ipsen, and James Foley. Educating cyber professionals: a view from academia, the private sector, and government. IEEE Security & Privacy, 10(2):50–53, March/ April 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kim:2021:ELS

[KKK^{+}21] Dongkwan Kim, Eunsoo Kim, Mingeun Kim, Yeongjin Jang, and Yongdae Kim. Enabling the large-scale emulation of Internet of Things firmware with heuristic workarounds. IEEE Security & Privacy, 19(6):26– 35, November/December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Kerschbaum:2023:PPM

[KL23] Florian Kerschbaum and Nils Lukas. Privacy-preserving machine learning [cryptography]. IEEE Security & Privacy, 21(6):90–94, November/ December 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Kott:2017:AMI

[KLL17] Alexander Kott, Jackson Ludwig, and Mona Lange. Assessing mission impact of cyberattacks: Toward a modeldriven paradigm. IEEE Security & Privacy, 15(5):65–74, September/October 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/05/
msp2017050065-abs.html.

Koblitz:2016:RWE

[KM16] Neal Koblitz and Alfred Menezes. A riddle wrapped in an enigma. IEEE Security & Privacy, 14(6):34– 42, November/December 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/06/
msp2016060034-abs.html.

Kardash:2022:PCG

[KM22] Adam Kardash and Suzanne Morin. The practices and challenges of generating nonidentifiable data. IEEE Security & Privacy, 20(1):113– 118, January/February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Karger:2011:LLB

[KMP^{+}11] Paul Karger, Suzanne McIntosh, Elaine Palmer, David Toll, and Samuel Weber. Lessons learned: Building the Caernarvon high-assurance operating system. IEEE Security & Privacy, 9(1):22–30, January/February 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Klipstein:2020:UIE

[KMP20] M. Klipstein, A. Minter, and J. Pittman. Understanding the information environment to win the next conflict without firing a shot. IEEE Security & Privacy, 18(1):39–45, January/ February 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Koppel:2013:CCA

[KN13] Benedikt Koppel and Stephan Neuhaus. Crypto Corner: Analysis of a hardware security module’s high-availability setting. IEEE Security & Privacy, 11(3):77–80, May/June 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kulyk:2017:NCF

[KNBV17] Oksana Kulyk, Stephan Neumann, Jurlind Budurushi, and Melanie Volkamer. Nothing comes for free: How much usability can you sacrifice for security? IEEE Security & Privacy, 15(3):24– 29, May/June 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/03/
msp2017030024-abs.html.

Knight:2015:ISC

[Kni15] John Knight. The importance of security cases: Proof is good, but not enough. IEEE Security & Privacy, 13(4):73– 75, July/August 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/04/
msp2015040073-abs.html.

Knijnenburg:2017:PCE

[Kni17] Bart P. Knijnenburg. Privacy?

I can’t even! Making a case for user-tailored privacy. IEEE Security & Privacy, 15(4):62– 67, July/August 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/04/
msp2017040062-abs.html.

Kristianto:2022:DPK

[KNL22] Edy Kristianto, Van-Linh Nguyen, and Po-Ching Lin.

Decentralized public-key infrastructure with blockchain in V2X communications: Promising or only euphoria? IEEE Security & Privacy, 20(4):40– 50, July/August 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Knutson:2007:BPS

[Knu07] Tina R. Knutson. Building privacy into software products and services. IEEE Security & Privacy, 5(3):72–74, May/June 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kolupaev:2008:CHV

[KO08] Aleksey Kolupaev and Juriy Ogijenko. CAPTCHAs: Humans vs. bots. IEEE Security & Privacy, 6(1):68–70, January/February 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kobeissi:2019:SRF

[Kob19] N. Kobeissi. Selfie’s reflections on formal verification for transport layer security 1.3: Largely opaque. IEEE Security & Privacy, 17(4):94–96, July/ August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Kohno:2022:END

[Koh22a] Tadayoshi Kohno. Excerpts from the New Dictionary of Cybersecurity, 2036. IEEE Security & Privacy, 20(3):8, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Kohno:2022:MPB

[Koh22b] Tadayoshi Kohno. Mx. President has a brain. IEEE Security & Privacy, 20(3):9–11, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Kohno:2022:ORP

[Koh22c] Tadayoshi Kohno. The our reality privacy policy. IEEE Security & Privacy, 20(6):4– 7, November/December 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Kohno:2022:S

[Koh22d] Tadayoshi Kohno. The Schuhmacher. IEEE Security & Privacy, 20(5):14–15, September/ October 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Kohno:2022:WLO

[Koh22e] Tadayoshi Kohno. Welcome to the latest off by one column. IEEE Security & Privacy, 20(6):5, November/December 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Kohno:2023:EUR

[Koh23a] Tadayoshi Kohno. In Earth until (ready). IEEE Security & Privacy, 21(2):7–8, March/ April 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Kohno:2023:YE

[Koh23b] Tadayoshi Kohno. In your eyes. IEEE Security & Privacy, 21(5):4–5, September/October 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Kosseff:2015:NLF

[Kos15] J. Kosseff. A new legal framework for online anonymity: California’s privacy-based approach. IEEE Security & Privacy, 13(6):66–70, November/ December 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kotzanikolaou:2008:DPD

[Kot08] Panayiotis Kotzanikolaou. Data privacy: Data retention and privacy in electronic communications. IEEE Security & Privacy, 6(5):46–52, September/ October 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kagal:2010:PPB

[KP10] Lalana Kagal and Joe Pato. Preserving privacy based on semantic policy tools. IEEE Security & Privacy, 8(4):25–30, July/August 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Krol:2015:EPN

[KP15] Kat Krol and Soren Preibusch. Effortless privacy negotiations. IEEE Security & Privacy, 13 (3):88–91, May/June 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/03/
msp2015030088-abs.html.

Kumar:2019:ETI

[KPM^{+}19] D. Kumar, R. Paccagnella, P. Murley, E. Hennenfent, J. Mason, A. Bates, and M. Bailey. Emerging threats in Internet of Things voice services. IEEE Security & Privacy, 17(4):18–24, July/ August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Kenny:2010:ESA

[KR10] J. Ryan Kenny and Craig Robinson. Embedded software assurance for configuring secure hardware. IEEE Security & Privacy, 8(5):20–26, September/October 2010. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Krawetz:2004:HFA

[Kra04] Neal Krawetz. The Honeynet files: Anti-honeypot technology. IEEE Security & Privacy, 2(1):76–79, January/ February 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2004/01/j1076abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/01/j1076.
pdf.

Krishnamurthy:2013:POS

[Kri13] Balachander Krishnamurthy.

Privacy and online social networks: can colorless green ideas sleep furiously? IEEE Security & Privacy, 11(3):14–20, May/ June 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kharraz:2018:PAR

[KRK18] Amin Kharraz, William Robertson, and Engin Kirda. Protecting against ransomware: A new line of research or restating classic ideas? IEEE Security & Privacy, 16(3):103– 107, May/June 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/03/
                                        msp2018030103-abs.html.

Kroll:2018:DSD

[Kro18] J. A. Kroll. Data science data governance [AI ethics]. IEEE Security & Privacy, 16(6):61– 70, November/December 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Keblawi:2006:ACC

[KS06] Feisal Keblawi and Dick Sullivan. Applying the common criteria in systems engineering. IEEE Security & Privacy, 4(2): 50–55, March/April 2006. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Karger:2008:VVM

[KS08a] Paul A. Karger and David R. Safford. Virtualization: I/O for virtual machine monitors:

Security and performance issues. IEEE Security & Privacy, 6(5):16–23, September/ October 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

King:2008:GEI

[KS08b] Samuel T. King and Sean W.

Smith. Guest Editors’ introduction: Virtualization and security: Back to the future. IEEE Security & Privacy, 6(5): 15, September/October 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kirlappos:2012:SEA

[KS12a] Iacovos Kirlappos and M. Angela Sasse. Security education against phishing: a modest proposal for a major rethink. IEEE Security & Privacy, 10 (2):24–32, March/April 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kumar:2012:MVA

[KS12b] Sanjeev Kumar and Sirisha Surisetty. Microsoft vs. Apple: Resilience against distributed denial-of-service attacks. IEEE Security & Privacy, 10(2):60– 64, March/April 2012. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kaur:2013:AAS

[KS13a] Sanmeet Kaur and Maninder Singh. Automatic attack signature generation systems: A review. IEEE Security & Privacy, 11(6):54–61, November/ December 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Koopman:2013:IEC

[KS13b] Philip Koopman and Christopher Szilagyi. Integrity in embedded control networks. IEEE Security & Privacy, 11(3):61– 63, May/June 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kshetri:2006:SEC

[Ksh06]  Nir Kshetri.  The sim-

ple economics of cybercrimes. IEEE Security & Privacy, 4 (1):33–39, January/February 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/33481/01588823.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        33481&arnumber=1588823.

Kshetri:2010:ECF

[Ksh10] Nir Kshetri. The economics of click fraud. IEEE Security & Privacy, 8(3):45–53, May/June 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kshetri:2014:CDP

[Ksh14] Nir Kshetri. China’s data privacy regulations: A tricky tradeoff between ICT’s productive utilization and cybercontrol. IEEE Security & Privacy, 12(4):38– 45, July/August 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/04/
msp2014040038-abs.html.

Kshetri:2015:ICL

[Ksh15] Nir Kshetri. India’s cybersecurity landscape: The roles of the private sector and publicprivate partnership. IEEE Security & Privacy, 13(3):16– 23, May/June 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/03/
msp2015030016-abs.html.

Kshetri:2017:ORS

[Ksh17]  Nir Kshetri.  An opinion

on the “Report on Securing and Growing the Digital Economy”. IEEE Security & Privacy, 15(1):80–85, January/February 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/01/
msp2017010080-abs.html.

Kohlbrenner:2020:BOT

[KSL^{+}20] D. Kohlbrenner, S. Shinde, D. Lee, K. Asanovic, and D. Song. Building open trusted execution environments. IEEE Security & Privacy, 18(5):47– 56, September/October 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

King:2021:CCF

[KST^{+}21] Samuel T. King, Nolen Scaife, Patrick Traynor, Zainul Abi Din, Christian Peeters, and Hari Venugopala. Credit card fraud is a computer security problem. IEEE Security & Privacy, 19(2):65–69, March/ April 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Kott:2020:DWI

[KT20] A. Kott and P. Theron. Doers, not watchers: Intelligent autonomous agents are a path to cyber resilience. IEEE Security & Privacy, 18(3):62–66, May/June 2020. ISSN 1558-4046.

Kaza:2018:HWC

[KTS18] Siddharth Kaza, Blair Taylor, and Kyle Sherbert. Hello, world! — code responsibly. IEEE Security & Privacy, 16(1):98–100, January/February 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/01/
                                        msp2018010098.html.

Kyriazanos:2019:ADM

[KTT19] D. M. Kyriazanos, K. G. Thanos, and S. C. A. Thomopoulos. Automated decision making in airport checkpoints: Bias detection toward smarter security and fairness. IEEE Security & Privacy, 17(2):8–16, March/April 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Kufel:2013:SEM

[Kuf13] Lukasz Kufel. Security event monitoring in a distributed systems environment. IEEE Security & Privacy, 11(1):36–43, January/February 2013. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kuhn:2016:LIT

[Kuh16]  Richard Kuhn.  Learn-

ing Internet-of-Things security “Hands-On”. IEEE Security & Privacy, 14(1):37–46, January/ February 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kuper:2005:SS

[Kup05] Peter Kuper. The state of security. IEEE Security & Privacy, 3(5):51–53, September/ October 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kuper:2006:WIF

[Kup06] Peter Kuper. A warning to industry—fix it or lose it. IEEE Security & Privacy, 4(2):56–60, March/April 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kargl:2014:ISD

[KvdHK^{+}14] Frank Kargl, Rens W. van der Heijden, Hartmut Konig, Alfonso Valdes, and Marc C. Dacier. Insights on the security and dependability of industrial control systems. IEEE Security & Privacy, 12(6):75– 78, November/December 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/06/
msp2014060075-abs.html.

Kaaniche:2011:IAD

[KvM11] Mohamed Kaaniche and Aad van Moorsel. It all depends, and increasingly so. IEEE Security & Privacy, 9(6):56– 57, November/December 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kerr:2014:ALE

[KvS14] Michael Kerr and Ron van Schyndel. Adapting law enforcement frameworks to address the ethical problems of CCTV product propagation. IEEE Security & Privacy, 12 (4):14–21, July/August 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/04/
msp2014040014-abs.html.

Kebbel-Wyen:2012:TAS

[KW12] Josh Kebbel-Wyen. Training an army of security ninjas. IEEE Security & Privacy, 10(6):91–93, November/ December 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kirkham:2013:PDS

[KWRK13] Tom Kirkham, Sandra Winfield, Serge Ravet, and Sampo Kellomäki. The personal data store approach to personal data security. IEEE Security & Privacy, 11(5):12–19, September/ October 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Kreutz:2018:KPS

[KYEV^{+}18] D. Kreutz, J. Yu, P. Esteves-Veríssimo, C. Magalhães, and F. M. V. Ramos. The KISS principle in software-defined networking: A framework for secure communications. IEEE Security & Privacy, 16(5):60– 70, September/October 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Kiayias:2017:EEV

[KZZ17] Aggelos Kiayias, Thomas Zacharias, and Bingsheng Zhang. An efficient E2E verifiable E-voting system without setup assumptions. IEEE Security & Privacy, 15(3):14– 23, May/June 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/03/
msp2017030014-abs.html.

Lu:2010:MSC

[LA10] H. Karen Lu and Asad M.

Ali. Making smart cards truly portable. IEEE Security & Privacy, 8(2):28–34, March/April 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ladd:2006:SPS

[Lad06] David Ladd. A software procurement and security primer. IEEE Security & Privacy, 4(6): 71–73, November/December 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Landwehr:2003:ESC

[Lan03]  Carl E. Landwehr.  From

the Editor: Security cosmology: Moving from Big Bang to worlds in collusion. IEEE Security & Privacy, 1(5):5, September/October 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/sp/2003/05/
                                        j5005.pdf.

Landau:2005:SWI

[Lan05a] Susan Landau. Security, wiretapping, and the Internet. IEEE Security & Privacy, 3(6): 26–33, November/December 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/33104/01556533.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        33104&arnumber=1556533.

Landwehr:2005:CPP

[Lan05b] C. Landwehr. Changing the puzzle pieces. IEEE Security & Privacy, 3(1):3–4, January/February 2005. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/30310/01392686.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
30310&arnumber=1392686&count=
17&index=0.

Landwehr:2006:SP

[Lan06] Carl E. Landwehr. Speaking of privacy. IEEE Security & Privacy, 4(4):4–5, July/ August 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2006/04/j4004.pdf.

Landwehr:2007:FTI

[Lan07a] Carl E. Landwehr. Food for thought: Improving the market for assurance. IEEE Security & Privacy, 5(3):3–4, May/ June 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Landwehr:2007:NCN

[Lan07b] Carl E. Landwehr. New challenges for the New Year. IEEE Security & Privacy, 5(1):3–4, January/February 2007. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Landwehr:2007:RTC

[Lan07c] Carl E. Landwehr. Revolution through competition? IEEE Security & Privacy, 5(6):3– 4, November/December 2007.

CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/sp/2007/06/
                                        msp2007060003.pdf.

Landau:2008:SPL

[Lan08a] Susan Landau. Security and privacy landscape in emerging technologies. IEEE Security & Privacy, 6(4):74–77, July/ August 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Landwehr:2008:ECA

[Lan08b] Carl E. Landwehr. From the Editors: Cybersecurity and artificial intelligence: From fixing the plumbing to smart water. IEEE Security & Privacy, 6(5): 3–4, September/October 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Landwehr:2008:S

[Lan08c] Carl E. Landwehr. Up scope. IEEE Security & Privacy, 6 (3):3–4, May/June 2008. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/sp/2008/03/
                                        msp2008030003.pdf.

Landau:2009:PNT

[Lan09a] Susan Landau. Perspectives: The NRC takes on data mining, behavioral surveillance, and privacy. IEEE Security & Privacy, 7(1):58–62, January/ February 2009. CODEN ????

ISSN 1540-7993 (print), 1558-4046 (electronic).

Landwehr:2009:ENG

[Lan09b] Carl E. Landwehr. From the Editors: a national goal for cyberspace: Create an open, accountable Internet. IEEE Security & Privacy, 7(3):3–4, May/June 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Landwehr:2010:DL

[Lan10a] Carl E. Landwehr. Drawing the line. IEEE Security & Privacy, 8(1):3–4, January/ February 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Landwehr:2010:S

[Lan10b] Carl E. Landwehr. Sailing away! IEEE Security & Privacy, 8(6):3–4, November/ December 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Langner:2011:SDC

[Lan11] Ralph Langner. Stuxnet: Dissecting a cyberwarfare weapon. IEEE Security & Privacy, 9(3): 49–51, May/June 2011. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Landau:2013:GEI

[Lan13a] Susan Landau. Guest Editor’s introduction: Politics, love, and death in a world of no privacy. IEEE Security & Privacy, 11(3):11–13, May/June 2013.

CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Landau:2013:MSS

[Lan13b] Susan Landau. Making sense from Snowden: What’s significant in the NSA surveillance revelations. IEEE Security & Privacy, 11(4):54–63, July/August 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Landau:2014:EET

[Lan14a] Susan Landau. Educating engineers: Teaching privacy in a world of open doors. IEEE Security & Privacy, 12(3):66–70, May/June 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Landau:2014:HMS

[Lan14b] Susan Landau. Highlights from making sense of Snowden, Part II: What’s significant in the NSA revelations. IEEE Security & Privacy, 12(1):62–64, January/February 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Landau:2014:SPF

[Lan14c] Susan Landau. Security and privacy: Facing ethical choices. IEEE Security & Privacy, 12 (4):3–6, July/August 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/04/
msp2014040003.html.

Landau:2015:WWS

[Lan15] Susan Landau. What was Samsung thinking? IEEE Security & Privacy, 13(3): 3–4, May/June 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/03/
msp2015030003.html.

Landau:2016:ILI

[Lan16] Susan Landau. Is it legal?

Is it right? The can and should of use. IEEE Security & Privacy, 14(5):3–5, September/October 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/05/
msp2016050003.html.

Lauter:2017:POL

[Lau17] Kristin Lauter. Postquantum opportunities: Lattices, homomorphic encryption, and supersingular isogeny graphs. IEEE Security & Privacy, 15(4):22– 27, July/August 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/04/
msp2017040022-abs.html.

Lawson:2009:SCA

[Law09] Nate Lawson. Side-channel attacks on cryptographic software. IEEE Security & Privacy, 7(6):65–68, November/ December 2009. CODEN ????

ISSN 1540-7993 (print), 1558-4046 (electronic).

Lawton:2010:NBa

[Law10a] George Lawton. News briefs. IEEE Security & Privacy, 8 (5):10–11, September/October 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lawton:2010:NBb

[Law10b] George Lawton. News briefs. IEEE Security & Privacy, 8 (6):7–8, November/December 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lawrence:2016:SDB

[Law16] Darren Lawrence. Security dialogues: Building better relationships between security and business. IEEE Security & Privacy, 14(3):82–87, May/June 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Liyanage:2016:OCS

[LAYG16] Madhusanka Liyanage, Ahmed Bux Abro, Mika Ylianttila, and Andrei Gurtov. Opportunities and challenges of softwaredefined mobile networks in network security. IEEE Security & Privacy, 14(4):34– 44, July/August 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2016/04/
                                        msp2016040034-abs.html.

Li:2004:LEI

[LB04] Hong-Lok Li and Stan Bush.

Letters to the editor: Interface: Usability, efficiency — or privacy?; does piracy increase sales? IEEE Security & Privacy, 2(3):8–9, May/June 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/03/j3008.
htm; j3008.pdf.

Leversage:2008:ESM

[LB08] David John Leversage and Eric James Byres. Estimating a system’s mean time-to-compromise. IEEE Security & Privacy, 6(1):52–60, January/ February 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Li:2007:CAS

[LBB07] Ninghui Li, Ji-Won Byun, and Elisa Bertino. A critique of the ANSI Standard on role-based access control. IEEE Security & Privacy, 5(6):41–49, November/December 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Larsen:2015:ASD

[LBF15] Per Larsen, Stefan Brunthaler, and Michael Franz. Automatic software diversity. IEEE Security & Privacy, 13(2):30– 37, March/April 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/sp/2015/02/
                                        msp2015020030-abs.html.

Linkov:2019:ARH

[LBF^{+}19] I. Linkov, F. Baiardi, M. Florin, S. Greer, J. H. Lambert, M. Pollock, J. Rickli, L. Roslycky, T. Seager, H. Thorisson, and B. D. Trump. Applying resilience to hybrid threats. IEEE Security & Privacy, 17 (5):78–83, September/October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Locasto:2009:BDR

[LBS09] Michael E. Locasto, Sergey Bratus, and Brian Schulte. Bickering in-depth: Rethinking the composition of competing security systems. IEEE Security & Privacy, 7(6):77– 81, November/December 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Li:2013:MSL

[LC13] Qing Li and Greg Clark. Mobile security: A look ahead. IEEE Security & Privacy, 11 (1):78–81, January/February 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lee:2021:DAS

[LCJ21] Yu-Tsung Lee, Haining Chen, and Trent Jaeger. Demystifying Android’s scoped storage defense. IEEE Security & Privacy, 19(5):16–25, September/ October 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Lee:2014:APP

[LDK^{+}14] Michael Z. Lee, Alan M. Dunn, Jonathan Katz, Brent Waters, and Emmett Witchel. Anon-Pass: Practical anonymous subscriptions. IEEE Security & Privacy, 12(3):20–27, May/ June 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Li:2022:BAN

[LDZ^{+}22] Shaofeng Li, Tian Dong, Benjamin Zi Hao Zhao, Minhui Xue, Suguo Du, and Haojin Zhu. Backdoors against natural language processing: A review. IEEE Security & Privacy, 20 (5):50–59, September/October 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Lenarcic:2003:DBT

[Len03]  John Lenarcic.  The di-

nosaur and the butterfly: a tale of computer ethics. IEEE Security & Privacy, 1 (5):61–63, September/October 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2003/05/j5061abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/05/j5061.
pdf.

Lesk:2003:CLR

[Les03a] Michael Lesk. Chicken Little and the recorded music crisis. IEEE Security & Privacy, 1(5):73–75, September/ October 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2003/05/j5073abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/05/j5073.
pdf.

Lesk:2003:CEC

[Les03b] Michael Lesk. Copyright enforcement or censorship: New uses for the DMCA? IEEE Security & Privacy, 1(2):67– 69, March/April 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j2067.pdf;   http://www.
computer.org/security/j2067abs.
htm.

Lesk:2003:CEE

[Les03c] Michael Lesk. Copyright extension: Eldred v. Ashcroft. IEEE Security & Privacy, 1(1):76–78, January/February 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j1076.pdf;   http://www.
computer.org/security/j1076abs.
htm.

Lesk:2003:FFI

[Les03d]  Michael Lesk.  Feist and

facts: If data is protected, will it be more or less available? IEEE Security & Privacy, 1(6):68–70, November/ December 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2003/06/j6068abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/06/j6068.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/06/j6068.
pdf.

Lesk:2003:GBU

[Les03e] Michael Lesk. The good, the bad, and the ugly: What might change if we had good DRM. IEEE Security & Privacy, 1(3): 63–66, May/June 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j3063.pdf.

Lesk:2004:BSS

[Les04a] Michael Lesk. Bigger share of a smaller pie. IEEE Security & Privacy, 2(4):62–64, July/August 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/04/j4062.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/04/j4062.
pdf.

Lesk:2004:DRC

[Les04b] Michael Lesk. Digital rights:

Copyright and creativity. IEEE Security & Privacy, 2(3):76– 78, May/June 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/sp/2004/03/j3076.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2004/03/j3076.
                                        pdf.

Lesk:2004:DRM

[Les04c] Michael Lesk. Digital rights: Micropayments: An idea whose time has passed twice? IEEE Security & Privacy, 2 (1):61–63, January/February 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/sp/2004/01/j1061abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2004/01/j1061.
                                        pdf.

Lesk:2004:SLC

[Les04d] Michael Lesk. Shedding light on creativity: The history of photography. IEEE Security & Privacy, 2(2):62–64, March/ April 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/sp/2004/02/j2062abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2004/02/j2062.
                                        pdf.

Lesk:2005:SBF

[Les05] M. Lesk. Salute the broadcast flag [digital protection for TV recording]. IEEE Security & Privacy, 3(3):84–87, May/June 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/31002/01439510.pdf;
http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
31002&arnumber=1439510&count=
20&index=17.

Lesk:2006:SIF

[Les06] Michael Lesk. Should indexing be fair use? The battle over Google book search. IEEE Security & Privacy, 4(2):80–83, March/April 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lesk:2007:NFL

[Les07a] Michael Lesk. The new front line: Estonia under cyberassault. IEEE Security & Privacy, 5(4):76–79, July/August 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lesk:2007:SKW

[Les07b] Michael Lesk. South Korea’s way to the future. IEEE Security & Privacy, 5(2):85–87, March/April 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lesk:2008:DRM

[Les08a] Michael Lesk. Digital rights management and individualized pricing. IEEE Security & Privacy, 6(3):76–79, May/June 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lesk:2008:FSI

[Les08b] Michael Lesk. Forum shopping on the Internet. IEEE Security & Privacy, 6(1):78–80, January/February 2008. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lesk:2009:SPEb

[Les09a] Michael Lesk. Security & privacy economics: Incentives to innovate: Improve the past or break with it? IEEE Security & Privacy, 7(5):78–81, September/October 2009. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lesk:2009:SPEa

[Les09b] Michael Lesk. Security & privacy economics: Reading over your shoulder. IEEE Security & Privacy, 7(3):78–81, May/ June 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lesk:2010:DLE

[Les10a] Michael Lesk. Do the Luddites ever win? IEEE Security & Privacy, 8(5):79–82, September/October 2010. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lesk:2010:SCN

[Les10b] Michael Lesk. Son of Carterfone: Network neutrality or regulation? IEEE Security & Privacy, 8(3):77–82, May/June 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lesk:2011:CE

[Les11a] Michael Lesk. Cybersecurity and economics. IEEE Security & Privacy, 9(6):76–79, November/December 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lesk:2011:RPP

[Les11b] Michael Lesk. Reading: From paper to pixels. IEEE Security & Privacy, 9(4):76–79, July/ August 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lesk:2011:SSB

[Les11c] Michael Lesk. Salmon, songs, and blankets: Creativity on the Northwest Coast. IEEE Security & Privacy, 9(3):78–81, May/June 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lesk:2011:WIW

[Les11d] Michael Lesk. What is information worth? IEEE Security & Privacy, 9(1):88–90, January/ February 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lesk:2012:CR

[Les12a] Michael Lesk. The clouds roll by. IEEE Security & Privacy, 10(3):84–87, May/June 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lesk:2012:GMM

[Les12b] Michael Lesk. Georgia on my mind. IEEE Security & Privacy, 10(4):88–90, July/ August 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lesk:2012:PP

[Les12c] Michael Lesk. The price of privacy. IEEE Security & Privacy, 10(5):79–81, September/ October 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lesk:2012:YMN

[Les12d] Michael Lesk. Your memory is now a vendor service. IEEE Security & Privacy, 10(1):88–90, January/February 2012. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lesk:2013:BDB

[Les13a] Michael Lesk. Big data, Big Brother, big money. IEEE Security & Privacy, 11(4):85–89, July/August 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lesk:2013:EMR

[Les13b] Michael Lesk. Electronic medical records: Confidentiality, care, and epidemiology. IEEE Security & Privacy, 11(6):19– 24, November/December 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lesk:2013:SPEa

[Les13c] Michael Lesk. Security & privacy economics: Privateers in cyberspace: Aargh! IEEE Security & Privacy, 11(3):81–84, May/June 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lesk:2013:SPEb

[Les13d] Michael Lesk. Security & privacy economics: The old is new again. IEEE Security & Privacy, 11(6):77–79, November/ December 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lesk:2014:CIW

[Les14a]  Michael Lesk.  Caller ID:

Whose privacy? IEEE Security & Privacy, 12(2):77–79, March/April 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lesk:2014:DCS

[Les14b] Michael Lesk. Does the cloud of surveillance have a silver lining? IEEE Security & Privacy, 12(3):78–81, May/June 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lesk:2014:SSD

[Les14c] Michael Lesk. Staffing for security: Don’t optimize. IEEE Security & Privacy, 12(4):71– 73, July/August 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/04/
msp2014040071-abs.html.

Lesk:2014:TV

[Les14d] Michael Lesk. Trust, but verify. IEEE Security & Privacy, 12(6):94–96, November/December 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/06/
msp2014060094-abs.html.

Lesk:2015:LC

[Les15a] M. Lesk. License creep. IEEE Security & Privacy, 13(6):85– 88, November/December 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lesk:2015:IAT

[Les15b] Michael Lesk. Ideas ahead of their time: Digital time stamping. IEEE Security & Privacy, 13(4):76–79, July/ August 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/04/
msp2015040076-abs.html.

Lesk:2015:SRH

[Les15c] Michael Lesk. Safety risks — human error or mechanical failure?: Lessons from railways. IEEE Security & Privacy, 13 (2):99–102, March/April 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/02/
msp2015020099-abs.html.

Levy:2003:COP

[Lev03a] Elias Levy. Crossover: Online pests plaguing the offline world. IEEE Security & Privacy, 1(6):71–73, November/ December 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2003/06/j6071abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/06/j6071.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/06/j6071.
pdf.

Levy:2003:PSS

[Lev03b] Elias Levy. Poisoning the software supply chain. IEEE Security & Privacy, 1(3):70– 73, May/June 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j3070.pdf.

Levy:2004:AZ

[Lev04a]  Elias Levy.  Approaching

zero. IEEE Security & Privacy, 2(4):65–66, July/ August 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/04/j4065.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/04/j4065.
pdf.

Levy:2004:CBT

[Lev04b] Elias Levy. Criminals become tech savvy. IEEE Security & Privacy, 2(2):65–68, March/ April 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/sp/2004/02/j2065abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2004/02/j2065.
                                        pdf.

Levy:2004:II

[Lev04c] Elias Levy. Interface illusions. IEEE Security & Privacy, 2(6):66–69, November/ December 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/sp/2004/06/j6066.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2004/06/j6066.
                                        pdf.

Levy:2005:WPG

[Lev05] E. Levy. Worm propagation and generic attacks. IEEE Security & Privacy, 3(2):63– 65, March/April 2005. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/30742/01423964.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        30742&arnumber=1423964&count=
                                        15&index=10.

Levy:2006:WCS

[Lev06] Elias Levy. Worst-case scenario. IEEE Security & Privacy, 4(5):71–73, September/ October 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lewis:2011:CTE

[Lew11] James A. Lewis. Cyberwar thresholds and effects. IEEE Security & Privacy, 9(5):23–29, September/October 2011. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Levine:2004:UHP

[LGO04] John G. Levine, Julian B.

Grizzard, and Henry L. Owen. Using honeynets to protect large enterprise networks. IEEE Security & Privacy, 2(6):73–75, November/ December 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/06/j6073.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/06/j6073.
pdf.

Levine:2006:DCK

[LGO06] John G. Levine, Julian B. Grizzard, and Henry L. Owen. Detecting and categorizing kernel-level rootkits to aid future detection. IEEE Security & Privacy, 4(1):24– 32, January/February 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/33481/01588822.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        33481&arnumber=1588822.

Lyda:2007:UEA

[LH07] Robert Lyda and James Hamrock. Using entropy analysis to find encrypted and packed malware. IEEE Security & Privacy, 5(2):40–45, March/April 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lipner:2023:IWS

[LH23] Steve Lipner and Michael Howard. Inside the Windows security push: a twenty-year retrospective. IEEE Security & Privacy, 21(2):24–31, March/ April 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Libicki:2005:RCG

[Lib05] Martin Libicki. Are RFIDs coming to get you? IEEE Security & Privacy, 3(6): 6, November/December 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/33104/01556529.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        33104&arnumber=1556529.

Linn:2005:TWU

[Lin05] J. Linn. Technology and Web user data privacy — a survey of risks and countermeasures. IEEE Security & Privacy, 3 (1):52–58, January/February 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/30310/01392701.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
30310&arnumber=1392701&count=
17&index=9.

Lin:2009:SRL

[Lin09] Herbert Lin. Special report:

Lifting the veil on cyber offense. IEEE Security & Privacy, 7(4):15–21, July/August 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lipner:2012:LVS

[LJZ12] Steve Lipner, Trent Jaeger, and Mary Ellen Zurko. Lessons from VAX/SVS for highassurance VM systems. IEEE Security & Privacy, 10(6):26– 35, November/December 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lubomski:2017:PEI

[LK17] Pawel Lubomski and Henryk Krawczyk. Practical evaluation of Internet systems’ security mechanisms. IEEE Security & Privacy, 15(1):32–40, January/February 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/01/
msp2017010032-abs.html.

Leon:2019:HBW

[LKL^{+}19] R. S. Leon, M. Kiperberg, A. A. Leon Zabag, A. Resh, A. Algawi, and N. J. Zaidenberg. Hypervisor-based white listing of executables. IEEE Security & Privacy, 17(5):58– 67, September/October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Locasto:2011:FBD

[LL11] Michael E. Locasto and Matthew C. Little. A failure-based discipline of trustworthy information systems. IEEE Security & Privacy, 9(4):71–75, July/ August 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ladabouche:2016:GIN

[LL16] Tina Ladabouche and Steve LaFountain. GenCyber: Inspiring the next generation of cyber stars. IEEE Security & Privacy, 14(5):84–86, September/October 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2016/05/
                                        msp2016050084-abs.html.

Lacoste:2023:TEE

[LL23] Marc Lacoste and Vincent Lefebvre. Trusted execution environments for telecoms: Strengths, weaknesses, opportunities, and threats. IEEE Security & Privacy, 21(3):37–46, May/June 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Lin:2016:SCU

[LLGJ16] Jingqiang Lin, Bo Luo, Le Guan, and Jiwu Jing. Secure computing using registers and caches: The problem, challenges, and solutions. IEEE Security & Privacy, 14(6):63–70, November/December 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/06/
msp2016060063-abs.html.

Liu:2018:TFC

[LLKA18] Jian Liu, Wenting Li, Ghassan O. Karame, and N. Asokan. Toward fairness of cryptocurrency payments. IEEE Security & Privacy, 16(3):81– 89, May/June 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/03/
msp2018030081-abs.html.

Lin:2014:RSC

[LLT14] Ying-Dar Lin, Chia-Yin Lee, and Hao-Chuan Tsai. Redefining security criteria for networking devices with case studies. IEEE Security & Privacy, 12(1):43–53, January/ February 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Liu:2011:SRM

[LLWC11] Debin Liu, Ninghui Li, XiaoFeng Wang, and L. Jean Camp. Security risk management using incentives. IEEE Security & Privacy, 9(6):20– 28, November/December 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Landau:2008:IPS

[LM08] Susan Landau and Deirdre K. Mulligan. I’m Pc01002/SpringPeeper/ ED288l.6; who are you? IEEE Security & Privacy, 6(2):13– 15, March/April 2008. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/sp/2008/02/
                                        msp2008020013.pdf.

Lisovich:2010:IPI

[LMW10] Mikhail A. Lisovich, Deirdre K. Mulligan, and Stephen B. Wicker. Inferring personal information from demandresponse systems. IEEE Security & Privacy, 8(1):11–20, January/February 2010. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Levieil:2008:CTC

[LN08] Eric Levieil and David Naccache. Cryptographic test correction. IEEE Security & Privacy, 6(2):69–71, March/April 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Locasto:2009:EHS

[Loc09] Michael E. Locasto. Education: Helping students own their own code. IEEE Security & Privacy, 7(3):53–56, May/June 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Laakkonen:2016:IPD

[LPJ^{+}16] Jussi Laakkonen, Janne Parkkila, Pekka Jappinen, Jouni Ikonen, and Ahmed Seffah. Incorporating privacy into digital game platform design: The what, why, and how. IEEE Security & Privacy, 14(4):22– 32, July/August 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/04/
msp2016040022-abs.html.

Ladisa:2023:JCS

[LPS^{+}23] Piergiorgio Ladisa, Serena Elisa Ponta, Antonino Sabetta, Matias Martinez, and Olivier Barais. Journey to the center of software supply chain attacks. IEEE Security & Privacy, 21(6):34–49, November/ December 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Lebeck:2018:AOS

[LRKR18] Kiron Lebeck, Kimberly Ruth, Tadayoshi Kohno, and Franziska Roesner. Arya: Operating system support for securely augmenting reality. IEEE Security & Privacy, 16(1):44–53, January/February 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/01/
msp2018010044.html.

Landwehr:2023:LBF

[LRW^{+}23] Carl Landwehr, Michael K. Reiter, Laurie Williams, Gene Tsudik, Trent Jaeger, Tadayoshi Kohno, and Apu Kapadia. Looking backwards (and forwards): NSF secure and trustworthy computing 20year retrospective panel transcription. IEEE Security & Privacy, 21(2):32–42, March/ April 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Landau:2005:OCS

[LS05a] S. Landau and M. R. Stytz. Overview of cyber security: a crisis of prioritization. IEEE Security & Privacy, 3(3):9–11, May/June 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/31002/01439495.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
31002&arnumber=1439495&count=
20&index=3.

Lekkas:2005:HRS

[LS05b] Dimitrios Lekkas and Diomidis Spinellis. Handling and reporting security advisories: a scorecard approach. IEEE Security & Privacy, 3(4):32–41, July/ August 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lee:2006:EPI

[LS06] Vincent C. S. Lee and Linyi Shao. Estimating potential IT security losses: An alternative quantitative approach. IEEE Security & Privacy, 4(6):44– 52, November/December 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Locasto:2008:HDW

[LS08] Michael E. Locasto and Angelos Stavrou. The hidden difficulties of watching and rebuilding networks. IEEE Security & Privacy, 6(2):79–82, March/April 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lala:2009:GEI

[LS09] Jaynarayan H. Lala and Fred B. Schneider. Guest Editors’ introduction: IT monoculture security risks and defenses. IEEE Security & Privacy, 7(1):12–13, January/ February 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Lindeman:2012:GIR

[LS12] Mark Lindeman and Philip B.

Stark. A gentle introduction to risk-limiting audits. IEEE Security & Privacy, 10(5):42–49, September/October 2012. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ling:2022:FLZ

[LSBB22] Chen Ling, Gianluca Stringhini, Utkucan Balcı, and Jeremy Blackburn. A first look at zoombombing. IEEE Security & Privacy, 20(2):22–30, March/April 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Lecuyer:2018:ESB

[LSG^{+}18] Mathias Lecuyer, Riley Spahn, Roxana Geambasu, Tzu-Kuo Huang, and Siddhartha Sen. Enhancing selectivity in big data. IEEE Security & Privacy, 16(1):34–42, January/February 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/01/
                                        msp2018010034.html.

Lenders:2015:GEC

[LTB15] Vincent Lenders, Axel Tanner, and Albert Blarer. Gaining an edge in cyberspace with advanced situational awareness. IEEE Security & Privacy, 13 (2):65–74, March/April 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/sp/2015/02/
                                        msp2015020065-abs.html.

Liao:2021:PPP

[LWL^{+}21] Song Liao, Christin Wilson, Cheng Long, Hongxin Hu, and Huixing Deng. Problematic privacy policies of voice assistant applications. IEEE Security & Privacy, 19(6):66– 73, November/December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Liu:2010:VMF

[LWLL10] Qian Liu, Chuliang Weng, Minglu Li, and Yuan Luo. An in-VM measuring framework for increasing virtual machine security in clouds. IEEE Security & Privacy, 8(6):56– 62, November/December 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Li:2022:CCB

[LXS22] Ying Li, Tong Xin, and Mikko Siponen. Citizens cybersecurity behavior: Some major challenges. IEEE Security & Privacy, 20(1):54–61, January/ February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Lyon:2024:TDS

[Lyo24] Gregory Lyon. Trust in data security protocols and knowledge of privacy and security technology. IEEE Security & Privacy, 22(2):29–37, March/ April 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Lysyanskaya:2007:AI

[Lys07] Anna Lysyanskaya. Authentication without identification. IEEE Security & Privacy, 5(3): 69–71, May/June 2007. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Li:2016:PSP

[LZ16] Qing-Yun Li and Lei Zhang. The public security and personal privacy survey: Biometric technology in Hong Kong. IEEE Security & Privacy, 14 (4):12–21, July/August 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2016/04/
                                        msp2016040012-abs.html.

McGraw:2007:SBTb

[MA07] Gary McGraw and Ross Anderson. Silver bullet talks with Ross Anderson. IEEE Security & Privacy, 5(4):10–13, July/ August 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McDonald:2012:IHS

[MA12] Jeffrey T. McDonald and Todd R. Andel. Integrating historical security jewels in information assurance education. IEEE Security & Privacy, 10(6):45–50, November/ December 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Memon:2015:CAT

[MA15a] A. M. Memon and A. Anwar. Colluding apps: Tomorrow’s mobile malware threat. IEEE Security & Privacy, 13(6):77– 81, November/December 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Modic:2015:ACE

[MA15b] David Modic and Ross Anderson. It’s all over but the crying: The emotional and financial impact of Internet fraud. IEEE Security & Privacy, 13(5):99– 103, September/October 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/csdl/mags/sp/2015/05/
msp2015050099-abs.html.

Mannan:2020:CLH

[MA20] M. Mannan and N. Asokan.

Confronting the limitations of hardware-assisted security. IEEE Security & Privacy, 18 (5):6–7, September/October 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

MacDonald:2006:CBT

[Mac06] Geraldine MacDonald. Crossborder transaction liability. IEEE Security & Privacy, 4(5): 7, September/October 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/comp/mags/sp/2006/05/
j5007.pdf.

Malkin:2023:CIE

[Mal23] Nathan Malkin. Contextual integrity, explained: a more usable privacy definition. IEEE Security & Privacy, 21(1):58– 65, January/February 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Mangard:2013:KSL

[Man13] Stefan Mangard. Keeping secrets on low-cost chips. IEEE Security & Privacy, 11(4):75– 77, July/August 2013. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Marin:2005:NSB

[Mar05]  Gerald A. Marin.  Net-

work security basics. IEEE Security & Privacy, 3(6): 68–72, November/December 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/33104/01556540.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        33104&arnumber=1556540.

Martin:2006:FSP

[Mar06] Luther Martin. Fitting square pegs into round holes. IEEE Security & Privacy, 4(5):64–66, September/October 2006. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Martin:2007:SCE

[Mar07] Keye Martin. Secure communication without encryption? IEEE Security & Privacy, 5(2): 68–71, March/April 2007. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Martin:2008:CCI

[Mar08] Luther Martin. Crypto corner: Identity-based encryption and beyond. IEEE Security & Privacy, 6(5):62–64, September/ October 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Martin:2010:XMA

[Mar10] Luther Martin. XTS: a mode of AES for encrypting hard disks. IEEE Security & Privacy, 8(3): 68–69, May/June 2010. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Martin:2013:IEC

[Mar13a] Dave Martin. Implementing effective controls in a mobile, agile, cloud-enabled enterprise. IEEE Security & Privacy, 11(1):13–14, January/ February 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Marx:2013:PPT

[Mar13b] Gary T. Marx. The public as partner? Technology can make us auxiliaries as well as vigilantes. IEEE Security & Privacy, 11(5):56–61, September/ October 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Margulies:2015:DGA

[Mar15a] Jonathan Margulies. A developer’s guide to audit logging. IEEE Security & Privacy, 13(3):84–86, May/June 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/03/
msp2015030084-abs.html.

Margulies:2015:GDO

[Mar15b] Jonathan Margulies. Garage door openers: An Internet of Things case study. IEEE Security & Privacy, 13(4):80– 83, July/August 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/04/
msp2015040080-abs.html.

Margulies:2015:SCB

[Mar15c]  Jonathan Margulies.  Se-

curing cloud-based applications, Part 1. IEEE Security & Privacy, 13(5):96–98, September/October 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/csdl/mags/sp/2015/05/
msp2015050096-abs.html.

Masiello:2009:PID

[Mas09] Betsy Masiello. Privacy interests: Deconstructing the privacy experience. IEEE Security & Privacy, 7(4):68–70, July/ August 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Massacci:2019:DAV

[Mas19] F. Massacci. Is “Deny access” a valid “Fail-Safe default” principle for building security in cyberphysical systems? IEEE Security & Privacy, 17(5):90– 93, September/October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Massacci:2024:HGV

[Mas24] Fabio Massacci. The holy grail of vulnerability predictions. IEEE Security & Privacy, 22(1):4–6, January/February 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Matthews:2012:OPI

[Mat12] Bronwen Matthews. Optimizing product improvement spending with third-party security consultants. IEEE Security & Privacy, 10(1):91–93, January/February 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Mathur:2023:RDT

[Mat23] Aditya P. Mathur. Reconfigurable digital twin to support research, education, and training in the defense of critical infrastructure. IEEE Security & Privacy, 21(4):51–60, July/ August 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Maxion:2020:RBL

[Max20] R. Maxion. Reproducibility:

Buy low, sell high. IEEE Security & Privacy, 18(6):33– 41, November/December 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Mayron:2010:SMC

[May10] Liam M. Mayron. Secure multimedia communications. IEEE Security & Privacy, 8(6):76– 79, November/December 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Maybury:2015:TAC

[May15a] Mark Maybury. Toward the assured cyberspace advantage: Air Force Cyber Vision 2025. IEEE Security & Privacy, 13(1):49–56, January/February 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/01/
msp2015010049-abs.html.

Mayron:2015:BAM

[May15b] Liam M. Mayron. Biometric authentication on mobile devices. IEEE Security & Privacy, 13(3):70– 73, May/June 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/03/
msp2015030070-abs.html.

Mirkovic:2012:TCD

[MB12] Jelena Mirkovic and Terry Benzel. Teaching cybersecurity with DeterLab. IEEE Security & Privacy, 10(1):73–76, January/February 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Margulies:2016:CYB

[MB16] Jonathan Margulies and Michael Berg. That certificate you bought could get you hacked. IEEE Security & Privacy, 14 (5):93–95, September/October 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL https:

//www.computer.org/csdl/
mags/sp/2016/05/msp2016050093-
abs.html.

Maiorca:2019:DIP

[MB19] D. Maiorca and B. Biggio. Digital investigation of PDF files: Unveiling traces of embedded malware. IEEE Security & Privacy, 17(1):63–71, January/ February 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Murdoch:2012:HCS

[MBA12] Steven J. Murdoch, Mike Bond, and Ross Anderson. How certification systems fail: Lessons from the Ware Report. IEEE Security & Privacy, 10(6):40–44, November/ December 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Munir:2022:SIS

[MBA^{+}22] Arslan Munir, Erik Blasch, Alexander Aved, Edward Paul Ratazzi, and Joonho Kong. Security issues in situational awareness: Adversarial threats and mitigation techniques. IEEE Security & Privacy, 20 (4):51–60, July/August 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Matias:2018:NNZ

[MBC^{+}18] P. Matias, P. Barbosa, T. N. C.

Cardoso, D. M. Campos, and D. F. Aranha. NIZKCTF: A noninteractive zero-knowledge capture-the-flag platform. IEEE Security & Privacy, 16(6):42– 51, November/December 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Maughan:2013:CVD

[MBLT13] Douglas Maughan, David Balenson, Ulf Lindqvist, and Zachary Tudor. Crossing the “Valley of death”: Transitioning cybersecurity research into practice. IEEE Security & Privacy, 11(2):14–23, March/ April 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Maiano:2024:HVM

[MBP^{+}24] Luca Maiano, Alexandra Benova, Lorenzo Papa, Mara Stockner, Michela Marchetti, Gianmarco Convertino, Giuliana Mazzoni, and Irene Amerini. Human versus machine: a comparative analysis in detecting artificial intelligence-generated images. IEEE Security & Privacy, 22 (3):77–86, May/June 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Minna:2021:USI

[MBR^{+}21] Francesco Minna, Agathe Blaise, Filippo Rebecchi, Balakrishnan Chandrasekaran, and Fabio Massacci. Understanding the security implications of Kubernetes networking. IEEE Security & Privacy, 19(5):46–56, September/ October 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2009:SOG

[MC09] Gary McGraw and Ming Chow.

Securing online games: Guest Editors’ introduction: Securing online games: Safeguarding the future of software security. IEEE Security & Privacy, 7(3): 11–12, May/June 2009. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Martin:2012:SIC

[MC12] Robert A. Martin and Steven M.

Christey. The software industry’s “Clean water act” alternative. IEEE Security & Privacy, 10(3):24–31, May/June 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Mazurczyk:2015:IHC

[MC15] Wojciech Mazurczyk and Luca Caviglione. Information hiding as a challenge for malware detection. IEEE Security & Privacy, 13(2):89– 93, March/April 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/02/
msp2015020089-abs.html.

McCarty:2003:AIT

[McC03a]  Bill McCarty.  Automated

identity theft. IEEE Security & Privacy, 1(5):89– 92, September/October 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/sp/2003/05/j5089abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2003/05/j5089.
                                        pdf.

McCarty:2003:HAR

[McC03b]  Bill McCarty.  The Hon-

eynet arms race. IEEE Security & Privacy, 1(6): 79–82, November/December 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/sp/2003/06/j6079abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2003/06/j6079.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2003/06/j6079.
                                        pdf.

McDaniel:2011:DPS

[McD11] Patrick McDaniel. Data provenance and security. IEEE Security & Privacy, 9(2):83–85, March/April 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McDaniel:2012:BCS

[McD12]  Patrick McDaniel.  Bloat-

ware comes to the smartphone. IEEE Security & Privacy, 10 (4):85–87, July/August 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGrath:2003:MEP

[McG03a] Dennis McGrath. Measuring the 4:11 effect: The power failure and the Internet. IEEE Security & Privacy, 1(5):16–18, September/ October 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2003/05/j5016abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/05/j5016.
pdf.

McGraw:2003:GDS

[McG03b] Gary McGraw. From the ground up: The DIMACS Software Security Workshop. IEEE Security & Privacy, 1(2):59– 66, March/April 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j2059.pdf;   http://www.
computer.org/security/j2059abs.

htm. See erratum [Ano03b].

McGraw:2004:SS

[McG04] Gary McGraw. Software security. IEEE Security & Privacy, 2(2):80–83, March/ April 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2004/02/j2080abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/02/j2080.
pdf.

McGraw:2006:ISBa

[McG06a] Gary McGraw. Interview: Silver Bullet speaks to Avi Rubin. IEEE Security & Privacy, 4(3): 11–13, May/June 2006. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2006:ISBc

[McG06b] Gary McGraw. Interview: Silver Bullet speaks to Marcus Ranum. IEEE Security & Privacy, 4(5):11–14, September/ October 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2006:ISBb

[McG06c] Gary McGraw. Interview: Silver Bullet speaks with Dan Geer. IEEE Security & Privacy, 4(4):10–13, July/August 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2006:SBS

[McG06d] Gary McGraw. Silver bullet speaks with Ed Felten. IEEE Security & Privacy, 4(6):10– 13, November/December 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2007:ISBa

[McG07a] Gary McGraw. Interview: Silver Bullet talks with Annie Antón. IEEE Security & Privacy, 5(5):8–11, September/ October 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2007:ISBb

[McG07b] Gary McGraw. Interview: Silver Bullet talks with Mikko Hyppönen. IEEE Security & Privacy, 5(6):8–11, November/ December 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2007:SBSb

[McG07c] Gary McGraw. Silver bullet speaks with Dorothy Denning. IEEE Security & Privacy, 5(2): 11–14, March/April 2007. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2007:SBSa

[McG07d] Gary McGraw. Silver bullet speaks with John Stewart. IEEE Security & Privacy, 5(1): 9–11, January/February 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2007:SBTa

[McG07e] Gary McGraw. Silver bullet talks with Becky Bace. IEEE Security & Privacy, 5(3):6–9, May/June 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2008:ISBd

[McG08a] Gary McGraw. Interview: Silver Bullet talks with Adam Shostack. IEEE Security & Privacy, 6(4):6–10, July/ August 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2008:ISBe

[McG08b]  Gary McGraw.  Interview:

Silver Bullet talks with Bill Cheswick. IEEE Security & Privacy, 6(5):7–11, September/October 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2008:ISBb

[McG08c]  Gary McGraw.  Interview:

Silver Bullet talks with Ed Amoroso. IEEE Security & Privacy, 6(2):6–9, March/ April 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2008/02/msp2008020006.
pdf.

McGraw:2008:ISBa

[McG08d]  Gary McGraw.  Interview:

Silver Bullet talks with Eugene Spafford. IEEE Security & Privacy, 6(1):10–15, January/February 2008. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/comp/mags/sp/2008/01/
msp2008010010.pdf.

McGraw:2008:ISBc

[McG08e]  Gary McGraw.  Interview:

Silver Bullet talks with Jon Swartz. IEEE Security & Privacy, 6(3):9–11, May/June 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2008:ISBf

[McG08f] Gary McGraw. Interview: Silver Bullet talks with Matt Bishop. IEEE Security & Privacy, 6(6):6–10, November/ December 2008. CODEN ????

ISSN 1540-7993 (print), 1558-4046 (electronic).

McGovern:2009:ISB

[McG09a] James McGovern. Interview:

Silver Bullet talks with Gary McGraw. IEEE Security & Privacy, 7(3):8–10, May/June 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2009:ISBd

[McG09b] Gary McGraw. Interview: Silver Bullet talks with Bob Blakley. IEEE Security & Privacy, 7(5):5–8, September/October 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2009:ISBe

[McG09c]  Gary McGraw.  Interview:

Silver Bullet talks with Fred Schneider. IEEE Security & Privacy, 7(6):5–7, November/ December 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2009:ISBa

[McG09d] Gary McGraw. Interview: Silver Bullet talks with Gunnar Peterson. IEEE Security & Privacy, 7(1):7–11, January/ February 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2009:ISBb

[McG09e] Gary McGraw. Interview: Silver Bullet talks with Jeremiah Grossman. IEEE Security & Privacy, 7(2):10–14, March/ April 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2009:ISBc

[McG09f] Gary McGraw. Interview: Silver bullet talks with Virgil Gligor. IEEE Security & Privacy, 7(4):11–14, July/August 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2010:ISBa

[McG10a] Gary McGraw. Interview: Silver Bullet talks with Christofer Hoff. IEEE Security & Privacy, 8(1):8–10, January/February 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2010:ISBb

[McG10b] Gary McGraw. Interview: Silver Bullet talks with Gillian Hayes. IEEE Security & Privacy, 8(2):5–7, March/April 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2010:ISBe

[McG10c] Gary McGraw. Interview: Silver Bullet talks with Greg Morrisett. IEEE Security & Privacy, 8(5):6–9, September/ October 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2010:ISBd

[McG10d] Gary McGraw. Interview: Silver Bullet talks with Richard Clarke. IEEE Security & Privacy, 8(4):5–11, July/August 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2010:ISBc

[McG10e] Gary McGraw. Interview: Silver Bullet talks with Steven Kent. IEEE Security & Privacy, 8(3):5–9, May/June 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2010:SBT

[McG10f] Gary McGraw. Silver bullet talks with Ivan Arce. IEEE Security & Privacy, 8(6):9– 13, November/December 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2011:ISBb

[McG11a] Gary McGraw. Interview: Silver Bullet talks with David Rice. IEEE Security & Privacy, 9(2):8–11, March/April 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2011:ISBd

[McG11b] Gary McGraw. Interview: Silver Bullet talks with Elinor Mills. IEEE Security & Privacy, 9(5):9–12, September/ October 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2011:ISBe

[McG11c]  Gary McGraw.  Interview:

Silver Bullet talks with Halvar Flake. IEEE Security & Privacy, 9(6):5–8, November/ December 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2011:ISBc

[McG11d] Gary McGraw. Interview: Silver Bullet talks with John Savage. IEEE Security & Privacy, 9(4):9–12, July/August 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2011:ISBa

[McG11e] Gary McGraw. Interview: Silver Bullet talks with Paul Kocher. IEEE Security & Privacy, 9(1):8–11, January/ February 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2011:SBT

[McG11f] Gary McGraw. Silver Bullet talks with Ralph Langner. IEEE Security & Privacy, 9(3): 9–14, May/June 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2012:ISBa

[McG12a] Gary McGraw. Interview: Silver Bullet talks with Deborah Frincke. IEEE Security & Privacy, 10(1):11–14, January/ February 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2012:ISBc

[McG12b] Gary McGraw. Interview: Silver Bullet talks with Giovanni Vigna. IEEE Security & Privacy, 10(3):9–11, May/June 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2012:ISBf

[McG12c] Gary McGraw. Interview: Silver Bullet talks with Howard Schmidt. IEEE Security & Privacy, 10(6):9–12, November/ December 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2012:ISBe

[McG12d] Gary McGraw. Interview: Silver Bullet talks with Kay Connelly. IEEE Security & Privacy, 10(5):5–7, September/ October 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2012:ISBb

[McG12e]  Gary McGraw.  Interview:

Silver Bullet talks with Neil Daswani. IEEE Security & Privacy, 10(2):11–14, March/ April 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2012:ISBd

[McG12f] Gary McGraw. Interview: Silver Bullet talks with Randy Sabett. IEEE Security & Privacy, 10(4):19–22, July/ August 2012. CODEN ????

ISSN 1540-7993 (print), 1558-4046 (electronic).

McGettrick:2013:TEC

[McG13a] Andrew McGettrick. Toward effective cybersecurity education. IEEE Security & Privacy, 11(6):66–68, November/ December 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2013:PCP

[McG13b]  Deven McGraw.  Point/

counterpoint: Privacy and security as enable, not barrier, to responsible health data uses. IEEE Security & Privacy, 11(6):42–44, November/ December 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2013:ISBd

[McG13c] Gary McGraw. Interview: Silver Bullet talks with Gary Warzala. IEEE Security & Privacy, 11(4):8–10, July/August 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2013:ISBa

[McG13d] Gary McGraw. Interview: Silver Bullet talks with Per-Olof Persson. IEEE Security & Privacy, 11(1):8–10, January/ February 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2013:ISBb

[McG13e] Gary McGraw. Interview: Silver Bullet talks with Steve Bellovin. IEEE Security & Privacy, 11(2):8–11, March/April 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2013:ISBc

[McG13f] Gary McGraw. Interview: Silver Bullet talks with Thomas Rid. IEEE Security & Privacy, 11(3):8–10, May/June 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2013:ISBf

[McG13g] Gary McGraw. Interview: Silver Bullet talks with W. Hord Tipton. IEEE Security & Privacy, 11(6):5–7, November/ December 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2013:ISBe

[McG13h] Gary McGraw. Interview: Silver Bullet talks with Wenyuan Xu. IEEE Security & Privacy, 11(5):8–10, September/ October 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2014:SBTe

[McG14a] Gary McGraw. Silver Bullet talks with Bart Miller. IEEE Security & Privacy, 12(5):6–8, September/October 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/05/
msp2014050006.html.

McGraw:2014:SBTa

[McG14b] Gary McGraw. Silver Bullet talks with Jon Callas. IEEE Security & Privacy, 12(1):6–8, January/February 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2014:SBTb

[McG14c] Gary McGraw. Silver Bullet talks with Matthew Green. IEEE Security & Privacy, 12 (2):7–10, March/April 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2014:SBTd

[McG14d] Gary McGraw. Silver Bullet talks with Nate Fick. IEEE Security & Privacy, 12(4):11– 13, July/August 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/04/
msp2014040011.html.

McGraw:2014:SBTf

[McG14e] Gary McGraw. Silver Bullet talks with the IEEE Center for Secure Design. IEEE Security & Privacy, 12(6):9– 12, November/December 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/06/
msp2014060009.html.

McGraw:2014:SBTc

[McG14f] Gary McGraw. Silver Bullet talks with Yoshi Kohno. IEEE Security & Privacy, 12(3):9–12, May/June 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2015:SBTe

[McG15a] Gary McGraw. Silver Bullet talks with Bart Preneel. IEEE Security & Privacy, 13(5):8–10, September/October 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/csdl/mags/sp/2015/05/
msp2015050008-abs.html.

McGraw:2015:SBTa

[McG15b] Gary McGraw. Silver Bullet talks with Brian Krebs. IEEE Security & Privacy, 13(1):7–11, January/February 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/01/
msp2015010007.html.

McGraw:2015:SBTd

[McG15c] Gary McGraw. Silver Bullet talks with Katie Moussouris. IEEE Security & Privacy, 13 (4):7–9, July/August 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/04/
msp2015040007.html.

McGraw:2015:SBTc

[McG15d] Gary McGraw. Silver Bullet talks with L. Jean Camp. IEEE Security & Privacy, 13 (3):5–7, May/June 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/sp/2015/03/
                                        msp2015030005.html.

McGraw:2015:SBTf

[McG15e] Gary McGraw. Silver Bullet talks with Steven M. Bellovin and Matthew Green. IEEE Security & Privacy, 13(6):12– 15, November/December 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2015:SBTb

[McG15f] Gary McGraw. Silver Bullet talks with Whitfield Diffie. IEEE Security & Privacy, 13 (2):7–10, March/April 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/sp/2015/02/
                                        msp2015020007.html.

McGraw:2016:SBTc

[McG16a] Gary McGraw. Silver Bullet talks with Jacob West. IEEE Security & Privacy, 14(3):7–10, May/June 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2016:SBTb

[McG16b] Gary McGraw. Silver Bullet talks with Jamie Butler. IEEE Security & Privacy, 14(2):8–10, March/April 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2016:SBTe

[McG16c] Gary McGraw. Silver Bullet talks with Jim Manico. IEEE Security & Privacy, 14(6):3– 5, November/December 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/06/
msp2016060003.html.

McGraw:2016:SBTd

[McG16d] Gary McGraw. Silver Bullet talks with Martin Hellman. IEEE Security & Privacy, 14 (4):7–11, July/August 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/04/
msp2016040007.html.

McGraw:2016:SBTa

[McG16e] Gary McGraw. Silver Bullet talks with Peiter (Mudge) Zatko. IEEE Security & Privacy, 14(1):7–10, January/February 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2017:SBTc

[McG17a] Gary McGraw. Silver Bullet talks with Kate Pearce. IEEE Security & Privacy, 15 (3):6–9, May/June 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/03/
                                        msp2017030006.html.

McGraw:2017:SBTd

[McG17b] Gary McGraw. Silver Bullet talks with Kelly Lum. IEEE Security & Privacy, 15(4):7– 10, July/August 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/04/
                                        msp2017040007.html.

McGraw:2017:SBTe

[McG17c] Gary McGraw. Silver Bullet talks with Ksenia Dmitrieva-Peguero. IEEE Security & Privacy, 15(5):7–9, September/October 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/05/
                                        msp2017050007.html.

McGraw:2017:SBTb

[McG17d] Gary McGraw. Silver Bullet talks with Lesley Carhart. IEEE Security & Privacy, 15 (2):8–10, March/April 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/02/
                                        msp2017020008.html.

McGraw:2017:SBTa

[McG17e] Gary McGraw. Silver Bullet talks with Marie Moe. IEEE Security & Privacy, 15(1):8–11, January/February 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/01/
msp2017010008.html.

McGraw:2017:SBTf

[McG17f] Gary McGraw. Silver Bullet talks with Nicole Perlroth. IEEE Security & Privacy, 15(6):7–9, November/December 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/06/
msp2017060007.html.

McGraw:2018:SBTf

[McG18a] G. McGraw. Silver Bullet talks with Kathleen Fisher [interview]. IEEE Security & Privacy, 16(6):4–9, November/ December 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2018:SBTe

[McG18b] G. McGraw. Silver Bullet talks with Tanya Janca. IEEE Security & Privacy, 16(5):7– 11, September/October 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

McGraw:2018:SBTc

[McG18c] Gary McGraw. Silver Bullet talks with Bruce Potter. IEEE Security & Privacy, 16(3):12– 14, May/June 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/03/
                                        msp2018030012.html.

McGraw:2018:SBTb

[McG18d] Gary McGraw. Silver Bullet talks with Craig Froelich. IEEE Security & Privacy, 16 (2):6–8, March/April 2018. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/02/
                                        msp2018020006.html.

McGraw:2018:SBTd

[McG18e] Gary McGraw. Silver Bullet talks with Nick Weaver. IEEE Security & Privacy, 16(4):7– 10, July/August 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/04/
                                        msp2018040007.html.

McGraw:2018:SBTa

[McG18f] Gary McGraw. Silver Bullet talks with Wafaa Mamilli. IEEE Security & Privacy, 16(1):6–9, January/February 2018. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL https:

                                        //www.computer.org/csdl/
                                        mags/sp/2018/01/msp2018010006.
                                        html.

McKinney:2007:VB

[McK07] David McKinney. Vulnerability bazaar. IEEE Security & Privacy, 5(6):69–73, November/ December 2007. CODEN ????

ISSN 1540-7993 (print), 1558-4046 (electronic).

McKinney:2008:NHV

[McK08] David McKinney. New hurdles for vulnerability disclosure. IEEE Security & Privacy, 6(2):76–78, March/April 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Matwyshyn:2010:ESV

[MCKS10] Andrea M. Matwyshyn, Ang Cui, Angelos D. Keromytis, and Salvatore J. Stolfo. Ethics in security vulnerability research. IEEE Security & Privacy, 8(2):67–72, March/April 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McLaughlin:2005:IHS

[McL05a] L. McLaughlin. Interview:

Holistic security. IEEE Security & Privacy, 3(3):6–8, May/ June 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/31002/01439494.pdf;
http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
31002&arnumber=1439494&count=
20&index=2.

McLaughlin:2005:AGP

[McL05b] Laurianne McLaughlin. From AWK to Google: Peter Weinberger talks search. IEEE Security & Privacy, 3(5):11–13, September/October 2005. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/comp/mags/sp/2005/05/
j5011.pdf.

McLaughlin:2005:WGR

[McL05c] Laurianne McLaughlin. Winning the game of risk: Neumann’s take on sound design. IEEE Security & Privacy, 3 (6):9–12, November/December 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/33104/01556531.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
33104&arnumber=1556531.

McLaughlin:2006:PZW

[McL06] Laurianne McLaughlin. Philip Zimmermann on What’s Next after PGP? IEEE Security & Privacy, 4(1):10–13, January/February 2006. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/33481/01588818.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
33481&arnumber=1588818.

McLaughlin:2008:PIC

[McL08] Peter McLaughlin. Privacy interests: Cross-border data flows and increased enforcement. IEEE Security & Privacy, 6(5):58–61, September/ October 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McLaughlin:2013:SCS

[McL13] Stephen McLaughlin. Securing control systems from the inside: A case for mediating physical behaviors. IEEE Security & Privacy, 11(4):82–84, July/August 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McKenzie:2008:UCI

[MCW08] Robin McKenzie, Malcolm Crompton, and Colin Wallis. Use cases for identity management in E-government. IEEE Security & Privacy, 6(2):51–57, March/April 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Mazurczyk:2017:RAD

[MCW17] Wojciech Mazurczyk, Luca Caviglione, and Steffen Wendzel. Recent advancements in digital forensics. IEEE Security & Privacy, 15(6):10–11, November/December 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/06/
msp2017060010.html.

Mazurczyk:2019:RAD

[MCW19] W. Mazurczyk, L. Caviglione, and S. Wendzel. Recent advancements in digital forensics, part 2. IEEE Security & Privacy, 17(1):7–8, January/ February 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Manes:2009:OLL

[MD09a] Gavin W. Manes and Elizabeth Downing. Overview of licensing and legal issues for digital forensic investigators. IEEE Security & Privacy, 7(2):45–48, March/April 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Massey:2009:SDN

[MD09b] Daniel Massey and Dorothy E. Denning. Securing the Domain Name System: Guest Editors’ introduction. IEEE Security & Privacy, 7(5):11–13, September/October 2009. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Mirkovic:2015:ECE

[MDD^{+}15] Jelena Mirkovic, Melissa Dark, Wenliang Du, Giovanni Vigna, and Tamara Denning. Evaluating cybersecurity education interventions: Three case studies. IEEE Security & Privacy, 13(3):63–69, May/June 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/03/
msp2015030063-abs.html.

Momenzadeh:2020:BPW

[MDR^{+}20] B. Momenzadeh, H. Dougherty, M. Remmel, S. Myers, and L. J. Camp. Best practices would make things better in the IoT. IEEE Security & Privacy, 18 (4):38–47, July/August 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

McDaniel:2010:GEW

[ME10] Patrick McDaniel and William Enck. Not so great expectations: Why application markets haven’t failed security. IEEE Security & Privacy, 8 (5):76–78, September/October 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Mead:2003:BF

[Mea03a]  Nancy R. Mead.  Build-

ing a foundation. IEEE Security & Privacy, 1(3): 14, May/June 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j3014.pdf.

Mead:2003:SFH

[Mea03b] Nancy R. Mead. SEHAS 2003:

The future of high-assurance systems. IEEE Security & Privacy, 1(5):68–72, September/ October 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2003/05/j5068abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/05/j5068.
pdf.

Meinel:2004:CTC

[Mei04] Carolyn P. Meinel. Cybercrime Treaty could chill research. IEEE Security & Privacy, 2(4):28–32, July/ August 2004. CODEN ????

ISSN 1540-7993 (print), 1558-4046 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/sp/2004/04/j4028.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2004/04/j4028.
                                        pdf.

Meier:2006:WAS

[Mei06] J. D. Meier. Web application security engineering. IEEE Security & Privacy, 4(4):16–24, July/August 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Meiklejohn:2018:TTO

[Mei18] Sarah Meiklejohn. Top ten obstacles along distributed ledgers path to adoption. IEEE Security & Privacy, 16(4):13– 19, July/August 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/04/
                                        msp2018040013-abs.html.

Mentens:2024:HSE

[Men24] Nele Mentens. Hardware security in the era of emerging device and system technologies. IEEE Security & Privacy, 22 (3):4–6, May/June 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Mailloux:2015:PEQ

[MGH^{+}15] Logan O. Mailloux, Michael R. Grimaila, Douglas D. Hodson, Gerald Baumgartner, and Colin McLaughlin. Performance evaluations of quantum key distribution system architectures. IEEE Security & Privacy, 13(1):30–40, January/February 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/01/
msp2015010030-abs.html.

McGraw:2007:OGS

[MH07] Gary McGraw and Greg Hoglund. Online games and security. IEEE Security & Privacy, 5(5):76–79, September/ October 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Mead:2013:BSP

[MH13] Nancy R. Mead and Thomas B.

Hilburn. Building security in: Preparing for a software security career. IEEE Security & Privacy, 11(6):80–83, November/December 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Majeed:2023:RSS

[MH23] Abdul Majeed and Seong Oun Hwang. Rectification of syntactic and semantic privacy mechanisms. IEEE Security & Privacy, 21(5):18–32, September/ October 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Momen:2019:DAP

[MHF19] N. Momen, M. Hatamian, and L. Fritsch. Did app privacy improve after the GDPR? IEEE Security & Privacy, 17(6):10– 20, November/December 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Matheu:2019:TCC

[MHRS19] S. N. Matheu, J. L. Hernandez-Ramos, and A. F. Skarmeta. Toward a cybersecurity certification framework for the Internet of Things. IEEE Security & Privacy, 17(3):66–76, May/June 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

McLean:2018:LE

[MHV18] John D. McLean, Cormac Herley, and P. C. Van Oorschot. Letter to the Editor. IEEE Security & Privacy, 16(3):6– 10, May/June 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/03/
msp2018030006.html.

Michael:2008:EGT

[Mic08a] Bret Michael. From the Editors: Are governments up to the task? IEEE Security & Privacy, 6(6):4–5, November/ December 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Michener:2008:CPM

[Mic08b] John R. Michener. Common permissions in Microsoft Windows Server 2008 and Windows Vista. IEEE Security & Privacy, 6(3):63–67, May/June 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Michael:2009:ECS

[Mic09] Bret Michael. From the editors: In clouds shall we trust? IEEE Security & Privacy, 7(5): 3, September/October 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Michaels:2010:BLS

[Mic10a] Bret Michaels. Balancing liberty, stability, and security. IEEE Security & Privacy, 8(3): 3, May/June 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Michiels:2010:OWB

[Mic10b] Wil Michiels. Opportunities in white-box cryptography. IEEE Security & Privacy, 8(1):64–67, January/February 2010. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Michael:2019:TAM

[Mic19] J. B. Michael. Trustworthiness of autonomous machines in armed conflict. IEEE Security & Privacy, 17(6):4–6, November/December 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Michael:2021:SPE

[Mic21] James Bret Michael. Security and privacy for edge artificial intelligence. IEEE Security & Privacy, 19(4):4–7, July/August 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Michael:2022:TMA

[Mic22] James Bret Michael. Taking a measured approach to investing in information infrastructure for attaining leadingedge trustworthy artificial intelligence. IEEE Security & Privacy, 20(5):4–6, September/October 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Miller:2003:LBL

[Mil03] Sandra Kay Miller. Legal battle looming for Internet protections acts. IEEE Security & Privacy, 1(1):10–12, January/February 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j1010.pdf.

Miller:2005:TPC

[Mil05] Ann Miller. Trends in process control systems security. IEEE Security & Privacy, 3(5):57–60, September/October 2005. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Miller:2011:MAD

[Mil11] Charlie Miller. Mobile attacks and defense. IEEE Security & Privacy, 9(4):68–70, July/ August 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Mabry:2007:USE

[MJF07] Frank J. Mabry, John R.

James, and Aaron J. Ferguson. Unicode steganographic exploits: Maintaining enterprise border security. IEEE Security & Privacy, 5(5):32–39, September/October 2007. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Massacci:2021:SCP

[MJP21] Fabio Massacci, Trent Jaeger, and Sean Peisert. SolarWinds and the challenges of patching: Can we ever stop dancing with the devil? IEEE Security & Privacy, 19(2):14–19, March/ April 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Matyas:2013:CBI

[MK13] Vashek Matyáš and Jiří Kůr.

Conflicts between intrusion detection and privacy mechanisms for wireless sensor networks. IEEE Security & Privacy, 11(5):73–76, September/ October 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Mohammed:2021:RSP

[MKF21] S. Mohammed, T.-H. Kim, and W. C. Fang. Requirements for security, privacy, and trust in the Internet of Things [Guest Editors’ introduction]. IEEE Security & Privacy, 19(1):8–10, January/February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

McGrath:2009:PIF

[MKG09] D. Kevin McGrath, Andrew Kalafut, and Minaxi Gupta. Phishing infrastructure fluxes all the way. IEEE Security & Privacy, 7(5):21–28, September/October 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Mitterhofer:2009:SSB

[MKKP09] Stefan Mitterhofer, Christopher Kruegel, Engin Kirda, and Christian Platzer. Serverside bot detection in massively multiplayer online games. IEEE Security & Privacy, 7(3): 29–36, May/June 2009. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Mullins:2007:HCD

[MLM^{+}07] Barry E. Mullins, Timothy H. Lacey, Robert F. Mills, Joseph E. Trechter, and Samuel D. Bass. How the cyber defense exercise shaped an information-assurance curriculum. IEEE Security & Privacy, 5(5):40–49, September/ October 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Martin:2022:FMS

[MLS22] William Martin, Patrick Lincoln, and William Scherlis. Formal methods at scale. IEEE Security & Privacy, 20(3):22– 23, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Mead:2005:PSS

[MM05] Nancy R. Mead and Gary Mc-Graw. A portal for software security. IEEE Security & Privacy, 3(4):75–79, July/August 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Montgomery:2006:TSR

[MM06] Doug Montgomery and Sandra Murphy. Toward secure routing infrastructures. IEEE Security & Privacy, 4(5):84–87, September/October 2006. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

MacKie-Mason:2009:SPE

[MM09a] Jeffrey K. MacKie-Mason. Security & privacy economics: Incentive-centered design for security. IEEE Security & Privacy, 7(4):80–83, July/August 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McDaniel:2009:SSS

[MM09b] Patrick McDaniel and Stephen McLaughlin. Secure systems: Security and privacy challenges in the Smart Grid. IEEE Security & Privacy, 7(3):75–77, May/June 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Moran:2010:PMP

[MM10] Tal Moran and Tyler Moore.

The Phish-Market Protocol:

Secure sharing between competitors. IEEE Security & Privacy, 8(4):40–45, July/August 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

MacKie-Mason:2011:ASW

[MM11] Jeffrey MacKie-Mason. All space will be public space. IEEE Security & Privacy, 9 (5):77–80, September/October 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

MacKie-Mason:2014:CWA

[MM14] Jeffrey MacKie-Mason. Can we afford privacy from surveillance? IEEE Security & Privacy, 12(5):86–89, September/October 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/05/
msp2014050086-abs.html.

Mulholland:2017:DCD

[MMB17] John Mulholland, Michele Mosca, and Johannes Braun. The day the cryptography dies. IEEE Security & Privacy, 15 (4):14–21, July/August 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/04/
msp2017040014-abs.html.

Mailloux:2016:PSS

[MMKP16] Logan O. Mailloux, Michael A. McEvilley, Stephen Khou, and John M. Pecarina. Putting the ‘Systems’ in security engineering: An examination of NIST Special Publication 800-160. IEEE Security & Privacy, 14 (4):76–80, July/August 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/04/
msp2016040076-abs.html.

Moreira:2024:SRA

[MMR24] Daniel Moreira, Sébastien Marcel, and Anderson Rocha. Synthetic realities and artificial intelligence-generated contents. IEEE Security & Privacy, 22(3):7–10, May/June 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Massacci:2021:DFE

[MN21] F. Massacci and C. N.

Ngo. Distributed financial exchanges: Security challenges and design principles. IEEE Security & Privacy, 19(1):54– 64, January/February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Mathieu:2008:VIS

[MNS08] Bertrand Mathieu, Saverio Niccolini, and Dorgham Sisalem. Voice over IP: SDRS: a Voiceover-IP spam detection and reaction system. IEEE Security & Privacy, 6(6):52–59, November/December 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Maimut:2012:LCR

[MO12] Diana Maimut and Khaled Ouafi. Lightweight cryptography for RFID tags. IEEE Security & Privacy, 10(2):76–79, March/April 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Murdock:2020:PHL

[MOG^{+}20] K. Murdock, D. Oswald, F. D. Garcia, J. Van Bulck, F. Piessens, and D. Gruss. Plundervolt: How a little bit of undervolting can create a lot of trouble. IEEE Security & Privacy, 18(5):28–37, September/ October 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Molotch:2013:ESD

[Mol13]  Harvey Molotch.  Every-

day security: Default to decency. IEEE Security & Privacy, 11(6):84–87, November/ December 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Moriarty:2011:IC

[Mor11] Kathleen M. Moriarty. Incident coordination. IEEE Security & Privacy, 9(6):71–75, November/December 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Mosca:2018:CEQ

[Mos18] M. Mosca. Cybersecurity in an era with quantum computers: Will we be ready? IEEE Security & Privacy, 16(5):38– 41, September/October 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Matthews:2017:SPE

[MOT^{+}17] Tara Matthews, Kathleen O’Leary, Anna Turner, Manya Sleeper, Jill Palzkill Woelfer, Martin Shelton, Cori Manthorne, Elizabeth F. Churchill, and Sunny Consolvo. Security and privacy experiences and practices of survivors of intimate partner abuse. IEEE Security & Privacy, 15(5):76–81, September/October 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/05/
msp2017050076.html.

McGraw:2004:SST

[MP04] Gary McGraw and Bruce Potter. Software security testing. IEEE Security & Privacy, 2(5):81–85, September/October 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/05/j5081.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/05/j5081.
pdf.

Morris:2007:WWY

[MP07] John Morris and Jon Peterson. Who’s watching you now? IEEE Security & Privacy, 5(1): 76–79, January/February 2007.

CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

McGovern:2010:QDC

[MP10] James McGovern and Gunnar Peterson. 10 quick, dirty, and cheap things to improve enterprise security. IEEE Security & Privacy, 8(2):83–85, March/April 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Massacci:2021:TLD

[MP21] Fabio Massacci and Ivan Pashchenko. Technical leverage: Dependencies are a mixed blessing. IEEE Security & Privacy, 19(3):58–62, May/June 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Mangaokar:2022:DMC

[MP22] Neal Mangaokar and Atul Prakash. Dispelling misconceptions and characterizing the failings of deepfake detection. IEEE Security & Privacy, 20 (2):61–67, March/April 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Marjanov:2022:MLS

[MPM22] Tina Marjanov, Ivan Pashchenko, and Fabio Massacci. Machine learning for source code vulnerability detection: What works and what isn’t there yet. IEEE Security & Privacy, 20(5):60– 76, September/October 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

McParland:2014:MSN

[MPS14] Chuck McParland, Sean Peisert, and Anna Scaglione. Monitoring security of networked control systems: It’s the physics. IEEE Security & Privacy, 12(6):32–39, November/December 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/06/
msp2014060032-abs.html.

Matyas:2003:TRU

[MR03] Václav Matyás, Jr. and Zdenek Riha. Toward reliable user authentication through biometrics. IEEE Security & Privacy, 1(3):45–49, May/June 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j3045.pdf;   http://www.
computer.org/security/j3045abs.
htm.

Maxion:2005:MFE

[MR05] R. A. Maxion and R. R. M.

Roberts. Methodological foundations: enabling the next generation of security. IEEE Security & Privacy, 3(2):54– 57, March/April 2005. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/30742/01423962.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=

                                        30742&arnumber=1423962&count=
                                        15&index=8.

Maler:2008:VIO

[MR08] Eve Maler and Drummond Reed. The Venn of identity: Options and issues in federated identity management. IEEE Security & Privacy, 6(2):16–23, March/April 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Maimut:2014:AET

[MR14] Diana Maimut and Reza Reyhanitabar. Authenticated encryption: Toward nextgeneration algorithms. IEEE Security & Privacy, 12(2):70– 72, March/April 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Moody:2022:CSP

[MR22] Dustin Moody and Angela Robinson. Cryptographic standards in the post-quantum era. IEEE Security & Privacy, 20(6):66–72, November/ December 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Mhaidli:2024:SML

[MRF^{+}24] Abraham Mhaidli, Shwetha Rajaram, Selin Fidan, Gina Herakovic, and Florian Schaub. Shockvertising, malware, and a lack of accountability: Exploring consumer risks of virtual reality advertisements and marketing experiences. IEEE Security & Privacy, 22(1):43–52, January/February 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Meza:2023:SVO

[MRO^{+}23] Andres Meza, Francesco Restuccia, Jason Oberg, Dominic Rizzo, and Ryan Kastner. Security verification of the Open-Titan hardware root of trust. IEEE Security & Privacy, 21 (3):27–36, May/June 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

McDaniel:2014:TSS

[MRS14] Patrick McDaniel, Brian Rivera, and Ananthram Swami. Toward a science of secure environments. IEEE Security & Privacy, 12(4):68– 70, July/August 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/04/
msp2014040068-abs.html.

Mack:2020:SMC

[MS20] H. Mack and T. Schroer. Security midlife crisis: Building security in a new world. IEEE Security & Privacy, 18(4):72– 74, July/August 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Menezes:2021:AES

[MS21a] Alfred Menezes and Douglas Stebila. The Advanced Encryption Standard: 20 years later. IEEE Security & Privacy, 19(6):98–102, November/ December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Menezes:2021:CC

[MS21b] Alfred Menezes and Douglas Stebila. Challenges in cryptography. IEEE Security & Privacy, 19(2):70–73, March/ April 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Menezes:2021:EES

[MS21c] Alfred Menezes and Douglas Stebila. End-to-end security: When do we have it? IEEE Security & Privacy, 19(4):60–64, July/August 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Malecha:2022:DFM

[MSeF^{+}22] Gregory Malecha, Gordon Stewart, Franti ek Farka, Jasper Haag, and Yoichi Hirai. Developing with formal methods at BedRock Systems, Inc. IEEE Security & Privacy, 20(3):33–42, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Massacci:2022:FFP

[MSM^{+}22] Fabio Massacci, Antonino Sabetta, Jelena Mirkovic, Toby Murray, Hamed Okhravi, Mohammad Mannan, Anderson Rocha, Eric Bodden, and Daniel E. Geer. Free as in freedom to protest? IEEE Security & Privacy, 20(5):16– 21, September/October 2022.

ISSN 1540-7993 (print), 1558-4046 (electronic).

Mell:2006:CVS

[MSR06] Peter Mell, Karen Scarfone, and Sasha Romanosky. Common vulnerability scoring system. IEEE Security & Privacy, 4(6):85–89, November/ December 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Meike:2009:ISS

[MSW09] Michael Meike, Johannes Sametinger, and Andreas Wiesauer. Internet security: Security in open source Web content management systems. IEEE Security & Privacy, 7(4): 44–51, July/August 2009. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Milankovich:2023:DSC

[MT23] Ákos Milánkovich and Katja Tuma. Delta security certification for software supply chains. IEEE Security & Privacy, 21(6):24–33, November/ December 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Mbaka:2024:RGE

[MT24] Winnie Mbaka and Katja Tuma. Role of gender in the evaluation of security decisions. IEEE Security & Privacy, 22 (2):38–48, March/April 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Melara:2023:VSS

[MTA23] Marcela S. Melara and Santiago Torres-Arias. A viewpoint on software supply chain security: Are we getting lost in translation? IEEE Security & Privacy, 21(6):55– 58, November/December 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Meland:2015:MRC

[MTS15] P. H. Meland, I. A. Tondel, and B. Solhaug. Mitigating risk with cyberinsurance. IEEE Security & Privacy, 13(6):38– 43, November/December 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Mulvenon:2005:TCS

[Mul05] James Mulvenon. Toward a cyberconflict studies research agenda. IEEE Security & Privacy, 3(4):52–55, July/August 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Mulligan:2014:EIT

[Mul14] Deirdre K. Mulligan. The enduring importance of transparency. IEEE Security & Privacy, 12(3):61–65, May/June 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Mussington:2021:CRU

[Mus21] D. Mussington. Calculating risks: Understanding the 2020 election experience and defining the “New Normal”. IEEE Security & Privacy, 19(1):74– 80, January/February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Massacci:2022:BPL

[MV22] Fabio Massacci and Silvia Vidor. Building principles for lawful cyber lethal autonomous weapons. IEEE Security & Privacy, 20(2):101–106, March/ April 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Michael:2009:IAD

[MVL09] Bret Michael, Jeffrey Voas, and Phil Laplante. It all depends: Cyberpandemics: History, inevitability, response. IEEE Security & Privacy, 7(1):63–67, January/February 2009. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Massacci:2023:SSC

[MW23] Fabio Massacci and Laurie Williams. Software supply chain security [Guest Editors introduction]. IEEE Security & Privacy, 21(6):8–10, November/December 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Mitra:2013:VCP

[MWM13] Sayan Mitra, Tichakorn Wongpiromsarn, and Richard M. Murray. Verifying cyberphysical interactions in safetycritical systems. IEEE Security & Privacy, 11(4):28–37, July/ August 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Mihelic:2022:SSD

[MŽ22] Anže Mihelič and Boštjan Žvanut. (In)secure smart device use among senior citizens. IEEE Security & Privacy, 20 (1):62–71, January/February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Naccache:2005:FF

[Nac05]  David Naccache.  Finding

faults. IEEE Security & Privacy, 3(5):61–65, September/ October 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Naccache:2016:FHE

[Nac16] David Naccache. Fully homomorphic encryption: Computations with a blindfold. IEEE Security & Privacy, 14(1):63– 67, January/February 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Nahra:2008:PIH

[Nah08] Kirk J. Nahra. Privacy interests: HIPAA security enforcement is here. IEEE Security & Privacy, 6(6):70–72, November/December 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Nance:2009:ETT

[Nan09] Kara Nance. Education: Teach them when they aren’t looking: Introducing security in CS1. IEEE Security & Privacy, 7(5):53–55, September/ October 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Narayanan:2013:WHCa

[Nar13a] Arvind Narayanan. What happened to the crypto dream?, Part 1. IEEE Security & Privacy, 11(2):75–76, March/ April 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Narayanan:2013:WHCb

[Nar13b] Arvind Narayanan. What happened to the crypto dream?, Part 2. IEEE Security & Privacy, 11(3):68–71, May/June 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Narayanan:2023:OLE

[NB23] Vikram Narayanan and Anton Burtsev. The opportunities and limitations of extended page table switching for finegrained isolation. IEEE Security & Privacy, 21(3):16–26, May/June 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Nance:2008:VMI

[NBH08] Kara Nance, Matt Bishop, and Brian Hay. Virtual machine introspection: Observation or interference? IEEE Security & Privacy, 6(5):32–37, September/October 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ni:2009:ACP

[NBLC09] Qun Ni, Elisa Bertino, Jorge Lobo, and Seraphin B. Calo. Access control: Privacy-aware role-based access control. IEEE Security & Privacy, 7(4):35–43, July/August 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Nestler:2019:NCP

[NCA19] V. Nestler, T. Coulson, and J. D. Ashley III. The NICE Challenge Project: Providing workforce experience before the workforce. IEEE Security & Privacy, 17(2):73–78, March/ April 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Nanayakkara:2023:WDC

[NH23] Priyanka Nanayakkara and Jessica Hullman. What’s driving conflicts around differential privacy for the U.S. Census. IEEE Security & Privacy, 21 (5):33–42, September/October 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Nicol:2005:MSS

[Nic05] David M. Nicol. Modeling and simulation in security evaluation. IEEE Security & Privacy, 3(5):71–74, September/ October 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Nicol:2020:PDC

[Nic20a] D. M. Nicol. In the Petri dish: Cybersecurity pushed to the edge. IEEE Security & Privacy, 18(3):4–5, May/June 2020. ISSN 1558-4046.

Nicol:2020:VUA

[Nic20b] D. M. Nicol. The value of useless academic research to the cyberdefense of critical infrastructures. IEEE Security & Privacy, 18(1):4–7, January/ February 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Nicol:2021:MIS

[Nic21a] D. M. Nicol. Message from IEEE S&P’s outgoing Editor in Chief. IEEE Security & Privacy, 19(1):10, January/ February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Nicol:2021:RTE

[Nic21b] David M. Nicol. The ransomware threat to energydelivery systems. IEEE Security & Privacy, 19(3):24–32, May/June 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Nissenbaum:2024:ASP

[Nis24] Helen Nissenbaum. AI safety:

a poisoned chalice? IEEE Security & Privacy, 22(2):94–96, March/April 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Nazario:2012:IIS

[NK12] Jose Nazario and John Kristoff. Internet infrastructure security. IEEE Security & Privacy, 10(4):24–25, July/ August 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Nikiforakis:2014:WCP

[NKJ^{+}14] Nick Nikiforakis, Alexandros Kapravelos, Wouter Joosen, Christopher Kruegel, Frank Piessens, and Giovanni Vigna. On the workings and current practices of Web-based device fingerprinting. IEEE Security & Privacy, 12(3):28–36, May/ June 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Neumann:2012:ISS

[NL12] Peter G. Neumann and Ulf Lindqvist. The IEEE Symposium on Security and Privacy is moving to San Francisco. IEEE Security & Privacy, 10(2):65– 66, March/April 2012. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Narayanan:2023:SPA

[NL23] Arvind Narayanan and Kevin Lee. Security policy audits: Why and how. IEEE Security & Privacy, 21(2):77–81, March/April 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Nichols:2007:MFD

[NP07] Elizabeth A. Nichols and Gunnar Peterson. A metrics framework to drive application security improvement. IEEE Security & Privacy, 5(2):88–91, March/April 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Neumann:2014:ISS

[NPS14] Peter G. Neumann, Sean Peisert, and Marvin Schaefer. The IEEE Symposium on Security and Privacy, in retrospect. IEEE Security & Privacy, 12(3):15–17, May/June 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Nautiyal:2022:UKC

[NRH^{+}22] Lata Nautiyal, Awais Rashid, Joseph Hallett, Ben Shreeve, Michael K., Chris E., and Catherine H. The United Kingdom’s Cyber Security Degree Certification program: a cyber security body of knowledge case study. IEEE Security & Privacy, 20(1):87–95, January/ February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Nair:2024:TMU

[NROS24] Vivek Nair, Louis Rosenberg, James F. O’Brien, and Dawn Song. Truth in motion: The unprecedented risks and opportunities of extended reality motion data. IEEE Security & Privacy, 22(1):24–32, January/ February 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Nguyen:2011:CIT

[NS11] Quyen L. Nguyen and Arun Sood. A comparison of intrusion-tolerant system architectures. IEEE Security & Privacy, 9(4):24–31, July/ August 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Nicol:2023:TCW

[NSA^{+}23] David M. Nicol, Gregory Shannon, Monika Akbar, Matt Bishop, Michael Chaney, and Matthew Luallen. Toward common weakness enumerations in industrial control systems. IEEE Security & Privacy, 21(4):84–93, July/ August 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Nour:2020:SPC

[NSLW20] B. Nour, K. Sharif, F. Li, and Y. Wang. Security and privacy challenges in informationcentric wireless Internet of Things networks. IEEE Security & Privacy, 18(2):35–45, March/April 2020. ISSN 1558-4046.

Nicol:2008:UGN

[NSSS08] David M. Nicol, William H. Sanders, Sankalp Singh, and Mouna Seri. Usable global network access policy for process control systems. IEEE Security & Privacy, 6(6):30–36, November/December 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Nguyen-Tuong:2018:XAC

[NTMD^{+}18] Anh Nguyen-Tuong, David Melski, Jack W. Davidson, Michele Co, William Hawkins, Jason D. Hiser, Derek Morris, Ducson Nguyen, and Eric Rizzi. Xandra: An autonomous cyber battle system for the cyber grand challenge. IEEE Security & Privacy, 16(2):42– 51, March/April 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/02/
msp2018020042-abs.html.

Otto:2007:CDH

[OAB07] Paul N. Otto, Annie I. Antón, and David L. Baumer. The ChoicePoint dilemma: How data brokers should handle the privacy of personal information. IEEE Security & Privacy, 5(5):15–23, September/ October 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Okhravi:2020:OGL

[OBS^{+}20] H. Okhravi, N. Burow, R. Skowyra, B. C. Ward, S. Jero, R. Kazan, and H. Shrobe. One giant leap for computer security. IEEE Security & Privacy, 18(4):8–19, July/August 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Older:2012:EAU

[OC12] Susan Older and Shiu-Kai Chin. Engineering assurance at the undergraduate level. IEEE Security & Privacy, 10(6):74– 77, November/December 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

ODonnell:2007:EMS

[O’D07] Adam J. O’Donnell. The evolutionary microcosm of stock spam. IEEE Security & Privacy, 5(1):70–72, January/ February 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

ODonnell:2008:WMA

[O’D08] Adam J. O’Donnell. When malware attacks (anything but Windows). IEEE Security & Privacy, 6(3):68–70, May/June 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Obrenovic:2012:IUC

[OdH12] Željko Obrenovic and Bart den Haak. Integrating user customization and authentication: The identity crisis. IEEE Security & Privacy, 10(5):82–85, September/October 2012. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Oehlert:2005:VAF

[Oeh05] P. Oehlert. Violating assumptions with fuzzing. IEEE Security & Privacy, 3(2):58– 62, March/April 2005. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/30742/01423963.pdf;
http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
30742&arnumber=1423963&count=
15&index=9.

OHagan:2024:VSI

[OGM^{+}24] Joseph O’Hagan, Jan Gugenheimer, Florian Mathis, Jolie Bonner, Richard Jones, and Mark McGill. A viewpoint on the societal impact of everyday augmented reality and the need for perceptual human rights. IEEE Security & Privacy, 22(1):64–68, January/ February 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Okhravi:2014:FFB

[OHBS14] Hamed Okhravi, Thomas Hobson, David Bigelow, and William Streilein. Finding focus in the blur of moving-target techniques. IEEE Security & Privacy, 12(2):16–26, March/ April 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ohm:2014:SSW

[Ohm14] Paul Ohm. Should sniffing Wi-Fi be illegal? IEEE Security & Privacy, 12(1):73–76, January/ February 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Obermeier:2023:ARM

[OJR^{+}23] Sebastian Obermeier, Thomas Jösler, Stephan Renggli, Maurus Unternährer, and Bernhard M. Hämmerli. Automating recovery in mixed operation technology/IT critical infrastructures. IEEE Security & Privacy, 21(5):43–54, September/October 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Oshri:2007:ISS

[OKH07] Ilan Oshri, Julia Kotlarsky, and Corey Hirsch. An information security strategy for networkable devices. IEEE Security & Privacy, 5(5):50–56, September/October 2007. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Okhravi:2021:CM

[Okh21] Hamed Okhravi. A cybersecurity moonshot. IEEE Security & Privacy, 19(3):8–16, May/ June 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Oluwatimi:2017:OMC

[OMB17] Oyindamola Oluwatimi, Daniele Midi, and Elisa Bertino. Overview of mobile containerization approaches and open research directions. IEEE Security & Privacy, 15(1):22–31, January/February 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/01/
msp2017010022-abs.html.

Orman:2015:MPB

[OP15] Hilarie Orman and Charles P.

Pfleeger. Mathematics and physics build a new future for secure communication [Guest Editors’ introduction]. IEEE Security & Privacy, 13(1):12– 13, January/February 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/01/
msp2015010012.html.

Oppliger:2007:PCM

[Opp07] Rolf Oppliger. Providing certified mail services on the Internet. IEEE Security & Privacy, 5(1):16–22, January/February 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Oppliger:2015:QRA

[Opp15] R. Oppliger. Quantitative risk analysis in information security management: A modern fairy tale. IEEE Security & Privacy, 13(6):18–21, November/ December 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Oppliger:2017:DAB

[Opp17] Rolf Oppliger. Disillusioning Alice and Bob. IEEE Security & Privacy, 15(5):82–84, September/October 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/05/
msp2017050082-abs.html.

Oppliger:2003:DED

[OR03] Rolf Oppliger and Ruedi Rytz. Digital evidence: Dream and reality. IEEE Security & Privacy, 1(5):44–48, September/ October 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2003/05/j5044abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/05/j5044.
pdf.

Oppliger:2005:DTC

[OR05] R. Oppliger and R. Rytz. Does trusted computing remedy computer security problems? IEEE Security & Privacy, 3(2):16–19, March/April 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/30742/01423956.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
30742&arnumber=1423956&count=
15&index=3.

OReilly:2017:ARD

[O’R17]  David O’Reilly.  Availabil-

ity of required data to support criminal investigations involving large-scale IP addresssharing technologies. IEEE Security & Privacy, 15(5):90–93, September/October 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/05/
msp2017050090-abs.html.

Orman:2003:MWF

[Orm03] Hilarie Orman. The Morris worm: a fifteen-year perspective. IEEE Security & Privacy, 1(5):35–43, September/ October 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2003/05/j5035abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/05/j5035.
pdf.

Ortega:2006:NBa

[Ort06a] Brandi Ortega. News briefs.

IEEE Security & Privacy, 4 (5):8–10, September/October 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2006/05/j5008.pdf.

Ortega:2006:NBb

[Ort06b] Brandi Ortega. News briefs.

IEEE Security & Privacy, 4 (6):6–9, November/December 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ortega:2007:NBa

[Ort07a] Brandi Ortega. News briefs.

IEEE Security & Privacy, 5(1): 12–15, January/February 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ortega:2007:NBb

[Ort07b] Brandi Ortega. News briefs.

IEEE Security & Privacy, 5(2):

7–10, March/April 2007. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ortega:2007:NBc

[Ort07c] Brandi Ortega. News briefs. IEEE Security & Privacy, 5(3): 10–12, May/June 2007. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ortega:2007:NBd

[Ort07d] Brandi Ortega. News briefs. IEEE Security & Privacy, 5(4): 6–9, July/August 2007. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ortega:2007:NBe

[Ort07e] Brandi Ortega. News briefs. IEEE Security & Privacy, 5 (5):12–14, September/October 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/sp/2007/05/msp2007050012.
                                        pdf.

Ortega:2007:NBf

[Ort07f] Brandi Ortega. News briefs. IEEE Security & Privacy, 5 (6):6–7, November/December 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/sp/2007/06/msp2007060006.
                                        pdf.

Ortega:2008:NBa

[Ort08a] Brandi Ortega. News briefs. IEEE Security & Privacy, 6(1):

6–8, January/February 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/comp/mags/sp/2008/01/
msp2008010006.pdf.

Ortega:2008:NBb

[Ort08b] Brandi Ortega. News briefs.

IEEE Security & Privacy, 6(2): 10–12, March/April 2008. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/comp/mags/sp/2008/02/
msp2008020010.pdf.

Ortega:2008:NBc

[Ort08c] Brandi Ortega. News briefs.

IEEE Security & Privacy, 6(3): 6–8, May/June 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ortega:2008:NBd

[Ort08d] Brandi Ortega. News briefs.

IEEE Security & Privacy, 6(4): 11–13, July/August 2008. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ortega:2008:NBe

[Ort08e] Brandi Ortega. News briefs.

IEEE Security & Privacy, 6 (5):12–13, September/October 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ortega:2009:NB

[Ort09a] Brandi Ortega. News briefs.

IEEE Security & Privacy, 7(2):

8–9, March/April 2009. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ortega:2009:NBS

[Ort09b] Brandi Ortega. News briefs: Shaking up the cybersecurity landscape. IEEE Security & Privacy, 7(1):5–6, January/ February 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ortlieb:2014:AVP

[Ort14] Martin Ortlieb. The anthropologist’s view on privacy. IEEE Security & Privacy, 12(3):85– 87, May/June 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

OKane:2011:OHM

[OSM11] Philip O’Kane, Sakir Sezer, and Kieran McLaughlin. Obfuscation: The hidden malware. IEEE Security & Privacy, 9(5):41–47, September/ October 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Osterweil:2020:CTU

[Ost20] E. Osterweil. A cybersecurity terminarch: Use it before we lose it. IEEE Security & Privacy, 18(4):67–70, July/ August 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Ortiz-Yepes:2016:RTA

[OY16] Diego A. Ortiz-Yepes. A review of technical approaches to realizing near-field communication mobile payments. IEEE Security & Privacy, 14(4):54– 62, July/August 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/04/
msp2016040054-abs.html.

Osterweil:2009:ICM

[OZ09] Eric Osterweil and Lixia Zhang. Interadministrative challenges in managing DNSKEYs. IEEE Security & Privacy, 7(5):44–51, September/October 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Petroni:2003:DMS

[PA03] Nick L. Petroni, Jr. and William A. Arbaugh. The dangers of mitigating security design flaws: a wireless case study. IEEE Security & Privacy, 1(1):28–36, January/February 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j1028.pdf;   http://www.
computer.org/security/j1028abs.
htm.

Palmer:2004:EBM

[Pal04] Charles C. Palmer. Editorial Board member profile: Can we win the security game? IEEE Security & Privacy, 2 (1):10–12, January/February 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/sp/2004/01/j1010abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2004/01/j1010.
                                        pdf.

Paul:2005:CLV

[Pau05] Nathanael Paul. A closer look at viruses and worms. IEEE Security & Privacy, 3 (4):7, July/August 2005. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/sp/2005/04/
                                        j4007.pdf.

Pauley:2010:CPT

[Pau10a] Wayne Pauley. Cloud provider transparency: An empirical evaluation. IEEE Security & Privacy, 8(6):32–39, November/December 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Paulson:2010:NBa

[Pau10b] Linda Dailey Paulson. News briefs. IEEE Security & Privacy, 8(3):11–13, May/June 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Paulson:2010:NBb

[Pau10c] Linda Dailey Paulson. News briefs. IEEE Security & Privacy, 8(4):12–14, July/August 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Payne:2004:RIS

[Pay04] Jeffery E. Payne. Regulation and information security: Can Y2K lessons help us? IEEE Security & Privacy, 2(2):58–61, March/April 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2004/02/j2058abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/02/j2058.
pdf.

Payer:2019:FHT

[Pay19] Matthias Payer. The fuzzing hype-train: How random testing triggers thousands of crashes. IEEE Security & Privacy, 17(1):78–82, January/ February 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Pincus:2004:BSS

[PB04] Jonathan Pincus and Brandon Baker. Beyond stack smashing: Recent advances in exploiting buffer overruns. IEEE Security & Privacy, 2(4):20–27, July/August 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/04/j4020.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/04/j4020.
pdf.

Pfleeger:2005:CSP

[PB05] S. L. Pfleeger and G. Bloom.

Canning SPAM: Proposed solutions to unwanted email.

IEEE Security & Privacy, 3 (2):40–47, March/April 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/30742/01423959.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        30742&arnumber=1423959&count=
                                        15&index=6.

Peisert:2007:SP

[PB07] Sean Peisert and Matt Bishop. I am a scientist, not a philosopher! IEEE Security & Privacy, 5(4):48–51, July/August 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Pai:2008:TCS

[PBW^{+}08] Sameer Pai, Sergio Bermudez, Stephen B. Wicker, Marci Meingast, Tanya Roosta, Shankar Sastry, and Deirdre K. Mulligan. Transactional confidentiality in sensor networks. IEEE Security & Privacy, 6(4): 28–35, July/August 2008. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Pfleeger:2010:WMS

[PC10] Shari Pfleeger and Robert Cunningham. Why measuring security is hard. IEEE Security & Privacy, 8(4):46–54, July/ August 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Park:2021:PTC

[PCK21] Sangchul Park, Gina J. Choi, and Haksoo Ko. Privacy in the time of COVID-19: Divergent paths for contact tracing and route-disclosure mechanisms in South Korea. IEEE Security & Privacy, 19(3):51–56, May/ June 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Pearson:2008:SIT

[PCMN08] Siani Pearson, Marco Casassa-Mont, and Manny Novoa. Securing information transfer in distributed computing environments. IEEE Security & Privacy, 6(1):34–42, January/ February 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Peeters:2007:CES

[PD07] Johan Peeters and Paul Dyson.

Cost-effective security. IEEE Security & Privacy, 5(3):85–87, May/June 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Parks:2011:PC

[PD11] Raymond C. Parks and David P.

Duggan. Principles of cyberwarfare. IEEE Security & Privacy, 9(5):30–35, September/ October 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Pauli:2012:FYC

[PE12] Josh Pauli and Patrick Engebretson. Filling your cyber operations training toolbox. IEEE Security & Privacy, 10(5):71–74, September/ October 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Pham:2014:BRS

[PEC^{+}14] Cuong Pham, Zachary J. Estrada, Phuong Cao, Zbigniew Kalbarczyk, and Ravishankar K. Iyer. Building reliable and secure Virtual Machines using architectural invariants. IEEE Security & Privacy, 12(5):82–85, September/October 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/05/
msp2014050082-abs.html.

Peel:2013:PCC

[Pee13]  Deborah C. Peel.  Point/

counterpoint: The consequences of the lack of privacy in today’s electronic health systems. IEEE Security & Privacy, 11(6):42–44, November/ December 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Peisert:2019:SED

[Pei19] S. Peisert. Some experiences in developing security technology that actually get used. IEEE Security & Privacy, 17(2):4–7, March/April 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Peisert:2020:IIC

[Pei20] S. Peisert. Isolating insecurely:

A call to arms for the security and privacy community during the time of COVID-19. IEEE Security & Privacy, 18(4):4–7, July/August 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Peisert:2021:RPP

[Pei21] S. Peisert. Reflections on the past, perspectives on the future [from the editors]. IEEE Security & Privacy, 19(1):4–7, January/February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Peisert:2022:UCS

[Pei22]  Sean Peisert.  Unsafe at

any clock speed: The insecurity of computer system design, implementation, and operation. IEEE Security & Privacy, 20(1):4–9, January/ February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Peisert:2023:FYI

[Pei23a] Sean Peisert. The first 20 years of IEEE Security & Privacy. IEEE Security & Privacy, 21(2):4–6, March/April 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Peisert:2023:SID

[Pei23b] Sean Peisert. On software infrastructure: Develop, prove, profit? IEEE Security & Privacy, 21(4):4–8, July/August 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Pearce:2018:TCM

[PEL^{+}18] Paul Pearce, Roya Ensafi, Frank Li, Nick Feamster, and Vern Paxson. Toward continual measurement of global network-level censorship. IEEE Security & Privacy, 16(1):24– 33, January/February 2018. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/01/
msp2018010024.html.

Perrone:2007:CCD

[Per07] Luiz Felipe Perrone. Could a caveman do it? The surprising potential of simple attacks. IEEE Security & Privacy, 5(6):74–77, November/ December 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Peterson:2006:IIM

[Pet06] Gunnar Peterson. Introduction to identity management risk metrics. IEEE Security & Privacy, 4(4):88–91, July/ August 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Petratos:2007:WIS

[Pet07] Pythagoras Petratos. Weather, information security, and markets. IEEE Security & Privacy, 5(6):54–57, November/ December 2007. CODEN ????

ISSN 1540-7993 (print), 1558-4046 (electronic).

Peterson:2009:BSS

[Pet09] Gunnar Peterson. Building security in: Service-oriented security indications for use. IEEE Security & Privacy, 7(2): 91–93, March/April 2009. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Peterson:2010:DTV

[Pet10] Gunnar Peterson. Don’t trust.

and verify: a security architecture stack for the cloud. IEEE Security & Privacy, 8(5):83–86, September/October 2010. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Petrlic:2019:GDP

[Pet19] R. Petrlic. The General Data Protection Regulation: From a data protection authority’s (technical) perspective. IEEE Security & Privacy, 17(6):31– 36, November/December 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Pfleeger:2004:BRG

[Pfl04] Shari Lawrence Pfleeger. Book reviews: a gift of impact: A Gift of Fire: The Social, Legal, and Ethical Issues for Computers and the Internet, by Sara Baase (Prentice-Hall 2002, ISBN 0-13-008215-5). IEEE Security & Privacy, 2(3): 7, May/June 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/sp/2004/03/j3007.
                                        htm.

Pfleeger:2006:EYW

[Pfl06] Shari Lawrence Pfleeger. Everything you wanted to know about privacy (but were afraid to ask). IEEE Security & Privacy, 4(3):5, May/June 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/sp/2006/03/
                                        j3005.pdf.

Pfleeger:2007:SL

[Pfl07] Shari Lawrence Pfleeger. Spooky lessons. IEEE Security & Privacy, 5(5):7, September/ October 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/sp/2007/05/
                                        msp2007050007.pdf.

Pfleeger:2009:BRS

[Pfl09] Shari Lawrence Pfleeger. Book reviews: Searching for you. IEEE Security & Privacy, 7(2): 15, March/April 2009. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Pfleeger:2010:CJD

[Pfl10] Charles P. Pfleeger. Crypto: Not just for the defensive team. IEEE Security & Privacy, 8(2): 63–66, March/April 2010. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Pfleeger:2012:KC

[Pfl12a] Shari Lawrence Pfleeger. A key to the castle. IEEE Security & Privacy, 10(3):3, May/June 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Pfleeger:2012:SMS

[Pfl12b] Shari Lawrence Pfleeger. Security measurement steps, missteps, and next steps. IEEE Security & Privacy, 10(4):5–9, July/August 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Pfleeger:2013:ESS

[Pfl13a] Shari Lawrence Pfleeger. Enlightened security: Shedding light on what works and why. IEEE Security & Privacy, 11(1):3–4, January/February 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Pfleeger:2013:FP

[Pfl13b] Shari Lawrence Pfleeger. Focus on policy. IEEE Security & Privacy, 11(6):3, November/ December 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Pfleeger:2013:OOR

[Pfl13c] Shari Lawrence Pfleeger. In our orbit: Ramsey theory: Learning about the needle in the haystack. IEEE Security & Privacy, 11(3):85–87, May/ June 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Pfleeger:2014:EMR

[Pfl14a] Shari Lawrence Pfleeger. Expanding to meet readers’ needs. IEEE Security & Privacy, 12(5):3–4, September/October 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/05/
msp2014050003.html.

Pfleeger:2014:EIS

[Pfl14b] Shari Lawrence Pfleeger. The eyes have it: Surveillance and how it evolved. IEEE Security & Privacy, 12(4):74– 79, July/August 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/04/
msp2014040074-abs.html.

Pfleeger:2014:TTT

[Pfl14c] Shari Lawrence Pfleeger. Technology, transparency, and trust. IEEE Security & Privacy, 12(6):3–5, November/December 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/06/
msp2014060003.html.

Pfleeger:2015:LLS

[Pfl15a] C. P. Pfleeger. Lesson learned: Security is inevitable. IEEE Security & Privacy, 13(6):22– 28, November/December 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Pfleeger:2015:SMH

[Pfl15b] S. L. Pfleeger. Spider-Man, hubris, and the future of security and privacy. IEEE Security & Privacy, 13(6):3–10, November/December 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Pfleeger:2015:LOD

[Pfl15c] Shari Lawrence Pfleeger. Learning from other disciplines. IEEE Security & Privacy, 13 (4):10–11, July/August 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/04/
msp2015040010.html.

Pfleeger:2015:LLO

[Pfl15d] Shari Lawrence Pfleeger. Lessons learned by our editorial board. IEEE Security & Privacy, 13 (6):16–17, November/December 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Pfleeger:2016:LST

[Pfl16a] Charles P. Pfleeger. Looking into software transparency. IEEE Security & Privacy, 14 (1):31–36, January/February 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Pfleeger:2016:SEG

[Pfl16b] Shari Lawrence Pfleeger. Software everywhere [Guest editors’ introduction]. IEEE Security & Privacy, 14(1):11, January/February 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Pastor-Galindo:2024:LLM

[PGNRV24] Javier Pastor-Galindo, Pantaleone Nespoli, and José A. Ruipérez-Valiente. Largelanguage-model-powered agentbased framework for misinformation and disinformation research: Opportunities and open challenges. IEEE Security & Privacy, 22(3):24–36, May/ June 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Pusey:2016:OCC

[PGP16] Portia Pusey, Mark Gondree, and Zachary Peterson. The outcomes of cybersecurity competitions and implications for underrepresented populations. IEEE Security & Privacy, 14(6):90–95, November/December 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/06/
msp2016060090-abs.html.

Power:2007:SBB

[PGT07] E. Michael Power, Jonathan Gilhen, and Roland L. Trope. Setting boundaries at borders: Reconciling laptop searches and privacy. IEEE Security & Privacy, 5(2):72–75, March/ April 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Provos:2003:HSI

[PH03] Niels Provos and Peter Honeyman. Hide and seek: An introduction to steganography. IEEE Security & Privacy, 1(3): 32–44, May/June 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j3032.pdf;   http://www.
computer.org/security/j3032abs.
htm.

Perrotta:2018:BBU

[PH18] Raffaello Perrotta and Feng Hao. Botnet in the browser: Understanding threats caused by malicious browser extensions. IEEE Security & Privacy, 16(4):66–81, July/ August 2018. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/04/
msp2018040066-abs.html.

Phillips:2019:BYG

[Phi19] A. M. Phillips. Buying your genetic self online: Pitfalls and potential reforms in DNA testing. IEEE Security & Privacy, 17(3):77–81, May/June 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Pencheva:2020:BCS

[PHR20] D. Pencheva, J. Hallett, and A. Rashid. Bringing cyber to school: Integrating cybersecurity into secondary school education. IEEE Security & Privacy, 18(2):68–74, March/ April 2020. ISSN 1558-4046.

Pretschner:2008:UCE

[PHS^{+}08] Alexander Pretschner, Manuel Hilty, Florian Schütz, Christian Schaefer, and Thomas Walter. Usage control enforcement: Present and future. IEEE Security & Privacy, 6(4): 44–53, July/August 2008. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Piessens:2023:TEA

[Pie23] Frank Piessens. Transient execution attacks. IEEE Security & Privacy, 21(5):79– 84, September/October 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Pfleeger:2012:GEI

[PIK12] Shari Lawrence Pfleeger, Cynthia Irvine, and Mischel Kwon. Guest Editors’ introduction. IEEE Security & Privacy, 10 (2):19–23, March/April 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Pallas:2024:PEP

[PKB^{+}24] Frank Pallas, Katharina Koerner, Isabel Barberá, Jaap-Henk Hoepman, Meiko Jensen, Nandita Rao Narla, Nikita Samarin, Max-R. Ulbricht, Isabel Wagner, Kim Wuyts, and Christian Zimmermann. Privacy engineering from principles to practice: a roadmap. IEEE Security & Privacy, 22 (2):86–92, March/April 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Peddinti:2015:USA

[PKBS15] Sai Teja Peddinti, Aleksandra Korolova, Elie Bursztein, and Geetanjali Sampemane. Understanding sensitivity by analyzing anonymity [Guest Editor’s introduction]. IEEE Security & Privacy, 13(2):14– 21, March/April 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/02/
msp2015020014-abs.html.

Phillips:2005:SSR

[PKK05] Ted Phillips, Ttom Karygiannis, and Rick Kuhn. Security standards for the RFID market. IEEE Security & Privacy, 3(6):85–89, November/December 2005. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/33104/01556544.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
33104&arnumber=1556544.

Peacock:2004:TPK

[PKW04] Alen Peacock, Xian Ke, and Matthew Wilkerson. Typing patterns: a key to user identification. IEEE Security & Privacy, 2(5):40–47, September/October 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

                                        http://csdl.computer.org/
                                        comp/mags/sp/2004/extras/
                                        j5040x1.pdf;   http://csdl.
                                        computer.org/comp/mags/sp/
                                        2004/extras/j5040x2.pdf;
                                         http://csdl.computer.org/
                                        dl/mags/sp/2004/05/j5040.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2004/05/j5040.
                                        pdf.

Park:2023:MPK

[PLK23] Soyeon Park, Sangho Lee, and Taesoo Kim. Memory protection keys: Facts, key extension perspectives, and discussions. IEEE Security & Privacy, 21(3):8–15, May/June 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Pfleeger:2007:IBC

[PLW07] Shari Lawrence Pfleeger, Martin Libicki, and Michael Webber. I’ll buy that! cybersecurity in the Internet marketplace. IEEE Security & Privacy, 5(3):25–31, May/June 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Peisert:2014:CGS

[PM14] Sean Peisert and Jonathan Margulies. Closing the gap on securing energy sector control systems [Guest Editors’ introduction]. IEEE Security & Privacy, 12(6):13–14, November/December 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/06/
msp2014060013.html.

Pujol:2021:EPM

[PM21] David Pujol and Ashwin Machanavajjhala. Equity and privacy: More than just a tradeoff. IEEE Security & Privacy, 19(6):93–97, November/ December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Padilha:2020:FEA

[PMA^{+}20] R. Padilha, C. Mazini Rodrigues, F. A. Andalo, G. Bertocco, Z. Dias, and A. Rocha. Forensic event analysis: From seemingly unrelated data to understanding. IEEE Security & Privacy, 18(6):23–32, November/December 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Peisert:2014:CSS

[PMB^{+}14] Sean Peisert, Jonathan Margulies, Eric Byres, Paul Dorey, Dale Peterson, and Zach Tudor. Control systems security from the front lines. IEEE Security & Privacy, 12(6):55– 58, November/December 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.

                                        org/csdl/mags/sp/2014/06/
                                        msp2014060055-abs.html.

Peisert:2014:DSC

[PMN^{+}14] Sean Peisert, Jonathan Margulies, David M. Nicol, Himanshu Khurana, and Chris Sawall. Designed-in security for cyberphysical systems. IEEE Security & Privacy, 12(5):9–12, September/October 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/sp/2014/05/
                                        msp2014050009.html.

Paulsen:2012:NCC

[PMNT12] Celia Paulsen, Ernest Mc-Duffie, William Newhouse, and Patricia Toth. NICE: Creating a cybersecurity workforce and aware public. IEEE Security & Privacy, 10(3):76–79, May/June 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Popli:2004:ACC

[Pop04] Ashish Popli. ACM Computer and Communication Security Conference. IEEE Security & Privacy, 2(2):46–47, March/ April 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/sp/2004/02/j2046abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2004/02/j2046.
                                        pdf.

Porras:2006:PEG

[Por06] Phillip A. Porras. Privacyenabled global threat monitoring. IEEE Security & Privacy, 4(6):60–63, November/ December 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Porras:2009:ATD

[Por09] Phillip Porras. Attack trends:

Directions in network-based security monitoring. IEEE Security & Privacy, 7(1):82–85, January/February 2009. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Potter:2009:HTT

[Pot09]  Bruce Potter.  High time

for trusted computing. IEEE Security & Privacy, 7(6):54– 56, November/December 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Potter:2010:MIY

[Pot10a] Bruce Potter. My information, your code. IEEE Security & Privacy, 8(2):51–53, March/April 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Potter:2010:NS

[Pot10b] Bruce Potter. Necessary but not sufficient. IEEE Security & Privacy, 8(5):57–58, September/October 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Potter:2010:TO

[Pot10c] Bruce Potter. Thinking operationally. IEEE Security & Privacy, 8(3):54–55, May/June 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Power:2007:DCP

[Pow07] E. Michael Power. Developing a culture of privacy: a case study. IEEE Security & Privacy, 5(6):58–60, November/ December 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Pfleeger:2006:WWW

[PP06a] Charles P. Pfleeger and Shari Lawrence Pfleeger. Why we won’t review books by hackers. IEEE Security & Privacy, 4(4):9, July/ August 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2006/04/j4009.pdf.

Popp:2006:CTT

[PP06b] Robert Popp and John Poindexter. Countering terrorism through information and privacy protection technologies. IEEE Security & Privacy, 4(6): 18–27, November/December 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Padilha:2015:CC

[PP15] Ricardo Padilha and Fernando Pedone. Confidentiality in the cloud. IEEE Security & Privacy, 13(1):57–60, January/February 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/01/
msp2015010057-abs.html.

Predd:2008:IBB

[PPHB08] Joel Predd, Shari Lawrence Pfleeger, Jeffrey Hunker, and Carla Bulford. Insiders behaving badly. IEEE Security & Privacy, 6(4):66–70, July/ August 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Prabhakar:2003:BRS

[PPJ03] Salil Prabhakar, Sharath Pankanti, and Anil K. Jain. Biometric recognition: Security and privacy concerns. IEEE Security & Privacy, 1(2):33– 42, March/April 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j2033.pdf;   http://www.
computer.org/security/j2033abs.
htm.

Parks:2008:VAC

[PR08] Raymond C. Parks and Edmond Rogers. Vulnerability assessment for critical infrastructure control systems. IEEE Security & Privacy, 6(6):37– 43, November/December 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Pettigrew:2012:MSS

[PR12a] James A. Pettigrew III and Julie J. C. H. Ryan. Making successful security decisions: a qualitative evaluation. IEEE Security & Privacy, 10(1):60– 68, January/February 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Prandini:2012:ROP

[PR12b] Marco Prandini and Marco Ramilli. Return-oriented programming. IEEE Security & Privacy, 10(6):84–87, November/December 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Pfleeger:2012:DPM

[PRB^{+}12] Shari Lawrence Pfleeger, Marc Rogers, Masooda Bashir, Kelly Caine, Deanna Caputo, Michael Losavio, and Sal Stolfo. Does profiling make us more secure? IEEE Security & Privacy, 10(4):10–15, July/ August 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Peddinti:2017:UAT

[PRC17] Sai Teja Peddinti, Keith W. Ross, and Justin Cappos. User anonymity on Twitter. IEEE Security & Privacy, 15(3):84– 87, May/June 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/03/
                                        msp2017030084-abs.html.

Prandini:2010:SHS

[PRCC10] Marco Prandini, Marco Ramilli, Walter Cerroni, and Franco Callegati. Splitting the HTTPS stream to attack secure Web connections. IEEE Security & Privacy, 8(6):80–84, November/December 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Preibusch:2015:VWS

[Pre15] Soren Preibusch. The value of Web search privacy. IEEE Security & Privacy, 13(5):24–32, September/October 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/csdl/mags/sp/2015/05/
msp2015050024-abs.html.

Portnoy:2009:WWC

[PRS09] Aaron Portnoy and Ali Rizvi-Santiago. Walking on water: a cheating case study. IEEE Security & Privacy, 7(3):20–22, May/June 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Peterson:2006:DMW

[PS06] Gunnar Peterson and John Steven. Defining misuse within the development process. IEEE Security & Privacy, 4(6):81– 84, November/December 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Pfleeger:2009:AIT

[PS09] Shari Lawrence Pfleeger and Salvatore J. Stolfo. Addressing the insider threat. IEEE Security & Privacy, 7(6):10– 13, November/December 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Polonetsky:2021:RPE

[PS21] Jules Polonetsky and Tim Sparapani. A review of the privacy-enhancing technologies software market. IEEE Security & Privacy, 19(6): 119–122, November/December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Park:2023:FDI

[PS23] Sungmi Park and Douglas Stebila. The future of digital investigation: Automated legal inference. IEEE Security & Privacy, 21(1):84–90, January/ February 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Piazzalunga:2007:SSM

[PSB^{+}07] Ugo Piazzalunga, Paolo Salvaneschi, Francesco Balducci, Pablo Jacomuzzi, and Cristiano Moroncelli. Security strength measurement for dongle-protected software. IEEE Security & Privacy, 5(6): 32–40, November/December 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Peisert:2021:PSI

[PSO^{+}21] Sean Peisert, Bruce Schneier, Hamed Okhravi, Fabio Massacci, Terry Benzel, Carl Landwehr, Mohammad Mannan, Jelena Mirkovic, Atul Prakash, and James Bret Michael. Perspectives on the SolarWinds incident. IEEE Security & Privacy, 19(2):7–13, March/April 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Power:2005:ASM

[PT05a] E. M. Power and R. L. Trope.

Averting security missteps in outsourcing. IEEE Security & Privacy, 3(2):70–73, March/ April 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/30742/01423966.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
30742&arnumber=1423966&count=
15&index=12.

Power:2005:ARG

[PT05b] E. Michael Power and Roland L.

Trope. Acting responsibly with geospatial data. IEEE Security & Privacy, 3(6): 77–80, November/December 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/33104/01556542.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
33104&arnumber=1556542.

Pfleeger:2007:GEI

[PTP07] Shari Lawrence Pfleeger, Roland L.

Trope, and Charles C. Palmer. Guest Editors’ introduction:

Managing organizational security. IEEE Security & Privacy, 5(3):13–15, May/June 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Puzanov:2016:AMN

[Puz16]  Anton Puzanov.  Analysis

and mitigation of NoSQL injections. IEEE Security & Privacy, 14(2):30–39, March/ April 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Perez:2008:VHB

[PvDS08] Ronald Perez, Leendert van Doorn, and Reiner Sailer. Virtualization and hardwarebased security. IEEE Security & Privacy, 6(5):24–31, September/October 2008. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Piessens:2024:SCA

[PvO24] Frank Piessens and Paul C. van Oorschot. Side-channel attacks: a short tour. IEEE Security & Privacy, 22(2):75–80, March/April 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Peisert:2017:OSC

[PW17] Sean Peisert and Von Welch. The open science cyber risk profile: The Rosetta Stone for open science and cybersecurity. IEEE Security & Privacy, 15(5):94–95, September/October 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/05/
msp2017050094-abs.html.

Poller:2012:EIC

[PWVT12] Andreas Poller, Ulrich Waldmann, Sven Vowe, and Sven Turpe. Electronic identity cards for user authentication — promise and practice. IEEE Security & Privacy, 10(1):46– 54, January/February 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Paterson:2006:LTT

[PY06] Kenneth G. Paterson and Arnold K. L. Yau. Lost in translation: Theory and practice in cryptography. IEEE Security & Privacy, 4(3):69–72, May/June 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Price:2018:HRD

[PZTE18] Benjamin Price, Michael Zhivich, Michael Thompson, and Chris Eagle. House rules: Designing the scoring algorithm for cyber grand challenge. IEEE Security & Privacy, 16(2):23– 31, March/April 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/02/
msp2018020023.html.

Quarterman:2003:UIG

[Qua03] John S. Quarterman. The ultimate in instant gratification. IEEE Security & Privacy, 1(6):56–58, November/ December 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/sp/2003/06/j6056abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2003/06/j6056.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2003/06/j6056.
                                        pdf.

Reeder:2005:S

[RA05] Robert W. Reeder and Fahd Arshad. Soups 2005. IEEE Security & Privacy, 3(5):47–50, September/October 2005. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Raaijmakers:2019:AIL

[Raa19] S. Raaijmakers. Artificial intelligence for law enforcement: Challenges and opportunities. IEEE Security & Privacy, 17 (5):74–77, September/October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Rodriguez:2020:CCI

[RAAB20] J. M. Rodriguez, B. J. Allison, C. W. Apsey, and T. M. Boudreau. Courseware as code: Instituting agile courseware collaboration. IEEE Security & Privacy, 18(6):59– 62, November/December 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Roman:2023:CPS

[RALS23] Rodrigo Roman, Cristina Alcaraz, Javier Lopez, and Kouichi Sakurai. Current perspectives on securing critical infrastructures supply chains. IEEE Security & Privacy, 21 (4):29–38, July/August 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Ranum:2016:SBT

[Ran16] Marcus Ranum. Silver Bullet talks with Gary Mc-Graw. IEEE Security & Privacy, 14(5):7–10, September/October 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/05/
msp2016050007.html.

Raskin:2011:YLE

[Ras11] Aza Raskin. Your life experiences, brought to you by Budweiser. IEEE Security & Privacy, 9(2):86–88, March/April 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ricci:2019:BBD

[RBB19] J. Ricci, I. Baggili, and F. Breitinger. Blockchain-based distributed cloud storage digital forensics: Where’s the beef? IEEE Security & Privacy, 17(1):34–42, January/ February 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Raynal:2004:HFPa

[RBBK04a] Frederic Raynal, Yann Berthier, Philippe Biondi, and Danielle Kaminsky. Honeypot forensics Part I: Analyzing the network. IEEE Security & Privacy, 2(4):72–78, July/ August 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/04/j4072.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/04/j4072.
pdf.

Raynal:2004:HFPb

[RBBK04b] Frederic Raynal, Yann Berthier, Philippe Biondi, and Danielle Kaminsky. Honeypot forensics, Part II: Analyzing the compromised host. IEEE Security & Privacy, 2(5):77–80, September/October 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/05/j5077.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/05/j5077.
pdf.

Reardon:2014:SDD

[RBC14] Joel Reardon, David Basin, and Srdjan Capkun. On secure data deletion. IEEE Security & Privacy, 12(3):37–44, May/June 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Rezgui:2003:PWF

[RBE03] Abdelmounaam Rezgui, Athman Bouguettaya, and Mohamed Y. Eltoweissy. Privacy on the Web: Facts, challenges, and solutions. IEEE Security & Privacy, 1(6): 40–49, November/December 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2003/06/j6040abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/06/j6040.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/06/j6040.
pdf.

Redmiles:2023:PCS

[RBK23] Elissa M. Redmiles, Mia M.

Bennett, and Tadayoshi Kohno. Power in computer security and privacy: a critical lens. IEEE Security & Privacy, 21 (2):48–52, March/April 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Ring:2004:TLS

[RC04] Sandra Ring and Eric Cole.

Taking a lesson from stealthy rootkits. IEEE Security & Privacy, 2(4):38–45, July/ August 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/04/j4038.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/04/j4038.
pdf.

Rubin:2006:CSE

[RC06] Bradley S. Rubin and Donald Cheung. Computer security education and research: Handle with care. IEEE Security & Privacy, 4(6):56–59, November/December 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ranganathan:2017:WRC

[RC17] Aanjhan Ranganathan and Srdjan Capkun. Are we really close? Verifying proximity in wireless systems. IEEE Security & Privacy, 15(3):52– 58, May/June 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/03/
                                        msp2017030052-abs.html.

Reinicke:2017:RDS

[RCK17] Bryan Reinicke, Jeffrey Cummings, and Howard Kleinberg. The right to digital self-defense. IEEE Security & Privacy, 15(4):68– 71, July/August 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/04/
                                        msp2017040068-abs.html.

Rieback:2006:RMT

[RCT06] Melanie R. Rieback, Bruno Crispo, and Andrew S. Tanenbaum. RFID malware: Truth vs. myth. IEEE Security & Privacy, 4(4):70–72, July/August 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Raykova:2012:USP

[RCV^{+}12] Mariana Raykova, Ang Cui, Binh Vo, Bin Liu, Tal Malkin, Steven M. Bellovin, and Salvatore J. Stolfo. Usable, secure, private search. IEEE Security & Privacy, 10(5):53–60, September/October 2012. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Rashid:2018:SCS

[RDC^{+}18] Awais Rashid, George Danezis, Howard Chivers, Emil Lupu, Andrew Martin, Makayla Lewis, and Claudia Peersman. Scoping the cyber security body of knowledge. IEEE Security & Privacy, 16(3):96– 102, May/June 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/03/
msp2018030096-abs.html.

Riazi:2019:DLP

[RDK19] M. S. Riazi, B. Darvish Rouani, and F. Koushanfar. Deep learning on private data. IEEE Security & Privacy, 17(6):54– 63, November/December 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Ruefle:2014:CSI

[RDM^{+}14] Robin Ruefle, Audrey Dorofee, David Mundie, Allen D.

Householder, Michael Murray, and Samuel J. Perl. Computer Security Incident Response Team development and evolution. IEEE Security & Privacy, 12(5):16–26, September/October 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/05/
msp2014050016-abs.html.

Ribalda:2010:MBS

[RdRdCG10] Ricardo Ribalda, Guillermo Gonzalez de Rivera, Angel de Castro, and Javier Garrido. A mobile biometric system-on-token system for signing digital transactions. IEEE Security & Privacy, 8(2):13–19, March/April 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Redmiles:2024:FMB

[Red24] Elissa M. Redmiles. Friction matters: Balancing the pursuit of perfect protection with target hardening. IEEE Security & Privacy, 22(1):76–75, January/February 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Reed:2016:CJT

[Ree16] Jeffrey H. Reed. A communications jamming taxonomy. IEEE Security & Privacy, 14(1):47–54, January/ February 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Renaud:2012:BNT

[Ren12] Karen Renaud. Blaming noncompliance is too convenient: What really causes information breaches? IEEE Security & Privacy, 10(3):57–63, May/June 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Rescorla:2005:FSH

[Res05] E. Rescorla. Is finding security holes a good idea? IEEE Security & Privacy, 3 (1):14–19, January/February 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/30310/01392694.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
30310&arnumber=1392694&count=
17&index=3.

Ressler:2016:MFV

[Res16] Eugene K. Ressler. Mettle fatigue: VW’s single-point-of-failure ethics. IEEE Security & Privacy, 14(1):12–30, January/ February 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Rosenberg:2006:TNR

[RH06] Timothy Rosenberg and Lance J.

Hoffman. Taking networks on the road: Portable solutions for security educators. IEEE Security & Privacy, 4 (1):57–60, January/February 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/33481/01588828.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        33481&arnumber=1588828.

Rohlf:2012:SCC

[RI12] Chris Rohlf and Yan Ivnitskiy. The security challenges of client-side just-in-time engines. IEEE Security & Privacy, 10 (2):84–86, March/April 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Reeder:2017:SSS

[RIC17] Robert W. Reeder, Iulia Ion, and Sunny Consolvo. 152 simple steps to stay safe online: Security advice for nontech-savvy users. IEEE Security & Privacy, 15(5):55–64, September/October 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/05/
                                        msp2017050055.html.

Rubin:2008:NRR

[RJ08] Aviel D. Rubin and David R. Jefferson. New research results for electronic voting. IEEE Security & Privacy, 6(3):12– 13, May/June 2008. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/sp/2008/03/
                                        msp2008030012.pdf.

Roesner:2024:SPM

[RK24] Franziska Roesner and Tadayoshi Kohno. Security and privacy in the metaverse. IEEE Security & Privacy, 22(1):7–9, January/February 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Ryoo:2015:AAS

[RKA15] Jungwoo Ryoo, R. Kazman, and P. Anand. Architectural analysis for security. IEEE Security & Privacy, 13(6):52– 59, November/December 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ruzomberka:2023:COB

[RLB^{+}23] Eric Ruzomberka, David J.

Love, Christopher G. Brinton, Arpit Gupta, Chih-Chun Wang, and H. Vincent Poor. Challenges and opportunities for Beyond-5G wireless security. IEEE Security & Privacy, 21(5):55–66, September/ October 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Razaq:2008:SGR

[RLS^{+}08] Abdul (Ali) Razaq, Wai Tong Luk, Kam Man Shum, Lee Ming Cheng, and Kai Ning Yung. Second-generation RFID. IEEE Security & Privacy, 6(4):21–27, July/August 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Rusman:2022:MEC

[RM22] Galina S. Rusman and Yulia A. Morozova. Measures to ensure cybersecurity of industrial enterprises: a legal perspective. IEEE Security & Privacy, 20 (4):23–28, July/August 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Roesner:2017:DAP

[Roe17] Franziska Roesner. Designing application permission models that meet user expectations. IEEE Security & Privacy, 15(1):75–79, January/February 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/01/
msp2017010075-abs.html.

Rogaway:2016:POP

[Rog16] Phillip Rogaway. Practiceoriented provable security and the social construction of cryptography. IEEE Security & Privacy, 14(6):10–17, November/December 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/06/
msp2016060010-abs.html.

Rios:2022:PIP

[RORL22] Ruben Rios, Jose A. Onieva, Rodrigo Roman, and Javier Lopez. Personal IoT privacy control at the edge. IEEE Security & Privacy, 20(1):23–32, January/February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Rosenblum:2007:WAC

[Ros07] David Rosenblum. What anyone can know: The privacy risks of social networking sites. IEEE Security & Privacy, 5(3): 40–49, May/June 2007. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Rostad:2008:LFF

[RØTM08] Lillian Røstad, Gunnar René Øie, Inger Anne Tøndel, and Per Håkon Meland. Learning by failing (and fixing). IEEE Security & Privacy, 6(4):54–56, July/August 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Roussev:2009:HDF

[Rou09] Vassil Roussev. Hashing and data fingerprinting in digital forensics. IEEE Security & Privacy, 7(2):49–55, March/April 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Rue:2009:CMB

[RP09] Rachel Rue and Shari Lawrence Pfleeger. Cybersecurity: Making the best use of cybersecurity economic models. IEEE Security & Privacy, 7(4):52–60, July/August 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ramilli:2010:ASN

[RP10] Marco Ramilli and Marco Prandini. Always the same, never the same. IEEE Security & Privacy, 8(2):73–75, March/ April 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Randell:2006:VTT

[RR06] Brian Randell and Peter Y. A. Ryan. Voting technologies and trust. IEEE Security & Privacy, 4(5):50–56, September/ October 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ryan:2008:RMP

[RR08] Julie J. C. H. Ryan and Daniel J. Ryan. Risk management: Performance metrics for information security risk management. IEEE Security & Privacy, 6(5):38–44, September/ October 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ryoo:2014:CSA

[RRAK14] Jungwoo Ryoo, Syed Rizvi, William Aiken, and John Kissell. Cloud security auditing: Challenges and emerging approaches. IEEE Security & Privacy, 12(6):68– 74, November/December 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/06/
msp2014060068-abs.html.

Rechberger:2006:NCW

[RRS06] Christian Rechberger, Vincent Rijmen, and Nicolas Sklavos. The NIST Cryptographic Workshop on Hash Functions. IEEE Security & Privacy, 4 (1):54–56, January/February 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/33481/01588827.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
33481&arnumber=1588827;
http://www.csrc.nist.gov/
pki/HashWorkshop/.

Robbins:2006:MIP

[RS06] Jim Robbins and John T. Sabo.

Managing information privacy: Developing a context for security and privacy standards convergence. IEEE Security & Privacy, 4(4):92–95, July/ August 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Reeder:2011:WPD

[RS11] Robert W. Reeder and Stuart Schechter. When the password doesn’t work: Secondary authentication for Websites. IEEE Security & Privacy, 9(2):43–49, March/April 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Rivest:2017:WEV

[RS17] Ronald L. Rivest and Philip B.

Stark. When is an election verifiable? IEEE Security & Privacy, 15(3):48– 50, May/June 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/03/
                                        msp2017030048-abs.html.

Roetteler:2018:QCC

[RS18] M. Roetteler and K. M. Svore. Quantum computing: Codebreaking and beyond. IEEE Security & Privacy, 16(5):22– 36, September/October 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Ruoti:2019:JJT

[RS19] S. Ruoti and K. Seamons. Johnny’s journey toward usable secure email. IEEE Security & Privacy, 17(6):72– 76, November/December 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Rosa:2013:MXI

[RSM13] Thiago Mattos Rosa, Altair Olivo Santin, and Andreia Malucelli. Mitigating XML injection 0-day attacks through strategy-based detection systems. IEEE Security & Privacy, 11(4):46–53, July/ August 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Raponi:2022:RTP

[RSOD22] Simone Raponi, Savio Sciancalepore, Gabriele Oligeri, and Roberto Di Pietro. Road traffic poisoning of navigation apps: Threats and countermeasures. IEEE Security & Privacy, 20(3):71–79, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Ryan:2015:EEV

[RST15] Peter Y. A. Ryan, Steve Schneider, and Vanessa Teague. End-to-end verifiability in voting systems, from theory to practice. IEEE Security & Privacy, 13(3):59– 62, May/June 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/03/
msp2015030059-abs.html.

Ronen:2018:IGN

[RSWO18] Eyal Ronen, Adi Shamir, Achi-Or Weingarten, and Colin O’Flynn. IoT goes nuclear: Creating a Zigbee chain reaction. IEEE Security & Privacy, 16(1):54–62, January/February 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/01/
msp2018010054.html.

Ryoo:2009:ESE

[RTL09] Jungwoo Ryoo, Angsana Techatassanasoontorn, and Dongwon Lee. Education: Security education using second life. IEEE Security & Privacy, 7(2):71–74, March/April 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Ruiu:2006:LIS

[Rui06] Dragos Ruiu. Learning from information security history. IEEE Security & Privacy, 4 (1):77–79, January/February 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/33481/01588833.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        33481&arnumber=1588833.

Ramon:2024:DDS

[RVG24] Meike Ramon, Matthew Vowels, and Matthew Groh. Deepfake detection in superrecognizers and police officers. IEEE Security & Privacy, 22(3):68–76, May/June 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Ray:2005:TAA

[RVK05] Helayne T. Ray, Raghunath Vemuri, and Hariprasad R. Kantubhukta. Toward an automated attack model for Red teams. IEEE Security & Privacy, 3(4):18–25, July/August 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Rahman:2021:DKS

[RW21] Akond Rahman and Laurie Williams. Different kind of smells: Security smells in infrastructure as code scripts.

IEEE Security & Privacy, 19 (3):33–41, May/June 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Ragsdale:2003:IAW

[RWD03] Dan Ragsdale, Don Welch, and Ron Dodge. Information assurance the West Point way. IEEE Security & Privacy, 1(5):64–67, September/ October 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2003/05/j5064abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/05/j5064.
pdf.

Ruan:2021:PCC

[RXJ^{+}21] Wenqiang Ruan, Mingxin Xu, Haoyang Jia, Zhenhuan Wu, LuShan Song, and Weili Han. Privacy compliance: Can technology come to the rescue? IEEE Security & Privacy, 19 (4):37–43, July/August 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Ryan:2003:TVS

[Rya03] Daniel J. Ryan. Two views on security software liability: Let the legal system decide. IEEE Security & Privacy, 1(1):70–72, January/February 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j1070.pdf;   http://www.

                                        computer.org/security/j1070abs.
                                        htm.

Ryan:2007:PGE

[Rya07] Julie J. C. H. Ryan. Plagiarism, graduate education, and information security. IEEE Security & Privacy, 5(5):62–65, September/October 2007. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Shrobe:2012:SWG

[SA12] Howard Shrobe and Daniel Adams. Suppose we got a doover: a revolution for secure computing. IEEE Security & Privacy, 10(6):36–39, November/December 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Stone:2004:NDB

[SAA04] Adam Stone, Benjamin Alfonsi, and Scott L. Andresen. News: The delicate balance: Security and privacy; corporate security under siege; news briefs. IEEE Security & Privacy, 2(4):12–13, July/August 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/sp/2004/04/j4012.
                                        htm.

Sambasivan:2019:TGE

[SAB^{+}19] N. Sambasivan, N. Ahmed, A. Batool, E. Bursztein, E. Churchill, L. Sanely Gaytan-Lugo, T. Matthews, D. Nemar, K. Thomas, and S. Consolvo. Toward gender-equitable privacy and security in South Asia. IEEE Security & Privacy, 17(4):71–77, July/ August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Sadeghi:2016:GFW

[Sad16] Ahmad-Reza Sadeghi. Games without frontiers: Whither information security and privacy? IEEE Security & Privacy, 14(1):3–5, January/ February 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Sadeghi:2017:SPM

[Sad17] Ahmad-Reza Sadeghi. Security and privacy more crucial than ever. IEEE Security & Privacy, 15(1):3–4, January/February 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/01/
msp2017010003.html.

Song:2015:DCG

[SAF15] Jia Song and J. Alves-Foss.

The DARPA Cyber Grand Challenge: A competitor’s perspective. IEEE Security & Privacy, 13(6):72–76, November/ December 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Sagarin:2013:BHT

[Sag13] Rafe Sagarin. Bio-hacking:

Tapping life’s code to deal with unpredictable risk. IEEE Security & Privacy, 11(4):93–95, July/August 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Sahinoglu:2005:SMP

[Sah05] M. Sahinoglu. Security meter: a practical decision-tree model to quantify risk. IEEE Security & Privacy, 3(3):18–24, May/ June 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/31002/01439498.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
31002&arnumber=1439498&count=
20&index=5.

Salka:2005:PLS

[Sal05] C. Salka. Programming languages and systems security. IEEE Security & Privacy, 3(3):80–83, May/June 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/31002/01439509.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
31002&arnumber=1439509&count=
20&index=16.

Sanders:2014:QSM

[San14] William H. Sanders. Quantitative security metrics: Unattainable Holy Grail or a vital breakthrough within our reach? IEEE Security & Privacy, 12(2):67–69, March/ April 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Sanderson:2021:BPH

[San21] Pollyanna Sanderson. Balancing public health and civil liberties: Privacy aspects of contact-tracing technologies. IEEE Security & Privacy, 19 (4):65–69, July/August 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Sasse:2007:REB

[Sas07] M. Angela Sasse. Red-eye blink, bendy shuffle, and the yuck factor: a user experience of biometric airport systems. IEEE Security & Privacy, 5(3): 78–81, May/June 2007. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Sasse:2015:SBP

[Sas15] Angela Sasse. Scaring and bullying people into security won’t work. IEEE Security & Privacy, 13(3):80– 83, May/June 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/03/
msp2015030080-abs.html.

Sasaki:2018:QKD

[Sas18] M. Sasaki. Quantum key distribution and its applications. IEEE Security & Privacy, 16 (5):42–48, September/October 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Savage:2017:CRH

[Sav17] Sara Renee Savage. Characterizing the risks and harms of linking genomic information to individuals. IEEE Security & Privacy, 15(5):14–19, September/October 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/05/
msp2017050014-abs.html.

Saydjari:2004:MSR

[Say04] O. Sami Saydjari. Multilevel security: Reprise. IEEE Security & Privacy, 2(5):64– 67, September/October 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/sp/
2004/05/j5064.htm;   http:
//csdl.computer.org/dl/mags/
sp/2004/05/j5064.pdf.

Saydjari:2008:SLC

[Say08]  O. Sami Saydjari.  Spot-

light: Launching into the cyberspace race: An interview with Melissa E. Hathaway. IEEE Security & Privacy, 6(6):11–17, November/ December 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Stytz:2006:DSS

[SB06] Martin R. Stytz and Sheila B. Banks. Dynamic software security testing. IEEE Security & Privacy, 4(3):77–79, May/June 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Singla:2019:HDL

[SB19] A. Singla and E. Bertino. How deep learning is making information security more intelligent. IEEE Security & Privacy, 17(3):56–65, May/June 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Stalla-Bourdillon:2021:MSD

[SB21] Sophie Stalla-Bourdillon. A maturity spectrum for data institutions. IEEE Security & Privacy, 19(5):90–94, September/October 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Silva:2018:EBC

[SBA^{+}18] P. Silva, T. Basso, N. Antunes, R. Moraes, M. Vieira, P. Simoes, and E. Montiero. A Europe–Brazil context for secure data analytics in the cloud. IEEE Security & Privacy, 16(6):52–60, November/ December 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Shoshitaishvili:2018:MPR

[SBB^{+}18] Yan Shoshitaishvili, Antonio Bianchi, Kevin Borgolte, Amat Cama, Jacopo Corbetta, Francesco Disperati, Audrey Dutcher, John Grosen, Paul Grosen, Aravind Machiry, Chris Salls, Nick Stephens, Ruoyu Wang, and Giovanni Vigna. Mechanical phish: Resilient autonomous hacking. IEEE Security & Privacy, 16 (2):12–22, March/April 2018. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/02/
msp2018020012-abs.html.

Stolfo:2011:MS

[SBE11] Sal Stolfo, Steven M. Bellovin, and David Evans. Measuring security. IEEE Security & Privacy, 9(3):60–65, May/June 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Steinke:2015:ICI

[SBF^{+}15] Julie Steinke, Balca Bolunmez, Laura Fletcher, Vicki Wang, Alan J. Tomassetti, Kristin M. Repchick, Stephen J. Zaccaro, Reeshad S. Dalal, and Lois E. Tetrick. Improving cybersecurity incident response team effectiveness using teamsbased research. IEEE Security & Privacy, 13(4):20– 29, July/August 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/04/
msp2015040020-abs.html.

Singh:2013:TTP

[SBM^{+}13] Indrajeet Singh, Michael Butkiewicz, Harsha V. Madhyastha, Srikanth V. Krishnamurthy, and Sateesh Addepalli. Twitsper: Tweeting privately. IEEE Security & Privacy, 11(3):46–50, May/ June 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Spencer:2004:LEI

[SC04] Brad Spencer and Michael A.

Caloyannides. Letters to the editor: Interface: Machine gun or blunderbuss?; usability or privacy redux; erratum. IEEE Security & Privacy, 2(4):7–8, July/August 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/04/j4007.
htm; j4007.pdf.

Sobiesk:2007:CFW

[SC07] Edward Sobiesk and Gregory Conti. The cost of free Web tools. IEEE Security & Privacy, 5(3):66–68, May/June 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Singh:2019:SID

[SC19] J. Singh and J. Cobbe. The security implications of data subject rights. IEEE Security & Privacy, 17(6):21–30, November/December 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Sherr:2005:SVW

[SCCB05] Michah Sherr, Eric Cronin, Sandy Clark, and Matt Blaze. Signaling vulnerabilities in wiretapping systems. IEEE Security & Privacy, 3(6): 13–25, November/December 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/33104/01556532.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
33104&arnumber=1556532.

Schneider:2003:LPM

[Sch03a]  Fred B. Schneider.  Least

privilege and more. IEEE Security & Privacy, 1(5): 55–59, September/October 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2003/05/j5055abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/05/j5055.
pdf.

Schneier:2003:AH

[Sch03b]  Bruce Schneier.  Airplane

hackers. IEEE Security & Privacy, 1(6):92, November/ December 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2003/06/j6092abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/06/j6092.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/06/j6092.
pdf.

Schneier:2003:GUP

[Sch03c] Bruce Schneier. Guilty until proven innocent? IEEE Security & Privacy, 1(3):88, 87, May/June 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j3088.pdf.

Schneier:2003:LFD

[Sch03d] Bruce Schneier. Locks and full disclosure. IEEE Security & Privacy, 1(2):88, March/April 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j2088.pdf;   http://www.
computer.org/security/j2088abs.
htm.

Schneier:2003:WAS

[Sch03e] Bruce Schneier. We are all security consumers. IEEE Security & Privacy, 1(1):104, January/February 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j1104.pdf;   http://www.
computer.org/security/j1104abs.
htm.

Schneider:2004:TSI

[Sch04a] Fred Schneider. Time out for station identification. IEEE Security & Privacy, 2(5):5, September/October 2004. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/comp/mags/sp/2004/05/
j5005.pdf;   http://csdl.
computer.org/dl/mags/sp/
2004/05/j5005.htm.

Schneider:2004:END

[Sch04b]  Fred B. Schneider.  From

the Editors: The next digital divide. IEEE Security & Privacy, 2(1):5, January/ February 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/sp/2004/01/j1005.pdf.

Schneier:2004:CTV

[Sch04c] Bruce Schneier. Clear text: Voting security and technology. IEEE Security & Privacy, 2(1):84, January/ February 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/sp/2004/01/j1084abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2004/01/j1084.
                                        pdf.

Schneier:2004:CPW

[Sch04d] Bruce Schneier. Customers, passwords, and Web sites. IEEE Security & Privacy, 2(4): 88, July/August 2004. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/sp/
                                        2004/04/j4088.htm;   http:
                                        //csdl.computer.org/dl/mags/
                                        sp/2004/04/j4088.pdf.

Schneier:2004:SC

[Sch04e]  Bruce Schneier.  Security

and compliance. IEEE Security & Privacy, 2(3):96, May/June 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/03/j3096.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/03/j3096.
pdf.

Schneier:2004:SSP

[Sch04f] Bruce Schneier. SIMS: Solution, or part of the problem? IEEE Security & Privacy, 2(5):88, September/October 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/05/j5088.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/05/j5088.
pdf.

Schechter:2005:TEM

[Sch05a] S. E. Schechter. Toward econometric models of the security risk from remote attacks. IEEE Security & Privacy, 3 (1):40–44, January/February 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/30310/01392698.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
30310&arnumber=1392698&count=
17&index=7.

Schneider:2005:IDW

[Sch05b] F. B. Schneider. It depends on what you pay. IEEE Security & Privacy, 3(3):3, May/June 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/31002/01439491.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        31002&arnumber=1439491&count=
                                        20&index=0.

Schneier:2005:AE

[Sch05c] B. Schneier. Authentication and expiration. IEEE Security & Privacy, 3(1):88, January/February 2005. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/30310/01392710.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        30310&arnumber=1392710&count=
                                        17&index=16.

Schneier:2005:ZS

[Sch05d]  Bruce Schneier.  The Zo-

tob storm. IEEE Security & Privacy, 3(6):96, November/December 2005. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/33104/01556546.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        33104&arnumber=1556546.

Schaffer:2006:WVB

[Sch06a] Gregory P. Schaffer. Worms and viruses and botnets, oh my!: Rational responses to emerging Internet threats.

IEEE Security & Privacy, 4(3): 52–58, May/June 2006. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneider:2006:HD

[Sch06b] Fred B. Schneider. Here be dragons. IEEE Security & Privacy, 4(3):3, May/June 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/comp/mags/sp/2006/03/
j3003.pdf.

Schneier:2006:UND

[Sch06c] Bruce Schneier. University networks and data security. IEEE Security & Privacy, 4(5):88, September/October 2006. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneider:2007:TCC

[Sch07a] Fred Schneider. Trusted computing in context. IEEE Security & Privacy, 5(2):4–5, March/April 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneider:2007:TSP

[Sch07b] Fred B. Schneider. Technology scapegoats and policy saviors. IEEE Security & Privacy, 5(5): 3–4, September/October 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/comp/mags/sp/2007/05/
msp2007050003.pdf.

Schneier:2007:DSI

[Sch07c] Bruce Schneier. The death of the security industry. IEEE Security & Privacy, 5(6): 88, November/December 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneier:2007:NCS

[Sch07d] Bruce Schneier. Nonsecurity considerations in security decisions. IEEE Security & Privacy, 5(3):88, May/June 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneider:2008:NNV

[Sch08a]  Fred B. Schneider.  Net-

work neutrality versus Internet trustworthiness? IEEE Security & Privacy, 6(4):3–4, July/ August 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneier:2008:HHB

[Sch08b] Bruce Schneier. How the human brain buys security. IEEE Security & Privacy, 6(4):80, July/August 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneider:2009:EAP

[Sch09a]  Fred B. Schneider.  From

the Editors: Accountability for perfection. IEEE Security & Privacy, 7(2):3–4, March/April 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneider:2009:LS

[Sch09b] Fred B. Schneider. Labelingin security. IEEE Security & Privacy, 7(6):3, November/ December 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneier:2009:CTA

[Sch09c] Bruce Schneier. Clear text: Architecture of privacy. IEEE Security & Privacy, 7(1):88, January/February 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneier:2009:CTS

[Sch09d] Bruce Schneier. Clear text: Security, group size, and the human brain. IEEE Security & Privacy, 7(4):88, July/August 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneider:2010:FFA

[Sch10a] Fred B. Schneider. Fumbling the future, again. IEEE Security & Privacy, 8(4):3, July/ August 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneier:2010:SFC

[Sch10b] Bruce Schneier. Security and function creep. IEEE Security & Privacy, 8(1):88, January/ February 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneier:2010:TSN

[Sch10c] Bruce Schneier. A taxonomy of social networking data. IEEE Security & Privacy, 8(4):88, July/August 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneier:2011:DC

[Sch11a] Bruce Schneier. Detecting cheaters. IEEE Security & Privacy, 9(2):96, 95, March/April 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneier:2011:ES

[Sch11b] Bruce Schneier. Empathy and security. IEEE Security & Privacy, 9(5):88, September/ October 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Schwartz:2011:PEA

[Sch11c] Paul M. Schwartz. Privacy, ethics, and analytics. IEEE Security & Privacy, 9(3):66–69, May/June 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneier:2012:HCT

[Sch12a] Bruce Schneier. How changing technology affects security. IEEE Security & Privacy, 10(2):104, March/April 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneier:2012:ISE

[Sch12b]  Bruce Schneier.  The im-

portance of security engineering. IEEE Security & Privacy, 10(5):88, September/October 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneider:2013:BR

[Sch13a] Fred B. Schneider. Breaking-in research. IEEE Security & Privacy, 11(2):3–4, March/April 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneider:2013:CEU

[Sch13b] Fred B. Schneider. Cybersecurity education in universities. IEEE Security & Privacy, 11(4):3–4, July/August 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneier:2013:IO

[Sch13c] Bruce Schneier. IT for oppression. IEEE Security & Privacy, 11(2):96, March/April 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneier:2013:LWT

[Sch13d] Bruce Schneier. Last word:

Trust in man/machine security systems. IEEE Security & Privacy, 11(5):96, September/ October 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneier:2014:FIR

[Sch14a] Bruce Schneier. The future of incident response. IEEE Security & Privacy, 12(5):96, September/October 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/05/
msp2014050096.html.

Schneier:2014:MS

[Sch14b] Bruce Schneier. Metadata = surveillance. IEEE Security & Privacy, 12(2):84, March/April 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneier:2015:SVM

[Sch15] Bruce Schneier. The security value of muddling through. IEEE Security & Privacy, 13 (2):108, March/April 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/02/
msp2015020108.html.

Schneier:2016:CHT

[Sch16a] Bruce Schneier. Cryptography is harder than it looks. IEEE Security & Privacy, 14(1):87– 88, January/February 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneier:2016:STF

[Sch16b] Bruce Schneier. Stop trying to fix the user. IEEE Security & Privacy, 14(5):96, September/October 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/05/
msp2016050096.html.

Schneier:2017:ITW

[Sch17a] Bruce Schneier. The Internet of Things will upend our industry. IEEE Security & Privacy, 15(2):108, March/April 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/02/
msp2017020108.html.

Schneier:2017:ISW

[Sch17b] Bruce Schneier. IoT security: What’s Plan B? IEEE Security & Privacy, 15(5):96, September/October 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/05/
msp2017050096.html.

Schneier:2018:CAA

[Sch18a] B. Schneier. Cryptography after the aliens land. IEEE Security & Privacy, 16(5):86– 88, September/October 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneier:2018:AIA

[Sch18b]  Bruce Schneier.  Artificial

intelligence and the attack/ defense balance. IEEE Security & Privacy, 16(2):96, March/April 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/02/
                                        msp2018020096.html.

Schneck:2019:CCN

[Sch19a] P. A. Schneck. Cybersecurity compliance is necessary but not sufficient: Bad guys don’t follow laws. IEEE Security & Privacy, 17(1):4–6, January/ February 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneier:2019:CPI

[Sch19b] Bruce Schneier. Cybersecurity for the public interest. IEEE Security & Privacy, 17(1):84– 83, January/February 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneck:2020:CDC

[Sch20a] P. A. Schneck. Cybersecurity during COVID-19. IEEE Security & Privacy, 18(6):4– 5, November/December 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneier:2020:HTC

[Sch20b] B. Schneier. Hacking the tax code. IEEE Security & Privacy, 18(5):79–80, September/ October 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneier:2020:TVP

[Sch20c] B. Schneier. Technologists vs. policy makers. IEEE Security & Privacy, 18(1):72–71, January/February 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneier:2021:WWI

[Sch21] Bruce Schneier. What will it take? IEEE Security & Privacy, 19(3):63–64, May/June 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Schmitt:2022:MDF

[Sch22a] Veronica Schmitt. Medical device forensics. IEEE Security & Privacy, 20(1):96–100, January/February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneier:2022:NPQ

[Sch22b] Bruce Schneier. NIST’s postquantum cryptography standards competition. IEEE Security & Privacy, 20(5):107– 108, September/October 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneier:2022:RHG

[Sch22c] Bruce Schneier. Robot hacking games. IEEE Security & Privacy, 20(1):119–120, January/ February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneier:2023:TAM

[Sch23] Bruce Schneier. Trustworthy AI means public AI [last word]. IEEE Security & Privacy, 21(6):95–96, November/ December 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Sun:2024:UMA

[SCHC24] Yeali S. Sun, Zhi-Kang Chen, Yi-Ting Huang, and Meng Chang Chen. Unleashing malware analysis and understanding with generative AI. IEEE Security & Privacy, 22(3):12–23, May/June 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Santin:2008:TBB

[SCM08] Altair O. Santin, Regivaldo G. Costa, and Carlos A. Maziero. A three-ballot-based secure electronic voting system. IEEE Security & Privacy, 6(3):14–21, May/June 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Stobert:2018:TAL

[SCMS18] E. Stobert, E. Cavar, L. Malisa, and D. Sommer. Teaching authentication as a life skill. IEEE Security & Privacy, 16 (5):82–85, September/October 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Salah:2013:UCC

[SCZ^{+}13] Khaled Salah, Jose M. Alcaraz Calero, Sherali Zeadally, Sameera Al-Mulla, and Mohammed Alzaabi. Using cloud computing to implement a security overlay network. IEEE Security & Privacy, 11(1):44– 53, January/February 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Sadeghi:2016:SPWa

[SD16a] Ahmad-Reza Sadeghi and Ghada Dessouky. Security & Privacy week interviews, Part 1. IEEE Security & Privacy, 14(5):58–67, September/October 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/05/
msp2016050058.html.

Sadeghi:2016:SPWb

[SD16b] Ahmad-Reza Sadeghi and Ghada Dessouky. Security & Privacy week interviews, part 2. IEEE Security & Privacy, 14(6):71–80, November/December 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/06/
msp2016060071-abs.html.

Sadeghi:2017:SPW

[SD17] Ahmad-Reza Sadeghi and Ghada Dessouky. Security & privacy week interviews, part 3. IEEE Security & Privacy, 15(1):66–74, January/February 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/01/
msp2017010066-abs.html.

Sherman:2017:ICC

[SDC^{+}17] Alan Sherman, Melissa Dark, Agnes Chan, Rylan Chong, Thomas Morris, Linda Oliva, John Springer, Bhavani Thuraisingham, Christopher Vatcher, Rakesh Verma, and Susanne Wetzel. INSuRE: Collaborating centers of academic excellence engage students in cybersecurity research. IEEE Security & Privacy, 15(4):72– 78, July/August 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/04/
msp2017040072-abs.html.

Stevens:2020:ILW

[SDKM20] R. Stevens, J. Dykstra, W. Knox Everette, and M. L. Mazurek. It lurks within: a look at the unexpected security implications of compliance programs. IEEE Security & Privacy, 18(6):51–58, November/ December 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Stark:2021:CTG

[SDO21] Emily Stark, Joe DeBlasio, and Devon O’Brien. Certificate transparency in Google Chrome: Past, present, and future. IEEE Security & Privacy, 19(6):112–118, November/December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Swart:2007:ESC

[SE07] Richard S. Swart and Robert F. Erbacher. Educating students to create trustworthy systems.

IEEE Security & Privacy, 5(3): 58–61, May/June 2007. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Spiekermann:2009:ACR

[SE09] Sarah Spiekermann and Sergei Evdokimov. Authentication: Critical RFID privacyenhancing technologies. IEEE Security & Privacy, 7(2):56–62, March/April 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Sood:2013:TCS

[SE13] Aditya K. Sood and Richard J.

Enbody. Targeted cyberattacks: A superset of advanced persistent threats. IEEE Security & Privacy, 11(1):54–61, January/February 2013. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Smith:2015:NMP

[SE15] Sean W. Smith and John S.

Erickson. Never mind Pearl Harbor — what about a Cyber Love Canal? IEEE Security & Privacy, 13(2):94– 98, March/April 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/02/
msp2015020094-abs.html.

Seacord:2006:SCC

[Sea06] Robert Seacord. Secure coding in C and C++: Of strings and integers. IEEE Security & Privacy, 4(1):74– 76, January/February 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/33481/01588832.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
33481&arnumber=1588832.

SeabraOliveira:2018:CPI

[SEKR18] D. Seabra Oliveira, J. Epstein, J. Kurose, and A. Rocha. Cybersecurity and privacy issues in Brazil: Back, now, and then [Guest Editors’ introduction]. IEEE Security & Privacy, 16(6):10–12, November/ December 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Seltzer:2011:EFC

[Sel11] Wendy Seltzer. Exposing the flaws of censorship by domain name. IEEE Security & Privacy, 9(1):83–87, January/ February 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Sendrier:2017:CBC

[Sen17] Nicolas Sendrier. Code-based cryptography: State of the art and perspectives. IEEE Security & Privacy, 15(4):44– 50, July/August 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/04/
msp2017040044-abs.html.

Shabtai:2010:SAP

[SFE10] Asaf Shabtai, Yuval Fledel, and Yuval Elovici. Securing Android-powered mobile devices using SELinux. IEEE Security & Privacy, 8(3):36–44, May/June 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Sheoran:2005:DSI

[SFH05] Pinny Sheoran, Oria Friesen, and Barbara J. Huffman de Belón. Developing and sustaining information assurance: The role of community colleges (part 1). IEEE Security & Privacy, 3(6): 61–63, November/December 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/33104/01556538.pdf;
http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
33104&arnumber=1556538.

Sheoran:2006:DSI

[SFH06] Pinny Sheoran, Oris Friesen, and Barbara J. Huffman de Belón. Developing and sustaining information assurance: The role of community colleges, part 2. IEEE Security & Privacy, 4(3):60–65, May/June 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Shabtai:2010:GAC

[SFK^{+}10] Asaf Shabtai, Yuval Fledel, Uri Kanonov, Yuval Elovici, Shlomi Dolev, and Chanan Glezer. Google Android: a comprehensive security assessment. IEEE Security & Privacy, 8(2):35–44, March/April 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Schrader:2018:PPA

[SG18] Dawn E. Schrader and Dipayan Ghosh. Proactively protecting against the singularity: Ethical decision making in AI. IEEE Security & Privacy, 16(3):56–63, May/June 2018. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/03/
msp2018030056-abs.html.

Schwarz:2020:HTE

[SG20] M. Schwarz and D. Gruss. How trusted execution environments fuel research on microarchitectural attacks. IEEE Security & Privacy, 18(5):18– 27, September/October 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Stone-Gross:2011:ABT

[SGCG^{+}11] Brett Stone-Gross, Marco Cova, Bob Gilbert, Richard Kemmerer, Christopher Kruegel, and Giovanni Vigna. Analysis of a botnet takeover. IEEE Security & Privacy, 9(1):64–72, January/February 2011. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Shi:2022:TMT

[SGSM22] Zhenpeng Shi, Kalman Graffi, David Starobinski, and Nikolay Matyunin. Threat modeling tools: a taxonomy. IEEE Security & Privacy, 20(4):29–39, July/August 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Souvignet:2020:LLL

[SHB20] T. R. Souvignet, T. Heckmann, and T. Bolle. From Lucky Luke to lock bits. IEEE Security & Privacy, 18(2):61–66, March/ April 2020. ISSN 1558-4046.

Spring:2021:TCC

[SHH^{+}21] Jonathan Spring, Eric Hatleback, Allen Householder, Art Manion, and Deana Shick. Time to change the CVSS? IEEE Security & Privacy, 19 (2):74–78, March/April 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Shostack:2022:YAS

[Sho22] Adam Shostack. 25 years in application security: Looking back, looking forward. IEEE Security & Privacy, 20(1):109– 112, January/February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Shostack:2023:NGE

[Sho23] Adam Shostack. Nothing is good enough: Fast and cheap are undervalued as influencers of security tool adoption. IEEE Security & Privacy, 21(1):78– 83, January/February 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Saydjari:2009:TTC

[SI09] O. Sami Saydjari and Cynthia E. Irvine. A tale of three cyber-defense workshops. IEEE Security & Privacy, 7(6): 60–64, November/December 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Simion:2015:RST

[Sim15] Emil Simion. The relevance of statistical tests in cryptography. IEEE Security & Privacy, 13(1):66–70, January/February 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/01/
msp2015010066-abs.html.

Simioni:2021:ITA

[Sim21]  Marco Simioni.  Investiga-

tive techniques for the deanonymization of hidden services. IEEE Security & Privacy, 19(2):60–64, March/ April 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Seigneur:2003:PRD

[SJ03] Jean-Marc Seigneur and Christian Damsgaard Jensen. Privacy recovery with disposable email addresses. IEEE Security & Privacy, 1(6): 35–39, November/December 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2003/06/j6035abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/06/j6035.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/06/j6035.
pdf.

Sherman:2019:OVB

[SJZG19] A. T. Sherman, F. Javani, H. Zhang, and E. Golaszewski. On the origins and variations of blockchain technologies. IEEE Security & Privacy, 17(1):72– 77, January/February 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Shankar:2004:COS

[SK04] K. S. Shankar and Helmut Kurth. Certifying open source—the Linux experience. IEEE Security & Privacy, 2(6):28–33, November/ December 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/06/j6028.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/06/j6028.
pdf.

Sidiroglou:2005:CNW

[SK05] Stelios Sidiroglou and Angelos D. Keromytis. Countering network worms through automatic patch generation. IEEE Security & Privacy, 3(6): 41–49, November/December 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/33104/01556535.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        33104&arnumber=1556535.

SaidElsayed:2021:DCN

[SLKJ21] M. Said Elsayed, N.-A. Le-Khac, and A. D. Jurcut. Dealing with COVID-19 network traffic spikes [cybercrime and forensics]. IEEE Security & Privacy, 19(1):90–94, January/ February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Strohmeier:2023:PAC

[SLM^{+}23] Martin Strohmeier, Mauro Leonardi, Sergei Markochev, Fabio Ricciato, Matthias Schäfer, and Vincent Lenders. In pursuit of aviation cybersecurity: Experiences and lessons from a competitive approach. IEEE Security & Privacy, 21(4):61– 73, July/August 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Slomovic:2014:PII

[Slo14] Anna Slomovic. Privacy issues in identity verification. IEEE Security & Privacy, 12(3):71– 73, May/June 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Sun:2018:TCC

[SLS18] X. Sun, P. Liu, and A. Singhal. Toward cyberresiliency in the context of cloud computing [resilient security]. IEEE Security & Privacy, 16(6):71– 75, November/December 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Skopik:2022:BSS

[SLW22a] Florian Skopik, Max Landauer, and Markus Wurzenberger. Blind spots of security monitoring in enterprise infrastructures: a survey. IEEE Security & Privacy, 20(6):18– 26, November/December 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Skopik:2022:OLD

[SLW22b] Florian Skopik, Max Landauer, and Markus Wurzenberger. Online log data analysis with efficient machine learning: A review. IEEE Security & Privacy, 20(3):80–90, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Schell:2010:MPK

[SLZ^{+}10] Roger Schell, Steve Lipner, Mary Ellen Zurko, Elaine R. Palmer, David Safford, Charles C. Palmer, and Carl E. Landwehr. In memoriam: Paul Karger. IEEE Security & Privacy, 8(6): 5, November/December 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Shannon:2004:SWW

[SM04] Colleen Shannon and David Moore. The spread of the witty worm. IEEE Security & Privacy, 2(4):46–50, July/August 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/04/j4046.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/04/j4046.
pdf.

Schneider:2011:DT

[SM11] Fred B. Schneider and Deirdre K. Mulligan. A doctrinal thesis. IEEE Security & Privacy, 9(4): 3–4, July/August 2011. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Smart:2023:CED

[Sma23] Nigel Smart. Computing on encrypted data. IEEE Security & Privacy, 21(4):94–98, July/ August 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Schwartz:2014:ETF

[SMCA14] Moses Schwartz, John Mulder, Adrian R. Chavez, and Benjamin A. Allan. Emerging techniques for field device security. IEEE Security & Privacy, 12(6):24–31, November/December 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/06/
msp2014060024-abs.html.

Sklavos:2004:CRC

[SMGK04] Nicolas Sklavos, Nikolay Moldovyan, Vladimir Gorodetsky, and Odysseas Koufopavlou. Conference reports: Computer network security: Report from MMM-ACNS. IEEE Security & Privacy, 2(1):49– 52, January/February 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2004/01/j1049abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/01/j1049.
pdf.

Smith:2003:FDS

[Smi03a] S. W. Smith. Fairy dust, secrets, and the real world. IEEE Security & Privacy, 1(1):89–93, January/February 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j1089.pdf;   http://www.
computer.org/security/j1089abs.
htm.

Smith:2003:HLH

[Smi03b] S. W. Smith. Humans in the loop: Human-computer interaction and security. IEEE Security & Privacy, 1(3):75– 79, May/June 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j3075.pdf.

Smith:2005:PSS

[Smi05a] S. W. Smith. Pretending that systems are secure. IEEE Security & Privacy, 3(6): 73–76, November/December 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/33104/01556541.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        33104&arnumber=1556541.

Smith:2005:TMS

[Smi05b] S. W. Smith. Turing is from Mars, Shannon is from Venus: computer science and computer engineering. IEEE Security & Privacy, 3(2):66– 69, March/April 2005. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/30742/01423965.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        30742&arnumber=1423965&count=
                                        15&index=11.

Smith:2011:RBA

[Smi11] Sean W. Smith. Room at the bottom: Authenticated encryption on slow legacy networks. IEEE Security & Privacy, 9(4):60–63, July/August 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Smith:2012:CLS

[Smi12a] Richard E. Smith. A contemporary look at Saltzer and Schroeder’s 1975 Design Principles. IEEE Security & Privacy, 10(6):20–25, November/ December 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Smith:2012:SCB

[Smi12b] Sean W. Smith. Security and cognitive bias: Exploring the role of the mind. IEEE Security & Privacy, 10(5):75–78, September/October 2012. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Schneider:2019:SBD

[SMJ^{+}19] M. Schneider, S. Matetic, A. Juels, A. Miller, and S. Capkun. Secure brokered delegation through DelegaTEE. IEEE Security & Privacy, 17 (4):43–52, July/August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Schiffman:2011:NBR

[SMJM11] Joshua Schiffman, Thomas Moyer, Trent Jaeger, and Patrick McDaniel. Networkbased root of trust for installation. IEEE Security & Privacy, 9(1):40–48, January/February 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Shoemaker:2020:TSA

[SMK20a] D. Shoemaker, N. R. Mead, and A. Kohnke. Teaching secure acquisition in higher education. IEEE Security & Privacy, 18(4):60–66, July/ August 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Sterlini:2020:GCE

[SMK^{+}20b] P. Sterlini, F. Massacci, N. Kadenko, T. Fiebig, and M. van Eeten. Governance challenges for European cybersecurity policies: Stakeholder views. IEEE Security & Privacy, 18(1):46–54, January/ February 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Sundaramurthy:2014:AAS

[SMO^{+}14] Sathya Chandran Sundaramurthy, John McHugh, Xinming Simon Ou, S. Raj Rajagopalan, and Michael Wesch. An anthropological approach to studying CSIRTs. IEEE Security & Privacy, 12(5):52–60, September/October 2014. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/05/
msp2014050052-abs.html.

Sang:2014:TAP

[SND14] Fernand Lone Sang, Vincent Nicomette, and Yves Deswarte. A tool to analyze potential I/O attacks against PCs. IEEE Security & Privacy, 12(2):60–66, March/April 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Snow:2005:FWI

[Sno05] B. Snow. Four ways to improve security. IEEE Security & Privacy, 3(3):65–67, May/ June 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/31002/01439505.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
31002&arnumber=1439505&count=
20&index=12.

Sorge:2010:LRC

[SNS10] Christoph Sorge, Saverio Niccolini, and Jan Seedorf. The legal ramifications of callfiltering solutions. IEEE Security & Privacy, 8(2):45–50, March/April 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Sherman:2019:CHC

[SOG^{+}19] A. T. Sherman, L. Oliva, E. Golaszewski, D. Phatak, T. Scheponik, G. L. Herman, D. S. Choi, S. E. Offenberger, P. Peterson, J. Dykstra, G. V. Bard, A. Chattopadhyay, F. Sharevski, R. Verma, and R. Vrecenar. The CATS hackathon: Creating and refining test items for cybersecurity concept inventories. IEEE Security & Privacy, 17(6):77– 83, November/December 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Solomon:2007:BPR

[Sol07] Dianne Solomon. Balancing privacy and risk in the Emessaging world. IEEE Security & Privacy, 5(5):72–75, September/October 2007. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Somayaji:2004:HWE

[Som04]  Anil Somayaji.  How to

win and evolutionary arms race. IEEE Security & Privacy, 2(6):70–72, November/ December 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/sp/2004/06/j6070.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2004/06/j6070.
                                        pdf.

Steven:2007:M

[SP07] John Steven and Gunnar Peterson. Metricon 2.0. IEEE Security & Privacy, 5(6):81– 83, November/December 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Steingruebl:2009:BSS

[SP09] Andy Steingruebl and Gunnar Peterson. Building security in: Software assumptions lead to preventable errors. IEEE Security & Privacy, 7(4):84–87, July/August 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Sasse:2014:PY

[SP14] M. Angela Sasse and Charles C. Palmer. Protecting you. IEEE Security & Privacy, 12(1):11– 13, January/February 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Spafford:2006:DIB

[Spa06] Eugene Spafford. Desert island books. IEEE Security & Privacy, 4(6):5, November/ December 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Spafford:2008:JPA

[Spa08] Eugene Spafford. James P.

Anderson: An information security pioneer. IEEE Security & Privacy, 6(1):9, January/February 2008. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/comp/mags/sp/2008/01/
msp2008010009.pdf.

Sassaman:2012:PPR

[SPB12] Len Sassaman, Meredith L.

Patterson, and Sergey Bratus. A patch for Postel’s Robustness Principle. IEEE Security & Privacy, 10(2):87–91, March/April 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Sherman:2019:PBL

[SPG^{+}19] A. T. Sherman, P. A. H. Peterson, E. Golaszewski, E. LaFemina, E. Goldschen, M. Khan, L. Mundy, M. Rather, B. Solis, W. Tete, E. Valdez, B. Weber, D. Doyle, C. O’Brien, L. Oliva, J. Roundy, and J. Suess. Project-based learning inspires cybersecurity students: A scholarship-for-service research study. IEEE Security & Privacy, 17(3):82–88, May/June 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Senior:2005:EVP

[SPH^{+}05] A. Senior, S. Pankanti, A. Hampapur, L. Brown, Ying-Li Tian, A. Ekin, J. Connell, Chiao Fe Shu, and M. Lu. Enabling video privacy through computer vision. IEEE Security & Privacy, 3(3):50–57, May/June 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/31002/01439502.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        31002&arnumber=1439502&count=
                                        20&index=9.

Spitzner:2003:HPT

[Spi03] Lance Spitzner. The Honeynet Project: Trapping the hackers. IEEE Security & Privacy, 1(2): 15–23, March/April 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://dlib.computer.
                                        org/sp/books/sp2003/pdf/
                                        j2015.pdf;   http://www.
                                        computer.org/security/j2015abs.
                                        htm.

Sasse:2014:HYP

[SPJ^{+}14] M. Angela Sasse, Charles C. Palmer, Markus Jakobsson, Sunny Consolvo, Rick Wash, and L. Jean Camp. Helping you protect you. IEEE Security & Privacy, 12(1):39–42, January/ February 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Sabetta:2024:KVO

[SPL^{+}24] Antonino Sabetta, Serena Elisa Ponta, Rocio Cabrera Lozoya, Michele Bezzi, Tommaso Sacchetti, Matteo Greco, Gerg Balogh, Péter Heged s, Rudolf Ferenc, Ranindya Paramitha, Ivan Pashchenko, Aurora Papotti, Ákos Milánkovich, and Fabio Massacci. Known vulnerabilities of open source projects: Where are the fixes? IEEE Security & Privacy, 22 (2):49–59, March/April 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Spring:2011:MCCa

[Spr11a] Jonathan Spring. Monitoring cloud computing by layer, part 1. IEEE Security & Privacy, 9(2):66–68, March/April 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Spring:2011:MCCb

[Spr11b] Jonathan Spring. Monitoring cloud computing by layer, part 2. IEEE Security & Privacy, 9(3):52–55, May/June 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Scaife:2019:MCL

[SPT19] N. Scaife, C. Peeters, and P. G. Traynor. More cowbell: Lessons learned in developing the skim reaper. IEEE Security & Privacy, 17(4):10–17, July/ August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Szekeres:2014:EWM

[SPWS14] Laszlo Szekeres, Mathias Payer, Lenx Tao Wei, and R. Sekar. Eternal war in memory. IEEE Security & Privacy, 12(3):45– 53, May/June 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Scott-Railton:2016:SHR

[SR16] John Scott-Railton. Security for the high-risk user: Separate and unequal. IEEE Security & Privacy, 14(2):79–87, March/April 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Smith:2003:FTH

[SRG03] S. W. Smith, Jothy Rosenberg, and Adam Golodner. A funny thing happened on the way to the marketplace. IEEE Security & Privacy, 1(6): 74–78, November/December 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/sp/2003/06/j6074abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2003/06/j6074.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2003/06/j6074.
                                        pdf.

Smith:2004:SSG

[SS04] Sean W. Smith and Eugene H. Spafford. Secure systems: Grand challenges in information security: Process and output. IEEE Security & Privacy, 2(1):69–71, January/ February 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2004/01/j1069abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/01/j1069.
pdf.

Sinclair:2005:TPT

[SS05] Sara Sinclair and S. W. Smith.

The TIPPI point: Toward trustworthy interfaces. IEEE Security & Privacy, 3(4):68–71, July/August 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Suchan:2006:SWL

[SS06] William Suchan and Edward Sobiesk. Strengthening the weakest link in digital protection. IEEE Security & Privacy, 4(6):78–80, November/ December 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Sinclair:2010:WWA

[SS10] Sara Sinclair and Sean W.

Smith. What’s wrong with access control in the real world? IEEE Security & Privacy, 8(4): 74–77, July/August 2010. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Sasse:2016:SUT

[SS16] M. Angela Sasse and Matthew Smith. The security–usability tradeoff myth [Guest Editors’ introduction]. IEEE Security & Privacy, 14(5):11–13, September/October 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2016/05/
                                        msp2016050011.html.

Siegel:2019:CPS

[SS19a] J. Siegel and S. Sarma. A cognitive protection system for the Internet of Things. IEEE Security & Privacy, 17(3):40–48, May/June 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Siegel:2019:UOC

[SS19b] J. E. Siegel and S. Sarma. Using open channels to trigger the invited, unintended consequences of the Internet of Things. IEEE Security & Privacy, 17(3):49–55, May/June 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Strobel:2022:DPT

[SS22] Martin Strobel and Reza Shokri. Data privacy and trustworthy machine learning. IEEE Security & Privacy, 20 (5):44–49, September/October 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Sasse:2016:DSU

[SSH^{+}16] M. Angela Sasse, Matthew Smith, Cormac Herley, Heather Lipford, and Kami Vaniea. Debunking security–usability tradeoff myths. IEEE Security & Privacy, 14(5):33–39, September/October 2016. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/05/
msp2016050033-abs.html.

Shim:2022:DLO

[SSH22] WooChul Shim, Hyejin Shin, and Yong Ho Hwang. On data licenses for open source threat intelligence. IEEE Security & Privacy, 20(4):8–22, July/August 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Sarieddine:2023:RTC

[SSJ^{+}23] Khaled Sarieddine, Mohammad Ali Sayed, Danial Jafarigiv, Ribal Atallah, Mourad Debbabi, and Chadi Assi. A real-time cosimulation testbed for electric vehicle charging and smart grid security. IEEE Security & Privacy, 21(4):74–83, July/August 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Sanzo:2021:DKC

[SSW21] Karen L. Sanzo, Jay Paredes Scribner, and Hongyi Wu. Designing a K-16 cybersecurity collaborative: CIPHER. IEEE Security & Privacy, 19(2):56– 59, March/April 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Slay:2006:CSE

[ST06] Jill Slay and Benjamin Turnbull. Computer security education and research in Australia. IEEE Security & Privacy, 4(5):60–63, September/ October 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Stolfo:2010:PPS

[ST10] Salvatore J. Stolfo and Gene Tsudik. Privacy-preserving sharing of sensitive information. IEEE Security & Privacy, 8(4):16–17, July/August 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Stallings:2020:HPI

[Sta20] W. Stallings. Handling of personal information and deidentified, aggregated, and pseudonymized information under the California Consumer Privacy Act. IEEE Security & Privacy, 18(1):61–64, January/ February 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Steven:2006:AES

[Ste06] John Steven. Adopting an enterprise software security framework. IEEE Security & Privacy, 4(2):84–87, March/ April 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Steven:2008:BSS

[Ste08] John Steven. Building security in: State of application assessment. IEEE Security & Privacy, 6(6):82–85, November/ December 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Steven:2010:TM

[Ste10] John Steven. Threat modeling.

IEEE Security & Privacy, 8(3): 83–86, May/June 2010. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Stevens:2011:MPD

[Ste11] Didier Stevens. Malicious PDF documents explained. IEEE Security & Privacy, 9(1):80–82, January/February 2011. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Stein:2013:ENS

[Ste13] Jeff Stein. The end of national security reporting? IEEE Security & Privacy, 11(4):64–68, July/August 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Steel:2015:APF

[Ste15] Graham Steel. Automated proof and flaw-finding tools in cryptography. IEEE Security & Privacy, 13(2):81– 83, March/April 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/02/
msp2015020081-abs.html.

Stolfo:2004:WAE

[Sto04] Salvatore J. Stolfo. Worm and attack early warning. IEEE Security & Privacy, 2(3):73– 75, May/June 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/03/j3073.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/03/j3073.
pdf.

Strembeck:2010:SDR

[Str10] Mark Strembeck. Scenariodriven role engineering. IEEE Security & Privacy, 8(1):28–35, January/February 2010. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Strigini:2012:RWI

[Str12] Lorenzo Strigini. Resilience: What is it, and how much do we want? IEEE Security & Privacy, 10(3):72–75, May/June 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Stytz:2003:CSW

[Sty03] Martin R. Stytz. The case for software warranties. IEEE Security & Privacy, 1(3):80– 82, May/June 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j3080.pdf.

Stytz:2004:BRH

[Sty04a] Martin R. Stytz. Book reviews: Hacking for understanding: Hacking: The Art of Exploitation, by Jon Erickson. IEEE Security & Privacy, 2 (2):8, March/April 2004. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/comp/mags/sp/2004/02/
j2008.pdf.

Stytz:2004:BRW

[Sty04b] Martin R. Stytz. Book reviews: Wireless world order [How Secure Is Your Wireless Network? Safeguarding Your Wi-Fi LAN by Lee Barken]; no need to fear [Beyond Fear: Thinking Sensibly About Security in an Uncertain World, by Bruce Schneier]. IEEE Security & Privacy, 2(1):20–21, January/ February 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2004/01/j1020abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/01/j1020.
pdf.

Stytz:2004:CDD

[Sty04c] Martin R. Stytz. Considering defense in depth for software applications. IEEE Security & Privacy, 2(1):72–75, January/ February 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/

                                        mags/sp/2004/01/j1072abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2004/01/j1072.
                                        pdf.

Stytz:2005:UB

[Sty05a]  M. R. Stytz.  Under the

Black Hat. IEEE Security & Privacy, 3(3):5, May/June 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/31002/01439493.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        31002&arnumber=1439493&count=
                                        20&index=1.

Stytz:2005:PPP

[Sty05b] Martin R. Stytz. Protecting personal privacy: Hauling down the Jolly Roger. IEEE Security & Privacy, 3(4):72–74, July/August 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Stytz:2007:WN

[Sty07a] Martin R. Stytz. What are the numbers? IEEE Security & Privacy, 5(6):12, November/December 2007. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/sp/2007/06/
                                        msp2007060012.pdf.

Stytz:2007:WEW

[Sty07b] Martin R. Stytz. Who are the experts, and what have they done for us lately? IEEE Security & Privacy, 5(6):78– 80, November/December 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Stytz:2008:BRS

[Sty08] Martin R. Stytz. Book reviews: The shape of crimeware to come. IEEE Security & Privacy, 6(5):5, September/ October 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Sunyaev:2016:DCC

[Sun16] Ali Sunyaev. Dynamic certification of cloud services: Trust, but verify! IEEE Security & Privacy, 14(2):66–71, March/ April 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Suri:2016:QTC

[Sur16]  Neeraj Suri.  Quantifiably

trusting the cloud: Putting metrics to work. IEEE Security & Privacy, 14(3):73–77, May/June 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Stevens:2019:ADT

[SVR^{+}19] R. Stevens, D. Votipka, E. M.

Redmiles, C. Ahern, and M. L. Mazurek. Applied digital threat modeling: It works. IEEE Security & Privacy, 17 (4):35–42, July/August 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Stytz:2003:CPC

[SW03a] Martin Stytz and James A. Whittaker. Caution: This product contains security code. IEEE Security & Privacy, 1 (5):86–88, September/October 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2003/05/j5086abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/05/j5086.
pdf.

Stytz:2003:SPS

[SW03b] Martin R. Stytz and James A. Whittaker. Software protection: Security’s last stand? IEEE Security & Privacy, 1(1): 95–98, January/February 2003. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j1095.pdf;   http://www.
computer.org/security/j1095abs.
htm.

Stark:2012:EBE

[SW12] Philip B. Stark and David Wagner. Evidence-based elections. IEEE Security & Privacy, 10(5):33–41, September/ October 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Sloan:2018:WAT

[SW18a] Robert H. Sloan and Richard Warner. When is an algorithm transparent? Predictive analytics, privacy, and public policy. IEEE Security & Privacy, 16(3):18–25, May/June 2018. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/03/
msp2018030018-abs.html.

Stahl:2018:EPA

[SW18b] Bernd Carsten Stahl and David Wright. Ethics and privacy in AI and big data: Implementing responsible research and innovation. IEEE Security & Privacy, 16(3):26– 33, May/June 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/03/
msp2018030026-abs.html.

Shi:2021:AVC

[SWCS21] Cong Shi, Yan Wang, Yingying Jennifer Chen, and Nitesh Saxena. Authentication of voice commands by leveraging vibrations in wearables. IEEE Security & Privacy, 19(6):83– 92, November/December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Skopik:2021:SGP

[SWL21] Florian Skopik, Markus Wurzenberger, and Max Landauer. The seven golden principles of effective anomaly-based intrusion detection. IEEE Security & Privacy, 19(5):36– 45, September/October 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Sheldon:2012:IWN

[SWYP12] Frederick T. Sheldon, John Mark Weber, Seong-Moo Yoo, and W. David Pan. The insecurity of wireless networks. IEEE Security & Privacy, 10(4):54–61, July/August 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Shamos:2012:RVS

[SY12] Michael Shamos and Alec Yasinsac. Realities of E-voting security. IEEE Security & Privacy, 10(5):16–17, September/ October 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Shanmugarasa:2022:APP

[SyPKZ22] Yashothara Shanmugarasa, Hye young Paik, Salil S. Kanhere, and Liming Zhu. Automated privacy preferences for smart home data sharing using personal data stores. IEEE Security & Privacy, 20(1):12–22, January/February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Sytz:2005:SAI

[Syt05] M. R. Sytz. Studying attacks to improve software defense. IEEE Security & Privacy, 3(1): 11, January/February 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/30310/01392692.pdf;
                                         http://ieeexplore.ieee.

org/xpls/abs_all.jsp?isnumber=
30310&arnumber=1392692&count=
17&index=1.

Schneider:2005:ITS

[SZ05] Fred B. Schneider and Lidong Zhou. Implementing trustworthy services using replicated state machines. IEEE Security & Privacy, 3(5):34–43, September/October 2005. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Shaunghe:2006:EPS

[SZ06] Peng Shaunghe and Han Zhen.

Enhancing PC security with a U-key. IEEE Security & Privacy, 4(5):34–39, September/ October 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Sood:2016:TDG

[SZ16] Aditya K. Sood and Sherali Zeadally. A taxonomy of domain-generation algorithms. IEEE Security & Privacy, 14 (4):46–53, July/August 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/04/
msp2016040046-abs.html.

Sadeghi:2017:ACIc

[SZ17a] Ahmad-Reza Sadeghi and Shaza Zeitouni. ACM CCS 2016 interview, part 3. IEEE Security & Privacy, 15(4):58– 61, July/August 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/04/
                                        msp2017040058.html.

Sadeghi:2017:ACIa

[SZ17b] Ahmad-Reza Sadeghi and Shaza Zeitouni. ACM CCS 2016 interviews, part 1. IEEE Security & Privacy, 15(2):88– 91, March/April 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/02/
                                        msp2017020088-abs.html.

Sadeghi:2017:ACIb

[SZ17c] Ahmad-Reza Sadeghi and Shaza Zeitouni. ACM CCS 2016 interviews, part 2. IEEE Security & Privacy, 15(3):70– 76, May/June 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/03/
                                        msp2017030070.html.

True:2023:PIA

[TA23] John True and Navid Asadizanjani. Physical inspection and attacks on electronics: an academic course for the hardware cybersecurity workforce. IEEE Security & Privacy, 21(2):63– 69, March/April 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Torres-Arias:2023:VKS

[TAGM23] Santiago Torres-Arias, Dan Geer, and John Speed Meyers.

A viewpoint on knowing software: Bill of materials quality when you see it. IEEE Security & Privacy, 21(6):50– 54, November/December 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Toth:2019:SSD

[TAP19] K. C. Toth and A. Anderson-Priddy. Self-sovereign digital identity: A paradigm shift for identity. IEEE Security & Privacy, 17(3):17–27, May/June 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Tschofenig:2019:CSM

[TB19] H. Tschofenig and E. Baccelli.

Cyberphysical security for the masses: A survey of the Internet protocol suite for Internet of Things security. IEEE Security & Privacy, 17(5):47– 57, September/October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Traynor:2017:FAS

[TBBR17] Patrick Traynor, Kevin Butler, Jasmine Bowers, and Bradley Reaves. FinTechSec: Addressing the security challenges of digital financial services. IEEE Security & Privacy, 15(5):85– 89, September/October 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2017/05/
msp2017050085-abs.html.

Tariang:2024:SIV

[TCC^{+}24] Diangarti Tariang, Riccardo Corvi, Davide Cozzolino, Giovanni Poggi, Koki Nagano, and Luisa Verdoliva. Synthetic image verification in the era of generative artificial intelligence: What works and what isn’t there yet. IEEE Security & Privacy, 22(3):37–49, May/ June 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Tian:2019:PAT

[TCJ^{+}19] H. Tian, Y. Chen, H. Jiang, Y. Huang, F. Nan, and Y. Chen. Public auditing for trusted cloud storage services. IEEE Security & Privacy, 17 (1):10–22, January/February 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Tsipenyuk:2005:SPK

[TCM05] Katrina Tsipenyuk, Brian Chess, and Gary McGraw. Seven pernicious kingdoms: a taxonomy of software security errors. IEEE Security & Privacy, 3(6):81– 84, November/December 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/33104/01556543.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        33104&arnumber=1556543.

Tsai:2006:S

[TE06] Janice Y. Tsai and Serge Egelman. Soups 2006. IEEE Security & Privacy, 4(6):53– 55, November/December 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Tene:2019:GYO

[TEG^{+}19] O. Tene, K. Evans, B. Gencarelli, G. Maldoff, and G. Zanfir-Fortuna. GDPR at year one: Enter the designers and engineers. IEEE Security & Privacy, 17(6):7–9, November/ December 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Telang:2015:PFD

[Tel15] Rahul Telang. Policy framework for data breaches. IEEE Security & Privacy, 13(1):77– 79, January/February 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/01/
msp2015010077-abs.html.

Thomsen:2012:LT

[TEN12] Dan Thomsen, Jeremy Epstein, and Peter G. Neumann. Lost treasures. IEEE Security & Privacy, 10(6):17– 19, November/December 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Tene:2016:MVU

[Ten16]  Omer Tene.  Microsoft v.

USA: Location of data and the law of the horse. IEEE Security & Privacy, 14(6):81– 85, November/December 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2016/06/
                                        msp2016060081-abs.html.

Talbot:2010:DC

[TFB10] Edward B. Talbot, Deborah Frincke, and Matt Bishop. Demythifying cybersecurity. IEEE Security & Privacy, 8(3): 56–59, May/June 2010. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Thorsteinson:2004:NSC

[TG04] Peter Thorsteinson and G. Gnana Arun Ganesh. .NET Security and Cryptography. The integrated .NET series from object innovations. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2004. ISBN 0-13-100851-X. xvii + 466 pp. LCCN ???? US$49.99.

Theoharidou:2007:CBK

[TG07] Marianthi Theoharidou and Dimitris Gritazalis. Common body of knowledge for information security. IEEE Security & Privacy, 5(2):64–67, March/ April 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Theofanos:2016:SUE

[TGC16] Mary Theofanos, Simson Garfinkel, and Yee-Yin Choong. Secure and usable enterprise authentication: Lessons from the field. IEEE Security & Privacy, 14(5):14–21, September/October 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/05/
msp2016050014-abs.html.

Tondel:2020:ISM

[TGS20] I. A. Tondel, M. Gilje Jaatun, and D. Soares Cruzes. IT security is from Mars, software security is from Venus. IEEE Security & Privacy, 18(4):48–54, July/August 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Trope:2013:EOD

[TH13] Roland L. Trope and Stephen J.

Humes. By executive order: Delivery of cyber intelligence imparts cyber responsibilities. IEEE Security & Privacy, 11 (2):63–67, March/April 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Thibadeau:2006:TCD

[Thi06] Robert Thibadeau. Trusted computing for disk drives and other peripherals. IEEE Security & Privacy, 4(5):26–33, September/October 2006. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Trappe:2015:LES

[THM15] Wade Trappe, Richard Howard, and Robert S. Moore. Lowenergy security: Limits and opportunities in the Internet of Things. IEEE Security & Privacy, 13(1):14–21, January/February 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/sp/2015/01/
                                        msp2015010014-abs.html.

Thompson:2005:APT

[Tho05]  H. H. Thompson.  Appli-

cation penetration testing. IEEE Security & Privacy, 3 (1):66–69, January/February 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/30310/01392704.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        30310&arnumber=1392704&count=
                                        17&index=11.

Thompson:2013:HEI

[Tho13] Hugh Thompson. The human element of information security. IEEE Security & Privacy, 11(1):32–35, January/ February 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Thompson:2018:EHC

[Tho18] Michael F. Thompson. Effects of a honeypot on the cyber grand challenge final event. IEEE Security & Privacy, 16 (2):37–41, March/April 2018. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/02/
                                        msp2018020037-abs.html.

Tian:2020:SUG

[THS20] Y. Tian, C. Herley, and S. Schechter. StopGuessing: Using guessed passwords to thwart online password guessing. IEEE Security & Privacy, 18(3):38–47, May/June 2020. ISSN 1558-4046.

Thompson:2018:ICL

[TI18] Michael F. Thompson and Cynthia E. Irvine. Individualizing cybersecurity lab exercises with labtainers. IEEE Security & Privacy, 16(2):91– 95, March/April 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/02/
msp2018020091-abs.html.

Takabi:2010:SPC

[TJA10] Hassan Takabi, James B. D.

Joshi, and Gail-Joon Ahn. Security and privacy challenges in cloud computing environments. IEEE Security & Privacy, 8(6):24–31, November/ December 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Tedeschi:2022:PPS

[TJS^{+}22] Pietro Tedeschi, Kang Eun Jeon, James She, Simon Wong, Spiridon Bakiras, and Roberto Di Pietro. Privacy-preserving and sustainable contact tracing using batteryless Bluetooth low-energy beacons. IEEE Security & Privacy, 20(3):91– 100, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Thimbleby:2015:HD

[TK15] H. Thimbleby and R. Koppel. The healthtech declaration. IEEE Security & Privacy, 13(6):82–84, November/ December 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Taylor:2021:CDS

[TKZ21] Blair Taylor, Sidd Kaza, and Paige A. Zaleppa. CLARK: A design science research project for building and sharing highquality cybersecurity curricula. IEEE Security & Privacy, 19 (5):72–76, September/October 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Taylor:2005:ASS

[TM05] D. Taylor and G. McGraw. Adopting a software security improvement program. IEEE Security & Privacy, 3(3):88– 91, May/June 2005. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/31002/01439511.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        31002&arnumber=1439511&count=
                                        20&index=18.

Tormo:2013:IMP

[TMGP13] Gines Dolera Tormo, Felix Gomez Marmol, Joao Girao, and Gregorio Martinez Perez.

Identity management — in privacy we trust: Bridging the trust gap in eHealth environments. IEEE Security & Privacy, 11(6):34–41, November/ December 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Turner:2023:RBT

[TMU23] Adam Brian Turner, Stephen McCombie, and Allon J. Uhlmann. Ransomware-bitcoin threat intelligence sharing using structured threat information expression. IEEE Security & Privacy, 21(3):47–57, May/ June 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Tomaszewski:2006:YSY

[Tom06] John P. Tomaszewski. Are you sure you had a privacy incident? IEEE Security & Privacy, 4(6):64–66, November/ December 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Tomb:2016:AVR

[Tom16] Aaron Tomb. Automated verification of real-world cryptographic implementations. IEEE Security & Privacy, 14 (6):26–33, November/December 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL https:

//www.computer.org/csdl/
mags/sp/2016/06/msp2016060026-
abs.html.

Torr:2005:DTM

[Tor05] Peter Torr. Demystifying the threat-modeling process. IEEE Security & Privacy, 3(5):66–70, September/October 2005. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Trope:2006:LLC

[TP06] Roland L. Trope and E. Michael Power. Lessons for laptops for the 18th Century. IEEE Security & Privacy, 4(4):64–68, July/August 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Trellue:2008:GEI

[TP08] Ron Trellue and Charles C. Palmer. Guest Editors’ introduction: Process control system security: Bootstrapping a legacy. IEEE Security & Privacy, 6(6):22–23, November/ December 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Theofanos:2011:GEI

[TP11] Mary Frances Theofanos and Shari Lawrence Pfleeger. Guest Editors’ introduction: Shouldn’t all security be usable? IEEE Security & Privacy, 9(2):12–17, March/April 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Tyagi:2015:PIU

[TPMT15] Rohit Tyagi, Tuhin Paul, B. S. Manoj, and B. Thanudas.

Packet inspection for unauthorized OS detection in enterprises. IEEE Security & Privacy, 13(4):60– 65, July/August 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/04/
msp2015040060-abs.html.

Trope:2007:CSD

[TPPM07] Roland L. Trope, E. Michael Power, Vincent I. Polley, and Bradford C. Morley. A coherent strategy for data security through data governance. IEEE Security & Privacy, 5(3): 32–39, May/June 2007. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Tene:2018:FFH

[TPS18] Omer Tene, Jules Polonetsky, and Ahmad-Reza Sadeghi. Five freedoms for the Homo Deus. IEEE Security & Privacy, 16(3):15–17, May/June 2018. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL https:

//www.computer.org/csdl/
mags/sp/2018/03/msp2018030015.
html.

Troutman:2009:CCGa

[TR09a] Justin Troutman and Vincent Rijmen. Crypto corner: Green cryptography: Cleaner engineering through recycling. IEEE Security & Privacy, 7(4): 71–73, July/August 2009. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Troutman:2009:CCGb

[TR09b] Justin Troutman and Vincent Rijmen. Crypto corner: Green cryptography: Cleaner engineering through recycling, part 2. IEEE Security & Privacy, 7 (5):64–65, September/October 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Traynor:2009:SSS

[Tra09] Patrick Traynor. Secure systems: Securing cellular infrastructure: Challenges and opportunities. IEEE Security & Privacy, 7(4):77–79, July/ August 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Trcek:2011:TMP

[Trč11] Denis Trček. Trust management in the pervasive computing era. IEEE Security & Privacy, 9(4):52–55, July/August 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Trope:2004:WC

[Tro04] Roland L. Trope. A warranty of cyberworthiness. IEEE Security & Privacy, 2(2): 73–76, March/April 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/sp/2004/02/j2073abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2004/02/j2073.
                                        pdf.

Trope:2005:DDF

[Tro05] R. L. Trope. Directors’ digital fiduciary duties. IEEE Security & Privacy, 3(1):78– 82, January/February 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/30310/01392707.pdf;
http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
30310&arnumber=1392707&count=
17&index=14.

Trope:2006:ITM

[Tro06] Roland L. Trope. Immaterial transfers with material consequences. IEEE Security & Privacy, 4(5):74–78, September/ October 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Troutman:2008:VMM

[Tro08] Justin Troutman. The virtues of mature and minimalist cryptography. IEEE Security & Privacy, 6(4):62–65, July/ August 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Tsang:2007:WCT

[Tsa07] Patrick P. Tsang. When cryptographers turn lead into gold. IEEE Security & Privacy, 5(2): 76–79, March/April 2007. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Tahaei:2023:EPD

[TVR23] Mohammad Tahaei, Kami Vaniea, and Awais Rashid. Embedding privacy into design through software developers: Challenges and solutions. IEEE Security & Privacy, 21 (1):49–57, January/February 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Tuma:2021:SPP

[TW21] Katja Tuma and Mathias Widman. Seven pain points of threat analysis and risk assessment in the automotive domain. IEEE Security & Privacy, 19(5):78–82, September/ October 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Trope:2008:DPH

[TWA08] Roland L. Trope, Monique Witt, and William J. Adams. Digital protection: Hardening the target. IEEE Security & Privacy, 6(5):77–81, September/October 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Turner:2015:BPO

[TWC^{+}15] Hamilton Turner, Jules White, Jaime A. Camelio, Christopher Williams, Brandon Amos, and Robert Parker. Bad parts: Are our manufacturing systems at risk of silent cyberattacks? IEEE Security & Privacy, 13 (3):40–47, May/June 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/03/
msp2015030040-abs.html.

Trestian:2021:PTC

[TXL^{+}21] Ramona Trestian, Guodong Xie, Pintu Lohar, Edoardo Celeste, Malika Bendechache, Rob Brennan, Evgeniia Jayasekera, Regina Connolly, and Irina Tal. Privacy in a time of COVID-19: How concerned are you? IEEE Security & Privacy, 19(5):26–35, September/ October 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Tziakouris:2018:CFC

[Tzi18] Giannis Tziakouris. Cryptocurrencies — a forensic challenge or opportunity for law enforcement? An INTER-POL perspective. IEEE Security & Privacy, 16(4):92– 94, July/August 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/04/
msp2018040092-abs.html.

Valasek:2012:PCE

[Val12]  Chris Valasek.  Primitive-

chaining exploits: a real-world example. IEEE Security & Privacy, 10(4):82–84, July/ August 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

vanOorschot:2019:SSS

[van19] P. C. van Oorschot. Software security and systematizing knowledge. IEEE Security & Privacy, 17(3):4–6, May/ June 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

vanOorschot:2020:BST

[van20] P. C. van Oorschot. Blockchains and stealth tactics for teaching security. IEEE Security & Privacy, 18(5):3–5, September/October 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Varadharajan:2009:HNT

[Var09] Vijay Varadharajan. On the horizon: a note on trustenhanced security. IEEE Security & Privacy, 7(3):57–59, May/June 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Varadharajan:2010:IF

[Var10] Vijay Varadharajan. Internet filtering. IEEE Security & Privacy, 8(4):62–65, July/August 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Vaudenay:2007:PT

[Vau07] Serge Vaudenay. E-passport threats. IEEE Security & Privacy, 5(6):61–64, November/ December 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Verissimo:2013:BWY

[VB13] Paulo Esteves Verissimo and Alysson Bessani. E-biobanking: What have you done to my cell samples? IEEE Security & Privacy, 11(6):62–65, November/December 2013. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Vaidya:2004:PPD

[VC04] Jaideep Vaidya and Chris Clifton. Privacy-preserving data mining: Why, how, and when. IEEE Security & Privacy, 2(6):19–27, November/ December 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/06/j6019.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/06/j6019.
pdf.

Viecco:2008:SSL

[VC08] Camilo Viecco and Jean Camp.

Secure systems: a life or death InfoSec subversion. IEEE Security & Privacy, 6(5):74–76, September/October 2008. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Vassilev:2007:PBW

[VdCA07] Apostol T. Vassilev, Bertrand du Castel, and Asad M. Ali. Personal brokerage of Web service access. IEEE Security & Privacy, 5(5):24–31, September/October 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

vandeGraaf:2017:LTT

[vdG17] Jeroen van de Graaf. Longterm threats to ballot privacy. IEEE Security & Privacy, 15(3):40–47, May/June 2017. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/03/
                                        msp2017030040-abs.html.

vanderKouwe:2020:BFU

[vdKHA^{+}20] E. van der Kouwe, G. Heiser, D. Andriesse, H. Bos, and C. Giuffrida. Benchmarking flaws undermine security research. IEEE Security & Privacy, 18(3):48–57, May/June 2020. ISSN 1558-4046.

Viega:2006:WAS

[VE06] John Viega and Jeremy Epstein. Why applying standards to Web services is not enough. IEEE Security & Privacy, 4(4): 25–31, July/August 2006. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Verdon:2006:SPS

[Ver06] Denis Verdon. Security policies and the software developer. IEEE Security & Privacy, 4(4): 42–49, July/August 2006. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Verbij:2016:NMS

[Ver16] Ruud Verbij. The navigation metaphor in security economics. IEEE Security & Privacy, 14(3):14–21, May/June 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Vetillard:2021:SCI

[Vét21] Eric Vétillard. Security certification: Is it box ticking? IEEE Security & Privacy, 19 (4):70–74, July/August 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Viega:2011:RC

[Vie11a] John Viega. Reality check.

IEEE Security & Privacy, 9(1): 3–4, January/February 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Viega:2011:TYT

[Vie11b] John Viega. Ten years of trustworthy computing: Lessons learned. IEEE Security & Privacy, 9(5):3–4, September/ October 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Viega:2012:CSP

[Vie12a] John Viega. Cloud security:

Not a problem. IEEE Security & Privacy, 10(4):3, July/ August 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Viega:2012:GB

[Vie12b] John Viega. Giving back. IEEE Security & Privacy, 10(6):3– 4, November/December 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Viega:2012:HA

[Vie12c] John Viega. Happy anniversary! IEEE Security & Privacy, 10(1):3–4, January/ February 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Viega:2012:TYH

[Vie12d] John Viega. Ten years on, how are we doing? (spoiler alert: We have no clue). IEEE Security & Privacy, 10(6):13– 16, November/December 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Vigna:2011:ICF

[Vig11] Giovanni Vigna. The 2010 International Capture the Flag Competition. IEEE Security & Privacy, 9(1):12–14, January/ February 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Vilardo:2004:OIS

[Vil04]  Mark F. Vilardo.  On-

line impersonation in securities scams. IEEE Security & Privacy, 2(3):82–85, May/June 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/03/j3082.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/03/j3082.
pdf.

Visaggio:2010:SMV

[Vis10] Corrado Visaggio. Session management vulnerabilities in today’s Web. IEEE Security & Privacy, 8(5):48–56, September/October 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Verma:2015:SAE

[VKM^{+}15] R. Verma, M. Kantarcioglu, D. Marchette, E. Leiss, and T. Solorio. Security analytics: Essential data analytics knowledge for cybersecurity professionals and students. IEEE Security & Privacy, 13(6):60– 65, November/December 2015. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Vidas:2018:CGS

[VLOS18] Timothy Vidas, Per Larsen, Hamed Okhravi, and Ahmad-Reza Sadeghi. Changing the game of software security. IEEE Security & Privacy, 16 (2):10–11, March/April 2018. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/02/
msp2018020010.html.

Verdon:2004:RAS

[VM04] Denis Verdon and Gary Mc-Graw. Risk analysis in software design. IEEE Security & Privacy, 2(4):79–84, July/August 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/04/j4079.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/04/j4079.
pdf.

Viega:2010:GEI

[VM10] John Viega and Bret Michael. Guest Editors’ introduction: Mobile device security. IEEE Security & Privacy, 8(2):11–12, March/April 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Vaidya:2021:SSE

[VM21] B. Vaidya and H. T. Mouftah. Security for shared electric and automated mobility services in smart cities. IEEE Security & Privacy, 19(1):24–33, January/ February 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Verissimo:2006:ITM

[VNC^{+}06] Paulo E. Veríssimo, Nuno F. Neves, Christian Cachin, Jonathan Poritz, David Powell, Yves Deswarte, Robert Stroud, and Ian Welch. Intrusion-tolerant middleware: The road to automatic security. IEEE Security & Privacy, 4(4):54–62, July/ August 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

vanOorschot:2020:USP

[vO20] P. C. van Oorschot. Untangling security and privacy. IEEE Security & Privacy, 18(2):4–6, March/April 2020. ISSN 1558-4046.

vanOorschot:2021:CSB

[vO21a] Paul C. van Oorschot. Coevolution of security’s body of knowledge and curricula. IEEE Security & Privacy, 19(5):83– 89, September/October 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

vanOorschot:2021:TUU

[vO21b] Paul C. van Oorschot. Toward unseating the unsafe C programming language. IEEE Security & Privacy, 19(2):4–6, March/April 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

vanOorschot:2022:SAS

[vO22a] Paul C. van Oorschot. Security as an artificial science, system administration, and tools. IEEE Security & Privacy, 20(6):74–78, November/December 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

vanOorschot:2022:VSS

[vO22b] Paul C. van Oorschot. A view of security as 20 subject areas in four themes. IEEE Security & Privacy, 20(1):102– 108, January/February 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

vanOorschot:2023:MEMb

[vO23a] Paul C. van Oorschot. Memory errors and memory safety: a look at Java and Rust. IEEE Security & Privacy, 21(3):62– 68, May/June 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

vanOorschot:2023:MEMa

[vO23b] Paul C. van Oorschot. Memory errors and memory safety: C as a case study. IEEE Security & Privacy, 21(2):70–76, March/April 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

vanOorschot:2019:ITS

[vS19] P. C. van Oorschot and S. W. Smith. The Internet of Things: Security challenges. IEEE Security & Privacy, 17(5):7– 9, September/October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Viega:2012:SED

[VT12] John Viega and Hugh Thompson. The state of embeddeddevice security (spoiler alert: It’s bad). IEEE Security & Privacy, 10(5):68–70, September/ October 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Vasserman:2012:SIM

[VVSL12a] Eugene Y. Vasserman, Krishna K. Venkatasubramanian, Oleg Sokolsky, and Insup Lee. Security and interoperablemedical-device systems, Part 2: Failures, consequences, and classification. IEEE Security & Privacy, 10(6):70–73, November/December 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Venkatasubramanian:2012:SIM

[VVSL12b] Krishna K. Venkatasubramanian, Eugene Y. Vasserman, Oleg Sokolsky, and Insup Lee. Security and interoperablemedical-device systems, part 1. IEEE Security & Privacy, 10 (5):61–63, September/October 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Villamor:2011:HUD

[VVYY11] Antonio Manuel Fernandez Villamor, Antonio Manuel Fernandez Villamor, Juan Yelmo, and Juan C. Yelmo. Helping users deal with digital threats: The Online User Supervision Architecture. IEEE Security & Privacy, 9(6):29–35, November/December 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Valente:2019:SSA

[VWC19] J. Valente, M. A. Wynn, and A. A. Cardenas. Stealing, spying, and abusing: Consequences of attacks on Internet of Things devices. IEEE Security & Privacy, 17(5):10– 21, September/October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

vanWyk:2005:BGB

[vWM05] Kenneth R. van Wyk and Gary McGraw. Bridging the gap between software development and information security. IEEE Security & Privacy, 3(5):75–79, September/October 2005. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

vanWyk:2006:EFS

[vWS06] Kenneth R. van Wyk and John Steven. Essential factors for successful software security awareness training. IEEE Security & Privacy, 4(5):80–83, September/October 2006. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Varian:2005:DDC

[VWW05] H. Varian, F. Wallenberg, and G. Woroch. The demographics of the do-not-call list [security of data]. IEEE Security & Privacy, 3 (1):34–39, January/February 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/30310/01392697.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        30310&arnumber=1392697&count=
                                        17&index=6.

vanderLinden:2019:BWY

[vZH^{+}19] D. van der Linden, A. Zamansky, I. Hadar, B. Craggs, and A. Rashid. Buddy’s wearable is not your buddy: Privacy implications of pet wearables. IEEE Security & Privacy, 17(3):28–39, May/June 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Woody:2007:COS

[WA07] Carol Woody and Christopher Alberts. Considering operational security risk during system development. IEEE Security & Privacy, 5(1):30–35, January/February 2007. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Weis:2011:SDS

[WAF11] Joel Weis and Jim Alves-Foss.

Securing database as a service: Issues and compromises. IEEE Security & Privacy, 9(6):49– 55, November/December 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Waidner:2016:HSP

[Wai16] Michael Waidner. HbbTV security and privacy: Issues and challenges. IEEE Security & Privacy, 14(3):61–67, May/June 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Wang:2018:ISP

[Wan18] Yang Wang. Inclusive security and privacy. IEEE Security & Privacy, 16(4):82– 87, July/August 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/04/
msp2018040082-abs.html.

Wash:2012:FS

[Was12] Rick Wash. Folk security. IEEE Security & Privacy, 10(6):88– 90, November/December 2012.

CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Wayner:2004:PCC

[Way04] Peter Wayner. The power of candy-coated bits. IEEE Security & Privacy, 2(2): 69–72, March/April 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2004/02/j2069abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/02/j2069.
pdf.

Wayman:2008:BIM

[Way08] James L. Wayman. Biometrics in identity management systems. IEEE Security & Privacy, 6(2):30–37, March/April 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Woods:2022:IRL

[WB22] Daniel W. Woods and Rainer Böhme. Incident response as a lawyers service. IEEE Security & Privacy, 20(2):68–74, March/April 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Wijesekera:2018:DRM

[WBT^{+}18] Primal Wijesekera, Arjun Baokar, Lynn Tsai, Joel Reardon, Serge Egelman, David Wagner, and Konstantin Beznosov. Dynamically regulating mobile application permissions. IEEE Security & Privacy, 16(1):64–71, January/February 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/01/
msp2018010064.html.

White:2004:AUF

[WC04] Greg White and Art Conklin.

The appropriate use of forceon-force cyberexercises. IEEE Security & Privacy, 2(4):33–37, July/August 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/04/j4033.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/04/j4033.
pdf.

Weber:2004:LED

[WCV^{+}04] Daniel Weber, Jean Camp, Tom Van Vleck, Bob Bruen, and James Whittaker. Letters to the Editors: Digital rights management; change the game?; no clear answers. IEEE Security & Privacy, 2(1):7–9, January/February 2004. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/comp/mags/sp/2004/01/
j1007.pdf.

Wiefling:2021:VYH

[WDL21] Stephan Wiefling, Markus Dürmuth, and Luigi Lo Iacono. Verify it’s you: How users perceive risk-based authentication. IEEE Security & Privacy, 19(6):47–57, November/ December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Weir:2023:DYS

[WDP23] Charles Weir, Anna Dyson, and Dan Prince. Do you speak cyber? Talking security with developers of health systems and devices. IEEE Security & Privacy, 21(1):27–36, January/ February 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Webb:2005:BSS

[Web05] Katharine W. Webb. Biometric security solutions. IEEE Security & Privacy, 3(5):7, September/October 2005. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/sp/2005/05/
                                        j5007.pdf.

Weis:2004:RPW

[Wei04]  Stephen A. Weis.  RFID

privacy workshop: Concerns, consensus, and questions. IEEE Security & Privacy, 2(2):48–50, March/April 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/sp/2004/02/j2048abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2004/02/j2048.
                                        pdf.

Weiss:2005:C

[Wei05] S. A. Weiss. Crypto 2004.

IEEE Security & Privacy, 3 (2):11–13, March/April 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/30742/01423953.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
30742&arnumber=1423953&count=
15&index=1.

Weis:2006:PET

[Wei06] Stephen A. Weis. Privacy enhancing technologies. IEEE Security & Privacy, 4(5):59, September/October 2006. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://csdl.computer.
org/comp/mags/sp/2006/05/
j5059.pdf.

Whittaker:2006:HTA

[WF06] James A. Whittaker and Richard Ford. How to think about security. IEEE Security & Privacy, 4(2):68–71, March/ April 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Whalen:2011:MDL

[Wha11a] Tara Whalen. Mobile devices and location privacy: Where do we go from here? IEEE Security & Privacy, 9(6):61– 62, November/December 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Whalen:2011:SIP

[Wha11b] Tara Whalen. Security as if people mattered. IEEE Security & Privacy, 9(4):64–67, July/August 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Whalen:2012:TPR

[Wha12] Tara Whalen. This time, it’s personal: Recent discussions on concepts of personal information. IEEE Security & Privacy, 10(1):77–79, January/ February 2012. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Williams:2009:STD

[WHD^{+}09] Daniel Williams, Wei Hu, Jack W. Davidson, Jason D. Hiser, John C. Knight, and Anh Nguyen-Tuong. Security through diversity: Leveraging virtual machine technology. IEEE Security & Privacy, 7(1): 26–33, January/February 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Willems:2007:TAD

[WHF07] Carsten Willems, Thorsten Holz, and Felix Freiling. Toward automated dynamic malware analysis using CWSandbox. IEEE Security & Privacy, 5(2):32–39, March/April 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Whittaker:2003:WSA

[Whi03] James Whittaker. Why secure applications are difficult to write. IEEE Security & Privacy, 1(2):81– 83, March/April 2003. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://dlib.computer.
org/sp/books/sp2003/pdf/
j2081.pdf;   http://www.
computer.org/security/j2081abs.
htm.

Whitty:2015:MMF

[Whi15] Monica T. Whitty. Massmarketing fraud: A growing concern. IEEE Security & Privacy, 13(4):84– 87, July/August 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/04/
msp2015040084-abs.html.

Williams:2016:AIT

[Wil16a] Julian Williams. Action, inaction, trust, and Cybersecurity’s common property problem. IEEE Security & Privacy, 14(1):82–86, January/ February 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Williams:2016:EIR

[Wil16b] Julian Williams. Economic impacts of rules- versus riskbased cybersecurity regulations for critical infrastructure providers. IEEE Security & Privacy, 14(3):52–60, May/ June 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Williams:2019:SLC

[Wil19] L. Williams. Science leaves clues. IEEE Security & Privacy, 17(5):4–6, September/ October 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Williams:2021:PWL

[Wil21] Laurie Williams. The people who live in glass houses are happy the stones weren’t thrown at them [from the Editors]. IEEE Security & Privacy, 19(3):4–7, May/June 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Williams:2022:TTH

[Wil22] Laurie Williams. Trusting trust: Humans in the software supply chain loop. IEEE Security & Privacy, 20(5):7– 10, September/October 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Williams:2024:NSS

[Wil24] Laurie Williams. Narrowing the software supply chain attack vectors: The SSDF is wonderful but not enough. IEEE Security & Privacy, 22(2):4–7, March/April 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Wing:2003:CAL

[Win03] Jeannette M. Wing. A call to action: Look beyond the horizon. IEEE Security & Privacy, 1(6):62–67, November/ December 2003. CODEN ????

ISSN 1540-7993 (print), 1558-4046 (electronic). URL http:

//csdl.computer.org/comp/
mags/sp/2003/06/j6062abs.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/06/j6062.
htm;   http://csdl.computer.
org/dl/mags/sp/2003/06/j6062.
pdf.

Wisniewski:2018:PPA

[Wis18] Pamela Wisniewski. The privacy paradox of adolescent online safety: A matter of risk prevention or risk resilience? IEEE Security & Privacy, 16 (2):86–90, March/April 2018. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2018/02/
msp2018020086-abs.html.

Walsh:2005:CSV

[WK05] T. J. Walsh and D. R. Kuhn.

Challenges in securing voice over IP. IEEE Security & Privacy, 3(3):44–49, May/June 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/31002/01439501.pdf;
 http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
31002&arnumber=1439501&count=
20&index=8.

Weigold:2008:RCA

[WKB08] Thomas Weigold, Thorsten Kramp, and Michael Baentsch. Remote client authentication. IEEE Security & Privacy, 6(4):

36–43, July/August 2008. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Weissman:2011:LLB

[WL11] Clark Weissman and Timothy Levin. Lessons learned from building a high-assurance crypto gateway. IEEE Security & Privacy, 9(1):31–39, January/February 2011. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Wang:2018:CDO

[WL18] Cliff Wang and Zhuo Lu. Cyber deception: Overview and the road ahead. IEEE Security & Privacy, 16(2):80– 85, March/April 2018. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2018/02/
                                        msp2018020080-abs.html.

Woods:2020:DIF

[WM20] D. W. Woods and T. Moore. Does insurance have a future in governing cybersecurity? IEEE Security & Privacy, 18(1):21–27, January/ February 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Wang:2020:DCP

[WMD20] Q.-H. Wang, S. M. Miller, and R. H. Deng. Driving cybersecurity policy insights from information on the Internet. IEEE Security & Privacy, 18(6):42– 50, November/December 2020.

ISSN 1540-7993 (print), 1558-4046 (electronic).

Wu:2023:CSD

[WMDX23] Jingwei Wu, Wanxi Mao, Yuxin Deng, and Jinghong Xu. Child safety and data privacy in smart homes with speakers: Comparison of privacy policies in the United States and China. IEEE Security & Privacy, 21 (5):6–17, September/October 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Williams:2010:PPN

[WMS10] Laurie Williams, Andrew Meneely, and Grant Shipley. Protection poker: The new software security “Game”. IEEE Security & Privacy, 8(3):14–20, May/June 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Weir:2022:ESS

[WMW22] Charles Weir, Sammy Migues, and Laurie Williams. Exploring the shift in security responsibility. IEEE Security & Privacy, 20(6):8–17, November/ December 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Wijngaards:2009:SDE

[WO09] Wouter C. A. Wijngaards and Benno J. Overeinder. Securing DNS: Extending DNS servers with a DNSSEC validator. IEEE Security & Privacy, 7(5):36–43, September/ October 2009. CODEN ????

ISSN 1540-7993 (print), 1558-4046 (electronic).

Wood:2023:HTS

[Woo23] Christopher A. Wood. Hot topics in security and privacy standardization at the IETF and beyond. IEEE Security & Privacy, 21(2):57–62, March/ April 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Weiss:2017:CEA

[WTML17] Richard Weiss, Franklyn Turbak, Jens Mache, and Michael E. Locasto. Cybersecurity education and assessment in EDURange. IEEE Security & Privacy, 15(3):90– 95, May/June 2017. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/sp/2017/03/
                                        msp2017030090-abs.html.

Wu:2010:KSP

[Wu10] Hui Jun Wu. Kernel service protection for client security. IEEE Security & Privacy, 8 (5):14–19, September/October 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

White:2010:CNH

[WWH10] Gregory B. White, Dwayne Williams, and Keith Harrison. The CyberPatriot National High School Cyber Defense Competition. IEEE Security & Privacy, 8(5):59–61, September/October 2010. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Wang:2006:CFC

[WYA06] WenJie Wang, Yufei Yuan, and Norm Archer. A contextual framework for combating identity theft. IEEE Security & Privacy, 4(2):30–38, March/ April 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Winkler:2018:PCD

[WZE18] S. Winkler, S. Zeadally, and K. Evans. Privacy and civilian drone use: The need for further regulation. IEEE Security & Privacy, 16(5):72–80, September/October 2018. ISSN 1540-7993 (print), 1558-4046 (electronic).

Xu:2016:ASP

[XL16] Hui Xu and Michael R. Lyu.

Assessing the security properties of software obfuscation. IEEE Security & Privacy, 14(5):80–83, September/October 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/05/
msp2016050080-abs.html.

Yannakogeorgos:2016:DCD

[Yan16] Panayotis A. Yannakogeorgos.

Designing cybersecurity into defense systems: An information economics approach. IEEE Security & Privacy, 14 (3):44–51, May/June 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Yavuz:2024:MLH

[Yav24] Can Yavuz. A multidisciplinary look at history and future of deepfake with Gartner Hype Cycle. IEEE Security & Privacy, 22(3):50–61, May/ June 2024. ISSN 1540-7993 (print), 1558-4046 (electronic).

Yang:2021:OCS

[YAZ21] Zheng Yang, Sridhar Adepu, and Jianying Zhou. Opportunities and challenges in securing critical infrastructures through cryptography. IEEE Security & Privacy, 19(5):57–65, September/October 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Yasinsac:2005:CAE

[YB05] A. Yasinsac and M. Burmester. Centers of academic excellence: a case study. IEEE Security & Privacy, 3(1):62– 65, January/February 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8013/30310/01392703.pdf;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        30310&arnumber=1392703&count=
                                        17&index=10.

Yasinsac:2008:DCR

[YB08] Alec Yasinsac and Matt Bishop. The dynamics of counting and recounting votes. IEEE Security & Privacy, 6(3): 22–29, May/June 2008. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Yao:2021:AFQ

[YB21] Danfeng Daphne Yao and Terry Benzel. ACSAC 2020: Furthering the quest to tackle hard problems and find practical solutions. IEEE Security & Privacy, 19(6):23–24, November/December 2021. ISSN 1540-7993 (print), 1558-4046 (electronic).

Yan:2004:PMS

[YBAG04] Jeff Yan, Alan Blackwell, Ross Anderson, and Alasdair Grant. Password memorability and security: Empirical results. IEEE Security & Privacy, 2(5):25–31, September/October 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/05/j5025.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/05/j5025.
pdf.

Yagemann:2021:MLS

[YCU^{+}21] Carter Yagemann, Pak Ho Chung, Erkam Uzun, Sai Ragam, Brendan Saltaformaggio, and Wenke Lee. Modeling large-scale manipulation in open stock markets. IEEE Security & Privacy, 19(6):58– 65, November/December 2021.

ISSN 1540-7993 (print), 1558-4046 (electronic).

Yan:2009:PCA

[YE09] Jeff Yan and Ahmad Salah El Ahmad. Pixel-count attacks: CAPTCHA security: a case study. IEEE Security & Privacy, 7(4):22–28, July/August 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Yee:2004:ASU

[Yee04] Ka-Ping Yee. Aligning security and usability. IEEE Security & Privacy, 2(5):48–55, September/October 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/sp/2004/05/j5048.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/sp/2004/05/j5048.
                                        pdf.

Yener:2019:CED

[YG19] B. Yener and T. Gal. Cybersecurity in the era of data science: Examining new adversarial models. IEEE Security & Privacy, 17(6):46– 53, November/December 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Yasinsac:2013:HTT

[YI13] Alec Yasinsac and Cynthia Irvine. Help! Is there a trustworthy-systems doctor in the house? IEEE Security & Privacy, 11(1):73–77, January/ February 2013. CODEN ????

ISSN 1540-7993 (print), 1558-4046 (electronic).

Yang:2016:TCP

[YK16] Baijian Justin Yang and Brian Kirk. Try-CybSI: A platform for trying out cybersecurity. IEEE Security & Privacy, 14 (4):74–75, July/August 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL https://www.computer.
org/csdl/mags/sp/2016/04/
msp2016040074-abs.html.

Yan:2009:ICO

[YR09] Jeff Yan and Brian Randell.

An investigation of cheating in online games. IEEE Security & Privacy, 7(3):37–44, May/June 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Yao:2022:BDF

[YRX^{+}22] Danfeng Daphne Yao, Sazzadur Rahaman, Ya Xiao, Sharmin Afrose, Miles Frantz, Ke Tian, Na Meng, Cristina Cifuentes, Yang Zhao, Nicholas Allen, Nathan Keynes, Barton P. Miller, Elisa Heymann, Murat Kantarcioglu, and Fahad Shaon. Being the developers friend: Our experience developing a high-precision tool for secure coding. IEEE Security & Privacy, 20(6):43– 52, November/December 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Yan:2008:BTR

[YZA08] Wei Yan, Zheng Zhang, and Nirwan Ansari. Basic training: Revealing packed malware. IEEE Security & Privacy, 6(5):65–69, September/ October 2008. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Zanero:2009:ATW

[Zan09] Stefano Zanero. Attack trends: Wireless malware propagation: a reality check. IEEE Security & Privacy, 7(5):70–74, September/October 2009. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Zatko:2016:RRS

[Zat16] Sarah Zatko. Rethinking the role of security in undergraduate education. IEEE Security & Privacy, 14(2):73–78, March/April 2016. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Zhivich:2009:SSR

[ZC09] Michael Zhivich and Robert K. Cunningham. Secure systems: The real cost of software errors. IEEE Security & Privacy, 7(2):87–90, March/April 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Zdrnja:2009:ATM

[Zdr09] Bojan Zdrnja. Attack trends: Malicious JavaScript insertion through ARP poisoning attacks. IEEE Security & Privacy, 7(3):72–74, May/June 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Zurko:2023:USP

[ZH23] Mary Ellen Zurko and Julie Haney. Usable security and privacy for security and privacy workers. IEEE Security & Privacy, 21(1):8–10, January/ February 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Zhang:2020:EPC

[ZHC^{+}20] F. Zhang, W. He, R. Cheng, J. Kos, N. Hynes, N. Johnson, A. Juels, A. Miller, and D. Song. The Ekiden platform for confidentiality-preserving, trustworthy, and performant smart contracts. IEEE Security & Privacy, 18(3):17–27, May/ June 2020. ISSN 1558-4046.

Zahan:2023:OSP

[ZKH^{+}23] Nusrat Zahan, Parth Kanakiya, Brian Hambleton, Shohanuzzaman Shohan, and Laurie Williams. OpenSSF scorecard: On the path toward ecosystemwide automated security metrics. IEEE Security & Privacy, 21(6):76–88, November/ December 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Zhao:2020:HMS

[ZL20] L. Zhao and D. Lie. Is hardware more secure than software? IEEE Security & Privacy, 18(5):8–17, September/ October 2020. ISSN 1540-7993 (print), 1558-4046 (electronic).

Zahan:2023:SBM

[ZLT^{+}23] Nusrat Zahan, Elizabeth Lin, Mahzabin Tamanna, William Enck, and Laurie Williams. Software bills of materials are required. are we there yet? IEEE Security & Privacy, 21 (2):82–88, March/April 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Zhang:2022:PCL

[ZOC^{+}22] Penghui Zhang, Adam Oest, Haehyun Cho, Zhibo Sun, R. C. Johnson, Brad Wardman, Shaown Sarker, Alexandros Kapravelos, Tiffany Bao, Ruoyu Wang, Yan Shoshitaishvili, Adam Doupé, and Gail-Joon Ahn. CrawlPhish: Large-scale analysis of clientside cloaking techniques in phishing. IEEE Security & Privacy, 20(2):10–21, March/ April 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Zonouz:2014:DIC

[ZRM14] Saman Zonouz, Julian Rrushi, and Stephen McLaughlin. Detecting industrial control malware using automated PLC code analytics. IEEE Security & Privacy, 12(6):40– 47, November/December 2014. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2014/06/
msp2014060040-abs.html.

Zou:2019:BMM

[ZS19] Y. Zou and F. Schaub. Beyond mandatory: Making data breach notifications useful for consumers. IEEE Security & Privacy, 17(2):67–72, March/ April 2019. ISSN 1540-7993 (print), 1558-4046 (electronic).

Zurko:2022:DRU

[Zur22] Mary Ellen Zurko. Disinformation and reflections from usable security. IEEE Security & Privacy, 20(3):4–7, May/June 2022. ISSN 1540-7993 (print), 1558-4046 (electronic).

Zurko:2023:USA

[Zur23] Mary Ellen Zurko. Unusable security for attackers [from the Editors]. IEEE Security & Privacy, 21(6):4–7, November/ December 2023. ISSN 1540-7993 (print), 1558-4046 (electronic).

Zhou:2015:DGW

[ZYG15] Zongwei Zhou, Miao Yu, and Virgil D. Gligor. Dancing with giants: Wimpy kernels for ondemand I/O isolation. IEEE Security & Privacy, 13(2):38– 46, March/April 2015. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

URL http://www.computer.
org/csdl/mags/sp/2015/02/
msp2015020038-abs.html.